A Complete Bibliography of Publications in North American Journal of Fisheries

Management: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.03

Title word cross-reference

r [MBS90]. _{s} [MP99]. F_{1} [HB91, Snu92]. × [KBF99, MMD91, STS^{+}96]. W_{r} [NM91a, PP97b]. W_{s} [BM91, KH97b, QGB98].

10-Inch [Hof95]. 13 [MSR92]. 15 [VHW^{+}96].

230-V [Sch92]. 254-mm [HLMP99].

356-mm [LKD96, SMTW98].

67 [SC91].

8 [ZDK99]. 8-Inch [Hof95].

Abiotic [AAMB97]. Abnormally [MSR92]. Abrasions [HTMP94].

Absence [Uph93, Uph97]. Absolute [PB94b]. Abundance [AFW98, ADS^{+}99, BB99, BLW^{+}97, BMNB92, BBM96, BW94, BM93, CHCC95, CB96, CBD97, CBJ93, CMP99, CBS93, DVG95, DFW91, DB94, FW99a, FH92, FS97, HSP^{+}95, HHY96, HWE96, Hou95, IT98, Jes94, Jow92, KHR98, LCGJ99, LCSA96a, MWL95, MB91, McH90, MM91, MHSH96, MP97, MNN91, MHW98, OC94, PJN97, Sal93, SM98, SBJ^{+}99, SG93, TS96, TB99, VPB95, VS95, VD95, WH97, YFM98, ZM98]. Abundances [VB95]. AC [HSCM96, HSS99, HW94, Sch92]. Access [Loc97, PHJ^{+}97, Wri92]. Acclimation [DW91]. Accounting [CH98]. Accuracy [BC94, FMC97, HMG92, IHK99, JQV98, LMR91, WSL96, WD95, WvdABD93]. Accurate [UHL97]. Acetic [PLL^{+}95]. Acid [BWG93, GZB98, PLL^{+}95]. Acid-Stressed [BWG93]. Acidified [CDM^{+}98]. Acoustic [CSCM99]. Act [Ant90, MW90]. Actions [Aas95]. Active [DVG95, Jes90a]. Activity [TH90]. Adaptation [SNS99]. Adaptive [WGR93, Wil99]. Additional [JW98]. adina [PBTS95]. Adipose [TB97a]. Adirondack [FW92, VKS93]. Adjustable [DMP91]. Adsorption [CKT96, DGDG91]. Adult [BMS97, CCQR96, CV94, DO98, DW95b, FHMT97, Fry98, GB95, HTMP94, HN95, HJ93, KB97, MPP94, RB95, RB99, SLM^{+}97, SFZ98, SES94, TG96, WH96]. Adults [BR91]. Advanced [BM99]. Advanced-Fingerling [BM99]. Advisories [CKB96]. Advisory [MB97]. Aerial [Hil97, JQV98, Kel93]. Affect [DGE^{+}97, MLSH^{+}99, WPE96]. Affecting [Fie92a, FT93, HBN^{+}98, HCFG96, MRK98, MW95, MB91, PB95, SO96, STS^{+}96, WH97]. Affects [HA95, PH96]. after [Ant90, BJ98, CCQR96, CHR93, DW97b, FW92, KGFR97, OS96, PP97a, SNF94, SGWN94, WB93, WP98a, YHN99]. Age [BF97, BDG99, BW94, CMP99, CS96c, DWS97, FW99a, FS97, Goo96a, Goo96b, HBJ95, HM91, HTM96, INJ97, KGW93, MRL93, MM91, Mic98a, MG99, MH94, MKK97, MKKG98, MWF94, NF99, PWW94, PJN97, PD91, RS97, RQ95, RS96b, SRL97, STS^{+}96, Tip91, TB99, WH96, WvdABD93]. Age-0 [BF97, BW94, CMP99, FW99a, FS97, HBJ95, INJ97, MRL93, MM91, Mic98a, MG99, MH94, PJN97, RS97, RS96b, SRL97, STS^{+}96, TB99]. Age-1 [SRL97]. Age-Specific [WH96]. Age-Structured [HTM96]. Aged [SMH92]. Ages [KH97a, PD91]. Aging [FMC97, LMR91, MRS98]. Aided [UHL97]. Air [BFJ90, SJ96]. Air-Lift [BFJ90]. Airboat [CCT99]. Aircraft [HP97]. Alabama [SGWN94, AGS^{+}99, DMS96, DMR95, DTR92, HD96, MBFD98, MS98, MOAS98, McH90, OAM98, RD91, SMMC96, SMTW98]. Alaska [BA93b, BFM99, Cla94, DO98, FB96, HOCSN99, HSS^{+}94, HG94, HJ99, JQV98, Leo97, OC94, RR92, SCKF98, TMJ^{+}94, ZG90, ZM98]. Alaskan [MW90]. Albemarle [Dor95, HWE96]. Alewife [BROV98]. Alewives [DRG^{+}92, Jes90b, Jes94, Rod95, RDT^{+}93, RDM^{+}96]. Algorithm [Han95]. Allocating [KM98]. Allocation [HSS^{+}94, Leg98]. Allocations [Edw91]. Allocative [GH94b]. Allopatric [Pow94]. Allowable [HCFG96, MR95]. Allozyme [GKAM94, JP92, VWT99]. along [CRS96, GH93, Jac95]. Altering [GSP93]. Alternative [Cri96, EGP97, HM93, Lan91, MHW98, WYL^{+}99]. Alternatives [JFW96].

America [BN95, DS90]. American [BRK91, BBM96, Cam92, DG93, HK97, HOCSN99, HBM91, HTB94, HWE96, JD92, JR95, KO93, KB97, Par92, RQ95, RBJ97]. Ammonia [Ber92]. Among [VD95, BNNS95, DCPP97, EHB^{+}93, FSBB97, GKAM94, MK91, MPL99, MM91, OL96, Pal95b, PJN97, TGM98, WD94]. Among- [VD95]. Amount [MT91]. Anadromous [BFM99, HTMP94, JNT92, Neg99a, PT93a, Wri92]. Anal [WvdABD93, WG95]. Analyses [Cam97, HA97, MBFD98, SF93]. Analysis [All97, BF93, CR97, EF90, GH94b, GHG97, HG93, HSS^{+}94, HG94, Jac96, JMBK92, JBHW95, JW98, Max99, MRS98, MMD91, ORM92, PMW98, Syl94, TW92, TBR96, WKG98]. Anchor [BG90, BSN95, BD99, Clu96, HARS^{+}99, MFG94, Muo92a, PLB96, PT93b, SC91, SBH96, TH95, WDM91]. Anchor-Tagged [PLB96]. Anchovy [JT93]. Anesthesia [JL98, SES94]. Anesthetic [AMC97, GZB98, PLL^{+}95]. Anesthetics [Pea98]. Angled [DB91, DMS^{+}94, PH94b]. Angler [BHY^{+}97, BM93, BMMF99, CW94, Col91b, DGE^{+}97, Fis97, JL92, Loc97, Mal95, Mil91, NA94, PLB96, PJM95, PIM95, PB94b, PT98, Qui92, RP99, RTB99, RW92, SK95, SS92, SGWN94, TR92, TBF93, TBR96]. Angler-Caught [SGWN94]. Anglers [CU90, CKB96, EHB^{+}93, ESB^{+}96, FT93, GP93, GKM93, HD92, KH98, MS92, Spe93, WD94, WRD98]. Angling [Bac95, BDBN97, BO98, BS91, Ded96, GITS99, MK91, Mor93, NA94, PTK^{+}97, RS97, RIT97, TMBH93, TR92]. Anna [Gra92]. Annapolis [SD91]. Annual [Wat92]. Annulus [CM99a, JDCS97]. Anodes [MT92]. Antibiotics [HBM91]. Antimycin [Gre91, RLM94, TB96]. Antimycin-Impregnated [RLM94]. Appalachian [CR97, DJO97, FD95, HSCM96]. Apparatus [FB93, SM96]. Applicability [SG97a]. Application [BM96, CCS^{+}99, CDM^{+}98, Fis97, GB99a, Han95, HSS^{+}94, HJ99, KWRH94, MR95, Ste99, WP98a, ZA97]. Applications [Mal95, MBS90]. Applied [FB96, Neg95, SBH96]. Applying [Joh95]. Approach [BSMM90, DW92, KEC93, LF99, SSA99, WD94]. Approaches [WYL^{+}99]. Aquatic [ANN91, BMNB92, BMNB93, BTT^{+}93, Kir92, MBK^{+}91, MRP99, ML94, NM91b, Rin96]. Aqueducts [Mue96]. Arctic [AKG93, BD99, Cla94, DW95b, DW97b, KKS95, MDI^{+}96, Roa99]. Area [BF98, Hil97]. Area-under-the-Curve [Hil97]. Areal [FB93]. Areas [SP96, WS91]. Argentina [BA93a]. Arguments [Edw91]. Arizona [BM94, MRP99, Mue96]. Arkansas [SJ98a]. Artificial [Ave96, BM97, FN92, FE95, Gra92, Gre95, LW91, MS96, MMK97, NA92b, PT93a, SB96, SS97c, SF96, TR98, WLJ91]. Artificially [MKG^{+}93]. Artisanal [Smi98]. Assemblages [Aad93, Fag98, JBH^{+}99, MR91, SL95]. Assembly [OB95]. Assess [MBHC93, MM94a, PP97b, SHS90]. Assessing [ANN91, Bac95, DVG95, PPQ94]. Assessment [BA93a, BLW^{+}97, BPL93, CLA95, DS90, JS95, KL91, LR97, MFP98, ML92, NM91a, NJ96, PR93, SCKF98, SMMC96, SJ98a, WBYO97, WWSF93a]. Assessments [JN95, Sta92, Ste99]. Assist [HMG^{+}96]. Associated [CBD97, CBC95, HCF^{+}96, KMHS95, SFZ98]. Associations [JJ98, Wic95]. Assurance [MRS98]. Asymmetry [Cam97, HA95, HA97]. Atlantic [AGG96, CRS96, SF96, AGG96, BJ98, BGJ^{+}98, BCCEJ96, CBJ93, CS97b, CSCM99, DS91, Han95, HN93, HJ93, JMD96, Lan91, LT98, Loc98, McM95, Moh97, OLMS94, OL96, PH94a, PH90, Pow94, RR93, Rot94, Sal93, SG93, TBR96, WL98]. Attachment [PMBM97]. Attacks [AAMB97]. Attempt [Jes90a]. Attitudes [GP93, HS93, WRD98]. Attraction [BKM99, WLJ91]. Attractors [RMV96]. Attributes [MKK97, TR92]. Attrition [TP94]. Auger [KWRH94]. Augmentation [CBB98]. Australia [SW97]. Automated [BC94]. Autumn [GR97]. Availability [Ber94, Lan91, Mic98a, PP97b, SW96, TLW96, ZS95]. Averaging [New97]. Avoid [DRG^{+}92]. Avoidance [Hog99]. Axial [WN99].

Back [KBF99, KH97a, SMH92]. Back-Calculation [KBF99, KH97a, SMH92]. Background [DW91]. Background-Color [DW91]. Backpack [HC94, KGFR97, OAM98]. Backwater [GPB95, KJS95]. Bacterial [SGWN94]. Bag [GSP93, MK97, WO91]. Bags [PK94]. Bait [MCC97, PT93a, RLM94, Sch96a]. Bait-Caught [Sch96a]. Baited [Stu91, TW97, TB97b]. Baitfish [FT93, Kir98, SMH95]. Baits [Muo92b, SB96]. Balance [TBF93]. Bank [BBM96]. Bar [Clu96, PC99]. Barb [MH97a, MMD^{+}93]. Barbed [SS97b, TW97, TB97b]. Barbless [TW97, TB97b]. Barkley [BMM91, BMB^{+}99]. Barrier [MSG99, PMN99, Swi99, VB95]. Barriers [TR98]. Based [BHY^{+}97, BF98, CR97, HM93, Mir93, MFP91, MKKR99, RS97, SC92, SJ98b, TR92, Uph97, WD95, WBPQ92, Wil99, ZM98, UHL97]. Basibranchial [LGS96]. Basin [BBW94b, GLLB97, JDL^{+}98, TB99, BF93, GHS^{+}97, JD92, PWW94, SF93, SSUL97, TLR97]. Basin-Scale [GLLB97]. Basins [RLT97, TLR97]. Basinwide [DJO97]. Bass [AGS^{+}99, BM94, BN95, BWF97, BMNB92, BO98, BJ98, BM91, BGS90, BS91, BWW95b, BKM99, BW94, BMM91, BKSM91a, BKSM91b, BM93, BM99, BMB^{+}99, BMMF99, Cob92, CMS^{+}99, CN94, DW92, DMR95, Dol91, DFW91, Dor95, DGE^{+}97, DG93, DTR92, ESW98, EDGR97, GPB95, GH94a, GSMG95, GC90, GW90, GW91, GBTD99, HM95, HGH97, HARS^{+}99, HW94, Hof95, INJ97, JN95, JL98, JNT92, JR98, KBM90, KWW90, KSS93, KWM93, KH95, LS90, Lee92, LW91, LKD96, MRL93, MWL95, MB98, Mar95, McH90, MD93, MC99, McN95, MM94b, MH94, MHSH96, MP97, MNN91, MMD91, Muo92a, Muo94, NNR99, Nel98, Nov90, OWSM99, PH96, PJM95, PTK^{+}97, PJN97, PVP99, RIO^{+}97, RR99, RS96a, RS97, RIT97, RB95, RB99, RM90, SDBF99, SB99, SMH92, SS97c, She91, SMTW98, SBH96]. Bass [SGWN94, SF96, SH92b, TSP93, TGM98, TGF92, UHP99, Uph93, Uph97, WDM91, WV96, WYL^{+}99, WN99, WB92b, WN97, WvdABD93, WRDY91, WGB96, WD94, WSOM98, WRD98, WF97, WW93, ZS95, ZHR99, vdAE90]. Bass-Club [Dol91]. Bass-Yellow [GW91]. Basses [SA98]. Bathymetric [JH93]. Bay [HL99a, MT91, Tay92, WPTW93, ADS^{+}99, CBJ93, FBR^{+}94, GH94b, Gre95, HN93, JNT92, JMBK92, KJ92, NN91, RR99, Uph97]. Bayesian [New97]. Bays [FW99a]. Beach [WW93]. Beaches [Leo97]. Bead [NF93]. Beam [DO98]. Beaver [Ell92, JRH92, WHR90]. Bed [JL92]. Beds [BAR^{+}92]. before [FW92]. Behavior [BO98, CCQR96, HM92, JJ97, KB97, Sor94]. Behavioral [MPL99]. Behaviors [GP93]. below [MRP99, SBT93, WB93]. Benchmark [Pow96]. Benefit [CW94, HSS^{+}94]. Benefit-Cost [HSS^{+}94]. Benefits [EM93, JBHW95, Pin94, PR93, SS97c]. Benign [OB93]. Benthic [KHP98, SH92a, WP98b]. Bering [CHGF98, PMW98, ZG90]. between [AFW98, AWR97, BBH97, BSN95, CRMH94, DMS96, EGP97, Fie92b, GP99, HM95, JDCS97, Jes94, LSG^{+}96, MS98, MC98, MLP95, MP97, MWC92, SS99, SLJ^{+}97, WW94, WP98a, WRD98, WHR90, ZS97, ZS95]. Beyond [Mar99b]. Bias [Stu91, TMBH93]. Bicarbonate [Pea98, PLL^{+}95]. Bight [SF96]. Bigmouth [VB95]. Binary [PK92b]. Binary-Coded [PK92b]. Biocharacteristics [WV96]. Bioeconomic [HTM96, JMBK92]. Bioenergetics [Neg95, PJM95, PBC98, ZHR99]. Biological [HTS95, Ove99]. Biologists [AS97]. Biology [Hoo91]. Biomass [GB99a, HMG^{+}96, MFP91, Wat92]. Biopsy [MLSH^{+}99]. Biota [MRP99]. Biotelemetry [BCAM99]. Biotic [LWS96]. Biscayne [ADS^{+}99]. Black [AHC98, AHP99, BK99, BS91, BKSM91a, CMS^{+}99, CS96b, CS99, DW92, GSMG95, GW95, KJS95, KGW93, LSS91, LW91, MBK^{+}91, MD91, MFBM90, NM91a, OWSM99, PW98, RD91, SB99, SA98, She91, SF91, SMTD95, SF96, WSOM98, WRD98, ZHR99, BB99]. Bladder [BCC93, CMS^{+}99, KBHM96, Lee92, SS97c]. Blocking [Swi99]. Blue [BHK94, Car99, Gui93, Jac95, JFY97, MP99, PBC98]. Blueback [HOCSN99, Jes90b, Jes94, NPP^{+}92]. Bluegill [BDBN97, BK99, CCG^{+}95, CMS95, DCPP97, EGP97, Ehl97, GW90, JCP97, KBF99, MMS^{+}98, PVP99, RP99, Sch99, UHP99]. Bluegills [KBF99, KJS95, McH90, Muo92b]. Boat [GG98, MT92, Ste99]. Boat-Mounted [MT92]. Bodies [CRC99]. Body [BWF97, BF97, Gab91, NM99, TBNB97, Tri93]. bonariensis [BA93a]. Bonavista [Tay92]. Bone [CM99b]. Bonneville [BR94, GWB^{+}91, HCO99]. Bonytail [TSKS99]. Boom [AS97, MT92]. Bootstrap [Sta92]. Border [ZHR99]. Boreal [CS97a]. borealis [RH97]. Bottom [AC92, EP94, SG93]. Bottom-Dwelling [AC92]. Boulder [MG97, SNF94]. Bow [EH93]. Bowfin [MMS^{+}98]. Box [SE98]. Boxes [GB96]. Brands [CS99]. Brazil [AAMB97]. Briefs [BC94, BG90, BM96, BS91, Cob92, CS97b, DW95b, DW97b, EH93, FE95, Goo90, Gre92, HN95, MMD^{+}93, MDI^{+}96, Mir93, Moh97, MTT90, PT93b, PB94b, PM90, Rin96, RBB96, Sch92, Sor94, Tip91, UHL97, WDM91, ZB94, ZD95]. Bristol [GH94b]. British [CCQR96, MR91, WBPQ92, HT92b, NWHM90, Ric94, WH94]. Brook [BN94, CR97, DW91, FW92, Gre91, HC94, JRH92, KMHS95, LH92, MDI^{+}96, NM99, NA92b, NA94, PP96, QKH^{+}94, Snu92, Sor94, SH92b, TR96, TR98, VKS93, WHR90]. Brown [Bor94, BSN95, CBH91, GB95, Gre92, GS93, HSS99, HJ94, Jow92, LH92, MBHC93, MK91, MTK92, NMK94, Nor99, Sal93, SNF94, SS99, TBN97, TBNB97, WW94]. Brunswick [Cam92, Jes90b, Loc98]. Brush [JL92]. Bucket [IH97]. Buffalo [TH95, VB95, SJ98a]. Bulbs [Kis93]. Bull [BM97, BTS95, Max99, RM96, RLT97, TS96, WH97]. Burbot [BPC91, BPL93, BCC93]. Bust [AS97]. By-Catch [HG93]. Bycatch [CDW97, CH98, CRMH94, CRS96, GC99a, PWB^{+}98, RC99, RFS95, RH97]. Bycatches [HD98]. Bypass [BWO97, KO93, KB97].

Cache [SS91]. Cage [SBF93]. Cage-Released [SBF93]. Caged [GR94, Rot94]. Calcein [BHK94, Moh97]. Calculated [Ove99]. Calculation [KBF99, KH97a, SMH92, Sta92]. California [Ber94, BMY94, KV96, KM96, MS92, Mat96, PWB^{+}98, SH92a, WBPQ92, YFM98]. Camera [GG98, IWTC91]. Can [GP99, WPE96]. Canada [Cam92, Coo90, SL90]. canadense [STS^{+}96]. Canadian [CP95, MFC^{+}92, PH90]. Canal [BC93]. Canals [ML94]. Canyon [MRP99]. Capacity [ZS95]. Cape [CHCC95, She91]. Captains [HS93]. Captive [MR96, MR97, RM97]. Capture [BPC91, BTS95, BC94, MKH91, Pin94, SH92b, TH99, TBN97, UHL97]. Captured [RM92, WV96]. Capturing [ECS98, SRL97]. Carbonic [GZB98]. Carcass [SCKF98]. Carcass-Derived [SCKF98]. Carolina [BWW95b, CRMH94, CS97b, Dor95, FD95, Gre95, GR97, HD98, HWE96, HRM99, Kir92, LW91, MKK97, Nel98, Pal95b, Par92, RM90, She91, SB94]. Carp [BC93, BTT^{+}93, BC91, CLA95, CHTC93, DGE^{+}97, Kir92, MSG99, MKK97, RLM94, VB95]. Carpsuckers [BDG99]. Carr [BMMF99]. Carriers [GG98]. Carrying [ZS95]. Case [GP93, GW95, HSS^{+}94, JW98, MS95, Pin94, RR99, SNS99, WD95]. Cast [SMH95]. Cataracts [KE91]. Catch [All97, Bac95, BHY^{+}97, BO98, BS91, BG95, BM93, CCS^{+}99, CCQR96, Car99, CDW97, CRMH94, Cob92, CBC95, DGFB99, DGE^{+}97, EDGR97, EP94, Fis97, GKM93, GWS96, HBJ95, HC92, HG93, HCFG96, HW94, HP92, KM98, Leg98, Loc97, MRL93, MWL95, MD93, Mor93, Nel98, OL90, OWJ93, PF99a, PJM95, PTK^{+}97, PHJ^{+}97, Ric94, RTB99, SB99, SS97a, SS97b, SMH95, SS97c, SL95, Smi90, Sta92, SB94, VBM^{+}95, ZM98, ZS97]. Catch-and-Release [BS91, Car99, Nel98, OL90, PJM95, PTK^{+}97, SS97a, SS97b, SS97c]. Catch-Curve [All97]. Catch-Depletion [MRL93]. Catch-Per-Unit-Effort [Sta92]. Catch-Survey [CCS^{+}99]. Catchability [Dwy90, KWW90]. Catchable [JBHW95, JW98, Mor93, SJ98a]. Catchable-Size [Mor93, SJ98a]. Catches [AH93, CMS95, Han95, HGH97, HN92, Mil91, PIM95, SC92]. Catching [Ell92]. Categories [QGB98]. Categorizing [MFP91]. Catfish [BPB^{+}97, Cun95, GH91, Jac95, JFY97, MWF94, MP99, PH96, PVA98, Pit97, Qui93, SWS94, SMMC96, SJ98a]. Catfishes [TH95]. Caught [CCC^{+}96, Ded96, Dor95, HM95, MH97a, MMD^{+}93, MM94b, Muo92b, NA92b, OWJ93, PMW98, PT93a, RFS95, Sch96a, SB96, SGWN94]. Causal [KH95]. Caused [Fry98, HSCM96, Sch96b]. Causes [FN92]. Cautions [BCCEJ97, CT97]. Cave [BKSM91a]. Cedar [RMV96]. Census [NRA97]. Center [JFB99]. Central [Ber94, BM94, FCP97, MCC97, MS95, OAM98, PBTS95, PTS95, YFM98]. Centrarchid [BNNS95, JJ97, SN92]. Centrarchid-Dominated [SN92]. Centrarchids [MM91]. Century [Edw99, Mar99b]. Certain [BBM96]. Cesium [Han90]. Cesium-137 [Han90]. Chains [NF93]. Chaleur [HL99a]. Challenge [BCAM99]. Chandalar [DO98]. Change [Edw99, MBD^{+}92]. Changes [BK99, BBM96, BMM91, CBC95, CHR93, Gab91, HSL90, HCFG96, JG99, JGF94, KLN97, LEW^{+}94, MJHG97, Pit97, TB99, YHW91]. Changing [WPE96]. Channel [BM97, BPB^{+}97, GH91, Jac95, JFY97, NJ96, ORM92, PH96, PVA98, Pit97, RCRR92, SWS94, SJ98a, WF93]. Channels [MH95]. Char [AKG93, KKS95]. Characteristics [Cam92, Col91b, DMS96, EE92, GLLB97, GW95, HM91, KBH97, KGW93, Mic98b, MS95, O’S95, OL96, PBTS95, RS96a, RIT97, SD93, SMTW98, SMTD95, TGM98, WNFK94, WB92b, WSOM98, WRD98]. Characterization [GR97, HD98, KMHS95, RC99]. Characterize [TR92]. Characterizing [Hal98]. Charter [Ste99]. Chautauqua [MFM99]. Check [OS96]. Checking [BC94]. Chemical [CU90, KH98]. Chesapeake [Uph97, CBJ93, RR99]. Chickamauga [MB91]. Chignik [RR92]. Children [CK99]. Chinook [AMJM96, BLLPD98, BA93b, BMS97, CCQR96, CCW^{+}97, CBB98, DW97a, EP94, Fry98, GKM93, GC99b, HL99b, HMG92, MLSH^{+}99, MM94a, NA92a, OWJ93, PF99b, PJMS99, Pol98, PM90, RCRR92, RQ95, RS96b, SBF93, SF93, TG96, WBYO97, WH94, YFM98, ZA97]. Chippewa [MSS98]. Chub [CV94, Hog99, TSKS99, WV94]. Chubsucker [ESW98]. Chum [DO98, LPT99, PMW98]. Circular [ESP92]. Clair [TH99]. Clarification [VHW^{+}96]. Class [LCSA96b, Mit91, MHW98, SB98]. Classes [CBS93, GKAM94, WP98a]. Classification [MFP91]. Classify [SJ96]. Classifying [RS95]. Claw [GB99b, HTMP94]. Cleaning [Sch96b, Wha91]. Cleithra [LMR91]. Climate [CHGF98]. Clips [CS99, TB97a]. Closed [WS91, WDR93]. Closing [SSM99]. Closures [CHCC95, SW97]. Clove [AMC97, Pea98]. Club [Dol91]. Clupeids [Pow94, SD91]. Coast [AHB95, CRS96, EP94, SL90, WPE96]. Coastal [CHCC95, CBB^{+}97, CHR93, FB96, Hou96, OHW96, Pal95b, PT93a, WS91, WG95, YHN99, vdAE90]. Coating [HT92a]. Cobalt [Han90]. Cobalt-60 [Han90]. Cobble [MG97, PK94]. Cod [Car99, CDW97, CSCM99, Han95, HSS^{+}94, HN93, SG93, WPE96, CHCC95]. Coded [BSS93a, BT93, DR97, GSC96, HG98, OB93, PK92b, PPQ94, PI95, UHL97]. Coded-Wire-Tagging [DR97]. Coefficient [Wat92]. Coffelt [VHW^{+}96]. Coho [BIDN98, BBW94b, BTAH95, BMY94, CBB^{+}97, Ell92, GKM93, HMG92, IB92, JSN90, NN91, NA92a, PF99b, PPQ94, RQ99, RSJB92, RR92, SBF93, SES94]. Coldwater [LWS96]. Cole [EE92]. colias [Car99]. Collapsible [ECS98]. Collect [MS96, SMH95]. Collected [GWS96, KGW93, NGW95, NW95]. Collecting [ASS91, Dew92, SM96]. Collection [DVG95, MMA^{+}99, MKK97, TWF^{+}97, TSKS99]. Colonization [BM95, BFM99]. Color [DW91]. Colorado [BC93, BR91, CV94, HM96, JBHW95, JW98, JG99, JR98, KBH97, KGFR97, MW95, MRP99, SNF94, SS91]. Columbia [CMP99, FW99b, GWB^{+}91, PB94a, RB90b, BRBV90, BF93, BAG96, CCQR96, CCW^{+}97, CBC95, DW97a, Fry98, GHS^{+}97, HT92b, MR91, MM94a, NWHM90, PWW94, Ric94, RB90a, RLT97, SF93, TSP93, TLR97, WBPQ92, WH94]. Combination [Fag98]. Combinations [PIM95]. Comment [AWR97, BCCEJ97, CT97, Cam97, FBB99, GSM93, MN93, Pol98, RK95, SCS^{+}95, TB97b, VHW^{+}96, WSF99]. Commercial [Bro99, CBS93, EP94, GNS97, HJ99, HL92, MFC^{+}92, MFBM90, Pit97, ZM98]. Commercially [OWJ93]. Common [DGE^{+}97, RLM94, VB95]. Commonly [Edw91]. Communication [CU90, Knu90]. Communities [CDM^{+}98, GW91, JLS^{+}92, ML94, PR95]. Community [BMNB93, BTT^{+}93, KLN97, Mue96, NNR99, Pal97, SC92]. Comparative [HCO99, JMD96, RQ99]. Comparing [Mir93, PP96, RTB99]. Comparison [AHP99, BR91, BDG99, CB96, Cun95, DW95a, DJO97, Dol91, DCPP97, DD97, Fag98, GHW91, GWS96, HM91, HN92, HHY96, Kis93, KV96, KGW93, LB99, MKH91, MKG^{+}93, MLP95, MC99, MBHJ95, MK99, MFG94, NRA97, NF99, NWHM90, OC94, OAM98, PK92a, PM97, PPQ94, PB94b, PC99, PS98, RSJB92, Rod95, RMV96, RS96b, SL95, SJ98b, SG97a, SB94, SSM99, TS96, VBM^{+}95, WB92a, WYL^{+}99, WFB97, WKG98]. Comparisons [CRMH94, Gre95, HBJ95, Hof95, HN95]. Compartment [MCCP91]. Compensated [DGFB99]. Competition [BLT95, Ehl97, WW94, WP98a]. Competitive [HTS95]. Complete [NRA97, TB97a]. Complex [WBMM97]. Composition [HHY96, Kir98, KV96, YHW91]. Compositions [WF93]. Computer [IWTC91, Joh95]. Computerized [HFS^{+}98]. Concealment [GS93]. Concentrating [RB99]. Concentration [MM94a]. Concentrations [JMG95, WN99]. Concrete [ML94]. Concrete-Lined [ML94]. Condition [ASB94, BN94, DMR95, Gab91, GC90, HM93, MWC92, SC91, TBNB97]. Conditions [BTAH95, DB94, Hal99]. Conducted [WBPQ92]. Conducting [SJ98b]. Conductivity [BTS95, HSS99, HW94, NM99]. Confidence [GB99a]. Configuration [MG97]. Confining [MSG99]. Confounding [ROB98]. Conjoint [TBR96]. Conne [WD95]. Connecticut [ESB^{+}96, WL98]. Conowingo [WB93]. Conroe [MBK^{+}91]. Consequences [Cri96, Goo96a, Goo96b, GH94b, Nor99]. Conservation [Ant90, MW90, SL90]. Considerations [HLDH98, WW93]. Considering [Goo96a]. Consistency [BG95]. Constant [SP96]. Constant-Flow [SP96]. Constraints [Aas95]. Constructed [Ave96]. Consumed [PF99b]. Consumption [CKB96, EHB^{+}93, ESB^{+}96, LEW^{+}94, MB97, PBC98, TLW96]. Consumptive [Jac96]. Container [Gre92]. Contaminated [OL90]. Content [CC98, NM99]. Contents [WV94]. Context [HD97, JNT92, Wah99]. Continuity [AGG96]. Continuous [APP98]. Contribution [BM99, HB98, Kay92, PJMS99, SN92]. Contributions [CBS93]. Control [BBSJ96, KEC93, MM91, MBB^{+}96, RO98]. Controlled [HB96]. Controlling [Kir92, RLM94]. Conversions [KM93]. Cooling [MD91]. Coolwater [GB95, Sch92]. Cooper [BWW95b, MKK97]. Coordinated [BSMM90]. Coordination [Pin94]. Coos [JNT92]. Coral [RQTW95]. Core [RC94]. Cormorant [MWH96]. Corrected [SC92]. Correction [WKG98]. Cortisol [MM94a]. Cost [DFW91, HSS^{+}94, JW98, Mar92, SWS94]. Cost-Effectiveness [DFW91, JW98, Mar92, SWS94]. Costs [JT93, JBHW95, LF99, PR93]. Count [Hil97, HRJP93, IWTC91, JH90, MK99, NRA97]. Counting [DS91, LB99, MCCP91]. Counts [BR91, FS97, GP99, HMG92, HRJP93, HA95, JQV98, Max99, PB94b, RM96]. Cover [GS93, GS95]. Coverage [MP97]. Coyote [JDL^{+}98]. Crab [CDW97, GH94b, Gui93, MM93]. Crabs [HJ99, Tay92]. Cradle [Lar95]. Crappie [AM95, All97, AHC98, AHP99, BK99, Col91a, GC90, Hal96, Hoo91, MS98, MJP91, MD91, MFBM90, Mit91, NM91a, PM97, PW98, SF91, SMTD95]. Crappies [Col91b, CS96b, CS99, Gab91, GW95, GWS96, HLMP99, Hal99, HM91, HB91, HBW94, IBS99, KJS95, KGW93, LSS91, MBK^{+}91, MOAS98, MB91, McH90, MSAM96, RD91, SK91, WO91]. Crayfish [Stu91]. Creek [BR94, BFM99, BMMF99, WP98a]. Creel [Dol91, ESB^{+}96, NRA97, WFB97]. Crew [HC92]. Criteria [BBH97, FBC97, FBB99, Uph97, WB92a, WSF99]. Critique [Edw91, MM93]. Crops [FW92]. Cross [ORM92]. Cross-Sectional [ORM92]. Crossflow [DG93]. Crude [NN91]. Crustacean [HKKR92]. Crypsis [DW91]. cucullatus [DW91]. Cui [ESB93]. Cui-ui [ESB93]. Culture [Pin94, SF91]. Culverts [CR90, Goo90, SE98]. Cumulative [JBH^{+}99]. Current [BNNS95, BG90, BMY94, KJS95, TBN97]. Currents [MR96]. Curve [All97, Hil97]. Curves [Chi91, WB92a]. Cutthroat [BR94, BT93, BTS95, BR91, DWS97, DVR97, Dwy90, DE95, DW95b, DW97b, FB96, GLLB97, GS93, HCO99, HHY96, Hou95, JDL^{+}98, KBH97, KH97b, LGS96, MDI^{+}96, PT93a, SRCIW96, SSUL97, STT^{+}90, WW94, WF93, WG95, YHN99]. Cyprinids [RM97].

Dace [GHG97]. Dahl [KBF99]. Daily [Col91a, SJ96, SK91]. Dakota [AFW98, Fie92a, Fie92b, FJ94, GW90, GW91, GW95, KGW93, MWP94, SSD93]. Dam [BJ98, CBB98, GWB^{+}91, Jes90a, KO93, MRP99, BFJ90, BAG96, Jes90b, KLN97, PC99, WB93]. Dams [MM94a, NPP^{+}92, PMN99]. Daphnia [TLW96]. Dark [SE98]. Dart [AQR95]. Darters [WP98b]. Data [BC94, Chi91, CMP99, Dol91, O’S95, ORM92, Ove99, Pal95a, UHL97, Uph93, ZM98]. Day [BFJ90, BAG96, DD97, PC99, TS96]. Days [DW97b]. Daytime [GP99]. DC [APP98, CV94, HW94]. Debris [CBB^{+}97, CHR93, FD95, HLDH98]. Decision [JT93]. Decisions [HSS^{+}94, HMG^{+}96]. Decline [BMY94, YFM98]. Declining [HES^{+}94]. Decompression [KBHM96]. Decrease [CDW97]. Decreases [BTAH95]. Deep [AC92]. Deepened [HOCSN99]. Deepwater [BAR^{+}92, BG94]. Deferred [MS95]. Deflating [BCC93]. Deflation [CMS^{+}99, Lee92, SS97c]. Degraded [Mat96]. Delaware [BW94, HLMP99, MJP91, Mar95]. Delineation [WBPQ92]. Demersal [MKH91, OC94]. Demographic [MBD^{+}92, WRD98]. Demography [JH93]. Denil [BKM99, HOCSN99]. Dense [Sch99]. Densities [MSA^{+}95]. Density [AHC98, BHY^{+}97, BF97, BCCEJ96, CA92, Cob92, EDGR97, Fie92b, Gra92, Kel93, MRL93, MD93, McM95, Mic99, MG99, MMK97, RM95, SLG98]. Density-Dependent [CA92]. Dependence [McG99]. Dependent [BROV98, CA92, Wal92]. Depletion [MRL93, MWL95, MHW98, TR96]. Depletion-Removal [TR96]. Deposition [BT95, PK94, ROB98, SHS90]. Depressurized [Lee92]. Depth [Mic98b, WLJ91]. Derive [Uph97]. Derived [Chi91, KKM93, SCKF98]. Describe [DO98]. Description [MSAM96, Qui93]. Desert [RHSB98]. Design [ADS^{+}99, DW92, DS91, HLDH98, MP92, PK94, WW93]. Designs [JH90, MC99]. DeSmet [HGMS94]. Detachable [CR90]. Detect [CCG^{+}95, YHW91]. Detectable [Moh97]. Detection [CBB98, HG98, SB98]. Deterioration [BHF^{+}99]. Determination [CS96c, Kel93, MSA^{+}95, TMJ^{+}94]. Determine [Kay92]. Determined [BBH97, PBL97]. Determining [HSP^{+}95, LB92, PIM95, Uph93]. Developing [GHG97]. Development [BWO97, LWS96, MLP95, MP99, RLM94, RCRR92, WLK98, WNFK94, WGM91]. Device [HJ99, SP96]. Devices [GHW91, KV96, RC99]. Devils [BTT^{+}93]. Diaries [CK99]. Diel [BM97, BLW^{+}97, BAG96, SNF94, ZDK99]. Diet [CBC95, HGMS94, MP97, PN91]. Diets [BBF99, Hal98, JMD96, SF96]. Differ [GP99]. Differences [EGP97, HM95, SS99, WRD98]. Different [Ded96, DCPP97, NA94]. Differential [BSN95, BKSM91a, BHM92, HP92]. Differential-Harvest [BKSM91a]. Differentiating [JDCS97]. Digested [HG92]. Dimension [Knu90]. Diminished [SNS99]. Dip [BT95]. Dip-Net [BT95]. Diploid [PM97, SSD93]. Dipnetting [TGF92]. Direct [BNNS95, PTK94, TBN97]. Discarded [RH97]. Discarding [HL92]. Discharge [HLD99, RM90]. Discharges [MSR92]. Discriminate [WYL^{+}99]. Discriminating [MC98]. Disease [CHF^{+}97]. Diseases [SGWN94]. Dispensing [SP96]. Dispersal [CA92, PP97a, WRDY91]. Displacement [RS96a]. Disposal [CBD97]. Dissolved [Hal99, JMG95, KJS95]. Distinction [MWC92]. Distinguish [HTB94, SMTD95]. Distribution [BLW^{+}97, BM95, FW99a, FBR^{+}94, INJ97, Jac95, PW98, RLT97, Smi90, TLR97, VBM^{+}95, VKS93, WH97, WF97, ZHR99]. Distributions [CS97b, JM93, PF99a]. Divers [GB99b]. Diversion [MMA^{+}99]. Diversity [WD94]. DNA [BWW95a, BBH92, GKAM94]. Do [Sta92]. Does [MLSH^{+}99, WD95]. Dolly [FB96]. Dolphin [Sak91]. Domestic [KKM93]. Dominated [SN92]. Dorsal [BDG99, SBH96]. Dover [JH93, WBPQ92]. Downstream [BMS97, GHS^{+}97, MCCP91, OL96]. Drainage [BR94, FBR^{+}94]. Drainage-Wide [BR94]. Drawdown [Pal97, RB95, VB95]. Dredge [CHTG92, CBD97, HL99b, SH92a]. Dreissena [TB99]. Drift [FH92, HJ94]. Drift-Feeding [HJ94]. Driving [Sak91]. Drop [Dew92]. Drum [GITS99, JDCS97]. Drums [MMD^{+}93, MT91, PD91]. Dummy [CSCM99, HT92a, MLP95]. Dungeness [HJ99]. Durability [RKHW98]. Duration [MHSH96]. during [BK99, BBM96, BTAH95, FJ94, GITS99, Goe91, HES^{+}94, KJS95, KH95, MKH91, MP97, MSR92, Neg99a, SD91, TGF92, WN97, WBPQ92]. Dwelling [AC92]. Dye [DR97]. Dynamics [BSS93b, Cla94, CS97a, GH91, MD91, MFBM90, Muo94, RD91, RB90a, Sch99, WPTW93, WDR93, Wil99].

Eagle [HGMS94]. Early [Ber92, Cad90, MSS98, SB98]. East [FCP97]. East-Central [FCP97]. Eastern [AFW98, CHGF98, OHW96, OC94, PMW98, SG93, ZG90]. Ecological [Wah99]. Ecology [GKDG97, RBJ97, WWTB98]. Econometric [SM94]. Economic [DBJW98, EM93, GH94b, HCFG96, JBHW95, JW98, Kir98, KEC93, Mil91, NMK94, Pin94, Ste99, SA93]. Economics [Edw91]. Ecosystem [GW95, RIO^{+}97]. Ecosystem-Specific [GW95]. Ecosystems [NM91b]. Ecotrophic [Wat92]. Editing [HFS^{+}98]. Effect [Aas95, BTT^{+}93, BS91, CBW97, DMF96, DW97b, ES92, GT90, HOCSN99, JCP97, KBM90, MB97, MBK^{+}91, MR95, Mit91, Mor93, MTT90, MT91, NN91, SS92, SW95]. Effective [CU90, PLL^{+}95, RR99]. Effectiveness [CCT99, Dew92, DFW91, GSP93, JW98, JFB99, MSG99, Mar92, MMS^{+}98, SWS94, Swi99, VB95]. Effects [APP98, All97, AMC97, BDBN97, BBW94a, Ber92, BF97, BHF^{+}99, BT95, BCC93, BN94, Cad90, CHCC95, CCG^{+}95, CM99b, CV94, DGDG91, DB94, DE95, EH93, FBR^{+}94, GSM93, Han90, HKKR92, Hof95, IB92, JBH^{+}99, JSN90, JLS^{+}92, JM93, KBHM96, LCSA96a, LCSA96b, Loc98, MN93, MBFD98, MFM99, MRP99, MG97, Mic98a, MR96, MR97, NM91b, Nor99, PJM95, Pet93, PVP99, Pow94, PN91, RP99, RS96a, RS97, RIT97, Rin96, RW92, RO98, RB95, RM97, SDBF99, SC91, SMTW98, SH92a, Sor94, STT^{+}90, SL90, SG97b, SW96, TLW96, TMBH93, TMJ^{+}94, TBNB97, TSKS99, UHP99, WN97, WO91, Wic95, WLC97]. Efficacy [BWG93, GZB98, IBS99, Kir92, MM94a]. Efficiencies [RS96b]. Efficiency [BA90, BC94, BKM99, CHTG92, HK97, HC92, MKH91, PC99]. Efficient [ADS^{+}99, EM93, HLBD96, UHL97]. Effluent [Wic95]. Effort [BF98, BM93, Cob92, EDGR97, GWS96, HRJP93, KM98, Mor93, Pal95b, PR95, PB94b, Ric94, SL95, Sta92, TBF93, TP94, ZM98]. Egg [BAR^{+}92, BT95, MKH91, MR96, Neg99a, Uph97]. Egg-Hatching [MR96]. Eggs [DFE93, DE95, Gre92, MS96, MKH91, PH94a, PK94, Roa99, RW92, SLM^{+}97, TG96, Uph93]. Eight [KGFR97]. Elastomer [HG98]. Electric [AC92, ASS91, Kis93, Pet93]. Electrical [NM99, SES94, Swi99, VB95]. Electroanesthetized [TG96]. Electrofisher [BF98, VHW^{+}96]. Electrofishing [APP98, BNNS95, BR91, BG95, BM93, CCT99, Cob92, Cun95, DMF96, DVG95, Dew92, DD97, EDGR97, Fag98, FB93, HSCM96, HSS99, HC92, HW94, HC94, KM98, KGFR97, KHR98, Lyo92, MWL95, MM94a, MD93, McM93, MFP98, MHSH96, MHW98, MR96, MR97, Nor99, PW94, Pal95a, PS98, RK95, Roa99, RM97, SB99, Sch92, SCS^{+}94, SCS^{+}95, TR96, TBNB97, TS96, TGF92]. Electrofishing-Induced [DMF96, RK95, SCS^{+}94, SCS^{+}95]. Electrophoretic [BBH97, HG92, MTT90]. Electroshock [CV94, DFE93, DE95, DW95b, DW97b]. Electroshocker [OAM98]. Electroshockers [MT92]. Electroshocking [CCC^{+}96, Sor94]. Elevated [Ber92]. Elevation [MJHG97]. Elimination [CBH91]. Embarrass [LS90]. Embayments [MSG99, PJN97]. Embryonic [WH96]. Embryos [MR97]. Emergence [JMG95, MJHG97, PH94a, ROB98, SS99, WF93, YHW91]. emergent [RW92]. Emigrating [Ell92, KWRH94]. Emigration [KRP^{+}96]. Empirical [QKH^{+}94]. Encounter [MS92, SC92]. End [WPE96]. Endangered [CWGT96, ESB93, GHG97, LSA99, RHSB98, RM97, TWF^{+}97]. England [Ant90, EM93, OLMS94, OL96]. Enhance [DS90, MFC^{+}92]. Enhanced [GBFT98]. Enhancement [BHF^{+}99, CBD97, HLDH98, JNT92, WB93]. Enhancements [MKKR99]. Enhancing [CR97, ML94]. Enlargement [STT^{+}90]. Enough [Sta92]. Entrainment [JSD94, NPP^{+}92]. Entry [ME93, SLG98, TP94]. Enumeration [BTS95, JS95]. Environment [RQTW95]. Environmental [AGG96, BMM91, DB94, Jes94, Mit91, WH98]. Epifauna [SF96]. Equalize [KM98]. Equation [KWM93, MP99, PHW93, QGB98, RBB96, WGB96, WGM91]. Equations [BCCEJ97, CT97, KH97b]. Equilibrium [MOAS98]. Equivalent [SLM^{+}97]. Erie [HN92, MR95, TB99]. Error [BC94, Cam97, HA97, SSA99]. Error-Checking [BC94]. Escanaba [HBN^{+}98, NRA97]. Escape [ESP92]. Escapement [HSM94, Hil97, JH90, MK99, SCKF98]. Esocids [SW95]. Establishing [HN95, KEC93]. Establishment [Mue96]. Estimate [ADS^{+}99, CKT96, HM96, MRL93, PK94, PD91, RSJB92, SK95, SLG98, TS96, YHW91]. Estimated [ESB^{+}96, HCFG96, IHK99, Lyo92, PMW98]. Estimates [BF98, BR91, BDG99, CBS93, Dor95, NH98, Nor99, PTK94, PC99, RF92, SLM^{+}97, SG93, TMBH93, TH90, VD95, WSL96, WD95, WvdABD93, WP98b, WKG98]. Estimating [BBW94b, BW94, CB96, DBJW98, EHB^{+}93, FH92, FS97, Gra92, GB98, HM91, HC92, HDM91, Hil97, HRJP93, IT98, JR95, LF99, MWL95, MC99, MM93, Mir93, MHSH96, MJHG97, MHW98, NF99, OC94, O’S95, PW94, ROB98, VPB95, ZB94]. Estimation [DS91, EDGR97, GB99a, GC90, HSM94, HMG^{+}96, JH90, Leo97, MK99, PHJ^{+}97]. Estimators [Loc97]. Estuarine [JFB99]. Estuary [OSS^{+}96]. Eurasian [UHP99]. Evaluate [CS97a, DGFB99, JP92, MOAS98]. Evaluating [CHCC95, JNT92, JMBK92, OAM98, SW97, Wah99].

Evaluation [AM95, BLT95, BHM98, BSS93a, BW96, BT93, BBSJ96, BCCEJ96, BMVV95, BWW95b, BKSM91a, BKSM91b, CDW97, CHTG92, CKT96, Col91a, CS96b, ESP92, Fie92a, FB96, GB96, GC90, GW91, HM96, HLMP99, HK97, HARS^{+}99, HB96, HB91, Hou96, HT92b, IWTC91, JL98, JMG95, JH90, KO93, LB92, Loc97, LKD96, MKG^{+}93, Mar95, MFC^{+}92, MDI^{+}96, Mit92, ML94, MH95, MK97, MP99, MBS90, NM91a, PF99a, PK92b, PBC98, PR93, RC99, SS97a, SMH92, SHS90, SS97c, TR96, TR98, VPB95, WLK98, WH98, WWSF93b, Wil99, WGM91]. Evaluations [Ave96, BIDN98, Sch92, TBR96]. Evergreen [JL92]. Evidence [BCAM99, Jud92, MSS98]. Examination [McM93]. Examine [CK99]. Exceptional [BSMM90]. Exemption [GBTD99]. Existing [JFB99]. Expanded [BCC93]. Expectations [BMMF99, SS92]. Experience [Fis97, SW95]. Experiment [WGR93, WKG98]. Experimental [BTAH95, HL99a, KB97, MP92, PM97]. Experiments [BROV98, GR94]. Explaining [JFW96]. Exploitation [LSS91, NA94, Par92, PLB96, PTS95, RO98]. Exploited [GH91, Muo94]. Export [HKKR92]. Exposed [NA94]. Exposure [BRK91]. Expressing [ZS97]. Extended [PC99, SL90]. Extension [FBR^{+}94]. Extinction [DVR97, SSUL97]. Extirpated [GB98].

F [SGWN94, KWW90]. Fabricated [RMV96]. Fabrication [MT92]. Facilities [JSD94, NMB96]. Facility [SFZ98]. Factor [MWC92, ROB98, WKG98]. Factors [AAMB97, AHC98, DMR95, Fie92a, FT93, GHS^{+}97, HBN^{+}98, JRH92, KH95, LH92, MRK98, MW95, MB91, PB95, RFS95, SRCIW96, Spe93, STT^{+}90, SO96, STS^{+}96, Wah99, WH97]. Failure [FN92, GS95, LGS96, NM99, SMTD95]. Fall [BMS97, DW97a, DO98, GC99b, Mar92, RB95, TG96]. Farm [Kir92]. Fast [Nor97]. FD [SC91]. FD-67 [SC91]. Feasibility [NMK94]. Features [INJ97, WHR90]. Fecundity [DWS97, Tri93]. Fed [WL98]. Feeding [HJ94, IB92, RIT97, RBJ97, SF91, SW95, WB93]. Female [BSN95]. Fence [MK99]. Fences [DS91]. Feral [KKM93]. Fidelity [WF97]. Field [CWGT96, DR97, GR94, Gre92, GHW91, HT92b, IWTC91, IHK99, KV96, MBHC93, MDI^{+}96, PLL^{+}95, Sch92, SSD93, TWF^{+}97]. Fielding [Cla94]. Fields [MR97, RM97]. Fill [NF93]. Fin [BDG99, CS96a, CS99, VWT99, WvdABD93, WG95]. Fine [CS97a, GB96]. Fine-Sediment [CS97a]. Finfish [CHCC95, RH97]. Fingerling [BM99, CBW97, Fie92b, HDM91, HN95, Kay92, LBB92, Mar92, ML92, Mit92, PK92a]. Fingerlings [CC98, Fie92a, PP97a, SF91]. Finnish [SNS99]. Fins [TB97a]. First [GKM93, GS95, HM96, HDM91, JDCS97, KSS93, Mar99b, Mic99, SMTD95, SO96]. First-Generation [SMTD95]. First-Winter [KSS93]. First-Year [HM96, HDM91, Mic99, SO96]. Fish [Aad93, AQR95, AAMB97, ASS91, AGG96, BRBV90, BLLPD98, BMNB93, BWO97, BC94, BFM99, BW94, Cad90, CP95, CBD97, CDM^{+}98, DW95a, DVG95, EHB^{+}93, ESB^{+}96, EB90, Fag98, Goo90, Gra92, GB98, GK94, HG92, HFS^{+}98, HCO99, HKKR92, HB96, HP97, HB91, IT98, JT93, JBH^{+}99, JLS^{+}92, JSD94, JFB99, KLN97, Kis93, LB99, LCSA96a, Lyo92, MB97, MBB^{+}96, Mic98a, Mic98b, MRS98, Mue96, MCCP91, NNR99, OAM98, OB95, Pal95b, Pal97, PJMS99, PK94, PIM95, PR95, PC99, Pow96, RS97, RHSB98, RMV96, RCVR91, SLM^{+}97, Sch92, SMH95, SLK94, SL95, SE98, SFZ98, SC92, TB99, WLJ91, WB93, Wha91, WH98, Wic95, UHP99]. Fisheries [AM95, AWR97, Bac95, BT90, BBM96, Bro99, CM96, CP95, CW94, Col91a, DS90, Edw99, GT90, Hal96, HR90, HTM96, HSS^{+}94, JFW96, Knu90, Lan91, LSG^{+}96, LS96, Mal95, MP92, MSA^{+}95, ME93, Mir99, NJ96, PJMS99, PWB^{+}98, PH90, Pin94, RP99, RR96, RM95, SB98, SS97b, SS97c, SSA99, Sta92, Ste99, SLJ^{+}97, SL90, Syl94, WDR93, Wri92, ZS97]. Fisherman [MK91]. Fishers [AS97]. Fishery [Ant90, BT95, BM99, CCS^{+}99, Cam92, CH98, CRMH94, CHM98, Coo90, Cri96, CHGF98, CBS93, Edw91, ESP92, GC99a, GP93, GH94b, GR97, HS93, HSL90, Han95, HL99a, HGH97, HG94, HJ99, HRM99, Hof95, HL92, Jac96, Jes90a, JMBK92, Joh95, MBFD98, MR95, MFC^{+}92, MW90, Mit91, MHS^{+}95, NMK94, PMW98, PJM95, PBTS95, PH90, Qui93, RC99, Ric94, RR99, RO98, RH97, SS97a, SD93, SH99, SMTW98, Smi98, SSA99, SW97, SLG98, SN92, WPE96, WB92b]. Fishery-Independent [CHM98]. Fishery-Related [PBTS95]. Fishes [AC92, BBW94a, CCT99, Dew92, FBC97, JJ98, KWRH94, KHP98, OSS^{+}96, Pea98, PR96, PG93, PMN99, Rin96, RHSB98, SO96, TH95]. Fishing [Aas95, AKG93, BN95, CHCC95, CK99, DBJW98, DCPP97, GP93, Goo96b, Gui93, GBTD99, HTS95, HD92, HD97, HS94, KH95, MB97, MBD^{+}92, OWSM99, PB94b, Sak91, SK95, SMMC96, SNS99, SS91, SS92, Ste99, SA93, TBF93, TH90, WSOM98, WRD98, WKG98]. Fishway [CR90, HK97, JH90]. Fishways [BKM99, HOCSN99]. Five [BHF^{+}99, Col91a, PT98, YHN99]. Fixed [BG90, DO98]. Fixed-Location [DO98]. Flag [Sak91]. Flaming [GH93, MJHG97]. Flathead [Cun95, MWF94, Qui93]. Fleet [TP94, WDR93]. Float [OB95]. Floating [MSAM96]. Flood [MM91]. Flooding [TMJ^{+}94]. Floodplain [RIO^{+}97]. Floods [RKHW98]. Florida [AHC98, AHP99, ADS^{+}99, Bro99, BGS90, CHTC93, DTR92, JFB99, KWW90, MFBM90, MHS^{+}95, MT91, MHN^{+}95]. Flow [Aad93, BWO97, BTAH95, CKT96, CBB98, DGFB99, HB96, LB92, LPT99, Leo97, MRP99, O’S95, SP96, WB92a, WB93]. Flows [DMP91, HHY96]. Floy [BSN95, MFG94, Muo92a, SC91]. Fluctuating [HA95, JR98, MW95]. Fluctuation [AGG96, CHGF98]. Fluctuations [CBJ93]. Fluorescent [BIDN98]. Fluvial [HR90]. Focus [BSMM90]. Following [BO98, JG99, JGF94, KLN97, RKHW98, TB99, BFM99, GKM93, MFBM90]. Food [JG99, JGF94, Kem96, PN91, RO98, TLW96]. Foods [MMK97]. Foot [JQV98]. Footgear [Wal92]. Forage [LEW^{+}94]. Forcing [BCCEJ97, CT97]. Forecasting [CBS93, JLS^{+}92, NWHM90, SM98, SSA99, ZM98]. Foreign [Sak91]. Foreign-Flag [Sak91]. Forest [CS97a, EH93]. Forested [Pet93]. Fork [GS93, GS95]. Form [DE95]. Formalin [MB90]. Formation [CM99a, OS96]. Four [BNNS95, Col91b, DW95a, FCP97, MB98, MBHJ95, NMB96, RP99]. Fourteen [JG99]. Fractional [TP94]. Fragmentation [DVR97]. Frame [Dew92]. Framework [RR96, SLK94]. Fraser [KBF99]. Freeze [CS99, RC94]. Freeze-Core [RC94]. Frequencies [DW95a, MHW98]. Frequency [BBH97, DRG^{+}92, Fry98, Hil97, NPP^{+}92, RDT^{+}93, RDM^{+}96, VPB95]. Fresh [PD91]. Freshwater [DGDG91, EHB^{+}93, HD97, McG99, SES94]. Fronts [AH93]. Fry [Bea95, BM95, BR91, CBW97, CA92, Fie92a, IB92, MKG^{+}93, MFC^{+}92, McM95, ML92, Mit92, Neg99a, Neg99b, OLMS94, PK92a, ROB98, RW92, RR92, WF93, WL98]. Full [SJ98b]. Full-Station [SJ98b]. Function [BMNB92, GSC96]. Fundy [HN93]. Funnel [Stu91]. Future [MBD^{+}92, RLT97]. Fuzzy [SLM^{+}97]. Fx [FSBB97]. Fyke [CCC^{+}96, Fag98, Goo90, KJ92, KHP98]. Fyke-Net [KJ92].

Gas [BRK91, KBHM96, Lee92]. Gastric [Hal98]. Gear [BMVV95, CHCC95, LS96, MR95, MHN^{+}95]. Gear-Related [LS96]. Gears [NGW95, VBM^{+}95]. Gene [LPT99, WBMM97]. Generalized [New97]. Generate [Sta92]. Generating [Rod95]. Generation [SMTD95]. Genetic [FCP97, KBH97, KKM93, KKS95, KMHS95, MKG^{+}93, WLC97]. Genetically [KSJ92, RQ99]. Genome [DTR92]. Geographic [FCP97, RQ95]. George [SGWN94]. Georgia [LSS91, MSG99, OHW96, SGWN94, WS91]. Ghost [Gui93]. Gila [Mue96]. Gill [GNS97, GWS96, HSS98, HN92, HLBD96, KJ92, MLSH^{+}99, MHN^{+}95, NW95, PF99a, PTK94, SC92, VPB95]. Gill-Net [HSS98, KJ92, MHN^{+}95, NW95, PTK94, SC92, VPB95]. Given [TB97a]. Gizzard [BMVV95, BHM92, CM99a, DMS96, HBJ95, Mic98a, Mic98b, Mic99, MG99]. Glacial [AFW98, NW95, TMJ^{+}94]. Glen [MRP99]. Global [HP97]. Goal [SCKF98]. Goals [BT90]. Goby [LSA99]. Gold [SH92a]. Golden [KV96, KM96, Mat96, KM96, Mat96]. Goldfish [Sor94]. Gonad [MLP95]. Good [ME93]. Goosefish [AHB95]. Gorge [GH93, MJHG97]. Gradient [Pal95a, TB96, WLK98]. Gradients [MG99]. Grand [Cam92, WPTW93]. Grande [SNF94]. Graphic [DW92]. Grappling [JFY97]. Grass [BC93, BTT^{+}93, BC91, CLA95, CHTC93, Kir92, MSG99, MKK97]. Gravel [Ave96, CKT96, NF93]. Grayling [BD99, Cla94, DW95b, DW97b, MDI^{+}96, Roa99]. Graysby [HS94]. Grazing [KM96, MS95]. Great [PBC98, KEC93, MC98, PH96, SRL97, TLR97]. Grebe [MWH96]. Green [KBF99, KJ92, MI98, JMBK92, KKM93, MH95]. Greenback [WW94]. Groundfish [Ant90, EM93, HL92, MW90, PWB^{+}98, SM94, Wal92]. Groundwater [GBFT98]. Grouper [Goo96a, HS94, RQTW95]. Grow [HCO99]. Growth [AHC98, BN95, BLLPD98, BW96, BPB^{+}97, BDG99, BHM92, CRC99, CA92, CS99, CSCM99, DMF96, DMR95, DCPP97, DW95b, DW97b, EGP97, Ehl97, Gab91, GC90, Hal99, HM91, HBW94, JM93, JPT93, JGF94, Kem96, KE91, KWW90, KRP^{+}96, LCGJ99, MBK^{+}91, MRK98, MLP95, Mic98a, Mic98b, Mic99, MKK97, MKKG98, NA94, OLMS94, PJM95, PPQ94, PD91, SWS94, SC91, Sch92, SO96, SW95, SW96, TLW96, TBNB97, TB99, VKS93, WB93, WL98]. Guidance [GWB^{+}91, PC99]. Guidelines [Wri92]. Gulf [GC99a, HL99a, OC94, PBTS95, CCS^{+}99, Clu96, HG93, HL92, MS96, MHR90, MKK^{+}98, MKKR99, NMK94, Pow94, PR93, RC99, RH97, vdAE90]. Guri [WWTB98]. Gut [WV94].

Habitat [Aad93, AWR97, BM94, BSPH90, BBW94a, BBW94b, Ber94, BC93, BHF^{+}99, Bin94, BR94, CCG^{+}95, CBD97, CHR93, DJO97, DVR97, FD95, FBC97, FBB99, FN92, FE95, GPB95, GC99b, HM92, HHY96, HLD99, HWH93, JBH^{+}99, JJ98, JR98, KLN97, Kem96, KBH97, KHR98, LSG^{+}96, LB92, MBHC93, MJP91, Mat96, MFP91, MSR92, MHS^{+}95, PB94a, RS95, RBJ97, SNF94, SLK94, SE98, SJ98b, SG97a, SLJ^{+}97, TGM98, WB92a, WSL96, WLK98, WSF99, Wil99, WHR90, YHN99, ZDK99, ZS95]. Habitat-Based [MFP91]. Habitats [BG94, CCT99, FB93, MR91, MM91, PS98, RIO^{+}97, RCRR92, Wri92]. Habits [JGF94]. Hake [HTM96]. Halibut [CDW97, CH98, Coo90, PWB^{+}98, RFS95]. Hand [JFY97, MP90, PK92b, RM92, SM96]. Hand-Held [MP90]. Hand-Jigging [RM92]. Hand-Tagging [PK92b]. Handling [Lar95, PT93b, SF91, TWF^{+}97]. Harbor [HK97]. Harbors [SBF93]. Hardness [HJ99, KBM90]. Harvest [AM95, BB99, BKSM91a, CBH91, Col91b, Cri96, EM93, GT90, Jes90a, JLS^{+}92, MR95, MFBM90, PJM95, Pit97, RD91, RTB99, SWS94, SSA99, WS91, WO91]. Harvestable [MWL95]. Harvesting [Pow94, UHP99]. Hatched [FH92]. Hatcheries [PBC98, WWSF93b]. Hatchery [BW96, BMS97, BW94, BBM92, CBC95, DFW91, Elr97, EE92, GITS99, HTB94, IB92, JDCS97, JSN90, JMD96, LPT99, Loc98, LF99, MC98, MW95, McM93, MPL99, MFG94, Neg99a, OS96, PJMS99, PG93, RQ99, RQTW95, RS96b, Sch96a, SF93, SBT93, Snu92, VS95, WV96, WW94, WRDY91, WWSF93a, WLC97]. Hatchery-Origin [IB92]. Hatchery-Produced [GITS99]. Hatchery-Reared [CBC95, DFW91, Elr97, HTB94, MPL99, OS96, PJMS99, PG93, RQTW95, Snu92, WV96]. Hatchery-Sustained [MW95]. Hatching [KSS93, MR96, PH94a]. Hatteras [She91]. Hauls [CB96]. Hawaii [TP94]. Hawaiian [ESP92]. Hayden [Mue96]. Head [NMB96, PMN99, SD91]. Headpond [HOCSN99]. Headwater [Bin94, CS97a, DWS97, Pet93, SRCIW96]. Health [CP95, CKB96, KH98, RHSB98]. Heavily [Muo94]. Held [MP90]. Henry [GS95]. Heritability [Tip91]. Heron [PBC98]. Herring [HOCSN99, HWE96, Jes90a, Jes90b, Jes94, NPP^{+}92, SHS90]. Heterogeneity [PJN97]. Heterogeneous [MD91]. Hierarchical [WH97]. High [ASB94, DRG^{+}92, NPP^{+}92, RDT^{+}93, RDM^{+}96, TB96]. High-Frequency [DRG^{+}92, NPP^{+}92, RDT^{+}93, RDM^{+}96]. High-Gradient [TB96]. High-Mountain [ASB94]. Highly [CMP99, DMP91].

Histocompatibility [WBMM97]. Historical [BMY94, GB98, HWE96, YFM98]. Histories [EGP97]. History [FCP97, GLLB97, Pow94, RR99]. Holding [RR93]. Holyoke [KO93]. Home [RS96a]. Homing [CCW^{+}97, JSN90, NN91, PVA98]. Hooded [DW91]. Hook [MHN^{+}95, OWJ93, SS97b]. Hook-and-Line [MHN^{+}95]. Hooking [BA93b, BO98, DB91, DMS^{+}94, LB96, MH97a, MMD^{+}93, Muo92b, NA92b, PH94b, Sch96a, SB96, TW92]. Hooks [Car99, MH97a, MMD^{+}93, TW97, TB97b]. Hoop [BPC91, HP92]. Hoopnetting [PS98]. Horizontal [SM96]. Hormone [BRC94, JPT93]. Hormone-Sterilized [JPT93]. Hourly [PB94b]. Housatonic [ESB^{+}96]. Hudson [FBR^{+}94, WRDY91]. Huff [BR94]. Huron [PM90, SBJ^{+}99, SC92]. Hybrid [BM91, HB91, HBW94, KWW90, NNR99, Neg99a, PM97, SMTD95]. Hybridization [BBH97]. Hybrids [FSBB97, KBF99, MMD91]. Hydroacoustic [BLW^{+}97, DW95a, JSD94, PC99]. Hydroacoustics [VD95]. Hydrodynamic [HR90]. Hydroelectric [DG93, GHS^{+}97, JSD94, MM94a, NMB96, SFZ98, SD91]. Hydrograph [HM93]. Hydrologic [O’S95]. Hydrology [MS98, SDBF99]. Hydropower [NPP^{+}92]. Hypothetical [RP99].

Ice [Ber94, DB91, DMS^{+}94, HK97, HLBD96, PH94b, TBF93]. Ichthyoplankton [Uph93]. Idaho [BLW^{+}97, BBM92, GBFT98, GS93, GS95, MBHC93, PB95, Sch96a, TS96]. Identification [BWW95a, HG92, SF93, SSA99, WBMM97]. Identifying [IBS99]. Identity [KKS95]. II [HWH93]. Iliamna [Leo97]. Illegal [GT90]. Illinois [BBH97, HB98, KSS93, RIO^{+}97, SBF93]. Illustrated [SSA99]. Image [MRS98]. Immersion [BHF^{+}99, BHK94, Moh97]. Immobilizing [SES94]. Impact [HTS95, HS94, MS95, NJ96, PTK^{+}97, Ste99, SA93]. Impacts [HCFG96, HB96, KM96, MJHG97, Pal97]. Impinged [Rod95]. Impingement [RDM^{+}96]. IMPLAN [Ste99]. Implant [BIDN98, BT93, BN94, FB96, MDI^{+}96, MFG94, PI95, SRCIW96, WG95]. Implanted [CSCM99, HT92a, MLP95, MLSH^{+}99]. Implanting [BRC94]. Implementation [RD91, RR96, SS97a]. Implications [BRBV90, BBW94b, BBH92, Bor94, BBM96, Bro99, CBD97, CBB98, CN94, FW99b, HR90, JJ97, KKS95, KM96, KMHS95, MBD^{+}92, NPP^{+}92, PBL97, PW98, RD91, RHSB98, RR92, Sch96a, Tri93]. Importance [BT90, HLD99, SCKF98, SF96, TR92]. Impounded [SA98]. Impoundment [CMP99, RMV96, SWS94, SN92, HKKR92]. Impoundments [BA90, CLA95, ESW98, GW91, HB91, HBW94, KSJ92, MB98, McH90, NNR99, Nov90, PW98]. Impregnated [RLM94]. Improve [Ave96, WD95]. Improved [DBJW98, MKH91, Wha91, Wic95]. Improvement [SE98]. Improving [BG95, KH98, NMK94, SG93]. Inch [Hof95]. Incidence [FN92]. Incidental [EP94, SMH95]. Incidentally [PMW98]. Incisions [SSM99]. Inclined [DMP91]. Inclined-Screen [DMP91]. Including [WDR93]. Incomplete [TB97a]. Increase [Aas95, CDW97, WB93, WWSF93b].

Increased [HJ93]. Increasing [CMS^{+}99, WS91]. Incremental [JBH^{+}99, WB92a]. Increments [SK91]. Incubation [BAR^{+}92, GBFT98, Gre92]. Independent [CHM98]. Index [GSM93, HWH93, Jes94, LWS96, MN93, MD93, MBS90, NM91a, SJ98b]. Indexes [NW95]. Indexing [CMP99]. Indiana [SM98]. Indicate [LGS96]. Indicating [LGS96]. Indices [ASB94, BA93a, CHM98, GC90, HSP^{+}95, Lan91, Mir93, SG97a]. Indirect [PTK94, RO98]. Individual [CR97, RS97]. Individual-Based [CR97, RS97]. Induce [Moh97]. Induced [DMF96, RK95, SCS^{+}94, SCS^{+}95, TBN97]. Industry [AS97, Kir98, MCC97]. Inequality [Bac95, Smi90]. Inexpensive [AQR95]. Inferences [HLD99]. Infestation [Sch96b]. Influence [BNNS95, BI99, DFE93, DW95b, Edw91, Edw99, GHS^{+}97, GLLB97, HT92a, HW94, LB96, Mic99, MP90, OWJ93, Roa99, SCKF98, TH90, WLJ91, WNFK94, WP98b]. Influencing [AAMB97, BF93, JRH92, LH92, RFS95, SRCIW96, Spe93, Wah99]. Information [KH98, SS92, Syl94]. Inhabiting [SRCIW96]. Initial [OWSM99, PG93]. Initially [CCC^{+}96]. Injection [DR97]. Injector [BRC94]. Injuries [BNNS95, TBN97]. Injury [DMF96, HSCM96, HSS99, HC94, KGFR97, KEC93, McM93, MFP98, RK95, SCS^{+}94, SCS^{+}95]. Inland [SNS99, SH92b]. Inlet [WGR93]. Inseason [ZM98]. Insertion [PMBM97]. Inshore [CRMH94, HRM99]. Insight [ZHR99]. Installation [BFM99]. Instantaneous [HRJP93, NRA97, PB94b]. Instream [CHR93, IB92, LB92, O’S95, RKHW98, WB92a]. Intake [RDT^{+}93, RDM^{+}96]. Integrated [AMJM96, BD99, Clu96, MMA^{+}99, MRS98, New97, PPQ94]. Integrity [LWS96]. Intensities [DCPP97]. Intensity [KEC93]. Intensive [AKG93]. Interaction [SG97b]. Interactions [MPL99]. Intercept [BWF97]. Interface [AWR97, LSG^{+}96, SLJ^{+}97]. Interior [KH97b, TLR97]. Internal [HARS^{+}99, SBH96, WDM91]. International [PH90]. Interpretation [FMC97, ORM92]. Interpreting [HGW94]. Interreplicate [Pal95a]. Interruption [BW96]. Interstice [WLJ91]. Intervals [GB99a]. Intragravel [CKT96, JMG95, Leo97, SM96]. Intraperitoneal [SG97b]. Intraperitoneally [CSCM99]. Intraspecific [WP98a]. Introduced [BM94, HT92b]. Introduction [GKDG97, JG99, JGF94, Mar99a, NNR99, Par95, RQTW95, TB99]. Introductions [HN95]. Introgression [CHF^{+}97, DTR92]. Intrusion [GB96]. Invasive [TSKS99]. Invertebrates [SH92a]. Investigation [WH97]. Investigations [Par95]. Investment [JCP97]. Ion [CKT96]. Iowa [LCGJ99, ML92, Mit92, PK92a]. Irregularly [KH97a]. Irrigation [BC93, JR98]. Island [HJ99, MR91, CBD97]. Islands [JJ98, ESP92]. Isolated [DWS97, Hou95]. Issue [SS97b].

Jaw [NH98]. Jigging [PH94b, RM92]. John [Jes90a, Jes90b, BFJ90, BAG96, PC99]. Johnstone [CCQR96]. Jordan [Swi99]. Juniata [JR95]. Jurisdiction [SL90]. Juvenile [APP98, AFW98, BM97, BRBV90, BLT95, BF93, BTS95, BCCEJ96, BTAH95, BFJ90, BAG96, BHK94, CBB^{+}97, CBC95, CS96b, CS99, CSCM99, Dew92, DG93, DW95b, DW97b, FW99b, GC99a, GWB^{+}91, GHS^{+}97, GS93, HFS^{+}95, HARS^{+}99, HMG92, JN95, Jes94, KRP^{+}96, MMA^{+}99, MLSH^{+}99, MM94a, MG97, MBHJ95, Mic99, MP97, NA92a, NGW95, PF99b, PK92b, PPQ94, PN91, RQ99, RCRR92, RB90a, RSJB92, RBJ97, RR92, RM90, RM92, Sal93, SFZ98, SD91, TSP93, WNFK94, WBYO97, ZA97, ZD95]. Juveniles [RM97, SLM^{+}97].

Kansas [Gab91, TGM98]. Karluk [SCKF98]. Kenai [BA93b]. Kendall [GHG97]. Kennedy [JFB99]. Kentucky [BMM91, BKSM91a, BKSM91b, BHM92, BM93, BM99, BMB^{+}99, BMMF99, Hal96, Hal99]. Kept [Mil91]. Kerr [WF97]. Kill [LB99]. King [CHM98, GH94b, HD98, PR93]. Klamath [RLT97, TLR97]. Knowledge [Smi98]. Known [PD91]. Known-Age [PD91]. Kodiak [HJ99]. Kokanee [BLT95, BLW^{+}97, GBFT98, GH93, JPT93, MW95, McG99, MJHG97, Par95, RM95]. Kokanees [MH95, PB95].

Laboratory [BC94, BROV98, SS97c]. Lacs [Goe91]. Lacustrine [Bor94]. Ladder [BFM99]. Laerdalselva [Sal93]. Lahontan [DVR97, JDL^{+}98]. Lake [BB99, BAR^{+}92, BTT^{+}93, BWG93, BSS93b, BGS90, BKSM91a, BM93, BM99, BMMF99, Cla94, CM99b, CBS93, DB91, ESW98, EF90, ES92, ESO93, Elr97, FHMT97, Fie92a, Fie92b, FB96, GNS97, GKAM94, HES^{+}94, HSP^{+}95, HBN^{+}98, HSS98, Hof95, HN95, HGW94, JDL^{+}98, Kem96, KE91, KKM93, KJ92, KH95, LMR91, LB96, Leo97, LEW^{+}94, MOAS98, MD99, MKG^{+}93, MFC^{+}92, MFM99, Mit92, MNN91, MMS^{+}98, Neg99b, NW95, NRA97, OS96, OSS^{+}96, PH94b, PHW93, PB94b, Rod95, SCKF98, SF98, SLE99, SBJ^{+}99, Snu92, STT^{+}90, TH99, Tri93, TB99, UHP99, VB95, WPTW93, AH93, BAR^{+}92, BBW94a, BLT95, Bea95, CA92, CKB96, CBS93, Ded96, DW95a, ES92, ESO93, Elr97, FJ94, FS97, GNS97, Gra92, Gre91, GLLB97, HSL90, HES^{+}94, HSP^{+}95, HSS98, HN92, HCF^{+}96, HGMS94, JLS^{+}92, Kem96]. Lake [KWRH94, KE91, KKM93, MBK^{+}91, MSG99, MR95, MD99, MC98, MKG^{+}93, MWP94, MFBM90, Mon94, MHS^{+}95, Neg95, NGW95, OS96, PB95, PJMS99, PMN99, PM90, RS96a, RDT^{+}93, RDM^{+}96, SBF93, SMH95, SF98, SM98, Smi98, SC92, Spe93, SO96]. Lakes [AHC98, AHP99, AFW98, BLW^{+}97, BPL93, BPB^{+}97, BWG93, BBF99, BMB^{+}99, CHTC93, CMS95, DCPP97, DTR92, Fag98, HT92b, JBH^{+}99, KJS95, KHP98, MRK98, MC99, ML92, PTS95, PT98, PW98, PVP99, RP99, RTB99, RR92, Sch99, SMMC96, Snu92, Tri93, WH98, BMM91, KEC93, MC98]. Lamprey [BG94, HK97, KEC93, MB90, PW94, Swi99, WP98a]. Lampreys [BSS93a, Han90, MSS98]. Lampricide [DB94, WP98a]. Landscape [INJ97]. Lane [Nor97]. Lapse [HSM94]. Large [BBW94b, Bor94, BM95, Car99, CBB^{+}97, CBD97, CHTC93, Col91a, Col91b, CHR93, HSS99, HJ94, HB96, HLDH98, HJ93, IT98, Jow92, KH95, Lar95, MFC^{+}92, MD93, MG99, Mir99, OB95, RIO^{+}97]. Large-Scale [BM95, HB96]. Largely [WW94]. Largemouth [AGS^{+}99, BN95, BMNB92, BGS90, BMM91, BM93, BM99, BMB^{+}99, BMMF99, Cob92, CN94, DGE^{+}97, DTR92, ESW98, EDGR97, GPB95, GH94a, GC90, GW90, GW91, HM95, HW94, INJ97, JN95, KWW90, KSS93, KH95, Lee92, MRL93, MWL95, MB98, Mar95, McH90, MD93, MC99, MM94b, MH94, MHSH96, MP97, Nov90, PJM95, PTK^{+}97, PJN97, PVP99, RIO^{+}97, RB95, RB99, SDBF99, SMH92, SS97c, SGWN94, SH92b, TGF92, UHP99, WN99, WN97, WD94]. Larvae [BG94, BC91, FAW98, FH92, JR95, KBM90, MB90, MR97, SLM^{+}97, WYL^{+}99]. Larval [AGS^{+}99, AFW98, BSS93a, BROV98, BHK94, JD92, Moh97, MMK97, PW94, PW98, SB98, WP98a, ZD95]. Lavage [Hal98]. Lawrence [FW99a, HL99a]. Lea [KBF99]. Learning [Joh95]. Lee [KBF99]. Length [AHB95, BWF97, CRC99, Chi91, CM99b, DW92, FAW98, Gab91, GK94, HLMP99, Hof95, KM93, KH97a, Lyo92, LKD96, MOAS98, Mar95, MHW98, MK97, Nov90, OL96, Pal95b, PF99a, PIM95, PWW94, PT98, QGB98, SMTW98, VBM^{+}95, VPB95, WO91]. Length-Class [MHW98]. Length-Scale [BWF97]. Lengths [SMH92]. Lentic [QKH^{+}94]. Leslie [MM93]. Lessons [WH94]. Lethal [Ber92]. Lethally [Moh97]. Level [KEC93]. Levels [NA94]. Liberalized [GSP93]. License [SS91]. Licenses [TP94]. Licensing [ME93]. Life [Ber92, Cad90, EGP97, FCP97, GKM93, GLLB97, Pow94]. Lift [BFJ90]. Lighted [Kis93]. Lightly [GH91]. Lights [NA92a]. Lightweight [SP96]. Likelihood [RF92]. Likely [RLT97]. Limestone [CDM^{+}98]. Liming [BWG93]. Limit [HLMP99, LKD96, MBFD98, SMTW98]. Limited [ME93, TP94]. Limited-Entry [ME93]. Limits [BLT95, CM96, Col91a, CM99b, DW92, GSP93, GH94b, Hof95, LEW^{+}94, MOAS98, Mar95, MK97, Nov90, PT98, PP96, Ric94, WO91]. Line [BM96, MHN^{+}95]. Linear [New97]. Lined [ML94]. Lining [PH94b]. Linking [Leg98]. Lipid [CC98, NM99]. Liquid [SP96]. Little [LPT99]. Littoral [BBW94a, Fag98, MM91]. Littoral-Zone [BBW94a]. Live [FJ94, Goe91, HCF^{+}96, IWTC91, IH97, KH95]. Live-Bucket [IH97]. Live-Release [FJ94, Goe91, HCF^{+}96, KH95]. Livestock [BRC94, KM96]. Living [Nor97]. Lobster [AS97, Cam92, ESP92, Kel93, MM93]. Local [LPT99]. Localized [WLC97]. Locating [HP97]. Location [DO98, GSC96]. Locations [BSS93a, BG90, TH93]. Log [PN91]. Logbooks [Sta92]. Logging [EH93, YHN99]. Logic [SLM^{+}97]. Long [Bin94, DMF96, Tay92, TBNB97, WDM91, WLC97]. Long-Term [Bin94, DMF96, Tay92, TBNB97, WDM91, WLC97]. Longevity [DWS97, GH94a, JPT93, SF98]. Longfin [Lan91]. Longitudinal [MG99]. Longline [TP94]. Loop [SBH96]. Lophodytes [DW91]. Loss [GSC96, Muo92a, NH98, PT93b, SBH96]. Losses [BBW94b, BSN95, JR95, SLM^{+}97]. Lotic [PS98]. Louis [OSS^{+}96, SLE99]. Louisiana [MKK^{+}98, HKKR92, KRP^{+}96, MHR90]. Louver [KB97]. Low [BTS95, Bor94, HSS99, HRM99, McM95, MSR92, NMB96, Pal95a, PMN99, Snu92, SD91, WLK98]. Low-Conductivity [BTS95, HSS99]. Low-Gradient [Pal95a, WLK98]. Low-Head [NMB96, PMN99, SD91]. Low-pH [Snu92]. Low-Profile [HRM99]. Lower [CCC^{+}96, BC91, CMP99, Fie92a, Fie92b, FW99b, HP92, MKKG98, PB94a, PLB96, PBL97, PVA98, PS98, RB90b, SSL96, WNFK94]. Lubbock [SD93]. Lure [OWJ93]. Lures [NA92b, PT93a].

Mackerel [CBJ93, HD98, PR93]. Mackerels [CHM98]. Macroeconomic [Edw99]. Macrohabitats [Bin94, Jac95]. Macroinvertebrate [GB99b]. Macroinvertebrates [Rin96]. Macrophyte [BAR^{+}92, BTT^{+}93, MBK^{+}91]. Macrophytes [GS95]. Mactaquac [Jes90a, Jes90b]. Made [VD95]. Madtom [FE95]. Magnetic [TH93]. Magnuson [Ant90, MW90]. Mail [RTB99]. Main [JFY97]. Maine [AS97, CCS^{+}99, EHB^{+}93, FT93, HL92, Kel93, Kir98, MB97, PH94a, RTB99, RH97, Ste99, TBF93, TBR96]. Mainstream [MB98]. Major [WBMM97]. Making [RCVR91]. Malaŵi [Smi98]. Male [BSN95, Ehl97, Han90, JCP97, RS97, Tay92]. Males [EGP97]. Mammal [HTMP94]. Manage [Wri92]. Management [Aas95, Ant90, ASB94, AWR97, BA93a, BT90, BRBV90, BBH92, Bin94, BR94, Bor94, BSMM90, BC94, BBM96, BG90, BM96, Bro99, BS91, BWW95b, BMB^{+}99, CM96, CP95, CR97, CH98, Cob92, CW94, CS97b, DW95b, DW97b, EH93, Fis97, FW99b, FE95, GP93, GSP93, Goo90, Gre92, GKDG97, GW95, HS93, HR90, HG93, HN95, Hoo91, HMG^{+}96, JH93, JFW96, JJ97, Jes90a, JMBK92, Joh95, KM96, Knu90, KMHS95, LSG^{+}96, Mal95, Mar99a, MMD^{+}93, MP92, MDI^{+}96, MW90, Mir93, Moh97, MKKR99, MTT90, Neg95, OL90, PBL97, PT93b, PB94b, PH90, Pin94, PM90, PP96, PR93, Qui92, RP99, RR96, RR99, Rin96, RHSB98, RBB96, RR92, SB98, Sch96a, Sch92, SE98, Sor94, SLJ^{+}97, SL90, Syl94, TBF93, TAL^{+}97, Tip91, Tri93, UHL97, Uph97, WDM91, WPTW93]. Management [WGR93, Wat92, Wic95, WD94, Wil99, ZB94, ZG90, ZD95]. Management-Oriented [Fis97, WD94]. Managing [AM95, GSMG95]. Manan [Cam92]. Mandatory [NRA97]. Manipulating [BW96, DS90]. Manipulations [CCG^{+}95]. Manitoba [FBR^{+}94]. Manufactured [Gre95]. Maps [IHK99]. Maracaibo [Mon94]. Margaret [BFM99, FB96]. Marine [HTMP94, NN91, Ste99, SA93]. Marine-Mammal [HTMP94]. Mark [CS99, DS91, EDGR97, IT98, MC99, MK99, Moh97, TH93, WD95]. Marked [BW94, CS99, KBF99, KSJ92, PT93b]. Markers [GKAM94, JP92]. Market [Syl94]. Marking [BIDN98, BW96, BTS95, BHK94, DR97, HM96, HBM91, IBS99, LT98, Neg99b]. Marks [CS96b]. Marlborough [Car99]. Marsh [HKKR92, KRP^{+}96]. Marshland [MHR90]. Maryland [DFW91]. masou [MK99]. Mass [BW96, BHK94]. Mass-Marking [BW96, BHK94]. Massachusetts [KO93, SA93]. Masu [MK99]. Maturation [MWF94, Tri93]. Maturity [AHB95, DWS97, Tip91]. Maximization [CHGF98]. Maximum [Coo90, GB99a, Goo96b, HMG^{+}96, PF99b, RF92, SJ96]. Maximum-Likelihood [RF92]. Mean [Mic98b]. Means [BWW95a, BPC91, BW96, WBMM97]. Measure [HFS^{+}95, HJ99, IWTC91, MC98, WF93]. Measured [BCCEJ96]. Measurements [GK94, MP90]. Measures [YHW91]. Measuring [DGFB99, JMG95, Mil91, NF93]. Mechanical [DFE93, UHP99]. Medicine [EH93]. Meet [BMMF99]. Melvern [Gab91]. Membership [GP93]. Mendota [JLS^{+}92]. Menhaden [MHR90, Pow94, RC99]. Menippe [PBTS95]. Mercury [NA92a, WN99]. Meredith [MK97]. Mergansers [DW91]. Meristic [HA95, She91]. Merrimack [BGJ^{+}98, JMD96, Rot94]. Mesh [CB96, Fag98, HP92, OAM98, PF99a, PK94, WPE96]. Meta [TW92]. Meta-Analysis [TW92]. Metamorphosis [MSS98]. Meters [BG90, MP90]. Method [AQR95, CS97a, CKT96, HLBD96, Hil97, IWTC91, JMG95, MRL93, MFM99, MM93, PIM95, SJ96, TH90, UHL97, VS95, WF93]. Methodology [WB92a]. Methods [DGFB99, HBJ95, HM91, HRJP93, JN95, Mal95, MC98, NWHM90, PW94, SJ98b]. Metolius [WLC97]. Mexico [GC99a, HG93, NMK94, PBTS95, PR93, vdAE90]. Michigan [BB99, FHMT97, GP93, HSL90, HSP^{+}95, HSS98, Jud92, KJ92, Loc97, MD99, MSS98, NMB96, SBF93, Sch99, SM98, Swi99, TH99, WPTW93]. Microhabitat [BM97, HJ94]. Micropterus [SB99]. Microstructure [HTB94]. Microtags [PK92b]. Mid [CBJ93, DW97a, GHS^{+}97, PWW94]. Mid-Atlantic [CBJ93]. Mid-Columbia [DW97a, GHS^{+}97, PWW94]. Midcoastal [Kel93]. Middle [MI98, SF96]. Midwater [KL91]. Migrants [MCCP91]. Migrating [BRBV90, Bea95, MPP94, ZA97]. Migration [AMJM96, BLLPD98, BF93, BGJ^{+}98, CCW^{+}97, DO98, EE92, GHS^{+}97, LS90, OL96, PVA98, RR93, Swi99]. Migrational [WNFK94]. Migratory [Par92]. Mille [Goe91]. Mills [KLN97]. Miltona [Spe93]. Milwaukee [KLN97]. Mine [RMV96]. Mini [Fag98]. Mini-Fyke [Fag98]. Miniature [BHM98]. Minimum [BM96, Col91a, Goo96a, HLMP99, Hof95, LKD96, MBFD98, O’S95, PP96, SMTW98, WB93]. Minimum-Length [LKD96]. Minimum-Size [Col91a, PP96]. Mining [SH92a]. Minisatellite [BWW95a]. Minnesota [CMS95, FBR^{+}94, Goe91, KH95, MC99, Neg95, O’S95, PTS95, PT98, PVP99, RP99, Spe93, TAL^{+}97]. Minnow [LCGJ99]. Minnow-Reared [LCGJ99]. Mississippi [GPB95, JJ98, KJS95, LBB92, MSR92, MKKG98, Pit97, PS98, ZDK99, JFY97, MM91, MKK^{+}98]. Missouri [BC91, SSUL97, Col91a, Col91b, HM93, KSJ92, Mic98b, MG99]. Mitigation [Sch96b]. Mitigative [MH95]. Mitochondrial [BBH92, GKAM94]. mm [HLMP99, LKD96, SMTW98]. Mobile [CHCC95, Cun95, EB90]. Model [BRBV90, HWH93, QKH^{+}94, RS97, SW97, SLG98, ZA97, ZS97]. Modeling [BA90, BBM92, CR97, Loc98, Neg95, SM94, Ste99]. Models [CCS^{+}99, HJ94, Joh95, Jow92, KBF99, MOAS98, New97, PBC98, SJ98b]. Moderate [BPL93]. Moderate-Size [BPL93]. Modification [JBH^{+}99, KWRH94, LB92]. Modified [CDW97, HK97, RC94]. Modules [RMV96]. Monitor [AMJM96]. Monitoring [GHG97, HFS^{+}98, MBHC93, Max99, MPP94, Smi98, SFZ98]. Monitors [NF93]. Mono [HN92]. Mono- [HN92]. Montana [GH91, SSL96, SRCIW96, SSUL97]. Morphologic [HR90]. Morphological [BBH97, EGP97]. Morphometric [MMD91, PBTS95, She91]. Mortalities [MMD^{+}93]. Mortality [BNNS95, BN95, BA93b, BO98, BGJ^{+}98, Bor94, BI99, Car99, CBW97, Ded96, DB91, Dor95, DG93, DMS^{+}94, FJ94, GKM93, Goe91, Goo96b, HSCM96, HM95, HCF^{+}96, JM93, KRP^{+}96, KH95, LS96, LB96, MH97a, McG99, McM93, MM94b, MKKG98, Muo92b, MHN^{+}95, Nel98, NA92b, OWJ93, OWSM99, PT93a, PH94b, PT93b, RFS95, Roa99, RH97, RM92, SWS94, Sch96a, SB96, Sch96b, Sch92, SBJ^{+}99, SGWN94, SD91, TW92, TW97, TB97b, WN97]. Motion [DGFB99]. Motion-Compensated [DGFB99]. Motives [WRD98]. Mountain [ASB94, HHY96, KHR98, LH92, MNN91, RBB96, TR96, TR98]. Mountains [KBH97]. Mounted [MT92]. Movement [BC93, CHTC93, HM92, LW91, MHR90, Mat96, MI98, PBL97, RB95, TR98, VKS93, ZDK99]. Movements [CN94, GB95, JR98, MJP91, MTK92, MSR92, Nor99, PMN99, RS96a, Tay92, WF97]. Moving [Nor97]. Mullet [Mon94]. Multifilament [HN92]. Multiple [HGH97, HBM91, MI98, Syl94]. Multiple-Objective [Syl94]. Multispecies [SSA99, SW97]. Multiuse [Qui93]. Multivariate [MMD91]. Muscle [WN99]. Muskellunge [BBF99, CRC99, CM99b, FW99a, FMC97, HM92, JM93, LCGJ99, Mar92, Mar99a, Mar99b, MFM99, SH99, SN92, SW96, TH93, Wah99]. Mussel [Sch96b]. Mussels [TB99]. mykiss [MPL99]. Mysis [CB96].

Narrow [LB99]. Nassau [RQTW95]. National [EH93]. Native [BBM92, DTR92, OSS^{+}96, TLR97]. Natural [BN95, BBH97, Bor94, BSMM90, Jud92, MC98, McG99, MPL99, MMK97, Muo92b, PT93a, PR96, PW98, Sal93, WF93, Wri92]. Naturalized [Neg99a]. Naturally [MKG^{+}93, RQ99, SF93]. Navigation [ZDK99]. Near [Elr97, Kel93, TSP93]. Nearshore [Par92, WPE96]. Nebish [Hof95]. Neosho [FE95]. Nest [BGS90]. Nesting [RS97]. Net [BT95, CB96, CMS95, Dew92, EB90, Fag98, Goo90, HSS98, KJ92, MSAM96, MHN^{+}95, NW95, PTK94, PC99, SC92, VPB95]. Nets [AHP99, CCC^{+}96, GNS97, GWS96, HN92, HLBD96, HP92, KHP98, MKH91, PF99a, SMH95, SB94]. Netted [AAMB97]. Networks [LR97]. Nevada [Ber94, MS95]. Newfoundland [BBM96, Tay92, WD95]. Night [DD97, GP99, TS96]. No [CBH91, Hof95, JFB99]. No-Harvest [CBH91]. No-Take [JFB99]. Non [FW92, Moh97, SBF93]. Non-Cage-Released [SBF93]. Non-Lethally [Moh97]. Non-trout [FW92]. Nonanadromous [TW92]. Noncompliance [PT98, SK95]. Nondeleterious [CS96a]. Nondisruptive [AQR95]. Nonlethal [HFS^{+}95, MLSH^{+}99, VWT99, WV94]. Nonresponse [TMBH93]. Nonsalmonid [Pea98]. Nontournament [WRD98]. Nonusers [SD93]. Nonwilderness [KBH97]. Normandy [FSBB97, SDBF99, SB99]. Norms [MS92]. North [Dor95, Gre95, HWE96, MWP94, Nel98, Par92, RM90, She91, AGG96, BN95, CRMH94, DS90, FD95, GR97, HRM99, JBH^{+}99, MCC97, OAM98, PBTS95, PTS95, RQ95].

North-Central [MCC97, OAM98, PBTS95, PTS95]. Northeast [AHB95, MTK92, NMB96]. Northeastern [SFZ98, WH98]. Northern [CCS^{+}99, CBC95, DMS^{+}94, FW99b, GSP93, GSMG95, JT93, KWW90, LMR91, MD99, MRK98, MMK97, NW95, PT93b, PTK94, PTS95, PT98, RB90a, RB95, RB99, RH97, SC91, SW97, TSP93, WS91, WH96, ZD95, BBF99, MRK98, SH92a]. Northwest [Lan91, PH90, RKHW98]. Northwestern [ESP92, FHMT97, FBR^{+}94, LCGJ99, Tri93]. Norway [HJ93, Sal93]. Nova [SD91, McN95]. Nuclear [KE91, Rod95]. Nuisance [Kir92]. Number [Pal95b, PIM95]. Nursery [FW99a, HKKR92]. Nutrients [SCKF98].

Oahe [Fie92a, Fie92b, FJ94]. Objective [Syl94]. Objectives [BT90]. Oblique [KL91]. Observations [Sta92, SJ96, Tay92]. Observed [Ded96]. Observer [Cam97, DGFB99, HA95, HA97, JQV98, RS95]. Occurrence [AHC98, AGS^{+}99, SS91]. Ocean [GKM93, LS96, Par92, WV96, AGG96]. Odonthestes [BA93a]. Off [Elr97, RCRR92, AHB95, BBM96, CHCC95, EP94, MR91, SB94, WPE96]. Off-Channel [RCRR92]. Ohio [HLMP99, MJP91, STS^{+}96]. Oil [AMC97, NN91, Pea98]. Okeechobee [MFBM90]. Oklahoma [SMH95, SFZ98, WF97]. Okoboji [ML92]. Oligotrophic [LMR91, WWTB98]. On-Site [RTB99]. Oncorhynchus [MPL99, MK99]. One [SSA99]. Onslow [Gre95]. Ontario [AH93, CRC99, CKB96, ES92, ESO93, Elr97, FCP97, FBR^{+}94, KWRH94, KE91, LMR91, MC98, MKG^{+}93, PMN99, Rod95, RDT^{+}93, RDM^{+}96, SJ98b, Tri93]. Open [MB97, Rod95, TBF93, Wri92]. Open-Access [Wri92]. Open-Water [MB97, Rod95, TBF93]. Opeongo [RS96a]. Operated [OC94]. Operator [MP90]. Opinion [GBTD99]. Opinions [RP99, TBF93]. Opossum [CB96]. Opportunity [CW94, JT93]. Optimizing [Hil97, MHSH96, PR95]. Optimum [CW94]. Options [BWW95b, TBF93]. Orange [BGS90]. Order [TS96]. Oregon [BTT^{+}93, CHR93, EE92, FN92, Hou95, Hou96, JNT92, JDL^{+}98, PWB^{+}98, WNFK94, WLC97]. Oreille [PB95]. Organization [BT90]. Organizational [WN97]. Organizations [GP93]. Oriented [Fis97, WD94]. Origin [BCCEJ97, CT97, IB92]. Origins [PMW98]. Orkla [HJ93]. Ossberger [DG93]. Other [STT^{+}90]. Otolith [HTB94, KGW93, PD91, SK91]. Otoliths [BHK94, CM99a, CS96b, HBM91, KH97a, LT98, Neg99b, SMH92, WvdABD93]. Otter [AHP99, CRMH94, Han95]. Outfalls [BWO97]. Outputs [VHW^{+}96]. Overexploited [RB90b]. Overview [BMVV95, EF92, McN95]. Overwinter [BF97, CC98, PPQ94]. Oviduct [PMBM97]. Oxygen [Hal99, JMG95, KJS95, LB96]. Oxytetracycline [BHK94, CS96b, CS99, IBS99, Kay92, KBF99]. Oxytetracycline-Marked [KBF99]. Oyster [CHTG92, OHW96]. Ozark [ZHR99]. Ozette [BLT95].

Pacific [CDW97, CH98, Coo90, HSM94, HSS^{+}94, PWB^{+}98, RFS95, RKHW98, SHS90, SL90, VWT99, WBMM97]. Paddlefish [FS97, GSC96, HD96, MWP94, MSR92, PF99a, PI95, SSL96, SRL97, SS97a, TH95, ZDK99]. Paid [FT93]. Paired [EF90]. Pallette [HN95]. Pallid [KM93]. Palmetto [JR98]. Pandalus [RH97]. Panfish [JL92]. Parameters [BBM96]. Parametric [SLM^{+}97]. Paraná [AAMB97]. Parapercis [Car99]. Parasites [RHSB98]. Parasitic [BSS93a]. Parasitic-Phase [BSS93a]. Parentage [WW94]. Part [CCW^{+}97]. Partial [DS91]. Partially [HG92]. Participating [ESB^{+}96]. Participation [Aas95, Fis97, MBD^{+}92, TMBH93]. Particular [AGG96]. Party [Ste99]. Pass [HSCM96, KHR98]. Passage [BJ98, BAG96, BKM99, Cad90, HK97, HOCSN99, HFS^{+}98, Jes90a, KB97, NMB96, SFZ98, SD91]. Passed [DG93]. Passing [MMA^{+}99, MM94a]. Passive [AMJM96, BD99, Clu96, DVG95, MKH91, MMA^{+}99, MPP94, New97, PPQ94]. Patent [CHTG92]. Pattern [PMW98, SF93]. Patterns [BW96, CBD97, CKB96, CK99, DO98, Ehl97, FCP97, Mat96, OHW96, Par92]. Patuxent [DFW91]. PC [UHL97]. PC-Based [UHL97]. PCBs [OL90]. Pearl [MKK^{+}98, MKKR99]. Pejerrey [BA93a]. Pelagic [DW95a, OB95]. Pellet [BRC94, LCGJ99]. Pellet-Reared [LCGJ99]. Penaeid [WDR93]. Pend [PB95]. Pennsylvania [JR95]. Penobscot [TBR96]. Peoria [BBH97, HB98]. Percentile [BM96]. Percentiles [GB99a]. Perceptions [SD93, TR92]. Perch [AFW98, BK99, BSS93b, FAW98, GW91, JMBK92, KBHM96, KJ92, MR95, NF99, SM98, WYL^{+}99, WGM91]. Percichthyid [NGW95]. Perforated [Gre92]. Performance [AMC97, CSCM99, HCO99, HRM99, RQ99, SSD93, TB96]. Performing [JL98]. Periodic [BG90, JH90]. Permanently [RC94]. Persistence [CS99, DGDG91, LSA99]. Perspective [WWSF93a]. Perspectives [AS97, ANN91, Qui92]. Pest [RO98]. pH [Ber92, Snu92]. PHABSIM [BCCEJ96]. Phase [BSS93a]. Phenology [SS99]. Phenotypes [MTT90]. Phenotypic [SMTD95]. Photochemicals [Kis93]. Photography [Kel93]. Phototaxis [ZD95]. Physical [FN92, GLLB97, RIT97]. Physiological [GITS99, RIT97]. Pickering [Rod95]. Pierre [BBM96]. Pike [DMS^{+}94, GSP93, LMR91, MRK98, MMK97, NW95, PT93b, PTK94, PTS95, PT98, RB95, RB99, SC91, ZD95]. Pikeminnow [FW99b]. Pink [ADS^{+}99, HG94, JQV98, Pin94, ZM98]. Pinniped [Fry98]. Pinniped-Caused [Fry98]. Pipe [SM96]. Piranha [AAMB97]. Piscicides [BBSJ96, SP96]. Piscivores [WWTB98]. Piscivorous [HG92]. Piscivory [BMNB92]. PIT [AMJM96, BRC94, CBB98, MPP94]. PIT-Tagged [CBB98]. Placement [CBB^{+}97, HLDH98, SNF94, WG95]. Plaice [BBM96, ZG90]. Plain [Pal95b]. Plains [PH96, SRL97]. Plan [MKKR99]. Plant [RDT^{+}93, RDM^{+}96]. Planted [PH94a]. Plasma [MM94a]. Plastic [PT93b]. Platte [HP92]. Point [Loc97, PHJ^{+}97]. Points [Ove99]. Policies [HG93]. Policy [CW94, HCFG96]. Pollock [CHGF98, HSS^{+}94, HCFG96, PMW98]. polymorpha [TB99]. Polypropylene [RMV96]. Ponds [BA90, DGDG91, DGE^{+}97, EDGR97, Ell92, GW90, JRH92, Kir92, KWW90, KRP^{+}96, Mar95, PM97, RR93, SSD93, WHR90]. Pool [BBH97, HB98, MSR92, ZDK99]. Pools [Ber94]. Pop [Dew92]. Population [AKG93, ADS^{+}99, BHY^{+}97, BF97, Bor94, BBM96, BSS93b, BN94, BWW95b, BM99, Cla94, Cob92, Col91b, CHF^{+}97, DS91, DMS96, EDGR97, Fis97, GH91, GSP93, GW95, HM96, HM91, HDM91, Hof95, HN95, Hou95, JCP97, KBH97, LR97, LCSA96a, MRL93, MMA^{+}99, MW95, MD91, MD93, MC99, Mic98b, MFBM90, MKK97, Muo94, Nor99, RD91, RB90a, RM95, RF92, SMTW98, TGM98, WPTW93, WB92b, WO91]. Populations [All97, ASB94, BN95, BDBN97, BK99, BPL93, BWG93, BMVV95, BMM91, CR97, Col91a, DW97a, DW95a, DWS97, DTR92, EB90, FCP97, GB98, GW90, JRH92, JS95, KKM93, Kis93, KGFR97, KMHS95, LSA99, LPT99, LGS96, MM93, NM91a, OAM98, PW94, PJMS99, Pit97, RSJB92, Sch99, SS99, UHP99, VBM^{+}95, VPB95, Wri92]. Portable [Lar95]. Portions [TLR97]. Position [JDCS97, MP90]. Positioning [HP97]. Possible [FBR^{+}94]. Postcards [ZB94]. Postflood [LSA99]. Postrelease [CMS^{+}99, SB96]. Poststocking [Loc98, PG93, WWSF93b]. Pot [GH94b]. Potamodromous [GKDG97]. Potamodromy [Nor97]. Potential [Bor94, Cla94, CMS^{+}99, EM93, FBR^{+}94, GB99a, HCFG96, HMG^{+}96, JNT92, MKK^{+}98, Pea98, PJM95, RB90a, SS97c, WWSF93b]. Pots [CDW97]. Poudre [SS91]. Powder [GH91]. Power [BG95, CCG^{+}95, GHG97, Max99, Pal95a, RDT^{+}93, RDM^{+}96, WKG98]. Powerhouse [GWB^{+}91]. Practices [PH90, Smi98]. Prairie [MFC^{+}92, NW95, SG97a]. Prawn [SW97]. Pre [RW92]. Pre-emergent [RW92]. Precision [GK94, HWH93, IHK99, JSD94, JQV98, KM98, MP90, NGW95, Ove99, SG93, VBM^{+}95, WSL96, WvdABD93]. Precontrol [HM93]. Predation [BRBV90, BLT95, Bea95, BJ98, BROV98, DW91, JNT92, JD92, JR95, MWP94, MWH96, OSS^{+}96, RB90a, RR92, SW95, SW96, TSP93]. Predator [MMS^{+}98]. Predators [MH94]. Predict [NM99, TMJ^{+}94]. Predicted [BN95]. Predicting [Cob92, Cri96, GSM93, HTS95, Lan91, MN93, PBC98, RS97]. Prediction [BHY^{+}97]. Predictor [BM93]. Predicts [KHR98]. Preference [BSPH90, Fis97, PIM95]. Preferences [BM94, BM97]. Preliminary [GR97, MT91]. Prepositioned [BF98, FB93]. Preseason [NWHM90]. Presence [Uph93, Uph97]. Present [FBR^{+}94]. Preservation [KKM93]. Preserved [FAW98, MB90]. Pressure [BHM98]. Pressure-Sensitive [BHM98]. Prevalence [HTMP94]. Prevent [KH97a]. Preventing [TR98]. Previous [KH97a]. Prey [ESW98, Hal96, HG92, JLS^{+}92, PP97b, SW96]. Prices [FT93]. Prince [Pin94]. Prior [OS96]. Private [Nov90]. Privatizing [JW98]. Probabalistic [ESB93]. Probabilistic [LR97]. Probability [GB98, SC92, Uph97]. Probability-Based [Uph97]. Procedure [JS95, LB92, PMBM97]. Procedures [LB99, SMH92, SL95, WN97]. Process [BT90]. Processes [HR90, JN95]. Processing [BC94]. Produced [GITS99]. Production [BLT95, BBW94b, Bor94, Edw99, JW98, MRS98, OLMS94, RS97, WD95, Wat92]. Production/Biomass [Wat92]. Productivity [Mic98b, Mic99, SCKF98]. Professional [SNS99]. Professionals [ANN91]. Profile [HRM99, SSL96]. Program [BKSM91b, SD93, UHL97, WLC97]. Programs [BBM92, KKS95]. Progress [SM94]. Progressive [HRJP93]. Projections [ESB93]. Projects [BHF^{+}99, HLDH98, MBB^{+}96]. Propagated [GKAM94, MKG^{+}93]. Propagating [LF99]. Propagation [CWGT96]. Propeller [Cad90]. Propeller-Type [Cad90]. Proper [MWC92]. Proportion [Mir93]. Proportion-Based [Mir93]. Proposed [BWF97, KWM93, KH97b, SLK94]. Protection [CP95]. Protocols [GHG97, MBHC93]. Protogynous [HS94]. Provide [GS95]. Providing [SS92]. Proximity [MS92]. Prudhoe [NN91]. Public [DTR92, JFW96, SMMC96]. Pulse [DMF96, DE95]. Pulsed [APP98, BNNS95, CV94, HW94, TBN97]. Pump [BFJ90, SM96]. Pumpkinseed [FCP97]. Pumpkinseeds [BF97]. Puncturing [KBHM96]. Pure [LGS96]. Purse [CCQR96]. Purse-Seine [CCQR96]. Put [HCO99, MSA^{+}95]. Put-and-Take [MSA^{+}95]. Put-Grow-and-Take [HCO99].

Quahog [WS91, WH96]. Qualitative [SG97a]. Quality [GPB95, HWH93, MRS98, Sch99, SM98]. Quality-Size [SM98]. Quantify [KWRH94]. Quantifying [GK94, SLK94]. Quantitative [Dew92, EB90, NJ96, SMTD95]. Question [FMC97]. Questions [BSPH90]. Quota [JMBK92]. Quotas [Leg98].

Radiation [Han90]. Radio [BHM98, BGS90, HP97, MLP95, MLSH^{+}99, PMBM97, PBL97]. Radio-Tagged [BGS90, HP97]. Radius [BWF97]. Rainbow [APP98, AMC97, BTAH95, BT95, BI99, CHF^{+}97, DMF96, Ded96, DW95b, FBR^{+}94, GZB98, GS95, HSCM96, HT92a, HDM91, Hog99, HGMS94, HT92b, JBHW95, JW98, JG99, JGF94, KL91, LGS96, LF99, MC98, MLP95, MDI^{+}96, MK91, McM93, MG97, MSA^{+}95, MWH96, MFG94, Neg99a, OB93, PBC98, RK95, Sch96a, SB96, SCS^{+}94, SCS^{+}95, SSD93, STT^{+}90, SG97b, SSM99, TLW96, TBN97, TBNB97, TSKS99, WFB97, WLC97]. Random [SK95]. Range [FBR^{+}94, RS96a]. Rapid [BTAH95, JS95]. Rate [BLLPD98, CBC95, DE95, EGP97, HJ93, Loc97, McG99, Mor93, MT91, PHJ^{+}97, SSA99, VBM^{+}95]. Rates [Aas95, BHY^{+}97, CLA95, CSCM99, DGE^{+}97, ESB^{+}96, FHMT97, GHS^{+}97, GR94, HC92, HW94, KGFR97, Loc98, MD93, NH98, PF99a, RTB99, SB99, SBH96, ZB94]. Rathbun [Mit92]. Rating [WLK98]. Ratio [DWS97, Wat92]. Ratios [ZM98]. Ray [BDG99, CS96a]. Rays [WvdABD93]. Razorback [CWGT96, MI98, MR96, MR97, TWF^{+}97, TSKS99]. Reach [DJO97, IHK99, MFP98, Pal95b, SMTW98, WB92b, ZS95]. Reading [UHL97]. Reared [CBC95, DFW91, Elr97, HTB94, LCGJ99, MPL99, OS96, PJMS99, PG93, RQ99, RQTW95, Snu92, WV96]. Rearing [BBW94b, MFM99, PB94a, WB92a]. Recall [TMBH93]. Recapture [DS91, EDGR97, HJ93, IT98, MC99, MK99, WD95]. Recently [FH92, JD92, JR95]. Recognition [AQR95]. Recolonization [LSA99]. Recolonized [WP98a]. Recommendations [BA93a, OHW96, ZG90]. Recommended [MR95, MKKR99]. Record [GH94a]. Recording [HSM94]. Recover [WV94]. Recoveries [EF90, New97]. Recovering [Mat96, Pow96]. Recovery [ESB93, MKK^{+}98, MKKR99, OB93, Pal97]. Recreation [Fis97]. Recreational [BT95, Bro99, GT90, GR97, SSL96, Ste99, SA93]. Recruit [Cla94, Jac96, ZG90]. Recruitment [All97, AGS^{+}99, Bor94, CHM98, Fie92b, HBN^{+}98, HBW94, JN95, Jes90b, MBK^{+}91, MS98, MB98, MC98, Mon94, OHW96, Ove99, RM90]. Rectangular [ESP92]. Red [GC99a, GITS99, GR94, Goo96a, GH94b, JDCS97, MMD^{+}93, MT91, PD91]. Redd [Max99, RM96]. Redds [HL99b, KV96]. Reduce [Car99, HB96, RB90a, VS95]. Reduced [Col91a, McH90]. Reduces [DW91]. Reducing [Cam97, HG93, HA97, NPP^{+}92, RK95, RR96, RDM^{+}96, SCS^{+}94, SCS^{+}95, VB95]. Reduction [Ber94, FAW98, GC99a, HB96, RC99, TP94]. Reductions [MRP99]. Reef [Gre95, LW91, RQTW95, SF96]. Reefs [ML94]. Reference [AGG96, Ove99]. Reflectance [HFS^{+}95]. Refuge [MD99]. Refugia [PR96]. Regarding [TBF93]. Regeneration [TB97a]. Region [CBJ93, MCC97, WBPQ92, YFM98, ZHR99]. Regional [Ste99]. Registry [Sak91]. Regression [BCCEJ97, BM96, CT97]. Regression-Line [BM96]. Regulated [HB96]. Regulation [BKSM91a, HGH97, PH90, Sch96a]. Regulations [BDBN97, CBH91, Cri96, GBTD99, JLS^{+}92, Pit97, Sak91, SK95, Wri92]. Regulatory [BMM91]. Rehabilitation [AKG93, BWG93, GPB95, JMBK92, MD99, MKG^{+}93]. Reintroduction [SLE99]. Related [AHC98, DMR95, HM91, KH95, LS96, MM94b, Pal97, PBTS95, RP99, RM90]. Relating [Cad90]. Relation [BLLPD98, BCCEJ96, INJ97, Mic98b, Sal93, ZS95]. Relations [AGS^{+}99, DMS96, Gab91, Jes94, KM93, MS98, MK91, TGM98, Tri93, WHR90]. Relationship [Aad93, AFW98, BWF97, CKB96, Fie92b, MP97, ZS97]. Relationships [AHB95, GW90, Jes90b, Pal95b]. Relative [CMP99, DFW91, DB94, FS97, GB95, GP93, HB98, HGW94, MH97b, MH94, MHSH96, MBS90, MWC92, NM91a, PP97b, SBT93, Snu92, SH92b, WvdABD93]. Release [BO98, BS91, CCQR96, Car99, CBC95, FJ94, GKM93, Goe91, HCF^{+}96, JSN90, KH95, Nel98, OS96, OL90, PJM95, PTK^{+}97, SS97a, SS97b, SS97c, SBT93]. Released [BA93b, BMS97, EE92, GNS97, GR94, HJ93, JD92, JR95, Mil91, MHN^{+}95, SBF93, SGWN94]. Releases [EF90]. Reliable [HC92]. Relicensing [SFZ98]. relicta [CB96]. Remnant [KKM93]. Remote [SP96]. Remotely [OC94]. Removal [BMNB93, CS96a, FW92, Gre91, KLN97, MBK^{+}91, PVP99, RF92, SL95, TR96]. Report [SM94, TMBH93]. Reporting [ZB94]. Reports [Dol91]. Represent [SLM^{+}97].

Representative [DJO97]. Reproducing [BWW95b]. Reproduction [Ave96, FSBB97, HS94, Jud92, Sal93, SDBF99]. Reproductive [DCPP97, EGP97, Ehl97, JJ97, JCP97, PTK^{+}97]. Requirements [EF90, HC92, LB92, O’S95, Pow96]. Rescue [Goo90]. Research [AWR97, HR90, LSG^{+}96, PR93, SLJ^{+}97]. Reservoir [BRBV90, BMNB93, BKSM91b, CBD97, CBC95, DMS96, FSBB97, Gab91, GB95, Hal96, HLMP99, Hal99, HDM91, IT98, IBS99, JG99, JR98, MWL95, MS98, MJP91, MW95, MB91, MD91, Mic98b, Mic99, MWH96, MJHG97, MK97, Pal97, PJN97, RD91, RB90a, RM90, SDBF99, SB99, SRL97, STT^{+}90, SGWN94, VBM^{+}95, VPB95, WF97, WWTB98, GH93, MJHG97]. Reservoirs [AGS^{+}99, BMVV95, BHM92, Col91a, Col91b, DMS96, DMR95, Dol91, DD97, HCO99, KSS93, LSS91, MH97b, MD93, MM91, Mic98b, MG99, MHSH96, Mir99, MKK97, STS^{+}96, TLW96, WO91]. Residency [MG97]. Resident [BRBV90, GP99, WLC97, Wri92, YHN99]. Residual [VS95]. Residues [KH98]. Resource [ANN91, HS93, PH90, SA98, Smi90]. Resources [BSMM90, MW90]. Response [BN95, BMNB93, BR94, BCCEJ97, BBM92, CBH91, CBB^{+}97, FBB99, GPB95, Hal99, HA97, MN93, NNR99, NA92a, Pit97, RDT^{+}93, SK95, SCS^{+}95, SLJ^{+}97, TW97]. Responses [Bin94, BMM91, CCG^{+}95, GITS99, Hog99, KJS95, McH90, MHS^{+}95, NPP^{+}92, RIT97]. Restocking [KKS95]. Restoration [BBW94b, HSP^{+}95, Hou96, MHS^{+}95, TBR96]. Restore [CDM^{+}98]. Restored [FE95]. Restraint [Lar95]. Restrict [PMN99]. Restrictions [AM95, RD91, SS97b]. Restructuring [CHR93]. Resulting [RM95]. Results [HL99a]. Retail [FT93]. Retention [BN94, BI99, BD99, Clu96, HG98, HT92a, MDI^{+}96, PI95, SRCIW96, TH95, WDM91]. Retrieval [BG90]. Return [MT91, NMK94]. Returning [Fry98, HTMP94, Jes90b]. Returns [BMS97, Coo90, CCC^{+}96, EE92, NWHM90, SBF93, SBT93, WFB97]. Reusable [OB95]. Revenue [CHGF98, HG94]. Review [BT90, Cad90, MP92, MBB^{+}96, TAL^{+}97, WBYO97]. Revised [WGB96]. Revision [KWM93]. Reward [MT91]. Rhodes [Mue96]. Richland [TSP93]. Richness [BF98, Lyo92]. Riffle [FE95]. Riffles [Ave96]. Rights [Pet93]. Rights-of-Way [Pet93]. Rio [SNF94]. Riparian [YHN99]. Ripe [MR96]. Rising [EB90]. Risk [DW91, DVR97, Knu90, KH98, NJ96, SSUL97]. River [AAMB97, BBW94b, BA93b, BF93, BDG99, DO98, DFW91, Dor95, ESB^{+}96, GH91, GHS^{+}97, GS93, GS95, HB96, HWE96, Jes90a, Jes90b, JD92, JR95, Jud92, KLN97, Loc98, MBFD98, McM95, MI98, MSR92, MSS98, MKK^{+}98, MH95, Nel98, OSS^{+}96, PVA98, PWW94, RIO^{+}97, RLT97, RM90, SSL96, SLE99, SA98, SSUL97, SNF94, SJ98a, SBT93, SMTW98, SS91, Swi99, TBR96, TMJ^{+}94, TH99, TLR97, WD95, WLC97, vdAE90, AMJM96, BRBV90, BBH97, BGJ^{+}98, BR91, BAG96, BC91, BMS97, BW94, BBM92, CCW^{+}97, CBC95, CBB98, CMP99, CV94, DW97a, FW99a, Fry98, GPB95, GWB^{+}91, GC99b, HTMP94, HB98, HTB94, HM93, HP92, HD96, HJ93, JFY97, Jes90a, JMD96, JJ98, KBH97, KJS95, LBB92, MB98, MM94a, MRP99, MSR92, MKKG98, MKKR99, New97, PB94a, PLB96, PBL97].

River [Pit97, Pol98, PS98, RIO^{+}97, RB90a, RB90b, RLT97, Rot94, Sal93, TSP93, TLR97, WNFK94, WBYO97, WB92b, WRDY91, WL98, ZA97, ZDK99]. River-Floodplain [RIO^{+}97]. Riverine [Bea95, Jac95, RO98, RBJ97, SA98]. Rivers [EE92, FH92, HJ94, Jow92, LS90, MP90, MCCP91, PK92a, TR92, VS95, FW99b, WGR93]. Roads [EH93]. Roanoke [Dor95, Nel98, RM90]. Robert [WF97]. Rockfish [MR91, Ric94]. Rockfishes [OC94]. Rocky [TR96, TR98]. Role [CP95]. Rotary [TMJ^{+}94]. Rotary-Screw [TMJ^{+}94]. Rotation [MS95]. Rotenone [BA90, DGDG91, Hog99]. Roundtail [CV94, WV94]. Route [CCW^{+}97]. Routes [MHR90]. Roving [PHJ^{+}97]. Royal [SD91]. Ruffe [BBSJ96, ECS98, OSS^{+}96]. Rule [BCAM99]. Run [BT93, BKSM91a, SLG98].

S [WF97]. S. [STS^{+}96]. Sablefish [RM92]. Sac [Neg99b]. Saginaw [Jud92]. Sagitta [JDCS97]. Sagittal [CS96b]. Saint [Jes90a, Jes90b]. Sakakawea [MWP94, FS97]. Salem [WP98a]. Salinity [KBM90]. Salmon [AMJM96, AGG96, BIDN98, BWW95a, BLT95, Bea95, BLW^{+}97, BLLPD98, BBW94b, BA93b, BJ98, BGJ^{+}98, BCCEJ96, BTAH95, BMY94, BMS97, CCQR96, CBB^{+}97, CCW^{+}97, CBB98, CHR93, DW97a, DO98, DS91, Ell92, EP94, Fry98, GHS^{+}97, GKM93, GC99b, HL99b, HSM94, HG94, HMG92, HJ93, IB92, JSN90, JMD96, JQV98, LS96, LPT99, Leo97, LT98, Loc98, MLSH^{+}99, MS92, MM94a, McG99, McM95, MK99, Moh97, NN91, NA92a, NWHM90, OLMS94, OL96, OWJ93, PMW98, PH94a, PF99b, PJMS99, PPQ94, Pin94, Pol98, PM90, PN91, RQ99, RCRR92, RSJB92, RQ95, RS96b, RR93, Rot94, RR92, Sal93, SBF93, SCKF98, SF93, SSA99, SLG98, SES94, TBR96, TG96, WGR93, WBYO97, WL98, WH94, WBMM97, YFM98, ZA97, ZM98]. Salmonid [ASB94, CKT96, FW99b, GWB^{+}91, HSL90, MMA^{+}99, TMJ^{+}94]. Salmonidae [Nor97]. Salmonids [BRBV90, BF93, BFJ90, BAG96, BFM99, CBC95, DB94, DR97, GKDG97, HFS^{+}95, HTMP94, IWTC91, JMG95, JNT92, Lar95, LR97, MFP98, Mor93, New97, RB90a, TSP93, TLR97, VWT99, WNFK94, YHW91]. Salmonine [AH93, JS95, Neg95]. Salt [Mue96]. Salt-Gila [Mue96]. Salvage [HB96]. Sample [BM96, DS91, EF90, GHW91, Lyo92, MFP98, Mir93, MHSH96, NW95, SLK94, VBM^{+}95]. Sampled [Fag98, KV96, Pal95b]. Sampler [GB99b]. Samplers [FH92]. Samples [BF98, Dew92, NGW95, SM96]. Sampling [AC92, AHP99, ADS^{+}99, BA90, BG94, BF98, BMVV95, BFJ90, CHTG92, CCT99, Cun95, DW95a, EB90, FB93, HBJ95, HC92, JN95, JSD94, KM98, Kis93, KHP98, MC99, MBHJ95, MSAM96, PW94, Pal95b, PR95, PW98, RC94, SB98, SHS90, SL95, Stu91, VPB95, VWT99]. Sanctuary [JFB99]. Santee [BWW95b, MKK97]. Santee-Cooper [MKK97]. Satisfaction [HD92, MK91, SS92, Spe93]. Satisfactory [MSA^{+}95]. Saturation [HSS98]. Sauger [HCF^{+}96, MBFD98, MWP94]. Saugers [BBH97, HB98, KH97a, PLB96, PBL97]. Saugeye [STS^{+}96]. Saugeyes [FSBB97]. Sawtooth [BLW^{+}97]. Scale [BWF97, BW96, BM95, DGFB99, DG93, FMC97, GLLB97, HB96, KGW93, MC98, PMW98, PD91, SF93, WP98b]. Scales [BDG99, LMR91, MFP98, OS96, WH97, WvdABD93, Wha91]. Scarcity [Smi90]. Scars [Fry98]. Scented [SB96]. Scheduling [HRJP93]. Schemes [Leg98]. Scorpaenidae [MR91]. Scotia [McN95, SD91]. Scotian [SG93]. Scour [HL99b, NF93]. Screen [DMP91, PC99]. Screw [TMJ^{+}94]. Scuba [GB99b, SHS90]. Sea [BSS93a, BG94, BT93, BS91, CMS^{+}99, DGFB99, Han90, HK97, KEC93, LW91, MB90, MSS98, PW94, She91, SF96, Swi99, WP98a, CHGF98, PMW98, ZG90]. Sea-Run [BT93]. Seabrook [SLM^{+}97]. Seam [SB94]. Season [ES92]. Seasonal [BAG96, CMS95, Gab91, GB95, Lan91, LS90, LW91, MTK92, MI98, Muo92b, NW95, PVA98, RS96a, RR93, SW97, WN99]. Seasons [WDR93]. Seatrout [MMD^{+}93, MHN^{+}95]. Second [GWB^{+}91, TS96]. Second-Order [TS96]. Section [BMVV95, GSMG95]. Sectional [ORM92]. Sections [BDG99, Jac95, KBH97, Mat96, SA98]. Sediment [Ave96, CS97a, DGDG91, GB96, GSM93, MN93, NJ96, ROB98]. Sediments [NM91b]. Segmentation [Fis97]. Seine [ASS91, CCQR96, Fag98, OAM98, WW93]. Seiners [Sak91]. Selected [MTT90, OWJ93, RTB99, WV96, WSL96]. Selecting [Wri92]. Selection [BM96, GH93, HM92, HD97, Mat96, MH95, SNF94, Wal92, YHW91, vdAE90]. Selective [JM93, LS96, TGF92]. Selectivities [KJ92]. Selectivity [JN95, MR95, PTK94, SC92]. Self [TMBH93]. Self-Report [TMBH93]. Semi [HKKR92]. Semi-impoundment [HKKR92]. Seminole [MSG99]. Seneca [GKAM94]. Sensitive [BHM98]. Separate [HN93]. Separation [Han95]. Sequencing [HLD99]. Sequential [Cri96, PPQ94]. Sequentially [BT93, UHL97]. Set [KHP98, PH94b]. Set-Lining [PH94b]. Setlines [TH99]. Setting [HLBD96]. Seven [PTS95, TLR97]. Several [BHM92]. Sewage [Wic95]. Sex [DWS97, Ehl97, ZM98]. Sex-Specific [Ehl97]. Sexual [AHB95, DWS97, Sor94]. Shad [AGS^{+}99, BRK91, BMVV95, BHM92, CM99a, DS90, DVG95, DMS96, DG93, HBJ95, Hal96, HK97, HOCSN99, HBM91, HTB94, HWE96, JD92, JR95, KO93, KB97, MBHJ95, Mic98a, Mic98b, Mic99, MG99, Par92, RBJ97, VBM^{+}95, VPB95, VD95]. Shallow [CCT99]. Shape [DMF96, WPE96]. Shaped [KH97a]. Shark [Bro99]. Shelf [OC94, SG93]. Shell [HJ99]. Shellfishing [WS91]. Shelter [MH94]. Shifts [ZHR99]. Shoal [BWG93, PM90]. Shoals [HJ93, SMTW98, WB92b]. Shock [DFE93]. Shore [Elr97, JMG95]. Shore-Spawned [JMG95]. Shoreline [BBW94a, GH93, INJ97, JBH^{+}99]. Short [BRK91, DW95b, GNS97, GITS99, HSCM96, MH97a, MMD^{+}93, McM93, Rin96, RH97, ZS95]. Short-Term [BRK91, DW95b, GNS97, GITS99, HSCM96, MH97a, MMD^{+}93, McM93, Rin96, RH97, ZS95]. Shortnose [CS97b]. Shovelnose [MKKG98, QGB98]. Shrimp [ADS^{+}99, CCS^{+}99, CB96, CRMH94, GC99a, GR97, HL99a, HD98, HG93, HRM99, HL92, KRP^{+}96, NMK94, PWB^{+}98, RH97, SB94]. Shrinkage [MB90]. Side [HN93]. Sierra [Ber94]. Silk [SSM99]. Silver [HTM96]. Silversides [SH92b]. Similar [RQ99]. Simple [HOCSN99, RCVR91, SJ96].

Simulated [BDBN97, BI99, LEW^{+}94, MBFD98]. Simulating [BGJ^{+}98]. Simulation [HTM96, HS94, Joh95, PJM95, SW97, WDR93]. Simulations [Wil99]. Simultaneously [MCCP91]. Single [KHR98, MH97a, MMD^{+}93, SJ96]. Single-Barb [MH97a, MMD^{+}93]. Single-Pass [KHR98]. Site [DG93, GH93, HD97, RTB99, SBT93, WB92a, WF97]. Site-Specific [WB92a]. Sites [BGS90, GBFT98, JSN90, MI98, MH95, PI95]. Situ [AQR95, GK94]. Situk [TMJ^{+}94]. Six [HARS^{+}99, VBM^{+}95]. Size [BLLPD98, BPL93, BF97, BROV98, BM96, Col91a, EF90, GSP93, Goo96a, GWS96, HM96, HDM91, JCP97, JRH92, JM93, KJ92, LH92, LCSA96a, MBFD98, McG99, MC99, MFM99, MM94b, MH94, MP97, Mor93, MMK97, MWF94, NGW95, OWJ93, PF99b, PTK94, PP96, RF92, RQ95, SM98, SJ98a, SLK94, SLG98, SW96, TS96, Tri93, WPE96, Wal92, WLJ91]. Size-Dependent [BROV98, Wal92]. Size-Related [MM94b]. Size-Selective [JM93]. Size-Specific [SW96]. Sized [NA92b]. Sizes [CB96, DB94, Goo96a, HP92, KSJ92, MM93, Mir93, PF99a, WFB97]. Skewed [CMP99]. Skimmer [CRMH94, HRM99]. Skin [HFS^{+}95]. Sliding [NF93]. Sliding-Bead [NF93]. Slope [HOCSN99]. Slopes [IHK99]. Slot [LEW^{+}94, Mar95, Nov90, PT98, PP96]. Small [BSS93a, BPL93, BTS95, BCCEJ96, CLA95, CS97a, DG93, DB94, ESW98, Fag98, FH92, GBFT98, GHW91, GW91, HCO99, HHY96, HB91, HBW94, IH97, JJ98, LB99, MRK98, MCCP91, MHW98, NNR99, Nor99, Nov90, OAM98, PB94b, PTS95, PT98, RF92, RSJB92, SWS94, SN92, TR96, TR98, VKS93]. Small-Mesh [Fag98, OAM98]. Small-Scale [DG93]. Smallmouth [BM94, BN95, BKM99, BKSM91b, HM95, HGH97, Hof95, KWM93, LS90, LKD96, McN95, PH96, PTK^{+}97, RS96a, RS97, SMTW98, TSP93, WB92b, ZS95]. Smelt [BT95, FBR^{+}94, JG99, JGF94, KL91]. Smith [MNN91]. Smolt [BBW94b, CCW^{+}97, DS91, DMP91, OLMS94, TMJ^{+}94, WD95]. Smoltification [HFS^{+}95]. Smolts [AMJM96, BIDN98, BLLPD98, BJ98, BGJ^{+}98, Ell92, HJ93, JMD96, MMA^{+}99, OL96, PWW94]. Smoltsand [MPL99]. Snaggers [SSL96]. Snake [FW99b, GS93, GS95, AMJM96, BMS97, CBB98, GC99b, HTMP94, New97, Pol98, WBYO97, ZA97]. Snapper [CMS^{+}99, GC99a, GR94]. Snorkeling [MHW98, TS96]. Snow [Tay92]. Soak [ZS97]. Social [Coo90, HD97, SS97b, SG97b]. Sockeye [BWW95a, BLT95, Bea95, BLW^{+}97, CCW^{+}97, Fry98, Leo97, McG99, NWHM90, PN91, RR92, SCKF98, WGR93]. Sodium [Pea98, PLL^{+}95]. Sole [JH93, WBPQ92]. Solutions [Moh97]. Some [IB92, MW95]. Sonar [DO98]. Soon [OS96]. Sound [DRG^{+}92, HWE96, NPP^{+}92, RDT^{+}93, RDM^{+}96, Dor95, Pin94]. Sounds [Car99]. South [AFW98, Fie92a, Fie92b, FJ94, BWW95b, CS97b, GS93, GW90, GW91, GW95, HD98, Kir92, KGW93, LW91, MKK97, Pal95b, SSD93, SB94]. Southeast [BFM99, FB96, JDL^{+}98, JQV98, TAL^{+}97]. Southeastern [LH92, Pal95a, Par92, WHR90, ZM98]. Southern [HL99a, CR97, CRS96, DJO97, HSCM96, MWH96, OLMS94, OL96, SBJ^{+}99, SJ98b]. Southwestern [MHR90, Rin96, SS99]. Space [JFB99]. Spacing [SLK94]. Spada [STT^{+}90]. Spaghetti [TH95]. Spanish [CBJ93, CHM98, HD98]. Sparse [KHR98]. Spatial [Ber94, BCCEJ96, DO98, Mic98a, PJN97, RM96, SB99, WP98b]. Spawn [SHS90]. Spawned [JMG95, MKG^{+}93]. Spawning [BAR^{+}92, CKT96, GBFT98, GH93, GC99b, Leo97, MBHC93, MKH91, MI98, MH95, PB94a, PM90, SF93, SS99, Swi99, Uph93]. Spawning-Site [GH93]. Special [BMVV95, GSMG95, HGH97, Sch96a]. Special-Regulation [Sch96a]. Specialization [Fis97]. Species [BNNS95, BHF^{+}99, BF98, FW92, GB98, HL92, Kir98, Lyo92, MKG^{+}93, Pal95b, SB99, SSA99, WH98, WBMM97]. Specific [Ehl97, GW95, SW96, WB92a, WH96]. Spent [KO93]. Sperm [TWF^{+}97]. Spherical [MT92]. Spinal [DMF96, KGFR97]. Spines [WvdABD93]. Splake [Snu92]. Split [DO98]. Split-Beam [DO98]. Sport [AH93, DS90, HSL90, HCO99, HB91, Knu90, KH98, MB97, Mic98a, Mic98b, Mit91, PJMS99, SMH95, SMTW98, SN92, WB92b]. Sport-Fisheries [Knu90]. Sportfish [CKB96]. Sportfishing [Aas95, HRJP93]. Spotted [DMR95, MMD^{+}93, MHN^{+}95, SDBF99, TGM98, WGB96]. spp [SB99]. Spreader [BWO97]. Spring [AH93, BLLPD98, Fry98, Mar92, SF93, ZA97]. Springs [GHG97]. Squawfish [CBC95, HM96, RB90a, TSP93]. Squid [Lan91]. St [FW99a, HL99a]. St. [BBM96, OSS^{+}96, SLE99, TH99]. Stability [GKAM94, SJ96]. Stages [Ber92, Cad90, MSR92, Neg99a]. Stainless [MT92]. Stake [JL92]. Stake-Bed [JL92]. Standard [BWF97, BM91, Chi91, KWM93, KH97b, MP99, PHW93, QGB98, RBB96, WB92a, WGB96, WGM91]. Standard-Weight [Chi91]. Standardizing [BG95]. Standing [FW92, WHR90]. Staples [SSM99]. State [AGS^{+}99, DMS96]. States [AHB95, MCC97, Mir99, Pal95a, Rin96, WPE96, WH98]. Statewide [MB97, TH90]. Station [Rod95, SJ98b, SLM^{+}97]. Stationary [Cun95]. Statistical [JSD94, LF99, Pal95a]. Status [BMY94, DW97a, HSP^{+}95, HD96, JDL^{+}98, MD99, MKK^{+}98, RLT97, SSUL97, TLR97, Uph93, WWTB98, YHN99]. Steel [MT92]. Steelhead [BBM92, CBB^{+}97, CA92, EE92, GHS^{+}97, HMG92, MPP94, MPL99, PWW94, SBT93, Tip91, TG96, VS95, WB92a]. Steep [CR90]. Steeppass [HOCSN99]. Stepped [KL91]. Stepped-Oblique [KL91]. Sterilized [JPT93]. Sterilizing [Han90]. Stizostedion [STS^{+}96]. Stock [BA93a, BWW95a, BPL93, GH91, HN93, Jac95, Jes90b, Jes94, Ove99, PR93, RB90b, Sta92, Uph93, WBPQ92, Wil99]. Stock-Recruitment [Jes90b]. Stocked [BPB^{+}97, BM99, CA92, ESO93, Elr97, HES^{+}94, HM92, HB98, HDM91, HBW94, IBS99, JM93, Jud92, Kay92, KE91, KSJ92, LCGJ99, Mar92, McM95, MWH96, MNN91, Mor93, MKK97, OLMS94, PB95, SWS94, Snu92, STS^{+}96, SS99, SW96, VKS93, WFB97, WL98]. Stocking [BM95, BKSM91b, BMMF99, CBH91, CLA95, CN94, DFW91, EF92, ES92, Fie92a, Fie92b, HDM91, IB92, JP92, JLS^{+}92, JBHW95, LBB92, LCSA96a, LCSA96b, Loc98, Mar99b, MKG^{+}93, MFC^{+}92, MFM99, ML92, MSA^{+}95, Mit92, PP97a, Sal93, SMMC96, SJ98a, Wah99, WWSF93a, WLC97]. Stockings [PK92a]. Stocks [AGG96, BBH92, EF92, Han95, HGW94, MC98, MH97b, Max99, NWHM90, Pow96, SF93, SNS99, WPTW93, WHR90]. Storage [MTT90, PN91, TWF^{+}97]. Strain [KKM93, LPT99]. Strains [Dwy90, HT92b, Neg99a, NA94, OL96, VKS93]. Strait [CCQR96]. Stranded [Goo90]. Stranding [BTAH95]. Strategic [CP95]. Strategies [CR97, DCPP97, Mar99a, Mar99b, SHS90]. Strategy [MKG^{+}93, ME93]. Stratification [LB96, WD95]. Stratified [DS91, NRA97]. Stream [Aad93, ANN91, ASB94, BM94, BM97, Ber94, Bin94, BCCEJ96, BN94, CBB^{+}97, CHR93, DB94, DR97, FB93, GBFT98, GP99, HSCM96, HLDH98, Hou96, IHK99, JQV98, KBH97, LB99, Lyo92, MBHC93, Mat96, MK91, MFP98, MFP91, MS95, Nor99, ORM92, OAM98, OL90, OLMS94, OL96, PH96, PR95, PR96, RHSB98, RS95, Sch96a, SL95, SH92a, SJ96, TS96, WSL96, WF93, Wri92, YHW91, ZS95]. Stream-Resident [GP99]. Streambed [RC94]. Streambeds [NF93]. Streamer [RCVR91]. Streamer-Type [RCVR91]. Streams [AC92, ASS91, Ave96, BTS95, CR97, CS97a, CDM^{+}98, DBJW98, EH93, FD95, FBC97, FN92, GB99a, GHW91, HSS99, HHY96, HMG^{+}96, IH97, JS95, KM96, KHR98, LH92, LB92, LWS96, LKD96, MPL99, Mor93, MHW98, O’S95, Pal95a, Pal95b, PH94a, Pet93, PMN99, RQ99, RF92, RSJB92, SRCIW96, SLK94, SG97a, SS99, SP96, TR96, TR98, TAL^{+}97, TB96, TGM98, VKS93, WLK98, Wat92, Wic95, WWSF93b, ZHR99]. Streamside [BR91]. Strength [LCSA96b, Mit91, SB98]. Stress [CBW97, GSM93, MN93]. Stressed [BWG93]. Stressors [WH98]. Strip [RMV96]. Striped [BO98, BJ98, BM91, BWW95b, BW94, DFW91, Dor95, DG93, HARS^{+}99, JL98, JNT92, KBM90, MNN91, MMD91, NNR99, Nel98, RR99, RIT97, RM90, SBH96, Uph93, Uph97, WDM91, WV96, WYL^{+}99, WvdABD93, WRDY91, WF97, WW93, vdAE90]. Strobe [NA92a]. Structural [ASB94, BA93a, GW90]. Structure [GSP93, GWS96, Jac95, JCP97, JRH92, LH92, NGW95, SC92, TS96, WO91]. Structured [HTM96]. Structures [BBW94a, FN92, Gra92, Hou96, JL92, NF99, RB99, RKHW98, SNF94, WLJ91]. Studies [Cad90, MKKR99, VWT99, WBYO97, WBPQ92]. Study [BMVV95, BTAH95, GP93, GSC96, JW98, MT91, MS95, WD95]. Stunted [AKG93]. Stunting [BDBN97]. Sturgeon [BB99, CS96a, KM93, MS96, MKKG98, MKK^{+}98, MKKR99, PB94a, QGB98, RB90b, SLE99, TH99]. Sturgeons [Clu96, CRS96, CS97b, CMP99, Hal98, Kem96]. Subadult [BMMF99]. Submersed [GS95, MRL93]. Submersible [OC94, PC99]. Subsequent [MR96]. Subspecies [DMR95]. Substrate [EH93, GHW91, KV96, MG97, PK94, WF93, WP98b, YHW91]. Substrates [MS96, YHW91]. Subyearling [BRK91, BMS97, CBB98]. Success [Fie92a, GBFT98, JP92, JL92, KSS93, LGS96, Loc98, MK91, MR96, PH94a, PTK^{+}97, Wah99, WF93]. Sucker [CWGT96, MR97, TWF^{+}97, TSKS99].

Suckers [BKM99, MI98, MR96]. Suction [HL99b, SH92a]. Suction-Dredge [SH92a]. Suitability [ESW98, FBC97, FBB99, SJ98b, WSF99]. Summary [GSMG95]. Summer [BBW94a, CBB98, Fry98, GP99, MJP91, WN97]. Sunfish [KBF99]. Superior [BSS93b, CA92, CBS93, GNS97, HES^{+}94, HSP^{+}95, HSS98, Neg95, OS96, OSS^{+}96, PJMS99, SF98, SLE99]. Supersaturation [BRK91]. Supplemental [CN94, Hal96, HDM91, IB92, LBB92, RIT97]. Supplementation [BBM92, WH94]. Supply [FT93, PN91]. Support [JFW96, SJ98b]. Surface [FS97, HBJ95, IHK99, Leo97]. Surgery [JL98]. Surgical [SSM99]. Surgically [MLP95]. Surrogates [ZB94]. Survey [ADS^{+}99, CCS^{+}99, DJO97, Dol91, GBTD99, JQV98, Lan91, Mal95, MCC97, NRA97, PIM95, TH90, Uph93, Wal92, ESB^{+}96]. Surveyed [Pal95b]. Surveys [IH97, JH90, Loc97, PHJ^{+}97, RTB99, RS95, SHS90, SG93, SJ98b, WW93]. Survival [BF97, BPB^{+}97, BCC93, BS91, CCQR96, CA92, CMS^{+}99, CS99, CC98, DMF96, DFW91, DFE93, DE95, ES92, ESO93, Elr97, FHMT97, FW99a, FW99b, GNS97, GR94, Gre92, HM96, HES^{+}94, HM92, HB98, HDM91, HBW94, JMG95, JSN90, KBM90, KBHM96, KE91, KSS93, KSJ92, LCGJ99, LSS91, Loc98, Mar92, MLP95, MFM99, McM95, MH94, MNN91, Neg99a, NA94, OLMS94, PB95, PH94a, PK94, PPQ94, PG93, ROB98, RW92, Rot94, Snu92, STS^{+}96, SW95, SW96, TG96, VKS93, WL98, WWSF93b, YHW91]. Susceptibility [CHF^{+}97, SH92b]. Suspend [GG98]. Suspended [GSM93, MN93, NM91b, NJ96]. Suspension [MFBM90]. Susquehanna [HTB94, JD92]. Sustainable [Goo96b]. Sustained [MW95]. Sutures [SSM99]. Swim [BCC93, CMS^{+}99, SS97c]. Swimming [AMC97, CSCM99]. Sympatric [SA98]. Symposium [EF92]. Synthesis [JBHW95, NJ96]. Synthetic [RB99]. System [BM95, BC94, BG90, BWW95b, GH91, GHS^{+}97, HFS^{+}98, HP97, HCF^{+}96, Jud92, Kem96, KO93, KB97, MMA^{+}99, MFP91, MRS98, MKK^{+}98, Ste99, WLK98, vdAE90].

T [Clu96]. T-Bar [Clu96]. Tag [BN94, EF90, GSC96, MDI^{+}96, MT91, New97, OB95, PT93b, PI95, RCVR91, TH93, ZB94]. Tag-Reporting [ZB94]. Tag-Return [MT91]. Tagged [BGS90, CCC^{+}96, HP97, KBHM96, MMA^{+}99, PLB96, Tay92, CBB98]. Tagging [AQR95, DR97, HARS^{+}99, PK92b, SRL97, WBPQ92]. Tags [AMJM96, BIDN98, BRC94, BSS93a, BT93, BSN95, BD99, Clu96, FB96, HG98, HARS^{+}99, MPP94, MFG94, Muo92a, NH98, OB93, PPQ94, PT93b, SC91, SRCIW96, SBH96, TH95, UHL97, WDM91, WG95]. Tahoe [BAR^{+}92, BBW94a]. Tailings [HL99b]. Tailwater [PH96]. Take [HCO99, JFB99, MSA^{+}95]. Taken [MP90]. Tampa [MT91]. Tandem [KHP98]. Tandem-Set [KHP98]. Targets [CM96]. Taupo [Ded96]. Taylorsville [BM93, BM99]. Technique [BG94, BM96, ES92, Hal98, HSM94, KL91, MWL95, RC94, RCVR91, WV94, Wha91]. Techniques [CWGT96, Ded96, DJO97, HM96, RSJB92, SF91]. Teeth [LGS96]. Telemetry [PBL97]. Temperate [JBH^{+}99, PR96]. Temperature [BI99, DGDG91, GB95, Hal99, KJS95, MK91, MTT90, PH96, vdAE90]. Temperatures [SJ96]. Temporal [Ber94, BCCEJ96, DO98, Hou95, Mic98a, RM96, SB99, WD95]. Tennessee [PLB96, PBL97, FSBB97, HD96, IBS99, Jac95, KMHS95, MBFD98, MB98, MB91, SDBF99, SB99, SMTW98, WB92b]. Term [BRK91, Bin94, DMF96, DW95b, GNS97, GITS99, HSCM96, MH97a, MMD^{+}93, McM93, Rin96, RH97, Tay92, TBNB97, WDM91, WLC97, ZS95]. Test [HWH93]. Testing [MBHC93]. Tests [FH92]. Tetracycline [HBM91]. Texas [Chi91, Dol91, DD97, KWW90, MBK^{+}91, MWF94, MK97, Muo94, MBD^{+}92, OWSM99, SMH95, SD93, WO91, WSOM98, WRD98]. Texoma [SMH95, DW95a, NGW95]. Their [Aad93, BR94, CCW^{+}97, CN94, Fie92a, GKM93, GS95, Mal95, Mic98b, SMTD95, SG97a, YHN99, GP93]. There [HN93]. Thermal [AH93, CBW97, LB96, LT98, Neg99b, PR96, SJ96]. These [Pow94]. Threadfin [BMVV95, HBJ95, Hal96]. Threatened [FE95]. Three [Ave96, DJO97, Dwy90, Edw91, Gre95, GKAM94, GHW91, HSCM96, HJ94, HCO99, HP92, HT92b, KGFR97, KSJ92, LSS91, MSG99, MHR90, MC99, NF99, NA94, OL96, PK92a, PF99a, PH94a, PI95, RSJB92, SB99, SG97a, WO91]. Three-Pass [HSCM96]. throughout [GS95]. Tidal [SD91]. Tidewater [LSA99]. Tiger [SN92, TH93]. Time [HSM94, MTT90, SLG98, ZA97, ZS97]. Time-Lapse [HSM94]. Times [ME93]. Timing [AMJM96, OL96, SLG98]. Tire [ML94]. Tissue [TSKS99, VWT99, WN99]. Tohopekaliga [MHS^{+}95]. Tolerance [BRK91]. Tolerances [WH98]. Tombigbee [Jac95]. Tongs [CHTG92]. Tongue [SB94]. Tool [CS96c, Neg95, SB98]. Tools [GSP93, Joh95]. Tooth [HTMP94]. Toronto [Wic95]. Total [DGFB99, FAW98, HCFG96, Leg98, NM99]. Tournament [Dol91, Goe91, MM94b, WRD98]. Tournament-Caught [MM94b]. Tournaments [FJ94, GBTD99, HM95, HCF^{+}96, KH95, OWSM99, SGWN94, WN97, WSOM98]. Towed [Lyo92]. Toxic [CU90]. Trade [SM94]. Traditional [Smi98]. Training [WWSF93b]. Traits [HA95, Neg99a, Pow94]. Transect [SLK94, SJ98b, VD95]. Transect-Based [SJ98b]. Transferability [FBC97, FBB99, WSF99]. Transmission [Pet93]. Transmitter [BHM98, PMBM97]. Transmitters [BI99, CSCM99, HT92a, MLP95, MLSH^{+}99, SG97b]. Transplanted [KKS95]. Transponder [AMJM96, BD99, Clu96, New97, PPQ94]. Transponders [MMA^{+}99]. Transportation [Pol98, WBYO97]. Transported [BMS97, CCW^{+}97]. Trap [AHP99, CMS95, CCC^{+}96, DMP91, ECS98, Ell92, GWS96, Kel93, KWRH94, MKH91, MSAM96, MCCP91, RS96b, ZS97]. Trap-Net [CMS95]. Trapping [Stu91]. Traps [Ave96, BPC91, Gui93, Kis93, ROB98, RM92, Stu91, TMJ^{+}94]. Travel [ZA97]. Traverse [WPTW93]. Trawl [AC92, CRMH94, GC99a, GR97, HS93, Han95, PMW98, PWB^{+}98, Ric94, RFS95, SG93, Sta92, WPE96, Wal92, WKG98]. Trawl-Caught [RFS95]. Trawlable [MR91]. Trawlers [HD98]. Trawling [CMP99, CCC^{+}96, KL91]. Trawls [AHP99, EP94, HRM99, MBHJ95]. Treatment [DW97b]. Treatments [DB94]. Treble [MMD^{+}93]. Tree [JL92]. Trees [BHF^{+}99, RMV96]. Trends [Bro99, HWE96, MH97b, PH90, RLT97, SH99]. Trial [HJ99]. Trials [DR97]. Tributaries [BC91, CA92, KWRH94]. Tributary [Gre91, Jud92, WL98]. Trip [HD92, Ric94]. Triploid [BC93, BTT^{+}93, CLA95, CHTC93, Kir92, MKK97, PM97, SSD93]. Troll [OWJ93]. Troll-Caught [OWJ93]. Trophic [AGS^{+}99, DMS96]. Trophies [LEW^{+}94]. Trophy [Jac96, NA92b]. Trophy-Sized [NA92b]. Tropical [WDR93, WWTB98]. Trough [Ell92]. Trout [APP98, AMC97, Ave96, BM97, BAR^{+}92, Ber94, BW96, Bin94, BR94, BT93, BTS95, BWG93, Bor94, BR91, BTAH95, BSN95, BN94, BI99, CBH91, CR97, CHF^{+}97, DMF96, DBJW98, Ded96, DB91, DW91, DWS97, DVR97, Dwy90, DFE93, DE95, DW95b, DW97b, EH93, EF90, ES92, ESO93, Elr97, FHMT97, FD95, FW92, FB96, GNS97, GB95, GZB98, GB99a, GP93, Gre92, Gre91, GLLB97, GKAM94, GS93, GS95, GP99, HSCM96, HSS99, HG98, HES^{+}94, HSP^{+}95, HSS98, HJ94, HT92a, HDM91, HCO99, HHY96, HN95, Hog99, HC94, Hou95, HGMS94, HGW94, HMG^{+}96, HT92b, JRH92, JBHW95, JW98, JDL^{+}98, Jow92, KBH97, KE91, KKM93, KV96, KM96, KGFR97, KMHS95, KH97b, KHR98, LH92, LGS96, LB96, LF99, LEW^{+}94, MD99, MC98, MBHC93, MKG^{+}93, MLP95, Mat96, Max99]. Trout [MDI^{+}96, MK91, McM93, MG97, MTK92, MSA^{+}95, MFP91, MWH96, Mor93, MFG94, MHW98, Neg99a, Neg99b, Nor99, NM99, NA92b, NA94, OS96, OL90, OB93, PT93a, PH94b, Pet93, PHW93, PBC98, PP96, QKH^{+}94, RK95, RM96, RLT97, RF92, RW92, Sal93, Sch96a, SS97b, SB96, SF98, SCS^{+}94, SCS^{+}95, SRCIW96, SSUL97, SNF94, SSD93, SBJ^{+}99, Snu92, Sor94, STT^{+}90, SS99, SG97b, SSM99, TLW96, TW92, TW97, TR96, TBN97, TBNB97, TR98, TAL^{+}97, TS96, Tri93, TB97b, TSKS99, VKS93, WFB97, WW94, Wat92, WH97, WF93, WG95, WWSF93a, WWSF93b, WLC97, WHR90, Wri92, YHN99]. Trucking [JSN90]. Tuna [Sak91]. Turbid [AC92]. Turbidity [DGE^{+}97]. Turbine [Cad90, NMB96, SFZ98, SD91]. Turbines [DG93]. Twenty [YHN99]. Twenty-Five [YHN99]. Two [AHP99, BSS93a, BKM99, CHCC95, DS91, GG98, Han95, HHY96, HT92b, JL98, KV96, KSS93, KSJ92, Mar95, McH90, MD93, MCCP91, MR96, NWHM90, PMN99, PVP99, Pow94, RHSB98, RM97, Sch99, SA98, SSA99, SB94, TLW96, VKS93, WFB97]. Two-Compartment [MCCP91]. Two-Sample [DS91]. Two-Seam [SB94]. Type [Cad90, HT92a, RCVR91]. Types [Aad93, JL98, MSG99, OWJ93, RS95]. Typology [HD92, Mir99].

U.S. [Edw99, ESO93, EP94, GC99a, HES^{+}94, HTM96, Sak91]. ui [ESB93]. Uinta [KBH97]. Ultimate [CRC99]. Ultrasonic [OB95]. Uncertainty [JMBK92, RR96, SLM^{+}97, WDR93]. Underestimation [RF92].

Understanding [WD94]. Underwater [AQR95, GG98, HMG92]. Unfed [WL98]. Unionid [Sch96b]. Unionids [Sch96b]. Unit [GWS96, KM98, Lyo92, Sta92]. United [Pal95a, Rin96, WH98, AHB95, MCC97, Mir99, WPE96]. Units [Gre95]. Untagged [MMA^{+}99]. Untrawlable [MR91]. Unweired [KRP^{+}96]. upon [Mit91]. Upper [SBT93, AAMB97, FW99a, GPB95, JJ98, KJS95, LBB92, MSR92, Pit97, SSUL97, Uph97, ZDK99]. Upper-River [SBT93]. Upstream [MCCP91, TR98]. Upwelling [GBFT98]. Urban [SD93]. Urbanization [Wic95]. USA [CRS96]. Use [AMJM96, AMC97, BBW94a, BC93, CS97a, DO98, DGFB99, FD95, FS97, Gre91, GS93, GHG97, Hal96, HS93, HTB94, HKKR92, HP97, IH97, JP92, JL92, JR98, Kay92, MRL93, MOAS98, MJP91, MT92, MFP91, MI98, MSR92, OB95, PK94, PP97b, PD91, PLL^{+}95, RCRR92, RBJ97, SLM^{+}97, SK95, SA98, Smi98, TR92, TH93, Uph97, WB92a, WP98b, ZDK99]. Used [BHF^{+}99, BGS90, Edw91, GHW91, GG98, GC99b, RIO^{+}97, SMH95, VBM^{+}95]. Usefulness [ASB94]. Users [SD93]. Uses [Ove99]. Using [BDG99, BW94, CK99, DS91, EDGR97, KBF99, LMR91, LR97, Max99, WS91]. Utah [GH93, HDM91, Hog99, KBH97, MWH96, MJHG97, MI98, TLW96]. Utility [Cla94, GSM93, MN93, RB99]. Utilization [LW91].

V [Sch92]. Vacuum [SM96]. Validate [PD91]. Validation [CM99a, KBF99, LWS96, SK91]. Valley [BLW^{+}97, YFM98]. Value [AM95, DBJW98, Jac96, Kir98, KH98, Mil91]. Value-per-Recruit [Jac96]. Values [ANN91, BT90, GP93]. Vancouver [MR91]. Vapor [NA92a]. Varden [FB96]. Variability [Ber94, Goo96b, HC92, MM91, RM96, RS95, VD95]. Variable [All97, DMP91, Goo96a]. Variables [BF93, GLLB97, Jes94, Mit91]. Variance [Pal95a, Sta92]. Variation [BWW95a, BBH92, CMS95, FCP97, Hou95, HA95, KKM93, MB98, Mic98a, NW95, RQ95, SB99, She91]. Variations [WN99]. Various [LEW^{+}94, NGW95, PIM95]. Varying [WF93]. Vegetated [CCT99, MMS^{+}98]. Vegetation [BMNB92, BMNB93, Dew92, Kir92, MRL93, MP97, PVP99, Sch99]. Vehicle [OC94]. Velocity [CKT96, MP90]. Vendace [SNS99]. Venezuela [Mon94]. Vents [ESP92]. Vermilion [CMS^{+}99]. Vermont [McM95, WL98]. versus [CW94, PH90, TP94]. Vertical [CB96]. Vessel [HS93]. Vessels [CCQR96]. Viability [LR97, WH96]. Vibert [GB96]. Video [GG98, HK97, HSM94, HFS^{+}98, IWTC91]. Virginia [Gra92, MNN91, SA98]. Visible [BIDN98, BT93, BN94, FB96, MDI^{+}96, MFG94, SRCIW96, WG95]. Visual [AQR95, BR91, FS97, SJ98b]. Visually [Gra92]. vitreum [STS^{+}96]. Volitional [RR93]. Volitionally [EE92]. Voltage [DE95]. Volumetric [DGFB99]. Vulnerability [NA94, SW95]. VVP [VHW^{+}96]. VVP-15 [VHW^{+}96].

W [MP99, MBS90]. Wading [RW92]. Walleye [Bac95, BHY^{+}97, Ber92, BBH92, CBW97, CC98, CHGF98, EF92, Fie92a, Fie92b, FJ94, FH92, Goe91, HSS^{+}94, HCFG96, HCF^{+}96, Jac96, JP92, JGF94, LCSA96a, LCSA96b, MH97b, MFC^{+}92, ML92, MWP94, Mit92, MBS90, MTT90, PK92a, PP97a, PMW98, PP97b, Qui92]. Walleyes [BBH97, BROV98, BHK94, CCC^{+}96, FSBB97, HBN^{+}98, JLS^{+}92, Jud92, Kay92, KSJ92, LBB92, MK97, NH98, PK92b]. Walter [SGWN94]. Warm [GHG97]. Warmwater [CLA95, FBC97, Kis93, LB92, NNR99, PR95, PR96, Sch92, SFZ98]. Warnings [CU90]. Was [WP98a]. Washington [BLT95, Bea95, BPB^{+}97, CBB^{+}97, FN92, PWW94, PWB^{+}98, STT^{+}90, TSP93]. Water [BHF^{+}99, CRC99, GPB95, HW94, IHK99, Leo97, MB97, PH96, PD91, Rod95, SM96, SJ96, TBF93]. Watermilfoil [UHP99]. Waters [BSMM90, ESO93, FW92, FBR^{+}94, GSMG95, GW95, HES^{+}94, HSP^{+}95, HSS98, HD96, KGW93, NN91, Neg95, Par92, SM98, SNS99, WV96]. Watershed [BFM99, O’S95]. Watersheds [DJO97, FCP97]. Waterway [Jac95, WNFK94]. Wave [DE95]. Way [Pet93]. Weakfish [MH97a]. Web [JG99, RO98]. Weight [AHB95, Chi91, GK94, KM93, KWM93, KH97b, MH97b, MP99, MBS90, MWC92, NM91a, PHW93, QGB98, RBB96, WGB96, WGM91]. Weights [BM91, HGW94, PP97b]. Weired [KRP^{+}96]. Weiss [MOAS98, RD91]. Wesche [CS97a]. West [EP94, McM95, WPE96]. Western [HN93, Sal93, SLE99, BSS93b, FN92, Hou95, TB99]. Westslope [DWS97, LGS96, SRCIW96, SSUL97, WF93]. Wetlands [MMK97]. Wetted [RC94]. White [BWF97, BM91, BKM99, Col91b, CMP99, Gab91, GC90, GWS96, Hal96, Hal99, HM91, KRP^{+}96, MBK^{+}91, MJP91, MB91, Mon94, MMD91, Muo92a, Muo94, NM91a, PB94a, PM97, RD91, RB90b, SMTD95, SK91, WYL^{+}99, WO91]. Whitefish [CBS93, RBB96, WPTW93]. Whitlock [GB96]. Whitlock-Vibert [GB96]. Wide [BR94]. Wild [CBH91, CHF^{+}97, DMF96, FB96, GITS99, HTB94, HC94, HJ93, IB92, JDCS97, JMD96, KGFR97, LPT99, MC98, MPL99, Mor93, NA92b, NA94, RS96b, SRL97, STT^{+}90, SS99, WW94, WG95]. Wilderness [FD95, KBH97, KM96, Mat96]. Wildfire [Rin96]. Willamette [WNFK94]. William [Pin94]. Winnebago [HCF^{+}96, Kem96, SO96]. Winter [BTAH95, EE92, GS93, GS95, Kir98, KJS95, KSS93, MG97, MH94, MP97, RIO^{+}97]. Wire [BSS93a, BT93, DR97, GSC96, HG98, OB93, PK92b, PPQ94, PI95, TH93, UHL97]. Wisconsin [PVA98, Ave96, BK99, BBF99, CM99b, Fag98, HSL90, HBN^{+}98, Hof95, HN95, HCF^{+}96, JLS^{+}92, KLN97, Kem96, LS90, LWS96, LKD96, Mar92, MRK98, Mar99b, MTK92, NRA97, PVA98, SH99, SO96, SS99, UHP99, WLK98]. Within [VD95, JFB99, ROB98, RLT97]. Within-Transect [VD95]. without [MLP95, SS91]. Wolf [LS90]. Wood [BHF^{+}99]. Woody [CBB^{+}97, CHR93, FD95, HLDH98]. Woolen [KLN97]. Work [HS93]. Workshop [GSMG95]. World [SM94]. Wounds [Fry98]. WS [KWM93, WGB96, WGM91]. Wyoming [GH91, Bin94, BR94, DBJW98, EH93, GB99a, GH93, HGMS94, HMG^{+}96, JRH92, KBH97, KH97a, LH92, MH97b, MJHG97, MH95, WWSF93a, WHR90].

Year [BHM92, CN94, CBS93, GKM93, GKAM94, HM96, HDM91, LCSA96b, Mic99, Mit91, SB98, SO96, SK91, WP98a, WRDY91]. Year-Class [LCSA96b, Mit91, SB98]. Year-Classes [CBS93, GKAM94, WP98a]. Yearling [BMS97, CN94, CBS93, HN95, JSN90, Mar92]. Years [Ant90, JG99, KGFR97, MW90, YHN99]. Yellow [AFW98, BK99, BSS93b, FAW98, GW91, JMBK92, KBHM96, KJ92, MR95, NF99, SM98, WGM91]. Yellowstone [SSL96, DW97b, Gre91, GLLB97]. Yield [Goo96b, LEW^{+}94, MOAS98, STT^{+}90, TMJ^{+}94, ZG90]. Yields [PH90]. York [GKAM94, MFM99, Pet93, WV96]. Young [BHM92, CN94, Kem96, Kis93, Mit91, SK91]. Young-of-the-Year [SK91, WRDY91]. Young-of-Year [BHM92, CN94].

Zealand [Car99, Ded96, HJ94, Jow92, MS92, TR92]. Zebra [Sch96b, TB99]. Zone [BBW94a].

References

Aadland:1993:SHT

[Aad93] Luther P. Aadland. Stream habitat types: Their fish assemblages and relationship to flow. North American Journal of Fisheries Management, 13(4):790–806, November 1993. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Agostinho:1997:AFI

[AAMB97] Carlos Sérgio Agostinho, Angelo Antonio Agostinho, Elineide Eugenio Marques, and Luis Mauricio Bini. Abiotic factors influencing piranha attacks on netted fish in the Upper Paraná River, Brazil. North American Journal of Fisheries Management, 17 (3):712–718, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Aas:1995:CSE

[Aas95] Øystein Aas. Constraints on sportfishing and effect of management actions to increase participation rates in fishing. North American Journal of Fisheries Management, 15(3):631–638, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Aadland:1992:ETS

[AC92] Luther P. Aadland and Chantel M. Cook. An electric trawl for sampling bottom-dwelling fishes in deep turbid streams. North American Journal of Fisheries Management, 12(3):652–656, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ault:1999:ESS

[ADS^{+}99] Jerald S. Ault, Guillermo A. Diaz, Steven G. Smith, Jiangang Luo, and Joseph E. Serafy. An efficient sampling survey design to estimate pink shrimp population abundance in biscayne Bay, Florida. North American Journal of Fisheries Management, 19 (3):696–712, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:1998:RBL

[AFW98] Marcy R. Anderson, Shannon J. Fisher, and David W. Willis. Relationship between larval and juvenile yellow perch abundance in Eastern South Dakota glacial lakes. North American Journal of Fisheries Management, 18(4):989–991, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Antonsson:1996:ECF

[AGG96] Thorolfur Antonsson, Gudni Gudbergsson, and Sigurdur Gudjonsson. Environmental continuity in fluctuation of fish stocks in the North Atlantic Ocean, with particular reference to Atlantic salmon. North American Journal of Fisheries Management, 16 (3):540–547, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:1999:RLB

[AGS^{+}99] M. S. Allen, J. C. Greene, F. J. Snow, M. J. Maceina, and D. R. DeVries. Recruitment of largemouth bass in Alabama reservoirs: Relations to trophic state and larval shad occurrence. North American Journal of Fisheries Management, 19(1):67–77, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Aultman:1993:STF

[AH93] Dana C. Aultman and James M. Haynes. Spring thermal fronts and salmonine sport catches in Lake Ontario. North American Journal of Fisheries Management, 13(3):502–510, August 1993.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Almeida:1995:LWR

[AHB95] Frank P. Almeida, Dana-Lyn Hartley, and Jay Burnett. Length– weight relationships and sexual maturity of goosefish off the Northeast Coast of the United States. North American Journal of Fisheries Management, 15(1):14–25, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:1998:FRB

[AHC98] M. S. Allen, M. V. Hoyer, and D. E. Canfield, Jr. Factors related to black crappie occurrence, density, and growth in Florida lakes. North American Journal of Fisheries Management, 18(4): 864–871, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:1999:CTN

[AHP99] M. S. Allen, M. M. Hale, and W. E. Pine III. Comparison of trap nets and otter trawls for sampling black crappie in two Florida lakes. North American Journal of Fisheries Management, 19(4): 977–983, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Amundsen:1993:RSP

[AKG93] P.-A. Amundsen, A. Klemetsen, and P. E. Grotnes. Rehabilitation of a stunted population of Arctic char by intensive fishing. North American Journal of Fisheries Management, 13(3):483– 491, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:1997:EVR

[All97] M. S. Allen. Effects of variable recruitment on catch-curve analysis for crappie populations. North American Journal of Fisheries Management, 17(1):202–205, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:1995:EVH

[AM95] M. S. Allen and L. E. Miranda. An evaluation of the value of harvest restrictions in managing crappie fisheries. North American Journal of Fisheries Management, 15(4):766–772, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:1997:UCO

[AMC97] W. Gary Anderson, R. Scott McKinley, and Maria Colavecchia. The use of clove oil as an anesthetic for rainbow trout and its effects on swimming performance. North American Journal of Fisheries Management, 17(2):301–307, May 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Achord:1996:UPI

[AMJM96] Stephen Achord, Gene M. Matthews, Orlay W. Johnson, and Douglas M. Marsh. Use of passive integrated transponder (PIT) tags to monitor migration timing of Snake River Chinook salmon smolts. North American Journal of Fisheries Management, 16(2): 302–313, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Angermeier:1991:ASV

[ANN91] Paul L. Angermeier, Richard J. Neves, and Larry A. Nielsen. Assessing stream values: Perspectives of aquatic resource professionals. North American Journal of Fisheries Management, 11 (1):1–10, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anthony:1990:NEG

[Ant90] Vaughn C. Anthony. The New England groundfish fishery after 10 years under the Magnuson Fishery Conservation and Management Act. North American Journal of Fisheries Management, 10(2):175–184, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ainslie:1998:EPC

[APP98] Barbara J. Ainslie, John R. Post, and Andrew J. Paul. Effects of pulsed and continuous DC electrofishing on juvenile rainbow trout. North American Journal of Fisheries Management, 18(4): 905–918, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Adkison:1995:INM

[AQR95] Milo D. Adkison, Thomas P. Quinn, and Otto C. Rutten. An inexpensive, nondisruptive method of in situ dart tagging for visual recognition of fish underwater. North American Journal of Fisheries Management, 15(2):507–511, May 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Acheson:1997:BTB

[AS97] James M. Acheson and Robert S. Steneck. Bust and then boom in the Maine lobster industry: Perspectives of fishers and biologists. North American Journal of Fisheries Management, 17(4): 826–847, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Austen:1994:USC

[ASB94] Douglas J. Austen, Dennis L. Scarnecchia, and Eric P. Bergersen. Usefulness of structural and condition indices in management of high-mountain stream salmonid populations. North American Journal of Fisheries Management, 14(4):681–691, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Angermeier:1991:ESC

[ASS91] Paul L. Angermeier, Roy A. Smogor, and Steven D. Steele. An electric seine for collecting fish in streams. North American Journal of Fisheries Management, 11(3):352–357, August 1991. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avery:1996:EST

[Ave96] E. L. Avery. Evaluations of sediment traps and artificial gravel riffles constructed to improve reproduction of trout in three Wisconsin streams. North American Journal of Fisheries Management, 16(2):282–293, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Auster:1997:CIB

[AWR97] Peter J. Auster, Les Watling, and Alison Rieser. Comment: The interface between fisheries research and habitat management. North American Journal of Fisheries Management, 17(2):591– 595, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bayley:1990:MSE

[BA90] Peter B. Bayley and Douglas J. Austen. Modeling the sampling efficiency of rotenone in impoundments and ponds. North American Journal of Fisheries Management, 10(2):202–208, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baigun:1993:SIS

[BA93a] Claudio R. M. Baigun and Richard O. Anderson. Structural indices for stock assessment of and management recommendations for pejerrey Odonthestes bonariensis in Argentina. North American Journal of Fisheries Management, 13(3):600–608, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bendock:1993:HMC

[BA93b] Terry Bendock and Marianna Alexandersdottir. Hooking mortality of Chinook salmon released in the Kenai River, Alaska. North American Journal of Fisheries Management, 13(3):540– 549, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baccante:1995:ACI

[Bac95] Dominic Baccante. Assessing catch inequality in walleye angling fisheries. North American Journal of Fisheries Management, 15 (3):661–665, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brege:1996:SDP

[BAG96] Dean A. Brege, Randall F. Absolon, and Ritchie J. Graves. Seasonal and diel passage of juvenile salmonids at John Day Dam on the Columbia River. North American Journal of Fisheries Management, 16(3):659–665, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beauchamp:1992:LTS

[BAR^{+}92] David A. Beauchamp, Brant C. Allen, Robert C. Richards, Wayne A. Wurtsbaugh, and Charles R. Goldman. Lake trout spawning in Lake Tahoe: Egg incubation in deepwater macrophyte beds. North American Journal of Fisheries Management, 12(3):442–449, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baker:1999:LSA

[BB99] Edward A. Baker and David J. Borgeson. Lake sturgeon abundance and harvest in Black Lake, Michigan, 1975–1999. North American Journal of Fisheries Management, 19(4):1080–1088, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bozek:1999:DMN

[BBF99] Michael A. Bozek, Thomas M. Burri, and Richard V. Frie. Diets of muskellunge in Northern Wisconsin lakes. North American Journal of Fisheries Management, 19(1):258–270, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Billington:1992:MIM

[BBH92] Neil Billington, Richard J. Barrette, and Paul D. N. Hebert. Management implications of mitochondrial DNA variation in walleye stocks. North American Journal of Fisheries Management, 12(2): 276–284, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Billington:1997:FNH

[BBH97] Neil Billington, Ronald C. Brooks, and Roy C. Heidinger. Frequency of natural hybridization between saugers and walleyes in the peoria pool of the Illinois River, as determined by morphological and electrophoretic criteria. North American Journal of Fisheries Management, 17(1):220–224, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Byrne:1992:MRN

[BBM92] Alan Byrne, T. C. Bjornn, and J. D. McIntyre. Modeling the response of native steelhead to hatchery supplementation programs in an Idaho River. North American Journal of Fisheries Management, 12(1):62–78, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowering:1996:CAC

[BBM96] W. R. Bowering, W. B. Brodie, and M. I. Morgan. Changes in abundance and certain population parameters of American plaice on St. Pierre Bank off Newfoundland during 1972–1994, with implications for fisheries management. North American Journal of Fisheries Management, 16(4):747–769, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boogaard:1996:EPC

[BBSJ96] Michael A. Boogaard, Terry D. Bills, James H. Selgeby, and David A. Johnson. Evaluation of piscicides for control of ruffe. North American Journal of Fisheries Management, 16(3):600– 607, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beauchamp:1994:SHU

[BBW94a] David A. Beauchamp, Earl R. Byron, and Wayne A. Wurtsbaugh. Summer habitat use by littoral-zone fishes in Lake Tahoe and the effects of shoreline structures. North American Journal of Fisheries Management, 14(2):385–394, May 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beechie:1994:ECS

[BBW94b] T. Beechie, E. Beamer, and L. Wasserman. Estimating Coho salmon rearing habitat and smolt production losses in a large river basin, and implications for habitat restoration. North American Journal of Fisheries Management, 14(4):797–811, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1991:GCL

[BC91] Danny J. Brown and Thomas G. Coon. Grass carp larvae in the Lower Missouri River and its tributaries. North American Journal of Fisheries Management, 11(1):62–66, February 1991. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beyers:1993:MHU

[BC93] Daniel W. Beyers and Clarence A. Carlson. Movement and habitat use of triploid grass carp in a Colorado irrigation canal. North American Journal of Fisheries Management, 13(1):141– 150, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bourque:1994:MBE

[BC94] C. Bourque and D. K. Cairns. Management briefs: Efficiency and accuracy of an automated data capture and error-checking system for laboratory fish processing. North American Journal of Fisheries Management, 14(3):650–655, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1999:ECR

[BCAM99] Richard S. Brown, Steven J. Cooke, W. Gary Anderson, and R. Scott McKinley. Evidence to challenge the “2% rule” for biotelemetry. North American Journal of Fisheries Management, 19(3):867–871, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bruesewitz:1993:EDE

[BCC93] Richard E. Bruesewitz, Daniel W. Coble, and Fred Copes. Effects of deflating the expanded swim bladder on survival of burbot. North American Journal of Fisheries Management, 13(2):346– 348, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bourgeois:1996:STE

[BCCEJ96] Gilles Bourgeois, Richard A. Cunjak, Daniel Caissie, and Nassir El-Jabi. A spatial and temporal evaluation of PHABSIM in relation to measured density of juvenile Atlantic salmon in a small stream. North American Journal of Fisheries Management, 16 (1):154–166, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bourgeois:1997:CFR

[BCCEJ97] Gilles Bourgeois, Richard A. Cunjak, Daniel Caissie, and Nassir El-Jabi. Cautions on forcing regression equations through the origin: Response to comment. North American Journal of Fisheries Management, 17(1):227–228, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buzby:1999:RAP

[BD99] Karen Buzby and Linda Deegan. Retention of anchor and passive integrated transponder tags by Arctic grayling. North American Journal of Fisheries Management, 19(4):1147–1150, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:1997:ESA

[BDBN97] T. Douglas Beard, Jr., Melissa T. Drake, James E. Breck, and Nancy A. Nate. Effects of simulated angling regulations on stunting in bluegill populations. North American Journal of Fisheries Management, 17(2):525–532, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Braaten:1999:CAG

[BDG99] Patrick J. Braaten, Matthew R. Doeringsfeld, and Christopher S. Guy. Comparison of age and growth estimates for river carpsuckers using scales and dorsal fin ray sections. North American Journal of Fisheries Management, 19(3):786–792, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beauchamp:1995:RPS

[Bea95] David A. Beauchamp. Riverine predation on sockeye salmon fry migrating to Lake Washington. North American Journal of Fisheries Management, 15(2):358–365, May 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergerhouse:1992:LEE

[Ber92] David L. Bergerhouse. Lethal effects of elevated pH and ammonia on early life stages of walleye. North American Journal of Fisheries Management, 12(2):356–366, May 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berg:1994:ISP

[Ber94] Neil H. Berg. Ice in stream pools in California’s Central Sierra Nevada: Spatial and temporal variability and reduction in trout habitat availability. North American Journal of Fisheries Management, 14(2):372–384, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berggren:1993:AVI

[BF93] Thomas J. Berggren and Margaret J. Filardo. An analysis of variables influencing the migration of juvenile salmonids in the Columbia River Basin. North American Journal of Fisheries Management, 13(1):48–63, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bernard:1997:EBS

[BF97] Gerry Bernard and Michael G. Fox. Effects of body size and population density on overwinter survival of age-0 pumpkinseeds. North American Journal of Fisheries Management, 17(2):581– 590, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowen:1998:SEE

[BF98] Zachary H. Bowen and Mary C. Freeman. Sampling effort and estimates of species richness based on prepositioned area electrofisher samples. North American Journal of Fisheries Management, 18(1):144–153, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brege:1990:ALP

[BFJ90] Dean A. Brege, Winston E. Farr, and Richard C. Johnsen. An air-lift pump for sampling juvenile salmonids at John Day Dam. North American Journal of Fisheries Management, 10(4): 481–483, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryant:1999:CWA

[BFM99] Mason D. Bryant, Brian J. Frenette, and Steven J. McCurdy. Colonization of a watershed by anadromous salmonids following the installation of a fish ladder in Margaret Creek, Southeast Alaska. North American Journal of Fisheries Management, 19 (4):1129–1136, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brege:1990:MBA

[BG90] Dean A. Brege and Albert E. Giorgi. Management briefs: An anchor system for periodic retrieval of current meters from fixed locations. North American Journal of Fisheries Management, 10 (1):114–115, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergstedt:1994:NTS

[BG94] Roger A. Bergstedt and Joseph H. Genovese. New technique for sampling sea lamprey larvae in deepwater habitats. North American Journal of Fisheries Management, 14(2):449–452, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burkhardt:1995:IEC

[BG95] Randy W. Burkhardt and Steve Gutreuter. Improving electrofishing catch consistency by standardizing power. North American Journal of Fisheries Management, 15(2):375–381, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:1998:SMM

[BGJ^{+}98] B. F. Blackwell, G. Gries, F. Juanes, K. D. Friedland, L. W. Stolte, and J. F. McKeon. Simulating migration mortality of Atlantic salmon smolts in the Merrimack River. North American Journal of Fisheries Management, 18(1):31–45, February 1998.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bruno:1990:NSU

[BGS90] Nicholas A. Bruno, Richard W. Gregory, and Harold L. Schramm, Jr. Nest sites used by radio-tagged largemouth bass in Orange Lake, Florida. North American Journal of Fisheries Management, 10(1):80–84, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bilby:1999:EIW

[BHF^{+}99] Robert E. Bilby, John T. Heffner, Brian R. Fransen, James W. Ward, and Peter A. Bisson. Effects of immersion in water on deterioration of wood from five species of trees used for habitat enhancement projects. North American Journal of Fisheries Management, 19(3):687–695, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brooks:1994:MMO

[BHK94] Ronald C. Brooks, Roy C. Heidinger, and Christopher C. Kohler. Mass-marking otoliths of larval and juvenile walleyes by immersion in oxytetracycline, calcein, or calcein blue. North American Journal of Fisheries Management, 14(1):143–150, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1992:DGY

[BHM92] Gerard L. Buynak, R. Scott Hale, and Bill Mitchell. Differential growth of young-of-year gizzard shad in several Kentucky reservoirs. North American Journal of Fisheries Management, 12 (3):656–662, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beeman:1998:ENM

[BHM98] John W. Beeman, Philip V. Haner, and Alec G. Maule. Evaluation of a new miniature pressure-sensitive radio transmitter. North American Journal of Fisheries Management, 18(2):458– 464, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:1997:PAC

[BHY^{+}97] T. Douglas Beard, Jr., Steven W. Hewett, Qing Yang, Ruth M. King, and Stephen J. Gilbert. Prediction of angler catch rates based on walleye population density. North American Journal of Fisheries Management, 17(3):621–627, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunnell:1999:ITM

[BI99] David B. Bunnell and J. Jeffery Isely. Influence of temperature on mortality and retention of simulated transmitters in rainbow trout. North American Journal of Fisheries Management, 19(1): 152–154, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bailey:1998:EVI

[BIDN98] R. E. Bailey, J. R. Irvine, F. C. Dalziel, and T. C. Nelson. Evaluations of visible implant fluorescent tags for marking Coho salmon smolts. North American Journal of Fisheries Management, 18 (1):191–196, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Binns:1994:LTR

[Bin94] N. Allen Binns. Long-term responses of trout and macrohabitats to habitat management in a Wyoming headwater stream. North American Journal of Fisheries Management, 14(1):87–98, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:1998:PAS

[BJ98] Bradley F. Blackwell and Francis Juanes. Predation on Atlantic salmon smolts by striped bass after dam passage. North American Journal of Fisheries Management, 18(4):936–939, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:1999:CBB

[BK99] T. Douglas Beard, Jr. and Jeffrey M. Kampa. Changes in bluegill, black crappie, and yellow perch populations in Wisconsin during 1967–1991. North American Journal of Fisheries Management, 19(4):1037–1043, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunt:1999:APE

[BKM99] Christopher M. Bunt, C. Katopodis, and R. S. McKinley. Attraction and passage efficiency of white suckers and smallmouth bass by two Denil fishways. North American Journal of Fisheries Management, 19(3):793–803, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1991:EDH

[BKSM91a] Gerard L. Buynak, Lewis E. Kornman, Albert Surmont, and Bill Mitchell. Evaluation of a differential-harvest regulation for black bass in Cave Run Lake, Kentucky. North American Journal of Fisheries Management, 11(3):277–284, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1991:ESB

[BKSM91b] Gerard L. Buynak, Lewis E. Kornman, Albert Surmont, and Bill Mitchell. Evaluation of a smallmouth bass stocking program in a Kentucky reservoir. North American Journal of Fisheries Management, 11(3):293–297, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beckman:1998:RFS

[BLLPD98] Brian R. Beckman, Donald A. Larsen, Beeda Lee-Pawlak, and Walton W. Dickhoff. Relation of fish size and growth rate to migration of spring Chinook salmon smolts. North American Journal of Fisheries Management, 18(3):537–546, August 1998. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beauchamp:1995:ECP

[BLT95] David A. Beauchamp, Mark G. Lariviere, and G. L. Thomas. Evaluation of competition and predation as limits to juvenile Kokanee and sockeye salmon production in Lake Ozette, Washington. North American Journal of Fisheries Management, 15 (1):193–207, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beauchamp:1997:HAA

[BLW^{+}97] David A. Beauchamp, Chris Luecke, Wayne A. Wurtsbaugh, Howard G. Gross, Phaedra E. Budy, Scott Spaulding, Robert Dillenger, and Chad P. Gubala. Hydroacoustic assessment of abundance and diel distribution of sockeye salmon and Kokanee in the Sawtooth Valley lakes, Idaho. North American Journal of Fisheries Management, 17(2):253–267, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1991:SWS

[BM91] Michael L. Brown and Brian R. Murphy. Standard weights (W_{s}) for striped bass, white bass, and hybrid striped bass. North American Journal of Fisheries Management, 11(3):451–467, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1993:ECP

[BM93] Gerard L. Buynak and Bill Mitchell. Electrofishing catch per effort as a predictor of largemouth bass abundance and angler catch in Taylorsville Lake, Kentucky. North American Journal of Fisheries Management, 13(3):630–633, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barrett:1994:HPI

[BM94] Paul J. Barrett and O. Eugene Maughan. Habitat preferences of introduced smallmouth bass in a Central Arizona stream. North American Journal of Fisheries Management, 14(1):112– 118, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bourgeois:1995:FDS

[BM95] C. E. Bourgeois and J. Murray. A fry distribution system for large-scale colonization stocking. North American Journal of Fisheries Management, 15(3):657–660, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1996:MBS

[BM96] Michael L. Brown and Brian R. Murphy. Management: Briefs selection of a minimum sample size for application of the regressionline–percentile technique. North American Journal of Fisheries Management, 16(2):427–432, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baxter:1997:DMP

[BM97] James S. Baxter and J. D. McPhail. Diel microhabitat preferences of juvenile bull trout in an artificial stream channel. North American Journal of Fisheries Management, 17(4):975–980, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1999:CSA

[BM99] Gerard L. Buynak and Bill Mitchell. Contribution of stocked advanced-fingerling largemouth bass to the population and fishery at Taylorsville Lake, Kentucky. North American Journal of Fisheries Management, 19(2):494–503, May 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1999:MLB

[BMB^{+}99] Gerard L. Buynak, Bill Mitchell, Don Bunnell, Bill McLemore, and Paul Rister. Management of largemouth bass at Kentucky and Barkley Lakes, Kentucky. North American Journal of Fisheries Management, 19(1):59–66, February 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1991:CLB

[BMM91] Gerard L. Buynak, William N. McLemore, and Bill Mitchell. Changes in largemouth bass populations at Kentucky and Barkley Lakes: Environmental or regulatory responses? North American Journal of Fisheries Management, 11(3):285–292, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1999:SSL

[BMMF99] Gerard L. Buynak, Bill Mitchell, Dan Michaelson, and Kevin Frey. Stocking subadult largemouth bass to meet angler expectations at Carr Creek Lake, Kentucky. North American Journal of Fisheries Management, 19(4):1017–1027, November 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettoli:1992:PLB

[BMNB92] Phillip W. Bettoli, Michael J. Maceina, Richard L. Noble, and Robert K. Betsill. Piscivory in largemouth bass as a function of aquatic vegetation abundance. North American Journal of Fisheries Management, 12(3):509–516, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettoli:1993:RRF

[BMNB93] Phillip W. Bettoli, Michael J. Maceina, Richard L. Noble, and Robert K. Betsill. Response of a reservoir fish community to aquatic vegetation removal. North American Journal of Fisheries Management, 13(1):110–124, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bugert:1997:ARS

[BMS97] Robert M. Bugert, Glen W. Mendel, and Paul R. Seidel. Adult returns of subyearling and yearling fall Chinook salmon released from a Snake River hatchery or transported downstream. North American Journal of Fisheries Management, 17(3):638–651, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:1995:SSS

[BMVV95] Jeff Boxrucker, Paul Michaletz, Michael J. Van Den Avyle, and Bruce Vondracek. Special section: Sampling shad in reservoirs: Overview of gear evaluation study for sampling gizzard shad and threadfin shad populations in reservoirs. North American Journal of Fisheries Management, 15(4):885–890, November 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1994:HDC

[BMY94] Larry R. Brown, Peter B. Moyle, and Ronald M. Yoshiyama. Historical decline and current status of Coho salmon in California. North American Journal of Fisheries Management, 14(2): 237–261, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryan:1994:VIT

[BN94] Roger D. Bryan and John J. Ney. Visible implant tag retention by and effects on condition of a stream population of brook trout. North American Journal of Fisheries Management, 14(1):216– 219, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beamesderfer:1995:GNM

[BN95] Raymond C. P. Beamesderfer and John A. North. Growth, natural mortality, and predicted response to fishing for largemouth bass and smallmouth bass populations in North America. North American Journal of Fisheries Management, 15(3):688–704, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bardygula-Nonn:1995:IPD

[BNNS95] L. G. Bardygula-Nonn, Robert Nonn, and J. Savitz. Influence of pulsed direct current electrofishing on mortality and injuries among four centrarchid species. North American Journal of Fisheries Management, 15(4):799–803, November 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettoli:1998:HMB

[BO98] Phillip W. Bettoli and Randall S. Osborne. Hooking mortality and behavior of striped bass following catch and release angling. North American Journal of Fisheries Management, 18(3):609– 615, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Borgstrom:1994:PLB

[Bor94] Reidar Borgstrøm. Production in a Lacustrine brown trout population with large recruitment potential and low natural mortality: Implications for management. North American Journal of Fisheries Management, 14(3):488–499, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonar:1997:SGC

[BPB^{+}97] Scott A. Bonar, John Pahutski, Bruce D. Bolding, Doug Fletcher, and Marc Divens. Survival and growth of channel catfish stocked in Washington lakes. North American Journal of Fisheries Management, 17(3):773–778, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bernard:1991:HTM

[BPC91] David R. Bernard, Gary A. Pearse, and Robert H. Conrad. Hoop traps as a means to capture burbot. North American Journal of Fisheries Management, 11(1):91–104, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bernard:1993:SAB

[BPL93] David R. Bernard, James F. Parker, and Robert Lafferty. Stock assessment of burbot populations in small and moderate-size lakes. North American Journal of Fisheries Management, 13(4): 657–675, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bozek:1991:CSV

[BR91] Michael A. Bozek and Frank J. Rahel. Comparison of streamside visual counts to electrofishing estimates of Colorado River cutthroat trout fry and adults. North American Journal of Fisheries Management, 11(1):38–42, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Binns:1994:RBC

[BR94] N. Allen Binns and Ronald Remmick. Response of Bonneville cutthroat trout and their habitat to drainage-wide habitat management at Huff Creek, Wyoming. North American Journal of Fisheries Management, 14(4):669–680, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beamesderfer:1990:MIM

[BRBV90] Raymond C. Beamesderfer, Bruce E. Rieman, Lewis J. Bledsoe, and Steven Vigg. Management implications of a model of predation by a resident fish on juvenile salmonids migrating through a Columbia River reservoir. North American Journal of Fisheries Management, 10(3):290–304, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergersen:1994:LHP

[BRC94] Eric P. Bergersen, Kevin B. Rogers, and Larry V. Conger. A livestock hormone pellet injector for implanting PIT tags. North American Journal of Fisheries Management, 14(1):224– 225, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Backman:1991:TSA

[BRK91] Thomas W. H. Backman, Robert M. Ross, and William F. Krise. Tolerance of subyearling American shad to short-term exposure to gas supersaturation. North American Journal of Fisheries Management, 11(1):67–71, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1999:TCR

[Bro99] Steven T. Brown. Trends in the commercial and recreational shark fisheries in Florida, 1980–1992, with implications for management. North American Journal of Fisheries Management, 19 (1):28–41, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brooking:1998:SDA

[BROV98] Thomas E. Brooking, Lars G. Rudstam, Mark H. Olson, and Anthony J. Vandevalk. Size-dependent alewife predation on larval walleyes in laboratory experiments. North American Journal of Fisheries Management, 18(4):960–965, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bugley:1991:MBE

[BS91] Karen Bugley and Gary Shepherd. Management briefs: Effect of catch-and-release angling on the survival of black sea bass. North American Journal of Fisheries Management, 11(3):468– 471, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Born:1990:EWA

[BSMM90] Stephen M. Born, William C. Sonzogni, Jeffrey Mayers, and J. Anderson Morton. The exceptional waters approach — a focus for coordinated natural resources management. North American Journal of Fisheries Management, 10(3):279–289, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brewin:1995:DLF

[BSN95] M. Kerry Brewin, Lucius L. Stebbins, and Joseph S. Nelson. Differential losses of floy anchor tags between male and female brown trout. North American Journal of Fisheries Management, 15(4): 881–884, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bartholow:1990:QHP

[BSPH90] John Bartholow, William Slauson, Bradford G. M. Parsons, and Wayne A. Hubert. Questions on habitat preference. North American Journal of Fisheries Management, 10(3):362–363, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergstedt:1993:ETL

[BSS93a] Roger A. Bergstedt, William D. Swink, and James G. Seelye. Evaluation of two locations for coded wire tags in larval and small parasitic-phase sea lampreys. North American Journal of Fisheries Management, 13(3):609–612, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bronte:1993:DYP

[BSS93b] Charles R. Bronte, James H. Selgeby, and Donald V. Swedberg. Dynamics of a yellow perch population in Western Lake Superior.

North American Journal of Fisheries Management, 13(3):511– 523, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barber:1990:IGO

[BT90] Willard E. Barber and John N. Taylor. The importance of goals, objectives, and values in the fisheries management process and organization: a review. North American Journal of Fisheries Management, 10(4):365–373, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blankenship:1993:EVI

[BT93] H. Lee Blankenship and Jack M. Tipping. Evaluation of visible implant and sequentially coded wire tags in sea-run cutthroat trout. North American Journal of Fisheries Management, 13(2): 391–394, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1995:ERD

[BT95] Russell W. Brown and William W. Taylor. Effects of a recreational dip-net fishery on rainbow smelt egg deposition. North American Journal of Fisheries Management, 15(1):165– 169, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bradford:1995:ESS

[BTAH95] Michael J. Bradford, Garth C. Taylor, J. Andrew Allan, and Paul S. Higgins. An experimental study of the stranding of juvenile Coho salmon and rainbow trout during rapid flow decreases under winter conditions. North American Journal of Fisheries Management, 15(2):473–479, May 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonneau:1995:CME

[BTS95] Joseph L. Bonneau, Russell E. Thurow, and Dennis L. Scarnecchia. Capture, marking, and enumeration of juvenile bull trout and cutthroat trout in small, low-conductivity streams. North American Journal of Fisheries Management, 15(3):563–568, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonar:1993:ETG

[BTT^{+}93] Scott A. Bonar, G. L. Thomas, Steven L. Th1Esfeld, Gilbert B. Pauley, and T. Brock Stables. Effect of triploid grass carp on the aquatic macrophyte community of Devils Lake, Oregon. North American Journal of Fisheries Management, 13(4): 757–765, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burton:1994:EAA

[BW94] William H. Burton and Stephen B. Weisberg. Estimating abundance of age-0 striped bass in the Delaware River by using marked hatchery fish. North American Journal of Fisheries Management, 14(2):347–354, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bigelow:1996:EGI

[BW96] Patricia E. Bigelow and Robert G. White. Evaluation of growth interruption as a means of manipulating scale patterns for massmarking hatchery trout. North American Journal of Fisheries Management, 16(1):142–153, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beck:1997:PSI

[BWF97] H. Denise Beck, David W. Willis, and Jay M. Francis. A proposed standard intercept for the white bass body length-scale radius relationship. North American Journal of Fisheries Management, 17(2):488–492, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Booth:1993:ESL

[BWG93] G. M. Booth, C. D. Wren, and J. M. Gunn. Efficacy of shoal liming for rehabilitation of lake trout populations in acid-stressed lakes. North American Journal of Fisheries Management, 13(4): 766–774, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bleyker:1997:DFS

[BWO97] Jeffrey S. Den Bleyker, Larry J. Weber, and A. Jacob Odgaard. Development of a flow spreader for fish bypass outfalls. North American Journal of Fisheries Management, 17(3):743–750, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:1995:SIS

[BWW95a] T. D. Beacham, R. E. Withler, and C. C. Wood. Stock identification of sockeye salmon by means of minisatellite DNA variation. North American Journal of Fisheries Management, 15(2): 249–265, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bulak:1995:EMO

[BWW95b] James S. Bulak, David S. Wethey, and Miller G. White III. Evaluation of management options for a reproducing striped bass population in the santee–Cooper system, South Carolina. North American Journal of Fisheries Management, 15(1):84–94, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Close:1992:DDD

[CA92] Tracy L. Close and Charles S. Anderson. Dispersal, densitydependent growth, and survival of stocked steelhead fry in Lake Superior tributaries. North American Journal of Fisheries Management, 12(4):728–735, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cada:1990:RSR

[Cad90] Glenn F. Cada. A review of studies relating to the effects of propeller-type turbine passage on fish early life stages. North American Journal of Fisheries Management, 10(4):418–426, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Campbell:1992:CAL

[Cam92] Alan Campbell. Characteristics of the American lobster fishery of Grand Manan, New Brunswick, Canada. North American Journal of Fisheries Management, 12(1):139–150, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Campbell:1997:CRO

[Cam97] W. B. Campbell. Comment: Reducing observer error in asymmetry analyses. North American Journal of Fisheries Management, 17(2):599, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carbines:1999:LHR

[Car99] G. D. Carbines. Large hooks reduce catch-and-release mortality of blue cod Parapercis colias in the marlborough sounds of New Zealand. North American Journal of Fisheries Management, 19 (4):992–998, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chipps:1996:CNM

[CB96] Steve R. Chipps and David H. Bennett. Comparison of net mesh sizes for estimating abundance of the opossum shrimp Mysis relicta from vertical hauls. North American Journal of Fisheries Management, 16(3):689–692, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cederholm:1997:RJC

[CBB^{+}97] C. J. Cederholm, R. E. Bilby, P. A. Bisson, T. W. Bumstead, B. R. Fransen, W. J. Scarlett, and J. W. Ward. Response of juvenile Coho salmon and steelhead to placement of large woody debris in a coastal Washington stream. North American Journal of Fisheries Management, 17(4):947–963, November 1997. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:1998:DPT

[CBB98] William P. Connor, Howard L. Burge, and David H. Bennett. Detection of PIT-Tagged subyearling Chinook salmon at a Snake River dam: Implications for summer flow augmentation. North American Journal of Fisheries Management, 18(3):530–536, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collis:1995:CCR

[CBC95] Ken Collis, Roy E. Beaty, and Bradford R. Crain. Changes in catch rate and diet of northern squawfish associated with the release of hatchery-reared juvenile salmonids in a Columbia River reservoir. North American Journal of Fisheries Management, 15 (2):346–357, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chipps:1997:PFA

[CBD97] Steve R. Chipps, David H. Bennett, and Thomas J. Dresser, Jr. Patterns of fish abundance associated with a dredge disposal Island: Implications for fish habitat enhancement in a large reservoir. North American Journal of Fisheries Management, 17(2): 378–386, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carline:1991:RWB

[CBH91] Robert F. Carline, Thomas Beard, Jr., and Bruce A. Hollender. Response of wild brown trout to elimination of stocking and to no-harvest regulations. North American Journal of Fisheries Management, 11(3):253–266, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chittenden:1993:FAS

[CBJ93] Mark E. Chittenden, Jr., Luiz R. Barbieri, and Cynthia M. Jones. Fluctuations in abundance of Spanish mackerel in Chesapeake Bay and the mid-Atlantic region. North American Journal of Fisheries Management, 13(3):450–458, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Curtis:1993:FCL

[CBS93] Gary L. Curtis, Charles R. Bronte, and James H. Selgeby. Forecasting contributions of lake whitefish year-classes to a Lake Superior commercial fishery from estimates of yearling abundance. North American Journal of Fisheries Management, 13(2):349– 352, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clapp:1997:ETS

[CBW97] David F. Clapp, Yamini Bhagwat, and David H. Wahl. The effect of thermal stress on walleye fry and fingerling mortality. North American Journal of Fisheries Management, 17(2):429–437, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Copeland:1998:OSL

[CC98] Timothy Copeland and Robert F. Carline. Overwinter survival and lipid content of walleye fingerlings. North American Journal of Fisheries Management, 18(2):383–390, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Czajkowski:1996:LRT

[CCC^{+}96] Stephen P. Czajkowski, Daniel W. Coble, Fred A. Copes, Ron Bruch, Kendall Kamke, and Paul J. Peeters. Lower returns of tagged walleyes initially caught by trawling than by trap nets, fyke nets, or electroshocking. North American Journal of Fisheries Management, 16(2):453–456, May 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carpenter:1995:RBH

[CCG^{+}95] S. R. Carpenter, P. Cunningham, S. Gafny, A. Munoz-Del-Rio, N. Nibbelink, M. Olson, T. Pellett, C. Storlie, and A. Trebitz. Responses of bluegill to habitat manipulations: Power to detect effects. North American Journal of Fisheries Management, 15 (3):519–527, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Candy:1996:ACS

[CCQR96] J. R. Candy, E. W. Carter, T. P. Quinn, and B. E. Riddell. Adult Chinook salmon behavior and survival after catch and release from purse-seine vessels in Johnstone Strait, British Columbia. North American Journal of Fisheries Management, 16(3):521– 529, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cadrin:1999:ACS

[CCS^{+}99] Steven X. Cadrin, Stephen H. Clark, Daniel F. Schick, Michael P. Armstrong, David McCarron, and Bruce Smith. Application of catch-survey models to the northern shrimp fishery in the Gulf of Maine. North American Journal of Fisheries Management, 19(2):551–568, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chick:1999:EAE

[CCT99] John H. Chick, Sean Coyne, and Joel C. Trexler. Effectiveness of airboat electrofishing for sampling fishes in shallow, vegetated habitats. North American Journal of Fisheries Management, 19 (4):957–967, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chapman:1997:HSC

[CCW^{+}97] D. Chapman, C. Carlson, D. Weitkamp, G. Matthews, J. Stevenson, and M. Miller. Homing in sockeye and Chinook salmon transported around part of their smolt migration route in the Columbia River. North American Journal of Fisheries Management, 17(1):101–113, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clayton:1998:ALR

[CDM^{+}98] Janet L. Clayton, Eric S. Dannaway, Raymond Menendez, Henry W. Rauch, John J. Renton, Sean M. Sherlock, and Peter E. Zurbuch. Application of limestone to Restore fish communities in acidified streams. North American Journal of Fisheries Management, 18(2):347–360, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlile:1997:EMC

[CDW97] David W. Carlile, Thomas A. Dinnocenzo, and Leslie J. Watson. Evaluation of modified crab pots to increase catch of Pacific cod and decrease bycatch of Pacific halibut. North American Journal of Fisheries Management, 17(4):910–928, November 1997. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1998:ABM

[CH98] William G. Clark and Steven R. Hare. Accounting for bycatch in management of the Pacific halibut fishery. North American Journal of Fisheries Management, 18(4):809–821, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cadrin:1995:EET

[CHCC95] Steven X. Cadrin, Arnold B. Howe, Steven J. Correia, and Thomas P. Currier. Evaluating the effects of two coastal mobile gear fishing closures on finfish abundance off Cape Cod. North American Journal of Fisheries Management, 15(2):300–315, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Currens:1997:ISD

[CHF^{+}97] Kenneth P. Currens, Alan R. Hemmingsen, Rod A. French, David V. Buchanan, Carl B. Schreck, and Hiram W. Li. Introgression and susceptibility to disease in a wild population of rainbow trout. North American Journal of Fisheries Management, 17(4): 1065–1078, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Criddle:1998:CFR

[CHGF98] Keith R. Criddle, Mark Herrmann, Joshua A. Greenberg, and Erika M. Feller. Climate fluctuation and revenue maximization in the Eastern Bering Sea fishery for walleye pollock. North American Journal of Fisheries Management, 18(1):1–10, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Childress:1991:SWC

[Chi91] W. Michael Childress. Standard-weight curves derived from Texas length and weight data. North American Journal of Fisheries Management, 11(3):480–482, August 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:1998:FIR

[CHM98] Mark R. Collins, Patrick J. Harris, and Philip P. Maier. Fisheryindependent recruitment indices for king and Spanish mackerels. North American Journal of Fisheries Management, 18(1):181– 186, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crispin:1993:CIH

[CHR93] Val Crispin, Robert House, and Dave Roberts. Changes in instream habitat, large woody debris, and salmon habitat after the restructuring of a coastal Oregon stream. North American Journal of Fisheries Management, 13(1):96–102, February 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clapp:1993:MTG

[CHTC93] David F. Clapp, Rue S. Hestand III, Boyd Z. Thompson, and Laurence L. Connor. Movement of triploid grass carp in large Florida lakes. North American Journal of Fisheries Management, 13(4):746–756, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chai:1992:EOS

[CHTG92] Ai-Ling Chai, Mark Homer, Chu-Fa Tsai, and Philippe Goulletquer. Evaluation of oyster sampling efficiency of patent tongs and an oyster dredge. North American Journal of Fisheries Management, 12(4):825–832, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connelly:1999:UDE

[CK99] Nancy A. Connelly and Barbara A. Knuth. Using diaries to examine Children’s fishing patterns. North American Journal of Fisheries Management, 19(4):1103–1107, November 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connelly:1996:SCP

[CKB96] Nancy A. Connelly, Barbara A. Knuth, and Tommy L. Brown. Sportfish consumption patterns of Lake Ontario anglers and the relationship to health advisories. North American Journal of Fisheries Management, 16(1):90–101, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clayton:1996:EIA

[CKT96] James L. Clayton, John G. King, and Russell E. Thurow. Evaluation of an ion adsorption method to estimate intragravel flow velocity in salmonid spawning gravel. North American Journal of Fisheries Management, 16(1):167–174, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1994:PDP

[Cla94] Robert A. Clark. Population dynamics and potential utility per recruit of Arctic grayling in Fielding Lake, Alaska. North American Journal of Fisheries Management, 14(3):500–515, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cassani:1995:ATG

[CLA95] John R. Cassani, Ernesto Lasso De La Vega, and Holly Allaire. An assessment of triploid grass carp stocking rates in small warmwater impoundments. North American Journal of Fisheries Management, 15(2):400–407, May 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clugston:1996:RBA

[Clu96] James P. Clugston. Retention of T-bar anchor tags and passive integrated transponder tags by Gulf sturgeons. North American Journal of Fisheries Management, 16(3):682–685, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caddy:1996:TLM

[CM96] J. F. Caddy and R. McGarvey. Targets or limits for management of fisheries? North American Journal of Fisheries Management, 16(3):479–487, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clayton:1999:VAF

[CM99a] Darien L. Clayton and Michael J. Maceina. Validation of annulus formation in gizzard shad otoliths. North American Journal of Fisheries Management, 19(4):1099–1102, November 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cornelius:1999:ELL

[CM99b] Richard R. Cornelius and Terry L. Margenau. Effects of length limits on muskellunge in Bone Lake, Wisconsin. North American Journal of Fisheries Management, 19(1):300–308, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Counihan:1999:IRA

[CMP99] Timothy D. Counihan, Allen I. Miller, and Michael J. Parsley. Indexing the relative abundance of age-0 white sturgeons in an impoundment of the Lower Columbia River from highly skewed trawling data. North American Journal of Fisheries Management, 19(2):520–529, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cross:1995:SVT

[CMS95] Timothy K. Cross, Michael C. McInerny, and Dennis H. Schupp. Seasonal variation in trap-net catches of bluegill in Minnesota lakes. North American Journal of Fisheries Management, 15(2): 382–389, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:1999:SBD

[CMS^{+}99] Mark R. Collins, John C. McGovern, George R. Sedberry, H. Scott Meister, and Renee Pardieck. Swim bladder deflation in black sea bass and vermilion snapper: Potential for increasing postrelease survival. North American Journal of Fisheries Management, 19(3):828–832, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Copeland:1994:MYY

[CN94] John R. Copeland and Richard L. Noble. Movements by youngof-year and yearling largemouth bass and their implications for supplemental stocking. North American Journal of Fisheries Management, 14(1):119–124, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coble:1992:MBP

[Cob92] Daniel W. Coble. Management briefs: Predicting population density of largemouth bass from electrofishing catch per effort. North American Journal of Fisheries Management, 12(3):650–652, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:1991:EMS

[Col91a] Michael A. Colvin. Evaluation of minimum-size limits and reduced daily limits on the crappie populations and fisheries in five large Missouri reservoirs. North American Journal of Fisheries Management, 11(4):585–597, November 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:1991:PCA

[Col91b] Michael A. Colvin. Population characteristics and angler harvest of white crappies in four large Missouri reservoirs. North American Journal of Fisheries Management, 11(4):572–584, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cook:1990:MSR

[Coo90] B. A. Cook. Maximum social returns for Canada’s Pacific halibut fishery. North American Journal of Fisheries Management, 10 (1):11–18, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carey:1995:FHP

[CP95] T. G. Carey and G. I. Pritchard. Fish health protection: a strategic role in Canadian fisheries management. North American Journal of Fisheries Management, 15(1):1–13, February 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clancy:1990:DFS

[CR90] Christopher G. Clancy and Donald R. Reichmuth. A detachable fishway for steep culverts. North American Journal of Fisheries Management, 10(2):244–246, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1997:IBM

[CR97] Mark E. Clark and Kenneth A. Rose. An individual-based modeling analysis of management strategies for enhancing brook trout populations in Southern Appalachian streams. North American Journal of Fisheries Management, 17(1):54–76, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Casselman:1999:GUL

[CRC99] John M. Casselman, Chris J. Robinson, and E. J. Crossman. Growth and ultimate length of muskellunge from Ontario water bodies. North American Journal of Fisheries Management, 19 (1):271–290, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Criddle:1996:PCA

[Cri96] Keith R. Criddle. Predicting the consequences of alternative harvest regulations in a sequential fishery. North American Journal of Fisheries Management, 16(1):30–40, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coale:1994:CSC

[CRMH94] J. Stuart Coale, Roger A. Rulifson, James D. Murray, and Robert Hines. Comparisons of shrimp catch and bycatch between a skimmer trawl and an otter trawl in the North Carolina inshore shrimp fishery. North American Journal of Fisheries Management, 14(4): 751–768, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:1996:BSA

[CRS96] Mark R. Collins, S. Gordon Rogers, and Theodore I. J. Smith. Bycatch of sturgeons along the Southern Atlantic Coast of the USA. North American Journal of Fisheries Management, 16 (1):24–29, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:1996:SFR

[CS96a] Mark R. Collins and Theodore I. J. Smith. Sturgeon fin ray removal is nondeleterious. North American Journal of Fisheries Management, 16(4):939–941, November 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Conover:1996:EOM

[CS96b] Greg A. Conover and Robert J. Sheehan. Evaluation of oxytetracycline marks in sagittal otoliths of juvenile black crappies.

North American Journal of Fisheries Management, 16(3):686– 688, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cyterski:1996:TAD

[CS96c] M. J. Cyterski and G. R. Spangler. A tool for age determination. North American Journal of Fisheries Management, 16(2):403– 412, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarke:1997:UWM

[CS97a] K. D. Clarke and D. A. Scruton. Use of the wesche method to evaluate fine-sediment dynamics in small boreal forest headwater streams. North American Journal of Fisheries Management, 17 (1):188–193, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:1997:MBD

[CS97b] Mark R. Collins and Theodore I. J. Smith. Management briefs: Distributions of shortnose and Atlantic sturgeons in South Carolina. North American Journal of Fisheries Management, 17(4): 995–1000, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Conover:1999:SGM

[CS99] Greg A. Conover and Robert J. Sheehan. Survival, growth, and mark persistence in juvenile black crappies marked with fin clips, freeze brands, or oxytetracycline. North American Journal of Fisheries Management, 19(3):824–827, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cote:1999:SPG

[CSCM99] D. Cote, D. A. Scruton, Lloyd Cole, and R. S. McKinley. Swimming performance and growth rates of juvenile Atlantic cod intraperitoneally implanted with dummy acoustic transmitters. North American Journal of Fisheries Management, 19(4):1137– 1141, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cade:1997:CCF

[CT97] Brian S. Cade and James W. Terrell. Comment: Cautions on forcing regression equations through the origin. North American Journal of Fisheries Management, 17(1):225–227, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cable:1990:ECT

[CU90] Ted T. Cable and Edward Udd. Effective communication of toxic chemical warnings to anglers. North American Journal of Fisheries Management, 10(4):382–387, November 1990. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cunningham:1995:CSM

[Cun95] Kenneth K. Cunningham. Comparison of stationary and mobile electrofishing for sampling flathead catfish. North American Journal of Fisheries Management, 15(2):515–517, May 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cowdell:1994:EPD

[CV94] Bryan R. Cowdell and Richard A. Valdez. Effects of pulsed DC electroshock on adult roundtail chub from the Colorado River in Colorado. North American Journal of Fisheries Management, 14 (3):659–660, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cole:1994:OFM

[CW94] Richard A. Cole and Frank A. Ward. Optimum fisheries management policy: Angler opportunity versus angler benefit. North American Journal of Fisheries Management, 14(1):22–33, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carmichael:1996:FPT

[CWGT96] Gary J. Carmichael, J. Holt Williamson, Owen T. Gorman, and Terrence R. Tiersch. Field propagation techniques for the endangered razorback sucker. North American Journal of Fisheries Management, 16(4):963–966, November 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dextrase:1991:HML

[DB91] Alan J. Dextrase and Helen E. Ball. Hooking mortality of lake trout angled through the ice. North American Journal of Fisheries Management, 11(3):477–479, August 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dubois:1994:ELT

[DB94] Robert B. Dubois and William H. Blust. Effects of lampricide treatments, relative to environmental conditions, on abundance and sizes of salmonids in a small stream. North American Journal of Fisheries Management, 14(1):162–169, February 1994. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dalton:1998:EEV

[DBJW98] Robert S. Dalton, Chris T. Bastian, James J. Jacobs, and Thomas A. Wesche. Estimating the economic value of improved trout fishing on Wyoming streams. North American Journal of Fisheries Management, 18(4):786–797, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drake:1997:CBR

[DCPP97] Melissa T. Drake, Julie E. Claussen, David P. Philipp, and Donald L. Pereira. A comparison of bluegill reproductive strategies and growth among lakes with different fishing intensities. North American Journal of Fisheries Management, 17(2):496–507, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dumont:1997:CDN

[DD97] Spencer C. Dumont and John A. Dennis. Comparison of day and night electrofishing in Texas reservoirs. North American Journal of Fisheries Management, 17(4):939–946, November 1997. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:1995:EEV

[DE95] William P. Dwyer and David A. Erdahl. Effects of electroshock voltage, wave form, and pulse rate on survival of cutthroat trout eggs. North American Journal of Fisheries Management, 15(3): 647–650, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dedual:1996:OMR

[Ded96] Michel Dedual. Observed mortality of rainbow trout caught by different angling techniques in Lake Taupo, New Zealand. North American Journal of Fisheries Management, 16(2):357–363, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dewey:1992:EDN

[Dew92] Michael R. Dewey. Effectiveness of a drop net, a pop net, and an electrofishing frame for collecting quantitative samples of juvenile fishes in vegetation. North American Journal of Fisheries Management, 12(4):808–813, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:1993:IEM

[DFE93] William P. Dwyer, Wade Fredenberg, and David A. Erdahl. Influence of electroshock and mechanical shock on survival of trout eggs. North American Journal of Fisheries Management, 13(4): 839–843, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorazio:1991:SHR

[DFW91] Robert M. Dorazio, Benjamin M. Florence, and Charles M. Wooley. Stocking of hatchery-reared striped bass in the Patuxent River, Maryland: Survival, relative abundance, and costeffectiveness. North American Journal of Fisheries Management, 11(3):435–442, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dubois:1993:MJA

[DG93] Robert B. Dubois and Steven P. Gloss. Mortality of juvenile American shad and striped bass passed through ossberger crossflow turbines at a small-scale hydroelectric site. North American Journal of Fisheries Management, 13(1):178–185, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dawson:1991:RPF

[DGDG91] V. K. Dawson, W. H. Gingerich, R. A. Davis, and P. A. Gilderhus. Rotenone persistence in freshwater ponds: Effects of temperature and sediment adsorption. North American Journal of Fisheries Management, 11(2):226–231, May 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drenner:1997:CCA

[DGE^{+}97] Ray W. Drenner, Kirsten L. Gallo, C. Michael Edwards, Kerri E. Rieger, and Eric D. Dibble. Common carp affect turbidity and angler catch rates of largemouth bass in ponds. North American Journal of Fisheries Management, 17(4):1010–1013, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorn:1999:MTC

[DGFB99] Martin W. Dorn, Sarah K. Gaichas, Shannon M. Fitzgerald, and Sally A. Bibb. Measuring total catch at sea: Use of a motioncompensated flow scale to evaluate observer volumetric methods. North American Journal of Fisheries Management, 19(4): 999–1016, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dolloff:1997:CBR

[DJO97] C. Andrew Dolloff, Holly E. Jennings, and Michael D. Owen. A comparison of basinwide and representative reach habitat survey techniques in three Southern Appalachian watersheds. North American Journal of Fisheries Management, 17(2):339–347, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dalbey:1996:EEP

[DMF96] Steven R. Dalbey, Thomas E. McMahon, and Wade Fredenberg. Effect of electrofishing pulse shape and electrofishing-induced spinal injury on long-term growth and survival of wild rainbow trout. North American Journal of Fisheries Management, 16 (3):560–569, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dubois:1991:ISS

[DMP91] Robert B. Dubois, John E. Miller, and Scott D. Plaster. An inclined-screen smolt trap with adjustable screen for highly variable flows. North American Journal of Fisheries Management, 11(2):155–159, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dicenzo:1995:FRG

[DMR95] Victor J. Dicenzo, Michael J. Maceina, and William C. Reeves. Factors related to growth and condition of the Alabama subspecies of spotted bass in reservoirs. North American Journal of Fisheries Management, 15(4):794–798, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dubois:1994:HMN

[DMS^{+}94] Robert B. Dubois, Terry L. Margenau, R. Scot Stewart, Paul K. Cunningham, and Paul W. Rasmussen. Hooking mortality of northern pike angled through ice. North American Journal of Fisheries Management, 14(4):769–775, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DiCenzo:1996:RBR

[DMS96] Victor J. DiCenzo, Michael J. Maceina, and Marc R. Stimpert. Relations between reservoir trophic state and gizzard shad population characteristics in Alabama reservoirs. North American Journal of Fisheries Management, 16(4):888–895, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daum:1998:UFL

[DO98] David W. Daum and Bruce M. Osborne. Use of fixed-location, split-beam sonar to describe temporal and spatial patterns of adult fall chum salmon migration in the Chandalar River, Alaska. North American Journal of Fisheries Management, 18(3):477– 486, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dolman:1991:CBC

[Dol91] William B. Dolman. Comparison of bass-club tournament reports and creel survey data from Texas reservoirs. North American Journal of Fisheries Management, 11(2):177–184, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorazio:1995:MES

[Dor95] Robert M. Dorazio. Mortality estimates of striped bass caught in Albemarle Sound and Roanoke River, North Carolina. North American Journal of Fisheries Management, 15(2):290–299, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dussault:1997:FTM

[DR97] Christian Dussault and Marco A. Rodríguez. Field trials of marking stream salmonids by dye injection and coded-wire-tagging. North American Journal of Fisheries Management, 17(2):451– 456, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunning:1992:AAH

[DRG^{+}92] Dennis J. Dunning, Quentin E. Ross, Paul Geoghegan, James J. Reichle, John K. Menezes, and John K. Watson. Alewives avoid high-frequency sound. North American Journal of Fisheries Management, 12(3):407–416, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Devries:1990:MSE

[DS90] Dennis R. Devries and Roy A. Stein. Manipulating shad to enhance sport fisheries in North America: An assessment. North American Journal of Fisheries Management, 10(2):209–223, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dempson:1991:UPC

[DS91] J. B. Dempson and D. E. Stansbury. Using partial counting fences and a two-sample stratified design for mark–recapture estimation of an Atlantic salmon smolt population. North American Journal of Fisheries Management, 11(1):27–37, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:1992:IFL

[DTR92] Rex A. Dunham, C. J. Turner, and W. C. Reeves. Introgression of the Florida largemouth bass genome into native populations in Alabama public lakes. North American Journal of Fisheries Management, 12(3):494–498, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Devries:1995:ASA

[DVG95] Dennis R. Devries, Michael J. Van Den Avyle, and Eugene R. Gilliland. Assessing shad abundance: Electrofishing with active and passive fish collection. North American Journal of Fisheries Management, 15(4):891–897, November 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:1997:HFE

[DVR97] Jason B. Dunham, Gary L. Vinyard, and Bruce E. Rieman. Habitat fragmentation and extinction risk of Lahontan cutthroat trout. North American Journal of Fisheries Management, 17(4):

1126–1133, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Donnelly:1991:BCA

[DW91] W. A. Donnelly and F. G. Whoriskey, Jr. Background-color acclimation of brook trout for crypsis reduces risk of predation by hooded mergansers Lophodytes cucullatus. North American Journal of Fisheries Management, 11(2):206–211, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dean:1992:BBL

[DW92] Jan Dean and Garland Wright. Black bass length limits by design: a graphic approach. North American Journal of Fisheries Management, 12(3):538–547, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Degan:1995:CFH

[DW95a] Donald J. Degan and Wayne Wilson. Comparison of four hydroacoustic frequencies for sampling pelagic fish populations in Lake Texoma. North American Journal of Fisheries Management, 15(4):924–932, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:1995:MBI

[DW95b] William P. Dwyer and Robert G. White. Management briefs: Influence of electroshock on short-term growth of adult rainbow trout and juvenile Arctic grayling and cutthroat trout. North American Journal of Fisheries Management, 15(1):148– 151, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauble:1997:SFC

[DW97a] D. D. Dauble and D. G. Watson. Status of fall Chinook salmon populations in the mid-Columbia River, 1948–1992. North American Journal of Fisheries Management, 17(2):283–300, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:1997:MBE

[DW97b] William P. Dwyer and Robert G. White. Management briefs: Effect of electroshock on juvenile Arctic grayling and Yellowstone cutthroat trout growth, 100 days after treatment. North American Journal of Fisheries Management, 17(1):174–177, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Downs:1997:ASM

[DWS97] Christopher C. Downs, Robert G. White, and Bradley B. Shepard. Age at sexual maturity, sex ratio, fecundity, and longevity of isolated headwater populations of westslope cutthroat trout. North American Journal of Fisheries Management, 17(1):85–92, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:1990:CTS

[Dwy90] William P. Dwyer. Catchability of three strains of cutthroat trout. North American Journal of Fisheries Management, 10(4): 458–461, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Espegren:1990:QSF

[EB90] Gregory D. Espegren and Eric P. Bergersen. Quantitative sampling of fish populations with a mobile rising net. North American Journal of Fisheries Management, 10(4):469–478, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:1998:CTC

[ECS98] Andrew J. Edwards, Gary D. Czypinski, and James H. Selgeby. A collapsible trap for capturing ruffe. North American Journal of Fisheries Management, 18(2):465–469, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:1997:EPD

[EDGR97] C. Michael Edwards, Ray W. Drenner, Kirsten L. Gallo, and Kerri E. Rieger. Estimation of population density of largemouth bass in ponds by using mark–recapture and electrofishing catch per effort. North American Journal of Fisheries Management, 17 (3):719–725, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:1991:CTE

[Edw91] Steven F. Edwards. A critique of three “economics” arguments commonly used to influence fishery allocations. North American Journal of Fisheries Management, 11(2):121–130, May 1991. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:1999:ICM

[Edw99] Steven F. Edwards. Influence of a century of macroeconomic change on U.S. fisheries production. North American Journal of Fisheries Management, 19(1):1–17, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evenson:1992:MCH

[EE92] M. D. Evenson and R. D. Ewing. Migration characteristics and hatchery returns of winter steelhead volitionally released from Cole Rivers Hatchery, Oregon. North American Journal of Fisheries Management, 12(4):736–743, November 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elrod:1990:SSR

[EF90] Joseph H. Elrod and Anthony Frank. Sample size requirements and analysis of tag recoveries for paired releases of lake trout. North American Journal of Fisheries Management, 10(2):196– 201, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ellison:1992:OSW

[EF92] Daryl G. Ellison and William G. Franzin. Overview of the symposium on walleye stocks and stocking. North American Journal of Fisheries Management, 12(2):271–275, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ehlinger:1997:MGR

[EGP97] Timothy J. Ehlinger, Mart R. Gross, and David P. Philipp. Morphological and growth rate differences between bluegill males of alternative reproductive life histories. North American Journal of Fisheries Management, 17(2):533–542, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eaglin:1993:MBE

[EH93] Gregory S. Eaglin and Wayne A. Hubert. Management briefs: Effects of logging and roads on substrate and trout in streams of the Medicine Bow National Forest, Wyoming. North American Journal of Fisheries Management, 13(4):844–846, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ebert:1993:ECF

[EHB^{+}93] Ellen S. Ebert, Natalie W. Harrington, Kevin J. Boyle, James W. Knight, and Russell E. Keenan. Estimating consumption of freshwater fish among Maine anglers. North American Journal of Fisheries Management, 13(4):737–745, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ehlinger:1997:MRC

[Ehl97] Timothy J. Ehlinger. Male reproductive competition and sexspecific growth patterns in bluegill. North American Journal of Fisheries Management, 17(2):508–515, May 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elliott:1992:TTC

[Ell92] Steven T. Elliott. A trough trap for catching Coho salmon smolts emigrating from Beaver ponds. North American Journal of Fisheries Management, 12(4):837–840, November 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elrod:1997:SHR

[Elr97] Joseph H. Elrod. Survival of hatchery-reared lake trout stocked near shore and off shore in Lake Ontario. North American Journal of Fisheries Management, 17(3):779–783, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:1993:PEB

[EM93] Steven F. Edwards and Steven A. Murawski. Potential economic benefits from efficient harvest of New England groundfish. North American Journal of Fisheries Management, 13(3):437–449, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Erickson:1994:ICC

[EP94] Daniel L. Erickson and Ellen K. Pikitch. Incidental catch of Chinook salmon in commercial bottom trawls off the U.S. West Coast. North American Journal of Fisheries Management, 14 (3):550–563, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elrod:1992:ESS

[ES92] Joseph H. Elrod and Clifford P. Schneider. Effect of stocking season and technique on survival of lake trout in Lake Ontario.

North American Journal of Fisheries Management, 12(1):131– 138, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Emlen:1993:PPR

[ESB93] John M. Emlen, Thomas A. Strekal, and Chester C. Buchanan. Probabalistic projections for recovery of the endangered cui-ui. North American Journal of Fisheries Management, 13(3):467– 474, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ebert:1996:ERF

[ESB^{+}96] Ellen S. Ebert, Steave H. Su, Timothy J. Barry, Michael N. Gray, and Natalie W. Harrington. Estimated rates of fish consumption by anglers participating in the Connecticut Housatonic River Creel Survey. North American Journal of Fisheries Management, 16(1):81–89, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elrod:1993:SLT

[ESO93] Joseph H. Elrod, Clifford P. Schneider, and David E. Ostergaard. Survival of lake trout stocked in U.S. waters of Lake Ontario. North American Journal of Fisheries Management, 13(4): 775–781, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Everson:1992:ERC

[ESP92] Alan R. Everson, Robert A. Skillman, and Jeffrey J. Polovina. Evaluation of rectangular and circular escape vents in the Northwestern Hawaiian Islands lobster fishery. North American Journal of Fisheries Management, 12(1):161–171, February 1992. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eberts:1998:SLC

[ESW98] Ronald C. Eberts, Jr., Victor J. Santucci, Jr., and David H. Wahl. Suitability of the lake chubsucker as prey for largemouth bass in small impoundments. North American Journal of Fisheries Management, 18(2):295–307, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fago:1998:CLF

[Fag98] Don Fago. Comparison of littoral fish assemblages sampled with a mini-fyke net or with a combination of electrofishing and smallmesh seine in Wisconsin lakes. North American Journal of Fisheries Management, 18(3):731–738, August 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisher:1998:TLR

[FAW98] Shannon J. Fisher, Marcy R. Anderson, and David W. Willis. Total length reduction in preserved yellow perch larvae. North American Journal of Fisheries Management, 18(3):739–742, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisher:1993:PAE

[FB93] William L. Fisher and Marshall E. Brown. A prepositioned areal electrofishing apparatus for sampling stream habitats. North American Journal of Fisheries Management, 13(4):807– 816, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frenette:1996:EVI

[FB96] Brian J. Frenette and Mason D. Bryant. Evaluation of visible implant tags applied to wild coastal cutthroat trout and Dolly Varden in Margaret Lake, Southeast Alaska. North American Journal of Fisheries Management, 16(4):926–930, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Freeman:1999:THS

[FBB99] Mary C. Freeman, Zachary H. Bowen, and Ken D. Bovee. Transferability of habitat suitability criteria: Response to comment. North American Journal of Fisheries Management, 19(2):626– 628, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Freeman:1997:THS

[FBC97] Mary C. Freeman, Zachary H. Bowen, and Johnie H. Crance. Transferability of habitat suitability criteria for fishes in warmwater streams. North American Journal of Fisheries Management, 17(1):20–31, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Franzin:1994:REP

[FBR^{+}94] William G. Franzin, Bruce A. Barton, Richard A. Remnant, Duncan B. Wain, and Shirley Jo Pagel. Range extension, present and potential distribution, and possible effects of rainbow smelt in Hudson Bay drainage waters of Northwestern Ontario, Manitoba, and Minnesota. North American Journal of Fisheries Management, 14(1):65–76, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fox:1997:GPG

[FCP97] Michael G. Fox, Julie E. Claussen, and David P. Philipp. Geographic patterns in genetic and life history variation in pumpkinseed populations from four East-Central Ontario watersheds. North American Journal of Fisheries Management, 17(2):543– 556, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flebbe:1995:TUW

[FD95] Patricia A. Flebbe and C. Andrew Dolloff. Trout use of woody debris and habitat in Appalachian wilderness streams of North Carolina. North American Journal of Fisheries Management, 15 (3):579–590, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fuselier:1995:MBA

[FE95] Linda Fuselier and David Edds. Management briefs: An artificial riffle as restored habitat for the threatened neosho madtom. North American Journal of Fisheries Management, 15(2):499– 503, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Franzin:1992:TDS

[FH92] W. G. Franzin and S. M. Harbicht. Tests of drift samplers for estimating abundance of recently hatched walleye larvae in small rivers. North American Journal of Fisheries Management, 12(2): 396–405, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fabrizio:1997:SRA

[FHMT97] Mary C. Fabrizio, Mark E. Holey, Patrick C. McKee, and Michael L. Toneys. Survival rates of adult lake trout in Northwestern Lake Michigan, 1983–1993. North American Journal of Fisheries Management, 17(2):413–428, May 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:1992:ESW

[Fie92a] David G. Fielder. Evaluation of stocking walleye fry and fingerlings and factors affecting their success in Lower Lake Oahe, South Dakota. North American Journal of Fisheries Management, 12(2):336–345, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:1992:RBW

[Fie92b] David G. Fielder. Relationship between walleye fingerling stocking density and recruitment in Lower Lake Oahe, South Dakota. North American Journal of Fisheries Management, 12(2):346– 352, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisher:1997:SAP

[Fis97] Mark R. Fisher. Segmentation of the angler population by catch preference, participation, and experience: a managementoriented application of recreation specialization. North American Journal of Fisheries Management, 17(1):1–10, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:1994:WMD

[FJ94] David G. Fielder and Bruce A. Johnson. Walleye mortality during live-release tournaments on Lake Oahe, South Dakota. North American Journal of Fisheries Management, 14(4):776– 780, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fitzgerald:1997:MSI

[FMC97] Thomas J. Fitzgerald, Terry L. Margenau, and Fred A. Copes. Muskellunge scale interpretation: The question of aging accuracy. North American Journal of Fisheries Management, 17(1):206– 209, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frissell:1992:ICP

[FN92] Christopher A. Frissell and Richard K. Nawa. Incidence and causes of physical failure of artificial habitat structures in streams of Western Oregon and Washington. North American Journal of Fisheries Management, 12(1):182–197, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fryer:1998:FPC

[Fry98] Jeffrey K. Fryer. Frequency of pinniped-caused scars and wounds on adult spring–summer Chinook and sockeye salmon returning to the Columbia River. North American Journal of Fisheries Management, 18(1):46–51, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fredericks:1997:USV

[FS97] James P. Fredericks and Dennis L. Scarnecchia. Use of surface visual counts for estimating relative abundance of age-0 paddlefish in Lake Sakakawea. North American Journal of Fisheries Management, 17(4):1014–1018, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fiss:1997:RAS

[FSBB97] Francis C. Fiss, Steven M. Sammons, Phillip W. Bettoli, and Neil Billington. Reproduction among saugeyes (Fx hybrids) and walleyes in Normandy Reservoir, Tennessee. North American Journal of Fisheries Management, 17(1):215–219, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frost:1993:FAB

[FT93] Frank O. Frost and Joan G. Trial. Factors affecting baitfish supply and retail prices paid by Maine anglers. North American Journal of Fisheries Management, 13(3):586–593, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flick:1992:SCB

[FW92] William A. Flick and Dwight A. Webster. Standing crops of brook trout in Adirondack waters before and after removal of non-trout species. North American Journal of Fisheries Management, 12 (4):783–796, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Farrell:1999:DAS

[FW99a] John M. Farrell and Robert G. Werner. Distribution, abundance, and survival of age-0 muskellunge in Upper St. Lawrence River nursery bays. North American Journal of Fisheries Management, 19(1):309–320, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Friesen:1999:MNP

[FW99b] Thomas A. Friesen and David L. Ward. Management of northern pikeminnow and implications for juvenile salmonid survival in the Lower Columbia and Snake Rivers. North American Journal of Fisheries Management, 19(2):406–420, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gabelhouse:1991:SCB

[Gab91] Donald W. Gabelhouse, Jr. Seasonal changes in body condition of white crappies and relations to length and growth in Melvern Reservoir, Kansas. North American Journal of Fisheries Management, 11(1):50–56, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garrett:1995:SMA

[GB95] James W. Garrett and David H. Bennett. Seasonal movements of adult brown trout relative to temperature in a coolwater reservoir. North American Journal of Fisheries Management, 15(2): 480–487, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garrett:1996:EFS

[GB96] James W. Garrett and David H. Bennett. Evaluation of fine sediment intrusion into whitlock-Vibert boxes. North American Journal of Fisheries Management, 16(2):448–452, May 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grogan:1998:EPH

[GB98] Christopher S. Grogan and John Boreman. Estimating the probability that historical populations of fish species are extirpated. North American Journal of Fisheries Management, 18(3):522– 529, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:1999:CIP

[GB99a] Ken Gerow and Canan Bilen. Confidence intervals for percentiles: An application to estimation of potential maximum biomass of trout in Wyoming streams. North American Journal of Fisheries Management, 19(1):149–151, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grandstaff:1999:CSM

[GB99b] E. Dale Grandstaff and Frank J. Bulow. The claw sampler — a macroinvertebrate sampler for scuba divers. North American Journal of Fisheries Management, 19(1):219–220, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garrett:1998:EIS

[GBFT98] James W. Garrett, David H. Bennett, Frank O. Frost, and Russell F. Thurow. Enhanced incubation success for Kokanee spawning in groundwater upwelling sites in a small Idaho stream. North American Journal of Fisheries Management, 18 (4):925–930, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:1999:EBT

[GBTD99] Christopher S. Guy, Matthew N. Burlingame, D. Mosher Thomas, and D. Nygren Douglas. Exemption of bass tournaments from fishing regulations: An opinion survey. North American Journal of Fisheries Management, 19(1):188–191, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gutreuter:1990:ECI

[GC90] Steve Gutreuter and W. Michael Childress. Evaluation of condition indices for estimation of growth of largemouth bass and white crappie. North American Journal of Fisheries Management, 10(4):434–441, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallaway:1999:RJR

[GC99a] Benny J. Gallaway and John G. Cole. Reduction of juvenile red snapper bycatch in the U.S. Gulf of Mexico shrimp trawl fishery. North American Journal of Fisheries Management, 19(2):342– 355, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Groves:1999:SHU

[GC99b] Phillip A. Groves and James A. Chandler. Spawning habitat used by fall Chinook salmon in the Snake River. North American Journal of Fisheries Management, 19(4):912–922, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Groves:1998:TCU

[GG98] Phillip A. Groves and Aaron P. Garcia. Two carriers used to suspend an underwater video camera from a boat. North American Journal of Fisheries Management, 18(4):1004–1007, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerhardt:1991:PDL

[GH91] David R. Gerhardt and Wayne A. Hubert. Population dynamics of a lightly exploited channel catfish stock in the Powder River System, Wyoming–Montana. North American Journal of Fisheries Management, 11(2):200–205, May 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gipson:1993:SSS

[GH93] Robert D. Gipson and Wayne A. Hubert. Spawning-site selection by Kokanee along the shoreline of Flaming Gorge Reservoir, Wyoming–Utah. North American Journal of Fisheries Management, 13(3):475–482, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Green:1994:LRL

[GH94a] David M. Green and Roy C. Heidinger. Longevity record for largemouth bass. North American Journal of Fisheries Management, 14(2):464–465, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Greenberg:1994:ACP

[GH94b] Joshua A. Greenberg and Mark Herrmann. Allocative consequences of pot limits in the Bristol Bay red king crab fishery: An economic analysis. North American Journal of Fisheries Management, 14(2):307–317, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gryska:1997:UPA

[GHG97] Andrew D. Gryska, Wayne A. Hubert, and Kenneth G. Gerow. Use of power analysis in developing monitoring protocols for the endangered Kendall warm springs dace. North American Journal of Fisheries Management, 17(4):1005–1009, November 1997. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Giorgi:1997:FID

[GHS^{+}97] A. E. Giorgi, T. W. Hillman, J. R. Stevenson, S. G. Hays, and C. M. Peven. Factors that influence the downstream migration rates of juvenile salmon and steelhead through the hydroelectric system in the mid-Columbia River Basin. North American Journal of Fisheries Management, 17(2):268–282, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grost:1991:FCT

[GHW91] Richard T. Grost, Wayne A. Hubert, and Thomas A. Wesche. Field comparison of three devices used to sample substrate in small streams. North American Journal of Fisheries Management, 11(3):347–351, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallman:1999:STP

[GITS99] E. A. Gallman, J. J. Isely, J. R. Tomasso, and T. I. J. Smith. Short-term physiological responses of wild and hatcheryproduced red drum during angling. North American Journal of Fisheries Management, 19(3):833–836, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gutreuter:1994:QPS

[GK94] Steve Gutreuter and Daniel J. Krzoska. Quantifying precision of in situ length and weight measurements of fish. North American Journal of Fisheries Management, 14(2):318–322, May 1994. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grewe:1994:SAM

[GKAM94] Peter M. Grewe, Charles C. Krueger, Charles F. Aquadro, and Bernie May. Stability of allozyme and mitochondrial DNA markers among three year-classes of lake trout propagated from Seneca Lake, New York. North American Journal of Fisheries Management, 14(3):467–474, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gresswell:1997:EMP

[GKDG97] Robert E. Gresswell, Jeff Kershner, Jason B. Dunham, and Robert E. Gresswell. Ecology and management of potamodromous salmonids: Introduction to ecology and management of potamodromous salmonids. North American Journal of Fisheries Management, 17(4):1027–1028, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gjernes:1993:MCC

[GKM93] T. Gjernes, A. R. Kronlund, and T. J. Mulligan. Mortality of Chinook and Coho salmon in their first year of ocean life following catch and release by anglers. North American Journal of Fisheries Management, 13(3):524–539, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gresswell:1997:IBS

[GLLB97] Robert E. Gresswell, William J. Liss, Gary L. Larson, and Patrick J. Bartlein. Influence of basin-scale physical variables on life history characteristics of cutthroat trout in Yellowstone Lake. North American Journal of Fisheries Management, 17(4): 1046–1064, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallinat:1997:STS

[GNS97] Michael P. Gallinat, Hock H. Ngu, and J. Dale Shively. Shortterm survival of lake trout released from commercial gill nets in Lake Superior. North American Journal of Fisheries Management, 17(1):136–140, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goeman:1991:WMD

[Goe91] Timothy J. Goeman. Walleye mortality during a live-release tournament on mille lacs, Minnesota. North American Journal of Fisheries Management, 11(1):57–61, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Good:1990:MBF

[Goo90] W. R. Good. Management briefs: a fyke net for rescue of fish stranded in culverts. North American Journal of Fisheries Management, 10(1):115–117, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goodyear:1996:MSR

[Goo96a] C. Phillip Goodyear. Minimum sizes for red grouper: Consequences of considering variable size at age. North American Journal of Fisheries Management, 16(3):505–511, August 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goodyear:1996:VFM

[Goo96b] C. Phillip Goodyear. Variability of fishing mortality by age: Consequences for maximum sustainable yield. North American Journal of Fisheries Management, 16(1):8–13, February 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gigliotti:1993:VBT

[GP93] Larry M. Gigliotti and R. Ben Peyton. Values and behaviors of trout anglers, and their attitudes toward fishery management, relative to membership in fishing organizations: a Michigan case study. North American Journal of Fisheries Management, 13 (3):492–501, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grost:1999:SCS

[GP99] Richard T. Grost and Linda Prendergast. Summer counts of stream-resident trout can differ between daytime and night. North American Journal of Fisheries Management, 19(3):837– 841, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gent:1995:LBR

[GPB95] Russell Gent, John Pitlo, Jr., and Tom Boland. Largemouth bass response to habitat and water quality rehabilitation in a backwater of the Upper Mississippi River. North American Journal of Fisheries Management, 15(4):784–793, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gitschlag:1994:FES

[GR94] Gregg R. Gitschlag and Maurice L. Renaud. Field experiments on survival rates of caged and released red snapper. North American Journal of Fisheries Management, 14(1):131–136, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griffith:1997:PCN

[GR97] David Griffith and Roger A. Rulifson. Preliminary characterization of the North Carolina autumn recreational shrimp trawl fishery. North American Journal of Fisheries Management, 17 (3):696–702, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graham:1992:VEF

[Gra92] Robert J. Graham. Visually estimating fish density at artificial structures in Lake Anna, Virginia. North American Journal of Fisheries Management, 12(1):204–212, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gresswell:1991:UAR

[Gre91] Robert E. Gresswell. Use of antimycin for removal of brook trout from a tributary of Yellowstone Lake. North American Journal of Fisheries Management, 11(1):83–90, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Greenberg:1992:MBF

[Gre92] Larry A. Greenberg. Management briefs: Field survival of brown trout eggs in a perforated incubation container. North American Journal of Fisheries Management, 12(4):833–835, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gregg:1995:CTM

[Gre95] Kurtis L. Gregg. Comparisons of three manufactured artificial reef units in onslow Bay, North Carolina. North American Journal of Fisheries Management, 15(2):316–324, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griffith:1993:UWC

[GS93] J. S. Griffith and R. W. Smith. Use of winter concealment cover by juvenile cutthroat and brown trout in the South Fork of the Snake River, Idaho. North American Journal of Fisheries Management, 13(4):823–830, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griffith:1995:FSM

[GS95] J. S. Griffith and R. W. Smith. Failure of submersed macrophytes to provide cover for rainbow trout throughout their first winter in the Henry’s Fork of the Snake River, Idaho. North American Journal of Fisheries Management, 15(1):42–48, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:1996:CWT

[GSC96] Christopher S. Guy, Randall D. Schultz, and Christopher P. Clouse. Coded wire tag loss from paddlefish: a function of study location. North American Journal of Fisheries Management, 16 (4):931–934, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gregory:1993:CUS

[GSM93] Robert S. Gregory, James A. Servizi, and Dennis W. Martens. Comment: Utility of the stress index for predicting suspended sediment effects. North American Journal of Fisheries Management, 13(4):868–873, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Green:1995:SSM

[GSMG95] David M. Green, Harold L. Schramm, Jr., Paul E. McKeown, and David M. Green. Special section: Managing black bass in northern waters: Managing black bass in northern waters: Summary of the workshop. North American Journal of Fisheries Management, 15(3):671–679, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goeman:1993:ELB

[GSP93] Timothy J. Goeman, Paul D. Spencer, and Rodney B. Pierce. Effectiveness of liberalized bag limits as management tools for altering northern pike population size structure. North American Journal of Fisheries Management, 13(3):621–624, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gigliotti:1990:EIH

[GT90] Larry M. Gigliotti and William W. Taylor. The effect of illegal harvest on recreational fisheries. North American Journal of Fisheries Management, 10(1):106–110, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guillory:1993:GFB

[Gui93] Vincent Guillory. Ghost fishing by blue crab traps. North American Journal of Fisheries Management, 13(3):459–466, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:1990:SRL

[GW90] Christopher S. Guy and David W. Willis. Structural relationships of largemouth bass and bluegill populations in South Dakota ponds. North American Journal of Fisheries Management, 10 (3):338–343, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:1991:ELB

[GW91] Christopher S. Guy and David W. Willis. Evaluation of largemouth bass-yellow perch communities in small South Dakota impoundments. North American Journal of Fisheries Management, 11(1):43–49, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:1995:PCB

[GW95] Christopher S. Guy and David W. Willis. Population characteristics of black crappies in South Dakota waters: a case for ecosystem-specific management. North American Journal of Fisheries Management, 15(4):754–765, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gessel:1991:JSG

[GWB^{+}91] Michael H. Gessel, John G. Williams, Dean A. Brege, Richard F. Krcma, and Donald R. Chambers. Juvenile salmonid guidance at the Bonneville Dam Second Powerhouse, Columbia River, 1983– 1989. North American Journal of Fisheries Management, 11 (3):400–412, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:1996:CCP

[GWS96] Christopher S. Guy, David W. Willis, and Randall D. Schultz. Comparison of catch per unit effort and size structure of white crappies collected with trap nets and gill nets. North American Journal of Fisheries Management, 16(4):947–951, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gelwicks:1998:ECA

[GZB98] Kevin R. Gelwicks, David J. Zafft, and Joe P. Bobbitt. Efficacy of carbonic acid as an anesthetic for rainbow trout. North American Journal of Fisheries Management, 18(2):432–438, May 1998.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1995:OVC

[HA95] Wayne A. Hubert and Charles B. Alexander. Observer variation in counts of meristic traits affects fluctuating asymmetry. North American Journal of Fisheries Management, 15(1):156– 158, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1997:RRO

[HA97] W. A. Hubert and C. B. Alexander. Response: Reducing observer error in asymmetry analyses. North American Journal of Fisheries Management, 17(2):599–600, May 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hale:1996:TSU

[Hal96] R. Scott Hale. Threadfin shad use as supplemental prey in reservoir white crappie fisheries in Kentucky. North American Journal of Fisheries Management, 16(3):619–632, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haley:1998:GLT

[Hal98] Nancy Haley. A gastric lavage technique for characterizing diets of sturgeons. North American Journal of Fisheries Management, 18(4):978–981, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hale:1999:GWC

[Hal99] R. Scott Hale. Growth of white crappies in response to temperature and dissolved oxygen conditions in a Kentucky reservoir. North American Journal of Fisheries Management, 19(2):591– 598, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:1990:SEC

[Han90] Lee H. Hanson. Sterilizing effects of cobalt-60 and cesium-137 radiation on male sea lampreys. North American Journal of Fisheries Management, 10(3):352–361, August 1990. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:1995:SCT

[Han95] J. Mark Hanson. Separation of catches of two Atlantic cod stocks in an otter trawl fishery by application of the EM algorithm. North American Journal of Fisheries Management, 15(2):277– 289, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henderson-Arzapalo:1999:ESI

[HARS^{+}99] Anne Henderson-Arzapalo, Paul Rago, Jorgen Skjeveland, Mike Mangold, Priscilla Washington, Jack Howe, and Timothy King. An evaluation of six internal anchor tags for tagging juvenile striped bass. North American Journal of Fisheries Management, 19(2):482–493, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hooe:1991:EHC

[HB91] Michael L. Hooe and D. Homer Buck. Evaluation of F_{1} hybrid crappies as sport fish in small impoundments. North American Journal of Fisheries Management, 11(4):564–571, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Higgins:1996:ELS

[HB96] P. S. Higgins and M. J. Bradford. Evaluation of a large-scale fish salvage to reduce the impacts of controlled flow reduction in a regulated river. North American Journal of Fisheries Management, 16(3):666–673, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heidinger:1998:RSC

[HB98] Roy C. Heidinger and Ronald C. Brooks. Relative survival and contribution of saugers stocked in the peoria pool of the Illinois River, 1990–1995. North American Journal of Fisheries Management, 18(2):374–382, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hale:1995:CCS

[HBJ95] R. Scott Hale, Gerard L. Buynak, and James R. Jackson. Catch comparisons of surface sampling methods for age-0 gizzard and threadfin shad. North American Journal of Fisheries Management, 15(4):862–870, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hendricks:1991:MMA

[HBM91] Michael L. Hendricks, T. R. Bender, Jr., and V. A. Mudrak. Multiple marking of American shad otoliths with tetracycline antibiotics. North American Journal of Fisheries Management, 11(2): 212–219, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:1998:FAR

[HBN^{+}98] Michael J. Hansen, Michael A. Bozek, Justine R. Newby, Steven P. Newman, and Michael D. Staggs. Factors affecting recruitment of walleyes in Escanaba Lake, Wisconsin, 1958– 1996. North American Journal of Fisheries Management, 18(4): 764–774, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hooe:1994:GSR

[HBW94] Michael L. Hooe, D. Homer Buck, and David H. Wahl. Growth, survival, and recruitment of hybrid crappies stocked in small impoundments. North American Journal of Fisheries Management, 14(1):137–142, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hardin:1992:VEC

[HC92] Scott Hardin and Laurence L. Connor. Variability of electrofishing crew efficiency, and sampling requirements for estimating reliable catch rates. North American Journal of Fisheries Management, 12(3):612–617, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hollender:1994:IWB

[HC94] Bruce A. Hollender and Robert F. Carline. Injury to wild brook trout by backpack electrofishing. North American Journal of Fisheries Management, 14(3):643–649, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoffman:1996:WSM

[HCF^{+}96] Gregory C. Hoffman, Daniel W. Coble, Richard V. Frie, Frederick A. Copes, Ronald M. Bruch, and Kendall K. Kamke. Walleye and sauger mortality associated with live-release tournaments on the Lake Winnebago System, Wisconsin. North American Journal of Fisheries Management, 16(2):364–370, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herrmann:1996:EEI

[HCFG96] Mark Herrmann, Keith R. Criddle, Erika M. Feller, and Joshua A. Greenberg. Estimated economic impacts of potential policy changes affecting the total allowable catch for walleye pollock. North American Journal of Fisheries Management, 16(4): 770–782, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hepworth:1999:CSF

[HCO99] Dale K. Hepworth, Charles B. Chamberlain, and M. J. Ottenbacher. Comparative sport fish performance of Bonneville cutthroat trout in three small put-grow-and-take reservoirs. North American Journal of Fisheries Management, 19(3):774–785, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holland:1992:FTS

[HD92] Stephen M. Holland and Robert B. Ditton. Fishing trip satisfaction: a typology of anglers. North American Journal of Fisheries Management, 12(1):28–33, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoxmeier:1996:SPA

[HD96] R. H. John Hoxmeier and Dennis R. Devries. Status of paddlefish in the Alabama waters of the Tennessee River. North American Journal of Fisheries Management, 16(4):935–938, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:1997:SCS

[HD97] Kevin M. Hunt and Robert B. Ditton. The social context of site selection for freshwater fishing. North American Journal of Fisheries Management, 17(2):331–338, May 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:1998:CKM

[HD98] Patrick J. Harris and John M. Dean. Characterization of king mackerel and Spanish mackerel bycatches of South Carolina shrimp trawlers. North American Journal of Fisheries Management, 18(2):439–453, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hepworth:1991:SSE

[HDM91] Dale K. Hepworth, Daniel J. Duffield, and Timothy Modde. Supplemental stocking for estimating population size and first-year survival of fingerling rainbow trout stocked in a Utah reservoir. North American Journal of Fisheries Management, 11(1):11–19, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:1994:DSL

[HES^{+}94] Michael J. Hansen, Mark P. Ebener, Richard G. Schorfhaar, Stephen T. Schram, Donald R. Schreiner, and James H. Selgeby. Declining survival of lake trout stocked during 1963–1986 in U.S. waters of Lake Superior. North American Journal of Fisheries Management, 14(2):395–402, May 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haner:1995:SRN

[HFS^{+}95] Philip V. Haner, Joyce C. Faler, Robin M. Schrock, Dennis W. Rondorf, and Alec G. Maule. Skin reflectance as a nonlethal measure of smoltification for juvenile salmonids. North American Journal of Fisheries Management, 15(4):814–822, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hatch:1998:CES

[HFS^{+}98] Douglas R. Hatch, Jeffery K. Fryer, Matthew Schwartzberg, David R. Pederson, and Andrew Wand. A computerized editing system for video monitoring of fish passage. North American Journal of Fisheries Management, 18(3):694–699, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartman:1992:EIP

[HG92] Kyle J. Hartman and David W. Garton. Electrophoretic identification of partially digested prey of piscivorous fish. North American Journal of Fisheries Management, 12(1):260–263, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hendrickson:1993:AMP

[HG93] Holly M. Hendrickson and Wade L. Griffin. An analysis of management policies for reducing shrimp by-catch in the Gulf of Mexico. North American Journal of Fisheries Management, 13(4): 686–697, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herrmann:1994:RAA

[HG94] Mark Herrmann and Joshua A. Greenberg. A revenue analysis of the Alaska pink salmon fishery. North American Journal of Fisheries Management, 14(3):537–549, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hale:1998:RDC

[HG98] R. Scott Hale and James H. Gray. Retention and detection of coded wire tags and elastomer tags in trout. North American Journal of Fisheries Management, 18(1):197–201, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayes:1997:MCS

[HGH97] Michael C. Hayes, Larry F. Gates, and Steven A. Hirsch. Multiple catches of smallmouth bass in a special regulation fishery. North American Journal of Fisheries Management, 17(1):182– 187, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1994:DEL

[HGMS94] Wayne A. Hubert, Robert D. Gipson, Robert A. McDowell, and Arthur C. Stewart. Diet of Eagle Lake rainbow trout in Lake DeSmet, Wyoming. North American Journal of Fisheries Management, 14(2):457–459, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1994:IRW

[HGW94] Wayne A. Hubert, Robert D. Gipson, and Roy A. Whaley. Interpreting relative weights of lake trout stocks. North American Journal of Fisheries Management, 14(1):212–215, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herger:1996:CHC

[HHY96] Lillian G. Herger, Wayne A. Hubert, and Michael K. Young. Comparison of habitat composition and cutthroat trout abundance at two flows in small mountain streams. North American Journal of Fisheries Management, 16(2):294–301, May 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:1997:OAC

[Hil97] Ryan A. Hill. Optimizing aerial count frequency for the areaunder-the-curve method of estimating escapement. North American Journal of Fisheries Management, 17(2):461–466, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hvidsten:1993:IRR

[HJ93] Nils Arne Hvidsten and Bjørn Ove Johnsen. Increased recapture rate of adult Atlantic salmon released as smolts into large shoals of wild smolts in the River Orkla, Norway. North American Journal of Fisheries Management, 13(2):272–276, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayes:1994:MML

[HJ94] J. W. Hayes and I. G. Jowett. Microhabitat models of large driftfeeding brown trout in three New Zealand rivers. North American Journal of Fisheries Management, 14(4):710–725, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hicks:1999:DMS

[HJ99] Dave Hicks and B. Alan Johnson. A device to measure shell hardness of Dungeness crabs and trial application in the Kodiak Island, Alaska, commercial fishery. North American Journal of Fisheries Management, 19(2):581–590, May 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hard:1997:VEP

[HK97] Alex Hard and Boyd Kynard. Video evaluation of passage efficiency of American shad and sea lamprey in a modified ice harbor fishway. North American Journal of Fisheries Management, 17 (4):981–987, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herke:1992:ESI

[HKKR92] William H. Herke, E. Eric Knudsen, Pamela A. Knudsen, and Barton D. Rogers. Effects of semi-impoundment of Louisiana marsh on fish and crustacean nursery use and export. North American Journal of Fisheries Management, 12(1):151–160, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Howell:1992:DCG

[HL92] W. Huntting Howell and Richard Langan. Discarding of commercial groundfish species in the Gulf of Maine shrimp fishery. North American Journal of Fisheries Management, 12(3):568– 580, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:1999:RES

[HL99a] J. Mark Hanson and M. Lanteigne. Results of an experimental shrimp fishery in Chaleur Bay, Southern Gulf of St. Lawrence. North American Journal of Fisheries Management, 19(3):713– 723, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:1999:SCS

[HL99b] Bret C. Harvey and Thomas E. Lisle. Scour of Chinook salmon redds on suction dredge tailings. North American Journal of Fisheries Management, 19(2):613–617, May 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:1996:EMS

[HLBD96] Tracy D. Hill, Sean T. Lynott, Scott D. Bryan, and Walter G. Duffy. An efficient method for setting gill nets under ice. North American Journal of Fisheries Management, 16(4): 960–962, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilderbrand:1999:HSI

[HLD99] Robert H. Hilderbrand, A. Dennis Lemly, and C. Andrew Dolloff. Habitat sequencing and the importance of discharge in inferences. North American Journal of Fisheries Management, 19(1):198– 202, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilderbrand:1998:DCL

[HLDH98] Robert H. Hilderbrand, A. Dennis Lemly, C. Andrew Dolloff, and Kelly L. Harpster. Design considerations for large woody debris placement in stream enhancement projects. North American Journal of Fisheries Management, 18(1):161–167, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hale:1999:EMM

[HLMP99] R. Scott Hale, Marty E. Lundquist, Randy L. Miller, and Raymond W. Petering. Evaluation of a 254-mm minimum length limit on crappies in Delaware Reservoir, Ohio. North American Journal of Fisheries Management, 19(3):804–814, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hammers:1991:CME

[HM91] Brad E. Hammers and L. Esteban Miranda. Comparison of methods for estimating age, growth, and related population characteristics of white crappies. North American Journal of Fisheries Management, 11(4):492–498, November 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:1992:MHS

[HM92] David A. Hanson and Terry L. Margenau. Movement, habitat selection, behavior, and survival of stocked muskellunge. North American Journal of Fisheries Management, 12(3):474–483, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hesse:1993:AHM

[HM93] Larry W. Hesse and Gerald E. Mestl. An alternative hydrograph for the Missouri River based on the precontrol condition. North American Journal of Fisheries Management, 13(2):360–366, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartley:1995:DMB

[HM95] Richard A. Hartley and John R. Moring. Differences in mortality between largemouth and smallmouth bass caught in tournaments. North American Journal of Fisheries Management, 15 (3):666–670, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haines:1996:EMT

[HM96] G. Bruce Haines and Timothy Modde. Evaluation of marking techniques to estimate population size and first-year survival of Colorado squawfish. North American Journal of Fisheries Management, 16(4):905–912, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hillman:1992:AUC

[HMG92] T. W. Hillman, J. W. Mullan, and J. S. Griffith. Accuracy of underwater counts of juvenile Chinook salmon, Coho salmon, and steelhead. North American Journal of Fisheries Management, 12 (3):598–603, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1996:EPM

[HMG^{+}96] Wayne A. Hubert, Todd D. Marwitz, Kenneth G. Gerow, N. Allen Binns, and Robert W. Wiley. Estimation of potential maximum biomass of trout in Wyoming streams to assist management decisions. North American Journal of Fisheries Management, 16(4): 821–829, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henderson:1992:CCM

[HN92] Bryan A. Henderson and Stephen J. Nepszy. Comparison of catches in mono- and multifilament gill nets in Lake Erie. North American Journal of Fisheries Management, 12(3):618–624, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:1993:TSS

[HN93] J. J. Hunt and J. D. Neilson. Is there a separate stock of Atlantic cod in the western side of the Bay of Fundy. North American Journal of Fisheries Management, 13(3):421–436, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoff:1995:MBC

[HN95] Michael H. Hoff and Steven P. Newman. Management briefs: Comparisons of fingerling and yearling lake trout introductions for establishing an adult population in Pallette Lake, Wisconsin. North American Journal of Fisheries Management, 15(4): 871–873, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haro:1999:ESH

[HOCSN99] Alex Haro, Mufeed Odeh, Theodore Castro-Santos, and John Noreika. Effect of slope and headpond on passage of American shad and blueback herring through simple Denil and deepened Alaska steeppass fishways. North American Journal of Fisheries Management, 19(1):51–58, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoff:1995:CEI

[Hof95] Michael H. Hoff. Comparisons of the effects of 10-inch, 8-inch, and no minimum length limits on the smallmouth bass population and fishery in Nebish Lake, Wisconsin. North American Journal of Fisheries Management, 15(1):95–102, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hogue:1999:ARR

[Hog99] Cheryl C. Hogue. Avoidance responses of rainbow trout and Utah chub to rotenone. North American Journal of Fisheries Management, 19(1):171–179, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hooe:1991:CBM

[Hoo91] Michael L. Hooe. Crappie biology and management. North American Journal of Fisheries Management, 11(4):483–484, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

House:1995:TVA

[Hou95] Robert House. Temporal variation in abundance of an isolated population of cutthroat trout in Western Oregon, 1981–1991. North American Journal of Fisheries Management, 15(1):33–41, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

House:1996:ESR

[Hou96] Robert House. An evaluation of stream restoration structures in a coastal Oregon stream, 1981–1993. North American Journal of Fisheries Management, 16(2):272–281, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holland:1992:DCH

[HP92] Richard S. Holland and Edward J. Peters. Differential catch by hoop nets of three mesh sizes in the Lower Platte River. North American Journal of Fisheries Management, 12(1):237– 243, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hockersmith:1997:UGP

[HP97] Eric E. Hockersmith and Bradley W. Peterson. Use of the global positioning system for locating radio-tagged fish from aircraft. North American Journal of Fisheries Management, 17(2):457– 460, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heede:1990:HFM

[HR90] Burchard H. Heede and John N. Rinne. Hydrodynamic and fluvial morphologic processes: Implications for fisheries management and research. North American Journal of Fisheries Management, 10(3):249–268, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoenig:1993:SCI

[HRJP93] John M. Hoenig, Douglas S. Robson, Cynthia M. Jones, and Kenneth H. Pollock. Scheduling counts in the instantaneous and progressive count methods for estimating sportfishing effort. North American Journal of Fisheries Management, 13(4): 723–736, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hines:1999:PLP

[HRM99] Kenneth L. Hines, Roger A. Rulifson, and James D. Murray. Performance of low-profile skimmer trawls in the inshore shrimp fishery of North Carolina. North American Journal of Fisheries Management, 19(2):569–580, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanna:1993:ATV

[HS93] Susan S. Hanna and Courtland L. Smith. Attitudes of trawl vessel captains about work, resource use, and fishery management. North American Journal of Fisheries Management, 13(2):367– 375, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huntsman:1994:SIF

[HS94] Gene R. Huntsman and William E. Schaaf. Simulation of the impact of fishing on reproduction of a protogynous grouper, the graysby. North American Journal of Fisheries Management, 14 (1):41–52, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Habera:1996:STM

[HSCM96] James W. Habera, Richard J. Strange, Bart D. Carter, and Stephen E. Moore. Short-term mortality and injury of rainbow trout caused by three-pass AC electrofishing in a Southern Appalachian stream. North American Journal of Fisheries Management, 16(1):192–200, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:1990:CWL

[HSL90] Michael J. Hansen, Paul T. Schultz, and Becky A. Lasee. Changes in Wisconsin’s Lake Michigan salmonid sport fishery, 1969–1985. North American Journal of Fisheries Management, 10(4):442– 457, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hatch:1994:EPS

[HSM94] Douglas R. Hatch, Matthew Schwartzberg, and Phillip R. Mundy. Estimation of Pacific salmon escapement with a time-lapse video recording technique. North American Journal of Fisheries Management, 14(3):626–635, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:1995:AID

[HSP^{+}95] Michael J. Hansen, Richard G. Schorfhaar, James W. Peck, James H. Selgeby, and William W. Taylor. Abundance indices for determining the status of lake trout restoration in Michigan waters of Lake Superior. North American Journal of Fisheries Management, 15(4):830–837, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herrick:1994:ABC

[HSS^{+}94] Samuel F. Herrick, Jr., Ivar Strand, Dale Squires, Morton Miller, Douglas Lipton, John Walden, and Stephen Freese. Application of benefit-cost analysis to fisheries allocation decisions: The case of Alaska walleye pollock and Pacific cod. North American Journal of Fisheries Management, 14(4):726–741, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:1998:GNS

[HSS98] Michael J. Hansen, Richard G. Schorfhaar, and James H. Selgeby. Gill-net saturation by lake trout in Michigan waters of Lake Superior. North American Journal of Fisheries Management, 18(4): 847–853, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Habera:1999:AEI

[HSS99] James W. Habera, Richard J. Strange, and Arnold M. Saxton. AC electrofishing injury of large brown trout in low-conductivity streams. North American Journal of Fisheries Management, 19 (1):120–126, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Helm:1992:ICT

[HT92a] William T. Helm and Harold M. Tyus. Influence of coating type on retention of dummy transmitters implanted in rainbow trout. North American Journal of Fisheries Management, 12(1):257– 259, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hume:1992:FET

[HT92b] Jeremy M. B. Hume and Kanji Tsumura. Field evaluation of two rainbow trout strains introduced into three British Columbia lakes. North American Journal of Fisheries Management, 12 (3):465–473, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hendricks:1994:UOM

[HTB94] Michael L. Hendricks, Donald L. Torsello, and Thomas W. H. Backman. Use of otolith microstructure to distinguish wild from hatchery-reared American shad in the Susquehanna River. North American Journal of Fisheries Management, 14(1):151– 161, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Helser:1996:ASB

[HTM96] Thomas E. Helser, Eric M. Thunberg, and Ralph K. Mayo. An age-structured bioeconomic simulation of U.S. silver hake fisheries. North American Journal of Fisheries Management, 16(4): 783–794, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harmon:1994:PMM

[HTMP94] Jerrel R. Harmon, Kenneth L. Thomas, Kenneth W. McIntyre, and Neil N. Paasch. Prevalence of marine-mammal tooth and claw abrasions on adult anadromous salmonids returning to the Snake River. North American Journal of Fisheries Management, 14(3):661–663, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayes:1995:PBI

[HTS95] Daniel B. Hayes, William W. Taylor, and Harold L. Schramm, Jr. Predicting the biological impact of competitive fishing. North American Journal of Fisheries Management, 15(2):457–472, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:1994:IWC

[HW94] Tracy D. Hill and David W. Willis. Influence of water conductivity on pulsed AC and pulsed DC electrofishing catch rates for largemouth bass. North American Journal of Fisheries Management, 14(1):202–207, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hightower:1996:HTA

[HWE96] Joseph E. Hightower, Anton M. Wicker, and Keith M. Endres. Historical trends in abundance of American shad and river herring in Albemarle Sound, North Carolina. North American Journal of Fisheries Management, 16(2):257–271, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hogle:1993:TPH

[HWH93] Jeffrey S. Hogle, Thomas A. Wesche, and Wayne A. Hubert. A test of the precision of the habitat quality index model II. North American Journal of Fisheries Management, 13(3):640–643, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Irvine:1992:SES

[IB92] J. R. Irvine and R. E. Bailey. Some effects of stocking Coho salmon fry and supplemental instream feeding on wild and hatchery-origin salmon. North American Journal of Fisheries Management, 12(1):125–130, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:1999:EIS

[IBS99] Daniel A. Isermann, Phillip W. Bettoli, and Steve M. Sammons. Efficacy of identifying stocked crappies in a Tennessee reservoir through oxytetracycline marking. North American Journal of Fisheries Management, 19(4):1122–1123, November 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isaak:1997:LBU

[IH97] Daniel J. Isaak and Wayne A. Hubert. A live-bucket for use in surveys of small streams. North American Journal of Fisheries Management, 17(4):1025–1026, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isaak:1999:APS

[IHK99] Daniel J. Isaak, Wayne A. Hubert, and Kirk L. Krueger. Accuracy and precision of stream reach water surface slopes estimated in the field and from maps. North American Journal of Fisheries Management, 19(1):141–148, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Irwin:1997:DAL

[INJ97] Elise R. Irwin, Richard L. Noble, and James R. Jackson. Distribution of age-0 largemouth bass in relation to shoreline landscape features. North American Journal of Fisheries Management, 17 (4):882–893, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isely:1998:EFA

[IT98] J. Jeffery Isely and Joseph R. Tomasso. Estimating fish abundance in a large reservoir by mark–recapture. North American Journal of Fisheries Management, 18(2):269–273, May 1998. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Irvine:1991:EMC

[IWTC91] J. R. Irvine, B. R. Ward, P. A. Teti, and N. B. F. Cousens. Evaluation of a method to count and measure live salmonids in the field with a video camera and computer. North American Journal of Fisheries Management, 11(1):20–26, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:1995:DSS

[Jac95] Donald C. Jackson. Distribution and stock structure of blue catfish and channel catfish in macrohabitats along riverine sections of the Tennessee–tombigbee waterway. North American Journal of Fisheries Management, 15(4):845–853, November 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:1996:TCV

[Jac96] Peter C. Jacobson. Trophy and consumptive value-per-recruit analysis for a walleye fishery. North American Journal of Fisheries Management, 16(1):75–80, February 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:1999:CEI

[JBH^{+}99] Martin J. Jennings, Michael A. Bozek, Gene R. Hatzenbeler, Edward E. Emmons, and Michael D. Staggs. Cumulative effects of incremental shoreline habitat modification on fish assemblages in North temperate lakes. North American Journal of Fisheries Management, 19(1):18–27, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1995:EBC

[JBHW95] Donn M. Johnson, Robert J. Behnke, David A. Harpman, and Richard G. Walsh. Economic benefits and costs of stocking catchable rainbow trout: a synthesis of economic analysis in Colorado. North American Journal of Fisheries Management, 15(1):26–32, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:1997:EPS

[JCP97] Martin J. Jennings, Julie E. Claussen, and David P. Philipp. Effect of population size structure on reproductive investment of male bluegill. North American Journal of Fisheries Management, 17(2):516–524, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1992:PRR

[JD92] James H. Johnson and David S. Dropkin. Predation on recently released larval American shad in the Susquehanna River Basin. North American Journal of Fisheries Management, 12(3):504– 508, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jenkins:1997:DBH

[JDCS97] Wallace E. Jenkins, Michael R. Denson, Mark R. Collins, and Theodore I. J. Smith. Differentiating between hatchery and wild red drum by the position of first annulus on sagitta. North American Journal of Fisheries Management, 17(4):1001–1004, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:1998:SLC

[JDL^{+}98] Kim K. Jones, Jeffrey M. Dambacher, Brad G. Lovatt, Andrew G. Talabere, and Wayne Bowers. Status of Lahontan cutthroat trout in the Coyote Lake Basin, Southeast Oregon. North American Journal of Fisheries Management, 18(2):308–317, May 1998. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jessop:1990:PHR

[Jes90a] B. M. Jessop. Passage and harvest of river herring at the Mactaquac Dam, Saint John River: An attempt at active fishery management. North American Journal of Fisheries Management, 10 (1):33–38, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jessop:1990:SRR

[Jes90b] B. M. Jessop. Stock-recruitment relationships of alewives and blueback herring returning to the Mactaquac Dam, Saint John River, New Brunswick. North American Journal of Fisheries Management, 10(1):19–32, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jessop:1994:RBS

[Jes94] B. M. Jessop. Relations between stock and environmental variables, and an index of abundance, for juvenile alewives and blueback herring. North American Journal of Fisheries Management, 14(3):564–579, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1999:EEE

[JFB99] Darlene R. Johnson, Nicholas A. Funicelli, and James A. Bohnsack. Effectiveness of an existing estuarine no-take fish sanctuary within the Kennedy space center, Florida. North American Journal of Fisheries Management, 19(2):436–453, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jakus:1996:EPS

[JFW96] Paul M. Jakus, J. Mark Fly, and J. Larry Wilson. Explaining public support for fisheries management alternatives. North American Journal of Fisheries Management, 16(1):41–48, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:1997:HGB

[JFY97] Donald C. Jackson, Jay M. Francis, and Qifeng Ye. Hand grappling blue catfish in the main channel of a Mississippi River. North American Journal of Fisheries Management, 17(4):1019– 1024, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1999:FWC

[JG99] Brett M. Johnson and John P. Goettl, Jr. Food Web changes over fourteen years following introduction of rainbow smelt into a Colorado reservoir. North American Journal of Fisheries Management, 19(3):629–642, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:1994:CWF

[JGF94] Mark S. Jones, John P. Goettl, Jr., and Stephen A. Flickinger. Changes in walleye food habits and growth following a rainbow smelt introduction. North American Journal of Fisheries Management, 14(2):409–414, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jessop:1990:EDP

[JH90] B. M. Jessop and C. J. Harvie. Evaluation of designs of periodic count surveys for the estimation of escapement at a fishway. North American Journal of Fisheries Management, 10(1):39–45, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:1993:BDM

[JH93] Larry D. Jacobson and John R. Hunter. Bathymetric demography and management of Dover sole. North American Journal of Fisheries Management, 13(3):405–420, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:1997:CRB

[JJ97] Martin J. Jennings and Martin J. Jennings. Centrarchid reproductive behavior: Centrarchid reproductive behavior: Implications for management. North American Journal of Fisheries Management, 17(2):493–495, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1998:HAS

[JJ98] Barry L. Johnson and Cecil A. Jennings. Habitat associations of small fishes around islands in the Upper Mississippi River. North American Journal of Fisheries Management, 18(2):327–336, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1992:PUA

[JL92] David L. Johnson and William E. Lynch, Jr. Panfish use of and angler success at evergreen tree, brush, and stake-bed structures. North American Journal of Fisheries Management, 12(1):222– 229, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:1998:ETT

[JL98] Cecil A. Jennings and Gregory L. Looney. Evaluation of two types of anesthesia for performing surgery on striped bass. North American Journal of Fisheries Management, 18(1):187–190, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1992:FEH

[JLS^{+}92] Brett M. Johnson, Chris Luecke, R. Scot Stewart, Michael D. Staggs, Stephen J. Gilbert, and James F. Kitchell. Forecasting effects of harvest regulations and stocking of walleyes on prey fish communities in Lake Mendota, Wisconsin. North American Journal of Fisheries Management, 12(4):797–807, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1993:GSS

[JM93] Brett M. Johnson and Terry L. Margenau. Growth and sizeselective mortality of stocked muskellunge: Effects on size distributions. North American Journal of Fisheries Management, 13 (3):625–629, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1992:EFR

[JMBK92] Barry L. Johnson, Scott R. Milliman, Richard C. Bishop, and James F. Kitchell. Evaluating fishery rehabilitation under uncertainty: a bioeconomic analysis of quota management for the Green Bay yellow perch fishery. North American Journal of Fisheries Management, 12(4):703–720, November 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1996:CDH

[JMD96] James H. Johnson, Joseph E. McKeon, and David S. Dropkin. Comparative diets of hatchery and wild Atlantic salmon smolts in the Merrimack River. North American Journal of Fisheries Management, 16(2):440–444, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jeric:1995:EMM

[JMG95] Randall J. Jeric, Timothy Modde, and Jeff M. Godfrey. Evaluation of a method for measuring intragravel dissolved oxygen concentrations and survival to emergence in shore-spawned salmonids. North American Journal of Fisheries Management, 15 (1):185–192, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:1995:SSM

[JN95] James R. Jackson and Richard L. Noble. Selectivity of sampling methods for juvenile largemouth bass in assessments of recruitment processes. North American Journal of Fisheries Management, 15(2):408–418, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1992:EES

[JNT92] James H. Johnson, Anthony A. Nigro, and Ray Temple. Evaluating enhancement of striped bass in the context of potential predation on anadromous salmonids in coos Bay, Oregon. North American Journal of Fisheries Management, 12(1):103–108, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1995:ACS

[Joh95] Barry L. Johnson. Applying computer simulation models as learning tools in fishery management. North American Journal of Fisheries Management, 15(4):736–747, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jowett:1992:MAL

[Jow92] Ian G. Jowett. Models of the abundance of large brown trout in New Zealand rivers. North American Journal of Fisheries Management, 12(3):417–432, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:1992:UAM

[JP92] Martin J. Jennings and David P. Philipp. Use of allozyme markers to evaluate walleye stocking success. North American Journal of Fisheries Management, 12(2):285–290, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnston:1993:LGH

[JPT93] N. T. Johnston, E. A. Parkinson, and K. Tsumura. Longevity and growth of hormone-sterilized Kokanee. North American Journal of Fisheries Management, 13(2):284–290, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:1998:OAP

[JQV98] Edgar L. Jones III, Terrance J. Quinn II, and Benjamin W. Van Alen. Observer accuracy and precision in aerial and foot survey counts of pink salmon in a Southeast Alaska stream. North American Journal of Fisheries Management, 18(4):832–846, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1995:ELP

[JR95] James H. Johnson and Neil H. Ringler. Estimating losses to predation of recently released American shad larvae in the Juniata River, Pennsylvania. North American Journal of Fisheries Management, 15(4):854–861, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:1998:PBM

[JR98] Mark S. Jones and Kevin B. Rogers. Palmetto bass movements and habitat use in a fluctuating Colorado irrigation reservoir. North American Journal of Fisheries Management, 18(3):640– 648, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1992:FIS

[JRH92] Shawn L. Johnson, Frank J. Rahel, and Wayne A. Hubert. Factors influencing the size structure of brook trout populations in Beaver ponds in Wyoming. North American Journal of Fisheries Management, 12(1):118–124, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:1995:RAP

[JS95] Michael L. Jones and Jason D. Stockwell. A rapid assessment procedure for the enumeration of salmonine populations in streams. North American Journal of Fisheries Management, 15(3):551– 562, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1994:SPH

[JSD94] Gary E. Johnson, John R. Skalski, and Don J. Degan. Statistical precision of hydroacoustic sampling of fish entrainment at hydroelectric facilities. North American Journal of Fisheries Management, 14(2):323–333, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1990:ESH

[JSN90] Steven L. Johnson, Mario F. Solazzi, and Thomas E. Nickelson. Effects on survival and homing of trucking hatchery yearling Coho salmon to release sites. North American Journal of Fisheries Management, 10(4):427–433, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:1993:OCD

[JT93] Larry D. Jacobson and Cynthia J. Thomson. Opportunity costs and the decision to fish for northern anchovy. North American Journal of Fisheries Management, 13(1):27–34, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jude:1992:ENR

[Jud92] David J. Jude. Evidence for natural reproduction by stocked walleyes in the Saginaw River Tributary System, Michigan. North American Journal of Fisheries Management, 12(2):386–395, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1998:EAC

[JW98] Donn M. Johnson and Richard G. Walsh. An economic analysis of the cost-effectiveness of privatizing additional production of catchable rainbow trout: a case study in Colorado. North American Journal of Fisheries Management, 18(1):168–174, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kayle:1992:UOD

[Kay92] Kevin A. Kayle. Use of oxytetracycline to determine the contribution of stocked fingerling walleyes. North American Journal of Fisheries Management, 12(2):353–355, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kynard:1997:PBA

[KB97] Boyd Kynard and Christopher Buerkett. Passage and behavior of adult American shad in an experimental louver bypass system. North American Journal of Fisheries Management, 17(3):734– 742, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klumb:1999:VDL

[KBF99] Robert A. Klumb, Michael A. Bozek, and Richard V. Frie. Validation of the Dahl–Lea and Fraser–Lee back-calculation models by using oxytetracycline-marked bluegills and bluegill × green sunfish hybrids. North American Journal of Fisheries Management, 19(2):504–514, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kershner:1997:PHG

[KBH97] Jeffrey L. Kershner, Catherine M. Bischoff, and Dona L. Horan. Population, habitat, and genetic characteristics of Colorado River cutthroat trout in wilderness and nonwilderness stream sections in the uinta mountains of Utah and Wyoming. North American Journal of Fisheries Management, 17(4):1134–1143, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keniry:1996:EDP

[KBHM96] Michael J. Keniry, Wayne A. Brofka, William H. Horns, and J. Ellen Marsden. Effects of decompression and puncturing the gas bladder on survival of tagged yellow perch. North American Journal of Fisheries Management, 16(1):201–206, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kane:1990:EHS

[KBM90] Andrew S. Kane, Richard O. Bennett, and Eric B. May. Effect of hardness and salinity on survival of striped bass larvae. North American Journal of Fisheries Management, 10(1):67–71, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kincaid:1991:GSS

[KE91] Harold L. Kincaid and Joseph H. Elrod. Growth and survival of stocked lake trout with nuclear cataracts in Lake Ontario. North American Journal of Fisheries Management, 11(3):429–434, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koonce:1993:EIL

[KEC93] Joseph F. Koonce, Randy L. Eshenroder, and Gavin C. Christie. An economic injury level approach to establishing the intensity of sea lamprey control in the Great Lakes. North American Journal of Fisheries Management, 13(1):1–14, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelly:1993:DLT

[Kel93] Kevin H. Kelly. Determination of lobster trap density near midcoastal Maine by aerial photography. North American Journal of Fisheries Management, 13(4):859–863, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kempinger:1996:HGF

[Kem96] James J. Kempinger. Habitat, growth, and food of young lake sturgeons in the Lake Winnebago System, Wisconsin.

North American Journal of Fisheries Management, 16(1):102– 114, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kocovsky:1997:SIR

[KGFR97] Patrick M. Kocovsky, Charles Gowan, Kurt D. Fausch, and Stephen C. Riley. Spinal injury rates in three wild trout populations in Colorado after eight years of backpack electrofishing. North American Journal of Fisheries Management, 17(2):308– 313, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kruse:1993:COS

[KGW93] Carter G. Kruse, Christopher S. Guy, and David W. Willis. Comparison of otolith and scale age characteristics for black crappies collected from South Dakota waters. North American Journal of Fisheries Management, 13(4):856–858, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kwak:1995:LBM

[KH95] Thomas J. Kwak and Mary G. Henry. Largemouth bass mortality and related causal factors during live-release fishing tournaments on a large Minnesota lake. North American Journal of Fisheries Management, 15(3):621–630, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krueger:1997:ISO

[KH97a] Kirk L. Krueger and Wayne A. Hubert. Irregularly shaped otoliths from saugers prevent back-calculation of length at previous ages in Wyoming. North American Journal of Fisheries Management, 17(3):769–772, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kruse:1997:PSW

[KH97b] Carter G. Kruse and Wayne A. Hubert. Proposed standard weight (W_{s}) equations for interior cutthroat trout. North American Journal of Fisheries Management, 17(3):784–790, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krieger:1998:IHR

[KH98] Douglas J. Krieger and John P. Hoehn. Improving health risk information for sport anglers: The value of information about chemical residues. North American Journal of Fisheries Management, 18(2):411–421, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krueger:1998:TSF

[KHP98] Kirk L. Krueger, Wayne A. Hubert, and Robert M. Price. Tandem-set fyke nets for sampling benthic fishes in lakes. North American Journal of Fisheries Management, 18(1):154– 160, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kruse:1998:SPE

[KHR98] Carter G. Kruse, Wayne A. Hubert, and Frank J. Rahel. Singlepass electrofishing predicts trout abundance in mountain streams with sparse habitat. North American Journal of Fisheries Management, 18(4):940–946, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kirk:1992:ETG

[Kir92] James P. Kirk. Efficacy of triploid grass carp in controlling nuisance aquatic vegetation in South Carolina farm ponds. North American Journal of Fisheries Management, 12(3):581–584, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kircheis:1998:SCE

[Kir98] Frederick W. Kircheis. Species composition and economic value of Maine’s winter baitfish industry. North American Journal of Fisheries Management, 18(1):175–180, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kissick:1993:CTL

[Kis93] Lee A. Kissick. Comparison of traps lighted by photochemicals or electric bulbs for sampling warmwater populations of young fish. North American Journal of Fisheries Management, 13(4): 864–867, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kraft:1992:FNG

[KJ92] Clifford E. Kraft and Barry L. Johnson. Fyke-net and gill-net size selectivities for yellow perch in green Bay, Lake Michigan.

North American Journal of Fisheries Management, 12(1):230– 236, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knights:1995:RBB

[KJS95] Brent C. Knights, Barry L. Johnson, and Mark B. Sandheinrich. Responses of bluegills and black crappies to dissolved oxygen, temperature, and current in backwater lakes of the Upper Mississippi River during winter. North American Journal of Fisheries Management, 15(2):390–399, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kincaid:1993:PGV

[KKM93] Harold L. Kincaid, Charles C. Krueger, and Bernie May. Preservation of genetic variation in the Green Lake strain lake trout derived from remnant domestic and feral populations. North American Journal of Fisheries Management, 13(2):318–325, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kircheis:1995:GIT

[KKS95] Frederick W. Kircheis, Irv Kornfield, and Seifu Seyoum. Genetic identity of transplanted Arctic char: Implications for restocking programs. North American Journal of Fisheries Management, 15 (1):54–59, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kirn:1991:SOM

[KL91] Richard A. Kirn and George W. Labar. Stepped-oblique midwater trawling as an assessment technique for rainbow smelt. North American Journal of Fisheries Management, 11(2):167–176, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kanehl:1997:CHF

[KLN97] Paul D. Kanehl, John Lyons, and John E. Nelson. Changes in the habitat and fish community of the Milwaukee River, Wisconsin, following removal of the woolen Mills Dam. North American Journal of Fisheries Management, 17(2):387–400, May 1997. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keenlyne:1993:LCL

[KM93] K. D. Keenlyne and S. J. Maxwell. Length conversions and length–weight relations for pallid sturgeon. North American Journal of Fisheries Management, 13(2):395–397, May 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knapp:1996:LGG

[KM96] Roland A. Knapp and Kathleen R. Matthews. Livestock grazing, golden trout, and streams in the Golden Trout Wilderness, California: Impacts and management implications. North American Journal of Fisheries Management, 16(4):805–820, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kershner:1998:ASE

[KM98] Mark W. Kershner and Elizabeth A. Marschall. Allocating sampling effort to equalize precision of electrofishing catch per unit effort. North American Journal of Fisheries Management, 18(4): 822–831, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kriegler:1995:GCT

[KMHS95] Frank J. Kriegler, Gary F. McCracken, James W. Habera, and Richard J. Strange. Genetic characterization of Tennessee brook trout populations and associated management implications. North American Journal of Fisheries Management, 15(4): 804–813, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knuth:1990:RCN

[Knu90] Barbara A. Knuth. Risk communication: a new dimension in sport-fisheries management. North American Journal of Fisheries Management, 10(4):374–381, November 1990. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kynard:1993:EBS

[KO93] Boyd Kynard and John O’L’a’y. Evaluation of a bypass system for spent American shad at Holyoke Dam, Massachusetts. North American Journal of Fisheries Management, 13(4):782– 789, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knudsen:1996:JWS

[KRP^{+}96] E. Eric Knudsen, Barton D. Rogers, Ronald F. Paille, William H. Herke, and James P. Geaghan. Juvenile white shrimp growth, mortality, and emigration in weired and unweired Louisiana marsh ponds. North American Journal of Fisheries Management, 16(3):640–652, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koppelman:1992:STS

[KSJ92] Jeffrey B. Koppelman, Kevin P. Sullivan, and Philip J. Jeffries, Jr. Survival of three sizes of genetically marked walleyes stocked into two Missouri impoundments. North American Journal of Fisheries Management, 12(2):291–298, May 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kohler:1993:LBH

[KSS93] Christopher C. Kohler, Robert J. Sheehan, and James J. Sweatman. Largemouth bass hatching success and first-winter survival in two Illinois reservoirs. North American Journal of Fisheries Management, 13(1):125–133, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knapp:1996:FCS

[KV96] Roland A. Knapp and Vance T. Vredenburg. A field comparison of the substrate composition of California golden trout redds sampled with two devices. North American Journal of Fisheries Management, 16(3):674–681, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kolander:1993:PRS

[KWM93] Todd D. Kolander, David W. Willis, and Brian R. Murphy. Proposed revision of the standard weight (WS) equation for smallmouth bass. North American Journal of Fisheries Management, 13(2):398–400, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kennen:1994:AMA

[KWRH94] Jonathan G. Kennen, Steven J. Wisniewski, Neil H. Ringler, and Harry M. Hawkins. Application and modification of an Auger trap to quantify emigrating fishes in Lake Ontario tributaries. North American Journal of Fisheries Management, 14 (4):828–836, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kleinsasser:1990:GCN

[KWW90] Leroy J. Kleinsasser, J. Holt Williamson, and B. G. Whiteside. Growth and catchability of northern, Florida, and F, hybrid largemouth bass in Texas ponds. North American Journal of Fisheries Management, 10(4):462–468, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lange:1991:ASI

[Lan91] Anne M. Lange. Alternative survey indices for predicting availability of longfin squid to seasonal Northwest Atlantic fisheries. North American Journal of Fisheries Management, 11(3):443– 450, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:1995:PRC

[Lar95] Larry L. Larson. A portable restraint cradle for handling large salmonids. North American Journal of Fisheries Management, 15 (3):654–656, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Layher:1992:MHE

[LB92] William G. Layher and Kenneth L. Brunson. A modification of the habitat evaluation procedure for determining instream flow requirements in warmwater streams. North American Journal of Fisheries Management, 12(1):47–54, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lee:1996:ITO

[LB96] Wade C. Lee and Eric P. Bergersen. Influence of thermal and oxygen stratification on lake trout hooking mortality. North American Journal of Fisheries Management, 16(1):175–181, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Labay:1999:CFK

[LB99] Andrew A. Labay and Dave Buzan. A comparison of fish kill counting procedures on a small, narrow stream. North American Journal of Fisheries Management, 19(1):209–214, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lajeone:1992:SSF

[LBB92] Larry J. Lajeone, Timothy W. Bowzer, and David L. Bergerhouse. Supplemental stocking of fingerling walleyes in the Upper Mississippi River. North American Journal of Fisheries Management, 12(2):307–312, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larscheid:1999:SGA

[LCGJ99] Joe Larscheid, Jim Christianson, Thomas Gengerke, and Wallace Jorgensen. Survival, growth, and abundance of pellet-reared and minnow-reared muskellunge stocked in Northwestern Iowa. North American Journal of Fisheries Management, 19(1):230– 237, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Li:1996:EWSa

[LCSA96a] Jingyin Li, Yosef Cohen, Dennis H. Schupp, and Ira R. Adelman. Effects of walleye stocking on population abundance and fish size. North American Journal of Fisheries Management, 16 (4):830–839, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Li:1996:EWSb

[LCSA96b] Jingyin Li, Yosef Cohen, Dennis H. Schupp, and Ira R. Adelman. Effects of walleye stocking on year-class strength. North American Journal of Fisheries Management, 16(4):840–850, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lee:1992:GBD

[Lee92] Dennis P. Lee. Gas bladder deflation of depressurized largemouth bass. North American Journal of Fisheries Management, 12(3): 662–664, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Legault:1998:LTC

[Leg98] Christopher M. Legault. Linking total catch quotas and allocation schemes. North American Journal of Fisheries Management, 18(2):454–457, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leonetti:1997:ESI

[Leo97] Frank E. Leonetti. Estimation of surface and intragravel water flow at sockeye salmon spawning beaches in Iliamna Lake, Alaska. North American Journal of Fisheries Management, 17(1):194– 201, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Luecke:1994:SCL

[LEW^{+}94] Chris Luecke, Thomas C. Edwards, Jr., Marvin W. Wengert, Jr., Steve Brayton, and Roger Schneidervin. Simulated changes in lake trout yield, trophies, and forage consumption under various slot limits. North American Journal of Fisheries Management, 14(1):14–21, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loomis:1999:SAE

[LF99] John Loomis and Peter Fix. A statistical approach to estimating costs of propagating hatchery rainbow trout. North American Journal of Fisheries Management, 19(1):110–119, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leary:1996:SBT

[LGS96] Robb E. Leary, William R. Gould, and George K. Sage. Success of basibranchial teeth in indicating pure populations of rainbow trout and failure to indicate pure populations of westslope cutthroat trout. North American Journal of Fisheries Management, 16(1):210–213, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larscheid:1992:FIS

[LH92] Joseph G. Larscheid and Wayne A. Hubert. Factors influencing the size structure of brook trout and brown trout in Southeastern Wyoming mountain streams. North American Journal of Fisheries Management, 12(1):109–117, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:1996:EMM

[LKD96] John Lyons, Paul D. Kanehl, and David M. Day. Evaluation of a 356-mm minimum-length limit for smallmouth bass in Wisconsin streams. North American Journal of Fisheries Management, 16 (4):952–957, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Laine:1991:AUS

[LMR91] Arnold O. Laine, Walter T. Momot, and Philip A. Ryan. Accuracy of using scales and cleithra for aging northern pike from an oligotrophic Ontario lake. North American Journal of Fisheries Management, 11(2):220–225, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lockwood:1997:ECR

[Loc97] Roger N. Lockwood. Evaluation of catch rate estimators from Michigan access point angler surveys. North American Journal of Fisheries Management, 17(3):611–620, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Locke:1998:MEP

[Loc98] Andrea Locke. Modeling the effects of poststocking survival rates on the success of stocking hatchery Atlantic salmon in a New Brunswick river. North American Journal of Fisheries Management, 18(3):547–560, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

LeClair:1999:LGF

[LPT99] Larry L. LeClair, Stevan R. Phelps, and Tim J. Tynan. Little gene flow from a hatchery strain of chum salmon to local wild populations. North American Journal of Fisheries Management, 19(2):530–535, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lee:1997:PVA

[LR97] Danny C. Lee and Bruce E. Rieman. Population viability assessment of salmonids by using probabilistic networks. North American Journal of Fisheries Management, 17(4):1144–1157, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Langhurst:1990:SMS

[LS90] Ross W. Langhurst and Dean L. Schoenike. Seasonal migration of smallmouth bass in the Embarrass and Wolf Rivers, Wisconsin. North American Journal of Fisheries Management, 10(2):224– 227, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lawson:1996:GRM

[LS96] Peter W. Lawson and David B. Sampson. Gear-related mortality in selective fisheries for ocean salmon. North American Journal of Fisheries Management, 16(3):512–520, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lafferty:1999:PPR

[LSA99] Kevin D. Lafferty, Camm C. Swift, and Richard F. Ambrose. Postflood persistence and recolonization of endangered tidewater goby populations. North American Journal of Fisheries Management, 19(2):618–622, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Langton:1996:IBF

[LSG^{+}96] Richard W. Langton, Robert S. Steneck, Vytenis Gotceitas, Francis Juanes, and Peter Lawton. The interface between fisheries research and habitat management. North American Journal of Fisheries Management, 16(1):1–7, February 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:1991:ESB

[LSS91] Scott C. Larson, Bruce Saul, and Steve Schleiger. Exploitation and survival of black crappies in three Georgia reservoirs. North American Journal of Fisheries Management, 11(4):604– 613, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Letcher:1998:TMA

[LT98] Benjamin H. Letcher and Timothy D. Terrick. Thermal marking of Atlantic salmon otoliths. North American Journal of Fisheries Management, 18(2):406–410, May 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Low:1991:SUM

[LW91] R. A. Low, Jr. and C. W. Waltz. Seasonal utilization and movement of black sea bass on a South Carolina artificial reef. North American Journal of Fisheries Management, 11(2):131–138, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:1996:DVI

[LWS96] John Lyons, Lizhu Wang, and Timothy D. Simonson. Development and validation of an index of biotic integrity for coldwater streams in Wisconsin. North American Journal of Fisheries Management, 16(2):241–256, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:1992:LSS

[Lyo92] John Lyons. The length of stream to sample with a towed electrofishing unit when fish species richness is estimated. North American Journal of Fisheries Management, 12(1):198–203, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Malvestuto:1995:ASM

[Mal95] Stephen P. Malvestuto. Angler survey methods and their applications in fisheries management. North American Journal of Fisheries Management, 15(1):238–240, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:1992:SCE

[Mar92] Terry L. Margenau. Survival and cost-effectiveness of stocked fall fingerling and spring yearling muskellunge in Wisconsin. North American Journal of Fisheries Management, 12(3):484–493, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:1995:ESL

[Mar95] Catherine C. Martin. Evaluation of slot length limits for largemouth bass in two Delaware ponds. North American Journal of Fisheries Management, 15(3):713–719, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:1999:ISM

[Mar99a] Terry L. Margenau. Introduction to strategies for muskellunge management. North American Journal of Fisheries Management, 19(1):221–222, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:1999:MSS

[Mar99b] Terry L. Margenau. Muskellunge stocking strategies in Wisconsin: The first century and beyond. North American Journal of Fisheries Management, 19(1):223–229, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matthews:1996:HSM

[Mat96] Kathleen R. Matthews. Habitat selection and movement patterns of California golden trout in degraded and recovering stream sections in the Golden Trout Wilderness, California. North American Journal of Fisheries Management, 16(3):579–590, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maxell:1999:PAM

[Max99] Bryce A. Maxell. A power analysis on the monitoring of bull trout stocks using redd counts. North American Journal of Fisheries Management, 19(3):860–866, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morkert:1990:SSL

[MB90] Sidney B. Morkert and Roger A. Bergstedt. Shrinkage of sea lamprey larvae preserved in formalin. North American Journal of Fisheries Management, 10(4):484–486, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDonough:1991:FAA

[MB91] Thomas A. McDonough and Johnny P. Buchanan. Factors affecting abundance of white crappies in Chickamauga Reservoir, Tennessee, 1970–1989. North American Journal of Fisheries Management, 11(4):513–524, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Macdonald:1997:ESS

[MB97] Hugh F. Macdonald and Kevin J. Boyle. Effect of a statewide sport fish consumption advisory on open-water fishing in Maine. North American Journal of Fisheries Management, 17(3):687– 695, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1998:VLB

[MB98] Michael J. Maceina and Phillip W. Bettoli. Variation in largemouth bass recruitment in four mainstream impoundments of the Tennessee River. North American Journal of Fisheries Management, 18(4):998–1003, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meronek:1996:RFC

[MBB^{+}96] Thomas G. Meronek, Patrick M. Bouchard, Edmund R. Buckner, Thomas M. Burri, Karen K. Demmerly, Daniel C. Hatleli, Robert A. Klumb, Stephen H. Schmidt, and Daniel W. Coble. A review of fish control projects. North American Journal of Fisheries Management, 16(1):63–74, February 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murdock:1992:IFD

[MBD^{+}92] Steve H. Murdock, Kenneth Backman, Robert B. Ditton, Md. Nazrul Hoque, and David Ellis. Implications of future demographic change for participation in fishing in Texas. North American Journal of Fisheries Management, 12(3):548–558, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1998:ASF

[MBFD98] Michael J. Maceina, Phillip W. Bettoli, Sean D. Finely, and Victor J. Dicenzo. Analyses of the sauger fishery with simulated effects of a minimum size limit in the Tennessee River of Alabama. North American Journal of Fisheries Management, 18 (1):66–75, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maret:1993:FTN

[MBHC93] T. R. Maret, T. A. Burton, G. W. Harvey, and W. H. Clark. Field testing of new monitoring protocols to assess brown trout spawning habitat in an Idaho stream. North American Journal of Fisheries Management, 13(3):567–580, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:1995:CFT

[MBHJ95] Paul Michaletz, Jeff Boxrucker, Scott Hale, and James R. Jackson. Comparison of four trawls for sampling juvenile shad. North American Journal of Fisheries Management, 15(4):918– 923, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1991:EAM

[MBK^{+}91] Michael J. Maceina, Phillip W. Bettoli, Wallace G. Klussmann, Robert K. Betsill, and Richard L. Noble. Effect of aquatic macrophyte removal on recruitment and growth of black crappies and white crappies in Lake Conroe, Texas. North American Journal of Fisheries Management, 11(4):556–563, November 1991. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1990:ERW

[MBS90] Brian R. Murphy, Michael L. Brown, and Timothy A. Springer. Evaluation of the relative weight (W_{r}) index, with new applications to walleye. North American Journal of Fisheries Management, 10(1):85–97, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marcogliese:1998:SMD

[MC98] Lucian A. Marcogliese and John M. Casselman. Scale methods for discriminating between Great Lakes stocks of wild and hatchery rainbow trout, with a measure of natural recruitment in Lake Ontario. North American Journal of Fisheries Management, 18 (2):253–268, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McInerny:1999:CTM

[MC99] Michael C. McInerny and Timothy K. Cross. Comparison of three mark–recapture sampling designs for estimating population size of largemouth bass in Minnesota lakes. North American Journal of Fisheries Management, 19(3):758–764, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meronek:1997:SBI

[MCC97] Thomas G. Meronek, Fred A. Copes, and Daniel W. Coble. A survey of the bait industry in the North-Central region of the United States. North American Journal of Fisheries Management, 17(3):703–711, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mullins:1991:TCF

[MCCP91] C. C. Mullins, P. L. Caines, D. Caines, and J. L. Peppar. A twocompartment fish trap for simultaneously counting downstream and upstream migrants in small rivers. North American Journal of Fisheries Management, 11(3):358–363, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McGurk:1999:SDN

[McG99] Michael D. McGurk. Size dependence of natural mortality rate of sockeye salmon and Kokanee in freshwater. North American Journal of Fisheries Management, 19(2):376–396, May 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McHugh:1990:RBC

[McH90] James J. McHugh. Responses of bluegills and crappies to reduced abundance of largemouth bass in two Alabama impoundments. North American Journal of Fisheries Management, 10(3):344– 351, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:1993:EEI

[McM93] Geoffrey A. McMichael. Examination of electrofishing injury and short-term mortality in hatchery rainbow trout. North American Journal of Fisheries Management, 13(2):229–233, May 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMenemy:1995:SAS

[McM95] James R. McMenemy. Survival of Atlantic salmon fry stocked at low density in the West River, Vermont. North American Journal of Fisheries Management, 15(2):366–374, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McNeill:1995:OSB

[McN95] Alan J. McNeill. An overview of the smallmouth bass in Nova Scotia. North American Journal of Fisheries Management, 15 (3):680–687, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McInerny:1991:DBC

[MD91] Michael C. McInerny and Donald J. Degan. Dynamics of a black crappie population in a heterogeneous cooling reservoir. North American Journal of Fisheries Management, 11(4):525– 533, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McInerny:1993:ECR

[MD93] Michael C. McInerny and Donald J. Degan. Electrofishing catch rates as an index of largemouth bass population density in two large reservoirs. North American Journal of Fisheries Management, 13(2):223–228, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:1999:SLT

[MD99] Charles P. Madenjian and Timothy J. DeSorcie. Status of lake trout rehabilitation in the northern refuge of Lake Michigan. North American Journal of Fisheries Management, 19(3):658– 669, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMahon:1996:MBF

[MDI^{+}96] Thomas E. McMahon, Steven R. Dalbey, Susan C. Ireland, James P. Magee, and Patrick A. Byorth. Management briefs: Field evaluation of visible implant tag retention by brook trout, cutthroat trout, rainbow trout, and Arctic grayling. North American Journal of Fisheries Management, 16(4):921–925, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:1993:GTL

[ME93] R. J. Miller and R. W. Elner. A good times licensing strategy for limited-entry fisheries. North American Journal of Fisheries Management, 13(2):387–390, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:1990:PDB

[MFBM90] Steven J. Miller, Donald D. Fox, Lawrence A. Bull, and T. D. McCall. Population dynamics of black crappie in Lake Okeechobee, Florida, following suspension of commercial harvest. North American Journal of Fisheries Management, 10(1):98–105, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mathias:1992:ESW

[MFC^{+}92] J. A. Mathias, W. G. Franzin, J. F. Craig, J. A. Babaluk, and J. F. Flannagan. Evaluation of stocking walleye fry to enhance a commercial fishery in a large, Canadian prairie lake. North American Journal of Fisheries Management, 12(2):299–306, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mourning:1994:CVI

[MFG94] Toby E. Mourning, Kurt D. Fausch, and Charles Gowan. Comparison of visible implant tags and floy anchor tags on hatchery rainbow trout. North American Journal of Fisheries Management, 14(3):636–642, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKeown:1999:ESS

[MFM99] Paul E. McKeown, John L. Forney, and Stephen R. Mooradian. Effects of stocking size and rearing method on muskellunge survival in Chautauqua Lake, New York. North American Journal of Fisheries Management, 19(1):249–257, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Modde:1991:UHB

[MFP91] Timothy Modde, Richard C. Ford, and Milton G. Parsons. Use of a habitat-based stream classification system for categorizing trout biomass. North American Journal of Fisheries Management, 11 (3):305–311, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:1998:EIS

[MFP98] Geoffrey A. McMichael, Anthony L. Fritts, and Todd N. Pearsons. Electrofishing injury to stream salmonids; injury assessment at the sample, reach, and stream scales. North American Journal of Fisheries Management, 18(4):894–904, November 1998. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:1997:ECB

[MG97] Kevin A. Meyer and J. S. Griffith. Effects of cobble–boulder substrate configuration on winter residency of juvenile rainbow trout. North American Journal of Fisheries Management, 17 (1):77–84, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:1999:LGA

[MG99] Paul H. Michaletz and Craig M. Gale. Longitudinal gradients in age-0 gizzard shad density in large Missouri reservoirs. North American Journal of Fisheries Management, 19(3):765–773, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1994:WSA

[MH94] L. E. Miranda and W. D. Hubbard. Winter survival of age-0 largemouth bass relative to size, predators, and shelter.

North American Journal of Fisheries Management, 14(4):790– 796, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mullner:1995:SSS

[MH95] Scott A. Mullner and Wayne A. Hubert. Selection of spawning sites by Kokanees and evaluation of mitigative spawning channels in the Green River, Wyoming. North American Journal of Fisheries Management, 15(1):174–184, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Malchoff:1997:STH

[MH97a] Mark H. Malchoff and Stephen W. Heins. Short-term hooking mortality of weakfish caught on single-barb hooks. North American Journal of Fisheries Management, 17(2):477–481, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marwitz:1997:TRW

[MH97b] Todd D. Marwitz and Wayne A. Hubert. Trends in relative weight of walleye stocks in Wyoming reservoirs. North American Journal of Fisheries Management, 17(1):44–53, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1995:MSS

[MHN^{+}95] Michael D. Murphy, Robert F. Heagey, Victor H. Neugebauer, Mark D. Gordon, and Jennifer L. Hintz. Mortality of spotted seatrout released from gill-net or hook-and-line gear in Florida. North American Journal of Fisheries Management, 15(4):748– 753, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marotz:1990:MGM

[MHR90] Brian L. Marotz, William H. Herke, and Barton D. Rogers. Movement of Gulf menhaden through three marshland routes in Southwestern Louisiana. North American Journal of Fisheries Management, 10(4):408–417, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moyer:1995:FRH

[MHS^{+}95] Edwin J. Moyer, Mikel W. Hulon, James J. Sweatman, Robert S. Butler, and Vincent P. Williams. Fishery responses to habitat restoration in Lake Tohopekaliga, Florida. North American Journal of Fisheries Management, 15(3):591–595, August 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1996:OES

[MHSH96] L. E. Miranda, W. D. Hubbard, S. Sangare, and T. Holman. Optimizing electrofishing sample duration for estimating relative abundance of largemouth bass in reservoirs. North American Journal of Fisheries Management, 16(2):324–331, May 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mullner:1998:SAD

[MHW98] Scott A. Mullner, Wayne A. Hubert, and Thomas A. Wesche. Snorkeling as an alternative to depletion electrofishing for estimating abundance and length-class frequencies of trout in small streams. North American Journal of Fisheries Management, 18 (4):947–953, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Modde:1998:UMS

[MI98] Timothy Modde and David B. Irving. Use of multiple spawning sites and seasonal movement by razorback suckers in the Middle Green River, Utah. North American Journal of Fisheries Management, 18(2):318–326, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:1998:ESF

[Mic98a] Paul H. Michaletz. Effects on sport fish growth of spatial and temporal variation in age-0 gizzard shad availability. North American Journal of Fisheries Management, 18(3):616–624, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:1998:PCG

[Mic98b] Paul H. Michaletz. Population characteristics of gizzard shad in Missouri reservoirs and their relation to reservoir productivity, mean depth, and sport fish growth. North American Journal of Fisheries Management, 18(1):114–123, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:1999:IRP

[Mic99] Paul H. Michaletz. Influence of reservoir productivity and juvenile density on first-year growth of gizzard shad. North American Journal of Fisheries Management, 19(3):842–847, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Milon:1991:MEV

[Mil91] J. Walter Milon. Measuring the economic value of anglers’ kept and released catches. North American Journal of Fisheries Management, 11(2):185–189, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1993:MBS

[Mir93] L. Esteban Miranda. Management briefs: Sample sizes for estimating and comparing proportion-based indices. North American Journal of Fisheries Management, 13(2):383–386, May 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1999:TFL

[Mir99] L. E. Miranda. A typology of fisheries in large reservoirs of the United States. North American Journal of Fisheries Management, 19(2):536–550, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitzner:1991:EEV

[Mit91] Larry Mitzner. Effect of environmental variables upon crappie young, year-class strength, and the sport fishery. North American Journal of Fisheries Management, 11(4):534–542, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitzner:1992:EWF

[Mit92] Larry Mitzner. Evaluation of walleye fingerling and fry stocking in Rathbun Lake, Iowa. North American Journal of Fisheries Management, 12(2):321–328, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Modde:1997:EIR

[MJHG97] Timothy Modde, Randall J. Jeric, Wayne A. Hubert, and Robert D. Gipson. Estimating the impacts of reservoir elevation changes on Kokanee emergence in Flaming Gorge Reservoir, Wyoming–Utah. North American Journal of Fisheries Management, 17(2):470–473, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Markham:1991:WCS

[MJP91] James L. Markham, David L. Johnson, and Raymond W. Petering. White crappie summer movements and habitat use in Delaware Reservoir, Ohio. North American Journal of Fisheries Management, 11(4):504–512, November 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:1991:RAS

[MK91] Geoffrey A. McMichael and Calvin M. Kaya. Relations among stream temperature, angling success for rainbow trout and brown trout, and fisherman satisfaction. North American Journal of Fisheries Management, 11(2):190–199, May 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munger:1997:ELB

[MK97] Charles R. Munger and Joseph E. Kraal. Evaluation of length and bag limits for walleyes in Meredith Reservoir, Texas. North American Journal of Fisheries Management, 17(2):438–445, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miyakoshi:1999:MRE

[MK99] Yasuyuki Miyakoshi and Satoshi Kudo. Mark–recapture estimation of escapement of masu salmon Oncorhynchus masou with a comparison to a fence count. North American Journal of Fisheries Management, 19(4):1108–1111, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsden:1993:GCN

[MKG^{+}93] J. Ellen Marsden, Charles C. Krueger, Peter M. Grewe, Harold L. Kincaid, and Bernie May. Genetic comparison of naturally spawned and artificially propagated Lake Ontario lake trout fry: Evaluation of a stocking strategy for species rehabilitation. North American Journal of Fisheries Management, 13(2):304–317, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsden:1991:ITP

[MKH91] J. Ellen Marsden, Charles C. Krueger, and Harry M. Hawkins. An improved trap for passive capture of demersal eggs during spawning: An efficiency comparison with egg nets. North American Journal of Fisheries Management, 11(3):364–368, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrow:1997:CAG

[MKK97] James V. Morrow, Jr., James P. Kirk, and K. Jack Killgore. Collection, age, growth, and population attributes of triploid grass carp stocked into the santee-Cooper reservoirs, South Carolina. North American Journal of Fisheries Management, 17(1):38–43, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrow:1998:SRP

[MKK^{+}98] James V. Morrow, Jr., James P. Kirk, K. Jack Killgore, Howard Rogillio, and Charles Knight. Status and recovery potential of Gulf sturgeon in the Pearl River System, Louisiana–Mississippi. North American Journal of Fisheries Management, 18(4):798– 808, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrow:1998:AGM

[MKKG98] James V. Morrow, Jr., James P. Kirk, K. Jack Killgore, and Steven G. George. Age, growth, and mortality of shovelnose sturgeon in the Lower Mississippi River. North American Journal of Fisheries Management, 18(3):725–730, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrow:1999:REG

[MKKR99] James V. Morrow, Jr., James P. Kirk, K. Jack Killgore, and Howard E. Rogillio. Recommended enhancements to the Gulf sturgeon recovery and management plan based on Pearl River studies. North American Journal of Fisheries Management, 19 (4):1117–1121, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McWilliams:1992:AWF

[ML92] R. H. McWilliams and Joe G. Larscheid. Assessment of walleye fry and fingerling stocking in the Okoboji Lakes, Iowa. North American Journal of Fisheries Management, 12(2):329–335, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:1994:ETR

[ML94] Gordon Mueller and Charles R. Liston. Evaluation of tire reefs for enhancing aquatic communities in concrete-lined canals. North American Journal of Fisheries Management, 14(3):616–625, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:1995:CSG

[MLP95] Steven W. Martin, John A. Long, and Todd N. Pearsons. Comparison of survival, gonad development, and growth between rainbow trout with and without surgically implanted dummy radio transmitters. North American Journal of Fisheries Management, 15(2):494–498, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinelli-Liedtke:1999:NGB

[MLSH^{+}99] T. L. Martinelli-Liedtke, R. S. Shively, G. S. Holmberg, M. B. Sheer, and R. M. Schrock. Nonlethal gill biopsy does not affect juvenile Chinook salmon implanted with radio transmitters. North American Journal of Fisheries Management, 19(3):856– 859, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meals:1991:VAA

[MM91] Keith O. Meals and L. Esteban Miranda. Variability in abundance of age-0 centrarchids among littoral habitats of flood control reservoirs in Mississippi. North American Journal of Fisheries Management, 11(3):298–304, August 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:1993:CLM

[MM93] R. J. Miller and R. K. Mohn. Critique of the Leslie method for estimating sizes of crab and lobster populations. North American Journal of Fisheries Management, 13(4):676–685, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mauls:1994:EEA

[MM94a] Alec G. Mauls and Matthew G. Mesa. Efficacy of electrofishing to assess plasma cortisol concentration in juvenile Chinook salmon passing hydroelectric dams on the Columbia River. North American Journal of Fisheries Management, 14(2):334–339, May 1994.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meals:1994:SRM

[MM94b] Keith O. Meals and L. Esteban Miranda. Size-related mortality of tournament-caught largemouth bass. North American Journal of Fisheries Management, 14(2):460–463, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsh:1999:DSS

[MMA^{+}99] Douglas M. Marsh, Gene M. Matthews, Stephen Achord, Thomas E. Ruehle, and Benjamin P. Sandford. Diversion of salmonid smolts tagged with passive integrated transponders from an untagged population passing through a juvenile collection system. North American Journal of Fisheries Management, 19(4):1142–1146, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muoneke:1991:MMA

[MMD91] Maurice I. Muoneke, O. Eugene Maughan, and Michael E. Douglas. Multivariate morphometric analysis of striped bass, white bass, and striped bass × white bass hybrids. North American Journal of Fisheries Management, 11(3):330–338, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matlock:1993:MBS

[MMD^{+}93] Gary C. Matlock, Lawrence W. McEachron, James A. Dailey, Philip A. Unger, and Peng Chai. Management briefs: Short-term hooking mortalities of red drums and spotted seatrout caught on single-barb and treble hooks. North American Journal of Fisheries Management, 13(1):186–189, February 1993. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrow:1997:DSF

[MMK97] James V. Morrow, Jr., Gary L. Miller, and K. Jack Killgore. Density, size, and foods of larval northern pike in natural and artificial wetlands. North American Journal of Fisheries Management, 17 (1):210–214, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mundahl:1998:EBP

[MMS^{+}98] Neal D. Mundahl, Christina Melnytschuk, Deena K. Spielman, Jason P. Harkins, Kate Funk, and Andrew M. Bilicki. Effectiveness of bowfin as a predator on bluegill in a vegetated lake. North American Journal of Fisheries Management, 18(2):286–294, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Macdonald:1993:USI

[MN93] D. D. Macdonald and C. P. Newcombe. Utility of the stress index for predicting suspended sediment effects: Response to comment. North American Journal of Fisheries Management, 13(4): 873–876, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moore:1991:SAS

[MNN91] Christopher M. Moore, Richard J. Neves, and John J. Ney. Survival and abundance of stocked striped bass in Smith Mountain Lake, Virginia. North American Journal of Fisheries Management, 11(3):393–399, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1998:UEY

[MOAS98] Michael J. Maceina, Ozcan Ozen, Micheal S. Allen, and Stephen M. Smith. Use of equilibrium yield models to evaluate length limits for crappies in Weiss Lake, Alabama. North American Journal of Fisheries Management, 18(4):854–863, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mohler:1997:MBI

[Moh97] Jerre W. Mohler. Management briefs: Immersion of larval Atlantic salmon in calcein solutions to induce a non-lethally detectable mark. North American Journal of Fisheries Management, 17(3):751–756, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Montao:1994:RWM

[Mon94] Orlando José Ferrer Montao. Recruitment of white mullet in Lake Maracaibo, Venezuela. North American Journal of Fisheries Management, 14(3):516–521, August 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moring:1993:EAE

[Mor93] John R. Moring. Effect of angling effort on catch rate of wild salmonids in streams stocked with catchable-size trout. North American Journal of Fisheries Management, 13(2):234–237, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Modde:1990:IOP

[MP90] Timothy Modde and Brian Platz. Influence of operator position on the precision of measurements taken with hand-held velocity meters in rivers. North American Journal of Fisheries Management, 10(2):247–248, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McAllister:1992:EDM

[MP92] Murdoch K. McAllister and Randall M. Peterman. Experimental design in the management of fisheries: a review. North American Journal of Fisheries Management, 12(1):1–18, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1997:RBV

[MP97] L. E. Miranda and L. L. Pugh. Relationship between vegetation coverage and abundance, size, and diet of juvenile largemouth bass during winter. North American Journal of Fisheries Management, 17(3):601–610, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muoneke:1999:DES

[MP99] Maurice I. Muoneke and Kevin L. Pope. Development and evaluation of a standard weight (w_{s}) equation for blue catfish. North American Journal of Fisheries Management, 19(3):878–879, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:1999:BIA

[MPL99] Geoffrey A. McMichael, Todd N. Pearsons, and Steven A. Leider. Behavioral interactions among hatchery-reared steelhead smoltsand wild Oncorhynchus mykiss in natural streams. North American Journal of Fisheries Management, 19(4):948–956, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCutcheon:1994:PMM

[MPP94] Clinton S. McCutcheon, Earl F. Prentice, and Donn L. Park. Passive monitoring of migrating adult steelhead with PIT tags. North American Journal of Fisheries Management, 14(1):220– 223, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matthews:1991:RSA

[MR91] Kathleen R. Matthews and Laura J. Richards. Rockfish (Scorpaenidae) assemblages of trawlable and untrawlable habitats off Vancouver Island, British Columbia. North American Journal of Fisheries Management, 11(3):312–318, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:1995:EGS

[MR95] Charles P. Madenjian and Philip A. Ryan. Effect of gear selectivity on recommended allowable harvest with application to the Lake Erie yellow perch fishery. North American Journal of Fisheries Management, 15(1):79–83, February 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muth:1996:ETE

[MR96] Robert T. Muth and Jack B. Ruppert. Effects of two electrofishing currents on captive ripe razorback suckers and subsequent egg-hatching success. North American Journal of Fisheries Management, 16(2):473–476, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muth:1997:EEF

[MR97] Robert T. Muth and Jack B. Ruppert. Effects of electrofishing fields on captive embryos and larvae of razorback sucker. North American Journal of Fisheries Management, 17(1):160– 166, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:1998:FAG

[MRK98] Terry L. Margenau, Paul W. Rasmussen, and Jeffrey M. Kampa. Factors affecting growth of northern pike in small Northern Wisconsin lakes. North American Journal of Fisheries Management, 18(3):625–639, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1993:UCD

[MRL93] Michael J. Maceina, Steven J. Rider, and Donny R. Lowery. Use of a catch-depletion method to estimate population density of age-0 largemouth bass in submersed vegetation. North American Journal of Fisheries Management, 13(4):847–851, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKinney:1999:EFR

[MRP99] T. McKinney, R. S. Rogers, and W. R. Persons. Effects of flow reductions on aquatic biota of the Colorado River below Glen Canyon Dam, Arizona. North American Journal of Fisheries Management, 19(4):984–991, November 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morison:1998:ISP

[MRS98] Alexander K. Morison, Simon G. Robertson, and David C. Smith. An integrated system for production fish aging: Image analysis and quality assurance. North American Journal of Fisheries Management, 18(3):587–598, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinson:1992:EPN

[MS92] Kristen S. Martinson and Bo Shelby. Encounter and proximity norms for salmon anglers in California and New Zealand. North American Journal of Fisheries Management, 12(3):559–567, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Myers:1995:IDR

[MS95] Thomas J. Myers and Sherman Swanson. Impact of deferred rotation grazing on stream characteristics in Central Nevada: a case study. North American Journal of Fisheries Management, 15(2):428–439, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marchant:1996:ASC

[MS96] Steve R. Marchant and Marvin K. Shutters. Artificial substrates collect Gulf sturgeon eggs. North American Journal of Fisheries Management, 16(2):445–447, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1998:RBR

[MS98] Michael J. Maceina and Marc R. Stimpert. Relations between reservoir hydrology and crappie recruitment in Alabama. North American Journal of Fisheries Management, 18(1):104– 113, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miko:1995:DSD

[MSA^{+}95] David A. Miko, Harold L. Schramm, Jr., Steven D. Arey, John A. Dennis, and Nancy E. Mathews. Determination of stocking densities for satisfactory put-and-take rainbow trout fisheries. North American Journal of Fisheries Management, 15(4):823– 829, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1996:DFT

[MSAM96] L. E. Miranda, M. S. Schorr, M. S. Allen, and K. O. Meals. Description of a floating trap net for sampling crappies. North American Journal of Fisheries Management, 16(2):457–460, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1999:ETB

[MSG99] Michael J. Maceina, Jeffery W. Slipke, and John M. Grizzle. Effectiveness of three barrier types for confining grass carp in embayments of Lake Seminole, Georgia. North American Journal of Fisheries Management, 19(4):968–976, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moen:1992:PMH

[MSR92] Clark T. Moen, Dennis L. Scarnecchia, and John S. Ramsey. Paddlefish movements and habitat use in Pool 13 of the Upper Mississippi River during abnormally low river stages and discharges. North American Journal of Fisheries Management, 12 (4):744–751, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morkert:1998:EEM

[MSS98] Sidney B. Morkert, William D. Swink, and James G. Seelye. Evidence for early metamorphosis of sea lampreys in the Chippewa River, Michigan. North American Journal of Fisheries Management, 18(4):966–971, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1991:PSE

[MT91] Michael D. Murphy and Ronald G. Taylor. Preliminary study of the effect of reward amount on tag-return rate for red drums in Tampa Bay, Florida. North American Journal of Fisheries Management, 11(3):471–474, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinez:1992:FSS

[MT92] Patrick J. Martinez and Kenneth F. Tiffan. Fabrication of stainless steel spherical anodes for use with boat-mounted boom electroshockers. North American Journal of Fisheries Management, 12(4):840–843, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyers:1992:SMB

[MTK92] Lee S. Meyers, Thomas F. Thuemler, and Greg W. Kornely. Seasonal movements of brown trout in Northeast Wisconsin. North American Journal of Fisheries Management, 12(3):433–441, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1990:MBE

[MTT90] Brian R. Murphy, David R. Terre, and Alan J. Temple. Management briefs: Effect of storage time and temperature on selected walleye electrophoretic phenotypes. North American Journal of Fisheries Management, 10(4):479–481, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:1996:EFC

[Mue96] Gordon Mueller. Establishment of a fish community in the Hayden–Rhodes and Salt-Gila Aqueducts, Arizona. North American Journal of Fisheries Management, 16(4):795–804, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muoneke:1992:LFA

[Muo92a] Maurice I. Muoneke. Loss of floy anchor tags from white bass. North American Journal of Fisheries Management, 12(4):819– 824, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muoneke:1992:SHM

[Muo92b] Maurice I. Muoneke. Seasonal hooking mortality of bluegills caught on natural baits. North American Journal of Fisheries Management, 12(3):645–649, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muoneke:1994:DHE

[Muo94] Maurice I. Muoneke. Dynamics of a heavily exploited Texas white bass population. North American Journal of Fisheries Management, 14(2):415–422, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Megrey:1990:AGR

[MW90] Bernard A. Megrey and Vidar G. Wespestad. Alaskan groundfish resources: 10 years of management under the Magnuson Fishery Conservation and Management Act. North American Journal of Fisheries Management, 10(2):125–143, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinez:1995:SFA

[MW95] Patrick J. Martinez and William J. Wiltzius. Some factors affecting a hatchery-sustained Kokanee population in a fluctuating Colorado reservoir. North American Journal of Fisheries Management, 15(1):220–228, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1992:PDB

[MWC92] Brian R. Murphy, David W. Willis, and W. Michael Childress. Proper distinction between relative weight and relative condition factor. North American Journal of Fisheries Management, 12 (3):665–666, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munger:1994:FCA

[MWF94] Charles R. Munger, Gene R. Wilde, and Billy J. Follis. Flathead catfish age and size at maturation in Texas. North American Journal of Fisheries Management, 14(2):403–408, May 1994. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Modde:1996:CGP

[MWH96] Timothy Modde, Anthony F. Wasowicz, and Dale K. Hepworth. Cormorant and grebe predation on rainbow trout stocked in a Southern Utah reservoir. North American Journal of Fisheries Management, 16(2):388–394, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1995:EHL

[MWL95] Michael J. Maceina, William B. Wrenn, and Donny R. Lowery. Estimating harvestable largemouth bass abundance in a reservoir with an electrofishing catch depletion technique. North American Journal of Fisheries Management, 15(1):103–109, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mero:1994:WSP

[MWP94] Steven W. Mero, David W. Willis, and Greg J. Power. Walleye and sauger predation on paddlefish in Lake Sakakawea, North Dakota. North American Journal of Fisheries Management, 14 (1):226–227, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nemeth:1992:RJC

[NA92a] Richard S. Nemeth and James J. Anderson. Response of juvenile Coho and Chinook salmon to strobe and mercury vapor lights. North American Journal of Fisheries Management, 12(4): 684–692, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nuhfer:1992:HMT

[NA92b] Andrew J. Nuhfer and Gaylord R. Alexander. Hooking mortality of trophy-sized wild brook trout caught on artificial lures. North American Journal of Fisheries Management, 12(3):634– 644, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nuhfer:1994:GSV

[NA94] Andrew J. Nuhfer and Gaylord R. Alexander. Growth, survival, and vulnerability to angling of three wild brook trout strains exposed to different levels of angler exploitation. North American Journal of Fisheries Management, 14(2):423–434, May 1994. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:1995:BMS

[Neg95] Mary T. Negus. Bioenergetics modeling as a salmonine management tool applied to Minnesota waters of Lake Superior. North American Journal of Fisheries Management, 15(1):60–78, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:1999:STN

[Neg99a] Mary T. Negus. Survival traits of naturalized, hatchery, and hybrid strains of anadromous rainbow trout during egg and fry stages. North American Journal of Fisheries Management, 19(4): 930–941, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:1999:TMO

[Neg99b] Mary T. Negus. Thermal marking of otoliths in lake trout sac fry. North American Journal of Fisheries Management, 19(1): 127–140, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nelson:1998:CRM

[Nel98] Kent L. Nelson. Catch-and-release mortality of striped bass in the Roanoke River, North Carolina. North American Journal of Fisheries Management, 18(1):25–30, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:1997:BAG

[New97] Ken Newman. Bayesian averaging of generalized linear models for passive integrated transponder tag recoveries from salmonids in the Snake River. North American Journal of Fisheries Management, 17(2):362–377, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nawa:1993:MSF

[NF93] Richard K. Nawa and Christopher A. Frissell. Measuring scour and fill of gravel streambeds with scour chains and sliding-bead monitors. North American Journal of Fisheries Management, 13 (3):634–639, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Niewinski:1999:CTS

[NF99] Brian C. Niewinski and C. Paola Ferreri. A comparison of three structures for estimating the age of yellow perch. North American Journal of Fisheries Management, 19(3):872–877, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neumann:1995:PSS

[NGW95] Robert M. Neumann, Christopher S. Guy, and David W. Willis. Precision and size structure of juvenile percichthyid samples collected with various gears from Lake Texoma. North American Journal of Fisheries Management, 15(4):956–962, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:1998:ELR

[NH98] Steven P. Newman and Michael H. Hoff. Estimates of loss rates of jaw tags on walleyes. North American Journal of Fisheries Management, 18(1):202–205, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newcombe:1996:CSS

[NJ96] Charles P. Newcombe and Jorgen O. T. Jensen. Channel suspended sediment and fisheries: a synthesis for quantitative assessment of risk and impact. North American Journal of Fisheries Management, 16(4):693–727, November 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neumann:1991:ERW

[NM91a] Robert M. Neumann and Brian R. Murphy. Evaluation of the relative weight (W_{r}) index for assessment of white crappie and black crappie populations. North American Journal of Fisheries Management, 11(4):543–555, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newcombe:1991:ESS

[NM91b] C. P. Newcombe and D. D. Macdonald. Effects of suspended sediments on aquatic ecosystems. North American Journal of Fisheries Management, 11(1):72–82, February 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Novinger:1999:FTB

[NM99] Douglas C. Novinger and Carlos Martinez Del Rio. Failure of total body electrical conductivity to predict lipid content of brook trout. North American Journal of Fisheries Management, 19(4): 942–947, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Navarro:1996:TPF

[NMB96] John E. Navarro, Dennis J. McCauley, and Andrew R. Blystra. Turbine passage at four low-head hydroelectric facilities in Northeast Michigan. North American Journal of Fisheries Management, 16(1):182–191, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nance:1994:FIE

[NMK94] James M. Nance, Eduardo X. Martinez, and Edward F. Klima. Feasibility of improving the economic return from the Gulf of Mexico brown shrimp fishery. North American Journal of Fisheries Management, 14(3):522–536, August 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nakatani:1991:EPB

[NN91] Roy E. Nakatani and Ahmad E. Nevissi. Effect of Prudhoe Bay crude oil on the homing of Coho salmon in marine waters. North American Journal of Fisheries Management, 11(2):160–166, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neal:1999:FCR

[NNR99] J. Wesley Neal, Richard L. Noble, and James A. Rice. Fish community response to hybrid striped bass introduction in small warmwater impoundments. North American Journal of Fisheries Management, 19(4):1044–1053, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Northcote:1997:PSL

[Nor97] T G. Northcote. Potamodromy in Salmonidae — living and moving in the fast lane. North American Journal of Fisheries Management, 17(4):1029–1045, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nordwall:1999:MBT

[Nor99] Fredrik Nordwall. Movements of brown trout in a small stream: Effects of electrofishing and consequences for population estimates. North American Journal of Fisheries Management, 19 (2):462–469, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Novinger:1990:SLL

[Nov90] Gary D. Novinger. Slot length limits for largemouth bass in small private impoundments. North American Journal of Fisheries Management, 10(3):330–337, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nestler:1992:RBH

[NPP^{+}92] John M. Nestler, Gene R. Ploskey, James Pickens, John Menezes, and Carl Schilt. Responses of blueback herring to high-frequency sound and implications for reducing entrainment at hydropower dams. North American Journal of Fisheries Management, 12(4): 667–683, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:1997:CSI

[NRA97] Steven P. Newman, Paul W. Rasmussen, and Lloyd M. Andrews. Comparison of a stratified, instantaneous count creel survey with a complete mandatory creel census on Escanaba Lake, Wisconsin. North American Journal of Fisheries Management, 17(2):321– 330, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neumann:1995:SVG

[NW95] Robert M. Neumann and David W. Willis. Seasonal variation in gill-net sample indexes for northern pike collected from a glacial prairie lake. North American Journal of Fisheries Management, 15(4):838–844, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Noakes:1990:CPF

[NWHM90] D. J. Noakes, D. W. Welch, M. Henderson, and E. Mansfield. A comparison of preseason forecasting methods for returns of two British Columbia sockeye salmon stocks. North American Journal of Fisheries Management, 10(1):46–57, February 1990. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Onorato:1998:CSM

[OAM98] Dave P. Onorato, Robert A. Angus, and Ken R. Marion. Comparison of a small-mesh seine and a backpack electroshocker for evaluating fish populations in a North-Central Alabama stream. North American Journal of Fisheries Management, 18(2):361– 373, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oven:1993:BRC

[OB93] Jackie Heinricher Oven and H. Lee Blankenship. Benign recovery of coded wire tags from rainbow trout. North American Journal of Fisheries Management, 13(4):852–855, November 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Osborne:1995:RUT

[OB95] Randall Osborne and Phillip W. Bettoli. A reusable ultrasonic tag and float assembly for use with large pelagic fish. North American Journal of Fisheries Management, 15(2):512–514, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

OConnell:1994:CRO

[OC94] Victoria M. O’Connell and David W. Carlile. Comparison of a remotely operated vehicle and a submersible for estimating abundance of demersal shelf rockfishes in the Eastern Gulf of Alaska. North American Journal of Fisheries Management, 14(1):196– 201, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

OBin:1996:REO

[OHW96] Francis X. O’B’i’n, Peter B. Heffernan, and Randal L. Walker. Recruitment of the Eastern oyster in coastal Georgia: Patterns and recommendations. North American Journal of Fisheries Management, 16(2):413–426, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Orciari:1990:CRM

[OL90] Robert D. Orciari and Gerald H. Leonard. Catch-and-release management of a trout stream contaminated with PCBs. North American Journal of Fisheries Management, 10(3):315–329, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Orciari:1996:LCS

[OL96] Robert D. Orciari and Gerald H. Leonard. Length characteristics of smolts and timing of downstream migration among three strains of Atlantic salmon in a Southern New England stream.

North American Journal of Fisheries Management, 16(4):851– 860, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Orciari:1994:SGS

[OLMS94] Robert D. Orciari, Gerald H. Leonard, Donald J. Mysling, and Eric C. Schluntz. Survival, growth, and smolt production of Atlantic salmon stocked as fry in a Southern New England stream. North American Journal of Fisheries Management, 14(3):588– 606, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olson-Rutz:1992:AIS

[ORM92] K. M. Olson-Rutz and C. B. Marlow. Analysis and interpretation of stream channel cross-sectional data. North American Journal of Fisheries Management, 12(1):55–61, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

OShea:1995:EMI

[O’S95] Daniel T. O’Shea. Estimating minimum instream flow requirements for Minnesota streams from hydrologic data and watershed characteristics. North American Journal of Fisheries Management, 15(3):569–578, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ogle:1996:CFS

[OS96] Derek H. Ogle and George R. Spangler. Check formation on the scales of hatchery-reared lake trout prior to and soon after release into Lake Superior. North American Journal of Fisheries Management, 16(4):896–904, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ogle:1996:PRN

[OSS^{+}96] Derek H. Ogle, James H. Selgeby, Jacqueline E. Saving, Raymond M. Newman, and Mary G. Henry. Predation on ruffe by native fishes of the St. Louis River Estuary, Lake Superior, 1989– 1991. North American Journal of Fisheries Management, 16(1): 115–123, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Overholtz:1999:PUB

[Ove99] W. J. Overholtz. Precision and uses of biological reference points calculated from stock recruitment data. North American Journal of Fisheries Management, 19(3):643–657, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Orsi:1993:ISH

[OWJ93] Joseph A. Orsi, Alex C. Wertheimer, and Herbert W. Jaenicke. Influence of selected hook and lure types on catch, size, and mortality of commercially troll-caught Chinook salmon. North American Journal of Fisheries Management, 13(4):709–722, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ostrand:1999:IMT

[OWSM99] Kenneth G. Ostrand, Gene R. Wilde, Dan W. Strickland, and Maurice I. Muoneke. Initial mortality in Texas black bass fishing tournaments. North American Journal of Fisheries Management, 19(4):1124–1128, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paller:1995:IVS

[Pal95a] Michael H. Paller. Interreplicate variance and statistical power of electrofishing data from low-gradient streams in the Southeastern United States. North American Journal of Fisheries Management, 15(3):542–550, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paller:1995:RAN

[Pal95b] Michael H. Paller. Relationships among number of fish species sampled, reach length surveyed, and sampling effort in South Carolina coastal plain streams. North American Journal of Fisheries Management, 15(1):110–120, February 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paller:1997:RRF

[Pal97] Michael H. Paller. Recovery of a reservoir fish community from drawdown related impacts. North American Journal of Fisheries Management, 17(3):726–733, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parker:1992:MPE

[Par92] Jerry A. Parker. Migratory patterns and exploitation of American shad in the nearshore ocean waters of Southeastern North Carolina. North American Journal of Fisheries Management, 12 (4):752–759, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1995:KII

[Par95] Vaughn L. Paragamian. Kokanee investigations: Introduction to Kokanee investigations. North American Journal of Fisheries Management, 15(1):173, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsley:1994:WSS

[PB94a] Michael J. Parsley and Lance G. Beckman. White sturgeon spawning and rearing habitat in the Lower Columbia River. North American Journal of Fisheries Management, 14(4):812– 827, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:1994:MBC

[PB94b] Rodney B. Pierce and Andrew G. Bindman. Management briefs: Comparison of absolute fishing effort and hourly instantaneous angler counts in a small lake. North American Journal of Fisheries Management, 14(2):447–448, May 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1995:FAS

[PB95] Vaughn L. Paragamian and Edward C. Bowles. Factors affecting survival of Kokanees stocked in Lake Pend Oreille, Idaho. North American Journal of Fisheries Management, 15(1):208– 219, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pitt:1998:EBM

[PBC98] William C. Pitt, David A. Beauchamp, and Michael R. Conover. Evaluation of bioenergetics models for predicting great blue Heron consumption of rainbow trout at hatcheries. North American Journal of Fisheries Management, 18(1):52–65, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pegg:1997:MSL

[PBL97] Mark A. Pegg, Phillip W. Bettoli, and James B. Layzer. Movement of saugers in the Lower Tennessee River determined by radio telemetry, and implications for management. North American Journal of Fisheries Management, 17(3):763–768, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:1995:FRM

[PBTS95] Harriet Perry, Walter Brehm, Christine Trigg, and Kenneth Stuck. Fishery-related morphometric characteristics of Menippe adina from the North-Central Gulf of Mexico. North American Journal of Fisheries Management, 15(3):639–646, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ploskey:1999:CHN

[PC99] Gene R. Ploskey and Thomas J. Carlson. Comparison of hydroacoustic and net estimates of fish guidance efficiency of an extended submersible bar screen at John Day Dam. North American Journal of Fisheries Management, 19(4):1066–1079, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prentice:1991:UKA

[PD91] John A. Prentice and Wilfred J. Dean, Jr. Use of known-age red drums to validate scale and otolith ages and to estimate growth in fresh water. North American Journal of Fisheries Management, 11(3):424–428, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peake:1998:SBC

[Pea98] S. Peake. Sodium bicarbonate and clove oil as potential anesthetics for nonsalmonid fishes. North American Journal of Fisheries Management, 18(4):919–924, November 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:1993:EET

[Pet93] Allen M. Peterson. Effects of electric transmission rights-of-way on trout in forested headwater streams in New York. North American Journal of Fisheries Management, 13(3):581–585, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:1999:EPL

[PF99a] Craig P. Paukert and William L. Fisher. Evaluation of paddlefish length distributions and catch rates in three mesh sizes of gill nets. North American Journal of Fisheries Management, 19(2): 599–603, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pearsons:1999:MSC

[PF99b] Todd N. Pearsons and Anthony L. Fritts. Maximum size of Chinook salmon consumed by juvenile Coho salmon. North American Journal of Fisheries Management, 19(1):165–170, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pitman:1993:IPS

[PG93] Veronica M. Pitman and Steve Gutreuter. Initial poststocking survival of hatchery-reared fishes. North American Journal of Fisheries Management, 13(1):151–159, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pinhorn:1990:CVI

[PH90] A. T. Pinhorn and R. G. Halliday. Canadian versus international regulation of Northwest Atlantic fisheries: Management practices, fishery yields, and resource trends, 1960–1986. North American Journal of Fisheries Management, 10(2):154–174, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pauwels:1994:SHE

[PH94a] Stanislas J. Pauwels and Terry A. Haines. Survival, hatching, and emergence success of Atlantic salmon eggs planted in three Maine streams. North American Journal of Fisheries Management, 14 (1):125–130, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Persons:1994:HML

[PH94b] Steven E. Persons and Steven A. Hirsch. Hooking mortality of lake trout angled through ice by jigging and set-lining. North American Journal of Fisheries Management, 14(3):664–668, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patton:1996:WTA

[PH96] Tim M. Patton and Wayne A. Hubert. Water temperature affects smallmouth bass and channel catfish in a tailwater stream on the Great Plains. North American Journal of Fisheries Management, 16(1):124–131, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pollock:1997:CRE

[PHJ^{+}97] Kenneth H. Pollock, John M. Hoenig, Cynthia M. Jones, Douglas S. Robson, and Colin J. Greene. Catch rate estimation for roving and access point surveys. North American Journal of Fisheries Management, 17(1):11–19, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Piccolo:1993:SWE

[PHW93] Jack J. Piccolo, Wayne A. Hubert, and Roy A. Whaley. Standard weight equation for lake trout. North American Journal of Fisheries Management, 13(2):401–404, May 1993. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pitman:1995:CWT

[PI95] Veronica M. Pitman and Jacob Isaac, Jr. Coded wire tag retention by paddlefish at three implant sites. North American Journal of Fisheries Management, 15(4):878–880, November 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Petering:1995:SMD

[PIM95] Raymond W. Petering, Gary L. Isbell, and Randy L. Miller. A survey method for determining angler preference for catches of various fish length and number combinations. North American Journal of Fisheries Management, 15(4):732–735, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pinkerton:1994:EMB

[Pin94] Evelyn Pinkerton. Economic and management benefits from the coordination of capture and culture fisheries: The case of Prince William Sound pink salmon. North American Journal of Fisheries Management, 14(2):262–277, May 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pitlo:1997:RUM

[Pit97] John Pitlo, Jr. Response of Upper Mississippi River channel catfish populations to changes in commercial harvest regulations. North American Journal of Fisheries Management, 17(4):

848–859, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:1995:BSP

[PJM95] William B. Perry, William A. Janowsky, and F. Joseph Margraf. A bioenergetics simulation of the potential effects of angler harvest on growth of largemouth bass in a catch-and-release fishery. North American Journal of Fisheries Management, 15(3):705– 712, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peck:1999:CHR

[PJMS99] James W. Peck, Thomas S. Jones, Wayne R. MacCallum, and Stephen T. Schram. Contribution of hatchery-reared fish to Chinook salmon populations and sport fisheries in Lake Superior. North American Journal of Fisheries Management, 19(1):155– 164, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phillips:1997:SHA

[PJN97] Jeff M. Phillips, James R. Jackson, and Richard L. Noble. Spatial heterogeneity in abundance of age-0 largemouth bass among reservoir embayments. North American Journal of Fisheries Management, 17(4):894–901, November 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1992:CWF

[PK92a] Vaughn L. Paragamian and Roger Kingery. A comparison of walleye fry and fingerling stockings in three rivers in Iowa. North American Journal of Fisheries Management, 12(2):313–320, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:1992:EHT

[PK92b] Mark S. Peterson and James P. Key. Evaluation of hand-tagging juvenile walleyes with binary-coded wire microtags. North American Journal of Fisheries Management, 12(4):814–818, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perkins:1994:DUM

[PK94] David L. Perkins and Charles C. Krueger. Design and use of mesh bags to estimate deposition and survival of fish eggs in cobble substrate. North American Journal of Fisheries Management, 14(4):866–869, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pegg:1996:AEA

[PLB96] Mark A. Pegg, James B. Layzer, and Phillip W. Bettoli. Angler exploitation of anchor-tagged saugers in the Lower Tennessee River. North American Journal of Fisheries Management, 16(1): 218–222, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prince:1995:SBA

[PLL^{+}95] Angela M. J. Prince, Susan E. Low, Timothy J. Lissimore, Ron E. Diewert, and Scott G. Hinch. Sodium bicarbonate and acetic acid: An effective anesthetic for field use. North American Journal of Fisheries Management, 15(1):170–172, February 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powell:1990:MBS

[PM90] M. J. Powell and M. Miller. Management briefs: Shoal spawning by Chinook salmon in Lake Huron. North American Journal of Fisheries Management, 10(2):242–244, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:1997:CTH

[PM97] Glenn R. Parsons and Keith Meals. Comparison of triploid hybrid crappie and diploid white crappie in experimental ponds. North American Journal of Fisheries Management, 17(3):803– 806, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peaks:1997:NPR

[PMBM97] Stephan Peaks, R. Scott McKinley, Toni A. Beddow, and Gerd Marmulla. New procedure for radio transmitter attachment: Oviduct insertion. North American Journal of Fisheries Management, 17(3):757–762, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porto:1999:LHB

[PMN99] L. M. Porto, R. L. McLaughlin, and D. L. G. Noakes. Low-head barrier dams restrict the movements of fishes in two Lake Ontario streams. North American Journal of Fisheries Management, 19 (4):1028–1036, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patton:1998:OCS

[PMW98] William S. Patton, Katherine W. Myers, and Robert V. Walker. Origins of chum salmon caught incidentally in the Eastern Bering Sea walleye pollock trawl fishery as estimated from scale pattern analysis. North American Journal of Fisheries Management, 18 (3):704–712, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Power:1991:ELS

[PN91] E. A. Power and T. G. Northcote. Effects of log storage on the food supply and diet of juvenile sockeye salmon. North American Journal of Fisheries Management, 11(3):413–423, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Policansky:1998:CTS

[Pol98] David Policansky. Comment: Transportation of Snake River Chinook salmon. North American Journal of Fisheries Management, 18(2):475–476, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powell:1994:LHT

[Pow94] Allyn B. Powell. Life history traits of two allopatric clupeids, Atlantic menhaden and Gulf menhaden, and the effects of harvesting on these traits. North American Journal of Fisheries Management, 14(1):53–64, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powers:1996:BRR

[Pow96] Joseph E. Powers. Benchmark requirements for recovering fish stocks. North American Journal of Fisheries Management, 16 (3):495–504, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Power:1996:CMS

[PP96] M. Power and G. Power. Comparing minimum-size and slot limits for brook trout management. North American Journal of Fisheries Management, 16(1):49–62, February 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:1997:DWF

[PP97a] Bradford G. Parsons and Donald L. Pereira. Dispersal of walleye fingerlings after stocking. North American Journal of Fisheries Management, 17(4):988–994, November 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porath:1997:UWR

[PP97b] Mark T. Porath and Edward J. Peters. Use of walleye relative weights (W_{r}) to assess prey availability. North American Journal of Fisheries Management, 17(3):628–637, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:1994:CSC

[PPQ94] N. Phil Peterson, Earl F. Prentice, and Thomas P. Quinn. Comparison of sequential coded wire and passive integrated transponder tags for assessing overwinter growth and survival of juvenile Coho salmon. North American Journal of Fisheries Management, 14(4):870–873, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powers:1993:ESA

[PR93] Joseph E. Powers and Victor R. Restrepo. Evaluation of stock assessment research for Gulf of Mexico king mackerel: Benefits and costs to management. North American Journal of Fisheries Management, 13(1):15–26, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:1995:OSE

[PR95] James T. Peterson and Charles F. Rabeni. Optimizing sampling effort for sampling warmwater stream fish communities. North American Journal of Fisheries Management, 15(3):528–541, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:1996:NTR

[PR96] James T. Peterson and Charles F. Rabeni. Natural thermal refugia for temperate warmwater stream fishes. North American Journal of Fisheries Management, 16(4):738–746, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pugh:1998:CEH

[PS98] Lawrence L. Pugh and Harold L. Schramm, Jr. Comparison of electrofishing and hoopnetting in lotic habitats of the Lower Mississippi River. North American Journal of Fisheries Management, 18(3):649–656, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pauley:1993:MAC

[PT93a] Gilbert B. Pauley and G. L. Thomas. Mortality of anadromous coastal cutthroat trout caught with artificial lures and natural bait. North American Journal of Fisheries Management, 13(2): 337–345, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:1993:MBT

[PT93b] Rodney B. Pierce and Cynthia M. Tomcko. Management briefs: Tag loss and handling mortality for northern pike marked with plastic anchor tags. North American Journal of Fisheries Management, 13(3):613–615, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:1998:ANS

[PT98] Rodney B. Pierce and Cynthia M. Tomcko. Angler noncompliance with slot length limits for northern pike in five small Minnesota lakes. North American Journal of Fisheries Management, 18(3):720–724, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:1994:IDE

[PTK94] Rodney B. Pierce, Cynthia M. Tomcko, and Todd D. Kolander. Indirect and direct estimates of gill-net size selectivity for northern pike. North American Journal of Fisheries Management, 14 (1):170–177, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Philipp:1997:ICR

[PTK^{+}97] David P. Philipp, C. Anna Toline, Mark F. Kubacki, David B. F. Philipp, and Frank J. S. Phelan. The impact of catch-and-release angling on the reproductive success of smallmouth bass and largemouth bass. North American Journal of Fisheries Management, 17(2):557–567, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:1995:ENP

[PTS95] Rodney B. Pierce, Cynthia M. Tomcko, and Dennis H. Schupp. Exploitation of northern pike in seven small North-Central Minnesota lakes. North American Journal of Fisheries Management, 15(3):601–609, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pellett:1998:SMH

[PVA98] Thomas D. Pellett, Gene J. Van Dyck, and Jean V. Adams. Seasonal migration and homing of channel catfish in the Lower Wisconsin River, Wisconsin. North American Journal of Fisheries Management, 18(1):85–95, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pothoven:1999:EVR

[PVP99] Steven A. Pothoven, Bruce Vondracek, and Donald L. Pereira. Effects of vegetation removal on bluegill and largemouth bass in two Minnesota lakes. North American Journal of Fisheries Management, 19(3):748–757, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pajos:1994:EPL

[PW94] Tiina A. Pajos and Jerry G. Weise. Estimating populations of larval sea lamprey with electrofishing sampling methods. North American Journal of Fisheries Management, 14(3):580–587, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pope:1998:LBC

[PW98] Kevin L. Pope and David W. Willis. Larval black crappie distribution: Implications for sampling impoundments and natural lakes. North American Journal of Fisheries Management, 18(2): 470–474, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pikitch:1998:PHB

[PWB^{+}98] Ellen K. Pikitch, John R. Wallace, Elizabeth A. Babcock, Daniel L. Erickson, Mark Saelens, and Geir Oddsson. Pacific halibut bycatch in the Washington, Oregon, and California groundfish and shrimp trawl fisheries. North American Journal of Fisheries Management, 18(3):569–586, August 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peven:1994:ALS

[PWW94] Charles M. Peven, Richard R. Whitney, and Kenneth R. Williams. Age and length of steelhead smolts from the mid-Columbia River Basin, Washington. North American Journal of Fisheries Management, 14(1):77–86, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:1998:SWE

[QGB98] Michael C. Quist, Christopher S. Guy, and Patrick J. Braaten. Standard weight (W_{s}) equation and length categories for shovelnose sturgeon. North American Journal of Fisheries Management, 18(4):992–997, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:1994:EML

[QKH^{+}94] Norman W. S. Quinn, Robert M. Korver, Frank J. Hicks, Brian P. Monroe, and Richard R. Hawkins. An empirical model of lentic brook trout. North American Journal of Fisheries Management, 14(4):692–709, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:1992:APW

[Qui92] Stephen P. Quinn. Angler perspectives on walleye management. North American Journal of Fisheries Management, 12(2):367– 378, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:1993:DMF

[Qui93] Stephen P. Quinn. Description of a multiuse fishery for flathead catfish. North American Journal of Fisheries Management, 13 (3):594–599, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rieman:1990:DNS

[RB90a] Bruce E. Rieman and Raymond C. Beamesderfer. Dynamics of a northern squawfish population and the potential to reduce predation on juvenile salmonids in a Columbia River reservoir. North American Journal of Fisheries Management, 10(2):228–241, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rieman:1990:WSL

[RB90b] Bruce E. Rieman and Raymond C. Beamesderfer. White sturgeon in the Lower Columbia River: Is the stock overexploited? North American Journal of Fisheries Management, 10(4):388– 396, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:1995:EFD

[RB95] Kevin B. Rogers and Eric P. Bergersen. Effects of a fall drawdown on movement of adult northern pike and largemouth bass. North American Journal of Fisheries Management, 15(3):596– 600, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:1999:USS

[RB99] Kevin B. Rogers and Eric P. Bergersen. Utility of synthetic structures for concentrating adult northern pike and largemouth bass. North American Journal of Fisheries Management, 19(4): 1054–1065, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:1996:MBS

[RBB96] Kevin B. Rogers, Lee C. Bergsted, and Eric E. Bergersen. Management briefs: Standard weight equation for mountain whitefish. North American Journal of Fisheries Management, 16(1): 207–209, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:1997:HUF

[RBJ97] Robert M. Ross, Randy M. Bennett, and James H. Johnson. Habitat use and feeding ecology of riverine juvenile American shad. North American Journal of Fisheries Management, 17(4): 964–974, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rood:1994:MFC

[RC94] Kenneth Rood and Michael Church. Modified freeze-core technique for sampling the permanently wetted streambed. North American Journal of Fisheries Management, 14(4):852–861, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rester:1999:CEB

[RC99] Jeffrey K. Rester and Richard E. Condrey. Characterization and evaluation of bycatch reduction devices in the Gulf menhaden fishery. North American Journal of Fisheries Management, 19 (1):42–50, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:1992:DCH

[RCRR92] Carl Richards, Phillip J. Cernera, Michael P. Ramey, and Dudley W. Reiser. Development of off-channel habitats for use by juvenile Chinook salmon. North American Journal of Fisheries Management, 12(4):721–727, November 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruiz-Campos:1991:STM

[RCVR91] Gorgonio Ruiz-Campos and Manuel M. Villalobos-Ramírez. A simple technique for making a streamer-type fish tag. North American Journal of Fisheries Management, 11(3):475–476, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reed:1991:PDB

[RD91] Jeffrey R. Reed and William D. Davies. Population dynamics of black crappies and white crappies in Weiss Reservoir, Alabama: Implications for the implementation of harvest restrictions. North American Journal of Fisheries Management, 11(4): 598–603, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:1996:RIA

[RDM^{+}96] Quentin E. Ross, Dennis J. Dunning, John K. Menezes, Mark J. Kenna, Jr., and Gary Tiller. Reducing impingement of alewives with high-frequency sound at a power plant intake on Lake Ontario. North American Journal of Fisheries Management, 16 (3):548–559, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:1993:RAH

[RDT^{+}93] Quentin E. Ross, Dennis J. Dunning, Richard Thorne, John K. Menezes, Gary W. Tiller, and John K. Watson. Response of alewives to high-frequency sound at a power plant intake on Lake Ontario. North American Journal of Fisheries Management, 13 (2):291–303, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riley:1992:UTP

[RF92] Stephen C. Riley and Kurt D. Fausch. Underestimation of trout population size by maximum-likelihood removal estimates in small streams. North American Journal of Fisheries Management, 12(4):768–776, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:1995:FIB

[RFS95] Laura J. Richards, Jeff Fargo, and Jon T. Schnute. Factors influencing bycatch mortality of trawl-caught Pacific halibut. North American Journal of Fisheries Management, 15(2):266–276, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:1997:STM

[RH97] Michael R. Ross and Steven R. Hokenson. Short-term mortality of discarded finfish bycatch in the Gulf of Maine fishery for northern shrimp Pandalus borealis. North American Journal of Fisheries Management, 17(4):902–909, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:1998:PFH

[RHSB98] Anthony T. Robinson, Philip P. Hines, Jeff A. Sorensen, and Scott D. Bryan. Parasites and fish health in a desert stream, and management implications for two endangered fishes. North American Journal of Fisheries Management, 18(3):599–608, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:1994:TLC

[Ric94] Laura J. Richards. Trip limits, catch, and effort in the British Columbia rockfish trawl fishery. North American Journal of Fisheries Management, 14(4):742–750, November 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rinne:1996:MBS

[Rin96] John N. Rinne. Management briefs: Short-term effects of wildfire on fishes and aquatic macroinvertebrates in the Southwestern United States. North American Journal of Fisheries Management, 16(3):653–658, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Raibley:1997:WHU

[RIO^{+}97] Paul T. Raibley, Kevin S. Irons, Timothy M. O’Hara, K. Douglas Blodgett, and Richard E. Sparks. Winter habitats used by largemouth bass in the Illinois River, a large river-floodplain ecosystem. North American Journal of Fisheries Management, 17(2): 401–412, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ridley:1997:ESF

[RIT97] Christine A. Ridley, J. Jeffery Isely, and Joseph R. Tomasso, Jr. Effects of supplemental feeding on physical characteristics and physiological responses to angling of striped bass. North American Journal of Fisheries Management, 17(3):791–794, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reynolds:1995:CRE

[RK95] J. B. Reynolds and A. L. Kolz. Comment: Reducing electrofishing-induced injury of rainbow trout. North American Journal of Fisheries Management, 15(4):963–965, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roper:1998:DPN

[RKHW98] Brett B. Roper, Deborah Konnoff, Dave Heller, and Ken Wieman. Durability of Pacific Northwest instream structures following floods. North American Journal of Fisheries Management, 18 (3):686–693, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rach:1994:DAI

[RLM94] Jeffrey J. Rach, James A. Luoma, and Leif L. Marking. Development of an antimycin-impregnated bait for controlling common carp. North American Journal of Fisheries Management, 14(2): 442–446, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rieman:1997:DSL

[RLT97] Bruce E. Rieman, Danny C. Lee, and Russell F. Thurow. Distribution, status, and likely future trends of bull trout within the Columbia River and Klamath River Basins. North American Journal of Fisheries Management, 17(4):1111–1125, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rulifson:1990:RJS

[RM90] Roger A. Rulifson and Charles S. Manooch III. Recruitment of juvenile striped bass in the Roanoke River, North Carolina, as related to reservoir discharge. North American Journal of Fisheries Management, 10(4):397–407, November 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rutecki:1992:MJS

[RM92] Thomas L. Rutecki and Theodore R. Meyers. Mortality of juvenile sablefish captured by hand-jigging and traps. North American Journal of Fisheries Management, 12(4):836–837, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rieman:1995:KPD

[RM95] Bruce E. Rieman and Meld A. Maiolie. Kokanee population density and resulting fisheries. North American Journal of Fisheries Management, 15(1):229–237, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rieman:1996:STV

[RM96] B. E. Rieman and J. D. McIntyre. Spatial and temporal variability in bull trout redd counts. North American Journal of Fisheries Management, 16(1):132–141, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruppert:1997:EEF

[RM97] Jack B. Ruppert and Robert T. Muth. Effects of electrofishing fields on captive juveniles of two endangered cyprinids. North American Journal of Fisheries Management, 17(2):314–320, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rold:1996:CCT

[RMV96] Robert E. Rold, Thomas S. McComish, and Donald E. Van Meter. A comparison of Cedar trees and fabricated polypropylene modules as fish attractors in a strip mine impoundment. North American Journal of Fisheries Management, 16(1):223– 227, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roell:1998:IEF

[RO98] Michael J. Roell and Donald J. Orth. Indirect effects of fishery exploitation and pest control in a riverine food Web. North American Journal of Fisheries Management, 18(2):337–346, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roach:1999:IEM

[Roa99] Stafford M. Roach. Influence of electrofishing on the mortality of Arctic grayling eggs. North American Journal of Fisheries Management, 19(4):923–929, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reiser:1998:SDW

[ROB98] Dudley W. Reiser, Alan Olson, and Keith Binkley. Sediment deposition within fry emergence traps: a confounding factor in estimating survival to emergence. North American Journal of Fisheries Management, 18(3):713–719, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rodgers:1995:CIO

[Rod95] David W. Rodgers. Comparison of impinged and open-water alewives at Pickering Nuclear Generating Station, Lake Ontario. North American Journal of Fisheries Management, 15(1):152– 155, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rottiers:1994:SCA

[Rot94] Donald V. Rottiers. Survival of caged Atlantic salmon in the Merrimack River. North American Journal of Fisheries Management, 14(2):355–361, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reed:1999:AOB

[RP99] Jeffrey R. Reed and Bradford G. Parsons. Angler opinions of bluegill management and related hypothetical effects on bluegill fisheries in four Minnesota lakes. North American Journal of Fisheries Management, 19(2):515–519, May 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:1995:GVS

[RQ95] Philip Roni and Thomas P. Quinn. Geographic variation in size and age of North American Chinook salmon. North American Journal of Fisheries Management, 15(2):325–345, May 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rhodes:1999:CPG

[RQ99] Justin S. Rhodes and Thomas P. Quinn. Comparative performance of genetically similar hatchery and naturally reared juvenile Coho salmon in streams. North American Journal of Fisheries Management, 19(3):670–677, August 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:1995:IHR

[RQTW95] Callum M. Roberts, Norman Quinn, John W. Tucker, Jr., and Peter N. Woodward. Introduction of hatchery-reared Nassau grouper to a coral reef environment. North American Journal of Fisheries Management, 15(1):159–164, February 1995. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruggerone:1992:PSS

[RR92] Gregory T. Ruggerone and Donald E. Rogers. Predation on sockeye salmon fry by juvenile Coho salmon in the Chignik Lakes, Alaska: Implications for salmon management. North American Journal of Fisheries Management, 12(1):87–102, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rottiers:1993:VMA

[RR93] Donald V. Rottiers and Lori A. Redell. Volitional migration of Atlantic salmon from seasonal holding ponds. North American Journal of Fisheries Management, 13(2):238–252, May 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rice:1996:FRI

[RR96] Jake C. Rice and Laura J. Richards. A framework for reducing implementation uncertainty in fisheries management. North American Journal of Fisheries Management, 16(3):488–494, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:1999:CHE

[RR99] R. Anne Richards and Paul J. Rago. A case history of effective fishery management: Chesapeake Bay striped bass. North American Journal of Fisheries Management, 19(2):356–375, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roper:1995:OVC

[RS95] Brett B. Roper and Dennis L. Scarnecchia. Observer variability in classifying habitat types in stream surveys. North American Journal of Fisheries Management, 15(1):49–53, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ridgway:1996:EDS

[RS96a] Mark S. Ridgway and Brian J. Shuter. Effects of displacement on the seasonal movements and home range characteristics of smallmouth bass in Lake Opeongo. North American Journal of Fisheries Management, 16(2):371–377, May 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roper:1996:CTE

[RS96b] Brett Roper and Dennis L. Scarnecchia. A comparison of trap efficiencies for wild and hatchery age-0 Chinook salmon. North American Journal of Fisheries Management, 16(1):214– 217, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ridgway:1997:PEA

[RS97] Mark S. Ridgway and Brian J. Shuter. Predicting the effects of angling for nesting male smallmouth bass on production of age-0 fish with an individual-based model. North American Journal of Fisheries Management, 17(2):568–580, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rodgers:1992:CTT

[RSJB92] Jeffrey D. Rodgers, Mario F. Solazzi, Steven L. Johnson, and Mary A. Buckman. Comparison of three techniques to estimate juvenile Coho salmon populations in small streams. North American Journal of Fisheries Management, 12(1):79–86, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roach:1999:CAC

[RTB99] Brian Roach, Joan Trial, and Kevin Boyle. Comparing 1994 angler catch and harvest rates from on-site and mail surveys on selected Maine lakes. North American Journal of Fisheries Management, 19(1):203–208, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:1992:EAW

[RW92] Bruce C. Roberts and Robert G. White. Effects of angler wading on survival of trout eggs and pre-emergent fry. North American Journal of Fisheries Management, 12(3):450–459, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Storey:1993:EIM

[SA93] David A. Storey and P. Geoffrey Allen. Economic impact of marine recreational fishing in Massachusetts. North American Journal of Fisheries Management, 13(4):698–708, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scott:1998:RUT

[SA98] Mark C. Scott and Paul L. Angermeier. Resource use by two sympatric black basses in impounded and riverine sections of the New River, Virginia. North American Journal of Fisheries Management, 18(2):221–235, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakagawa:1991:URT

[Sak91] Gary T. Sakagawa. Are U.S. regulations on tuna–dolphin fishing driving U.S. seiners to foreign-flag registry? North American Journal of Fisheries Management, 11(3):241–252, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saltveit:1993:AJA

[Sal93] S. J. Saltveit. Abundance of juvenile Atlantic salmon and brown trout in relation to stocking and natural reproduction in the River Laerdalselva, Western Norway. North American Journal of Fisheries Management, 13(2):277–283, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stender:1994:CCT

[SB94] B. W. Stender and C. A. Barans. Comparison of the catch from tongue and two-seam shrimp nets off South Carolina. North American Journal of Fisheries Management, 14(1):178– 195, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schisler:1996:PHM

[SB96] George J. Schisler and Eric P. Bergersen. Postrelease hooking mortality of rainbow trout caught on scented artificial baits. North American Journal of Fisheries Management, 16(3):570– 578, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:1998:LSF

[SB98] Steve M. Sammons and Phillip W. Bettoli. Larval sampling as a fisheries management tool: Early detection of year-class strength. North American Journal of Fisheries Management, 18(1):137– 143, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:1999:STV

[SB99] Steve M. Sammons and Phillip W. Bettoli. Spatial and temporal variation in electrofishing catch rates of three species of black bass (Micropterus spp.) from Normandy Reservoir, Tennessee. North American Journal of Fisheries Management, 19(2):454– 461, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Savitz:1993:RCR

[SBF93] J. Savitz, L. G. Bardygula, and G. Funk. Returns of cage-released and non-cage-released Chinook and Coho salmon to Illinois harbors of Lake Michigan. North American Journal of Fisheries Management, 13(3):550–557, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sprankle:1996:LRD

[SBH96] Kenneth Sprankle, John Boreman, and Jay B. Hestbeck. Loss rates for dorsal loop and internal anchor tags applied to striped bass. North American Journal of Fisheries Management, 16(2): 461–464, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sitar:1999:LTM

[SBJ^{+}99] Shawn P. Sitar, James R. Bence, James E. Johnson, Mark P. Ebener, and William W. Taylor. Lake trout mortality and abundance in Southern Lake Huron. North American Journal of Fisheries Management, 19(4):881–900, November 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slaney:1993:RHS

[SBT93] P. A. Slaney, L. Berg, and A. F. Tautz. Returns of hatchery steelhead relative to site of release below an upper-river hatchery. North American Journal of Fisheries Management, 13(3):558– 566, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scheirer:1991:EFF

[SC91] Jeffrey W. Scheirer and Daniel W. Coble. Effects of floy FD-67 anchor tags on growth and condition of northern pike. North American Journal of Fisheries Management, 11(3):369–373, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spangler:1992:LHF

[SC92] George R. Spangler and John J. Collins. Lake Huron fish community structure based on gill-net catches corrected for selectivity and encounter probability. North American Journal of Fisheries Management, 12(3):585–597, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schneider:1992:MBF

[Sch92] James C. Schneider. Management briefs: Field evaluations of 230-v AC electrofishing on mortality and growth of warmwater and coolwater fish. North American Journal of Fisheries Management, 12(1):253–256, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:1996:HMB

[Sch96a] D. J. Schill. Hooking mortality of bait-caught rainbow trout in an Idaho trout stream and a hatchery: Implications for specialregulation management. North American Journal of Fisheries Management, 16(2):348–356, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schloesser:1996:MUM

[Sch96b] Don W. Schloesser. Mitigation of unionid mortality caused by zebra mussel infestation: Cleaning of unionids. North American Journal of Fisheries Management, 16(4):942–946, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schneider:1999:DQB

[Sch99] James C. Schneider. Dynamics of quality bluegill populations in two Michigan lakes with dense vegetation. North American Journal of Fisheries Management, 19(1):97–109, February 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmidt:1998:ICD

[SCKF98] Dana C. Schmidt, Stan R. Carlson, Gary B. Kyle, and Bruce P. Finney. Influence of carcass-derived nutrients on sockeye salmon productivity of Karluk Lake, Alaska: Importance in the assessment of an escapement goal. North American Journal of Fisheries Management, 18(4):743–763, November 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharber:1994:REI

[SCS^{+}94] N. G. Sharber, S. W. Carothers, J. P. Sharber, J. C. De Vos, Jr., and D. A. House. Reducing electrofishing-induced injury of rainbow trout. North American Journal of Fisheries Management, 14(2):340–346, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharber:1995:REI

[SCS^{+}95] N. G. Sharber, S. W. Carothers, J. P. Sharber, J. C. De Vos, Jr., and D. A. House. Reducing electrofishing-induced injury of rainbow trout: Response to comment. North American Journal of Fisheries Management, 15(4):965–968, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stokesbury:1991:MJC

[SD91] Kevin D. E. Stokesbury and Michael J. Dadswell. Mortality of juvenile clupeids during passage through a tidal, low-head hydroelectric turbine at Annapolis Royal, Nova Scotia. North American Journal of Fisheries Management, 11(2):149–154, May 1991. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:1993:CPU

[SD93] Harold L. Schramm, Jr. and John A. Dennis. Characteristics and perceptions of users and nonusers of an urban fishery program in Lubbock, Texas. North American Journal of Fisheries Management, 13(2):210–216, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:1999:ERH

[SDBF99] Steve M. Sammons, Lawrence G. Dorsey, Phillip W. Bettoli, and Francis C. Fiss. Effects of reservoir hydrology on reproduction by largemouth bass and spotted bass in Normandy Reservoir, Tennessee. North American Journal of Fisheries Management, 19(1):78–88, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slawski:1998:FHI

[SE98] Thomas M. Slawski and Timothy J. Ehlinger. Fish habitat improvement in box culverts: Management in the dark? North American Journal of Fisheries Management, 18(3):676–685, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sterritt:1994:EAI

[SES94] David A. Sterritt, Steven T. Elliott, and Artwin E. Schmidt. Electrical anesthesia for immobilizing adult Coho salmon in freshwater. North American Journal of Fisheries Management, 14(2): 453–456, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smeltzer:1991:CHF

[SF91] John F. Smeltzer and Stephen A. Flickinger. Culture, handling, and feeding techniques for black crappie fingerlings. North American Journal of Fisheries Management, 11(4):485–491, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schwartzberg:1993:IHN

[SF93] Matthew Schwartzberg and Jeffrey K. Fryer. Identification of hatchery and naturally spawning stocks of Columbia Basin spring Chinook salmon by scale pattern analyses. North American Journal of Fisheries Management, 13(2):263–271, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steimle:1996:IAR

[SF96] Frank W. Steimle and William Figley. The importance of artificial reef epifauna to black sea bass diets in the Middle Atlantic Bight. North American Journal of Fisheries Management, 16(2): 433–439, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schram:1998:LLS

[SF98] Stephen T. Schram and Mary C. Fabrizio. Longevity of Lake Superior lake trout. North American Journal of Fisheries Management, 18(3):700–703, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sorenson:1998:TPJ

[SFZ98] Kent M. Sorenson, William L. Fisher, and Alexander V. Zale. Turbine passage of juvenile and adult fish at a warmwater hydroelectric facility in Northeastern Oklahoma: Monitoring associated with relicensing. North American Journal of Fisheries Management, 18(1):124–136, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1993:IPA

[SG93] Stephen J. Smith and Stratis Gavaris. Improving the precision of abundance estimates of Eastern Scotian Shelf Atlantic cod from bottom trawl surveys. North American Journal of Fisheries Management, 13(1):35–47, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stauffer:1997:CTQ

[SG97a] Joseph C. Stauffer and Robert M. Goldstein. Comparison of three qualitative habitat indices and their applicability to prairie streams. North American Journal of Fisheries Management, 17 (2):348–361, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swanberg:1997:EIT

[SG97b] Tim R. Swanberg and David R. Geist. Effects of intraperitoneal transmitters on the social interaction of rainbow trout. North American Journal of Fisheries Management, 17(1):178– 181, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steeger:1994:BDM

[SGWN94] Thomas M. Steeger, John M. Grizzle, Ken Weathers, and Mike Newman. Bacterial diseases and mortality of angler-caught largemouth bass released after tournaments on Walter F. George Reservoir, Alabama–Georgia. North American Journal of Fisheries Management, 14(2):435–441, May 1994. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Somer:1992:ESD

[SH92a] William L. Somer and Thomas J. Hassler. Effects of suctiondredge gold mining on benthic invertebrates in a Northern California stream. North American Journal of Fisheries Management, 12(1):244–252, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stoeckel:1992:RSI

[SH92b] Joseph N. Stoeckel and Roy C. Heidinger. Relative susceptibility of inland and brook silversides to capture by largemouth bass. North American Journal of Fisheries Management, 12(3):499– 503, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simonson:1999:TWM

[SH99] Timothy D. Simonson and Steven W. Hewett. Trends in Wisconsin’s muskellunge fishery. North American Journal of Fisheries Management, 19(1):291–299, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepherd:1991:MMV

[She91] Gary Shepherd. Meristic and morphometric variation in black sea bass north of Cape Hatteras, North Carolina. North American Journal of Fisheries Management, 11(2):139–148, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schweigert:1990:ESS

[SHS90] J. F. Schweigert, C. W. Haegele, and M. Stocker. Evaluation of sampling strategies for scuba surveys to assess spawn deposition by Pacific herring. North American Journal of Fisheries Management, 10(2):185–195, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stoneman:1996:SMC

[SJ96] Christine L. Stoneman and Michael L. Jones. A simple method to classify stream thermal stability with single observations of daily maximum water and air temperatures. North American Journal of Fisheries Management, 16(4):728–737, November 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Siegwarth:1998:ASC

[SJ98a] G. L. Siegwarth and J. E. Johnson. Assessment of stocking catchable-size channel catfish in the Buffalo River, Arkansas. North American Journal of Fisheries Management, 18(1):96–103, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanfield:1998:CFS

[SJ98b] Les W. Stanfield and Michael L. Jones. A comparison of fullstation visual and transect-based methods of conducting habitat surveys in support of habitat suitability index models for Southern Ontario. North American Journal of Fisheries Management, 18(3):657–675, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sweatman:1991:VDO

[SK91] James J. Sweatman and Christopher C. Kohler. Validation of daily otolith increments for young-of-the-year white crappies. North American Journal of Fisheries Management, 11(4):499– 503, November 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:1995:URR

[SK95] D. I. Schill and P A. Kline. Use of random response to estimate angler noncompliance with fishing regulations. North American Journal of Fisheries Management, 15(4):721–731, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stocker:1990:EEJ

[SL90] M. Stocker and B. M. Leaman. Effects of extended jurisdiction on management and conservation of Canada’s Pacific Coast fisheries. North American Journal of Fisheries Management, 10(2):144– 153, May 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simonson:1995:CCP

[SL95] Timothy D. Simonson and John Lyons. Comparison of catch per effort and removal procedures for sampling stream fish assemblages. North American Journal of Fisheries Management, 15(2):419–427, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schram:1999:RLS

[SLE99] Stephen T. Schram, John Lindgren, and Lori M. Evrard. Reintroduction of lake sturgeon in the St. Louis River, Western Lake Superior. North American Journal of Fisheries Management, 19 (3):815–823, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Springborn:1998:TDM

[SLG98] Robert Richard Springborn, Nickolas D. Lampsakis, and Vincent F. Gallucci. A time density model to estimate run size and entry timing in a salmon fishery. North American Journal of Fisheries Management, 18(2):391–405, May 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steneck:1997:RIB

[SLJ^{+}97] Robert S. Steneck, Richard W. Langton, Francis Juanes, Vytenis Gotceitas, and Peter Lawton. Response: The interface between fisheries research and habitat management. North American Journal of Fisheries Management, 17(2):596–598, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simonson:1994:QFH

[SLK94] Timothy D. Simonson, John Lyons, and Paul D. Kanehl. Quantifying fish habitat in streams: Transect spacing, sample size, and a proposed framework. North American Journal of Fisheries Management, 14(3):607–615, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saila:1997:EAE

[SLM^{+}97] Saul B. Saila, Ernesto Lorda, J. Dale Miller, Ronald A. Sher, and W. Huntting Howell. Equivalent adult estimates for losses of fish eggs, larvae, and juveniles at Seabrook Station with use of fuzzy logic to represent parametric uncertainty. North American Journal of Fisheries Management, 17(4):811–825, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schrank:1994:EMW

[SM94] W. E. Schrank and R. L. Mazany. Econometric modeling of world trade in groundfish: Progress report. North American Journal of Fisheries Management, 14(2):291–296, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saiki:1996:VHP

[SM96] Michael K. Saiki and Barbara A. Martin. Vacuum hand pump apparatus for collecting water samples from a horizontal intragravel pipe. North American Journal of Fisheries Management, 16(1): 231–232, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shroyer:1998:FAQ

[SM98] Steven M. Shroyer and Thomas S. McComish. Forecasting abundance of quality-size yellow perch in Indiana waters of Lake Michigan. North American Journal of Fisheries Management, 18 (1):19–24, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:1992:EPB

[SMH92] Harold L. Schramm, Jr., Stephen P. Malvestuto, and Wayne A. Hubert. Evaluation of procedures for back-calculation of lengths of largemouth bass aged by otoliths. North American Journal of Fisheries Management, 12(3):604–608, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schorr:1995:ICS

[SMH95] Mark S. Schorr, Michael R. Meador, and Loren G. Hill. Incidental catch of sport fish in cast nets used to collect baitfish in Lake Texoma, Oklahoma–Texas. North American Journal of Fisheries Management, 15(1):142–147, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1990:RSI

[Smi90] Courtland L. Smith. Resource scarcity and inequality in the distribution of catch. North American Journal of Fisheries Management, 10(3):269–278, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1998:UTP

[Smi98] Lance W. Smith. Use of traditional practices and knowledge in monitoring a Lake Malaŵi artisanal fishery. North American Journal of Fisheries Management, 18(4):982–988, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shaner:1996:ACS

[SMMC96] Brian L. Shaner, Michael J. Maceina, James J. McHugh, and Stanley F. Cook. Assessment of catfish stocking in public fishing lakes in Alabama. North American Journal of Fisheries Management, 16(4):880–887, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1995:FQP

[SMTD95] Stephen M. Smith, Michael J. Maceina, Vincent H. Travnichek, and Rex A. Dunham. Failure of quantitative phenotypic characteristics to distinguish black crappie, white crappie, and their first-generation hybrid. North American Journal of Fisheries Management, 15(1):121–125, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slipke:1998:EMM

[SMTW98] Jeffrey W. Slipke, Michael J. Maceina, Vincent H. Travnichek, and Kenneth C. Weathers. Effects of a 356-mm minimum length limit on the population characteristics and sport fishery of smallmouth bass in the shoals reach of the Tennessee River, Alabama. North American Journal of Fisheries Management, 18(1):76–84, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Storck:1992:CTM

[SN92] Ted W. Storck and Dennis L. Newman. Contribution of tiger muskellunge to the sport fishery of a small, centrarchiddominated impoundment. North American Journal of Fisheries Management, 12(1):213–221, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shuler:1994:DHS

[SNF94] Scot W. Shuler, R. Barry Nehring, and Kurt D. Fausch. Diel habitat selection by brown trout in the Rio Grande River, Colorado, after placement of boulder structures. North American Journal of Fisheries Management, 14(1):99–111, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sipponen:1999:APF

[SNS99] Matti Sipponen, Hannu Niittykangas, and Hannu Salo. The adaptation of professional fishing to diminished vendace stocks: The case of Finnish inland waters. North American Journal of Fisheries Management, 19(3):737–747, August 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snucins:1992:RSH

[Snu92] E. J. Snucins. Relative survival of hatchery-reared lake trout, brook trout, and F_{1} splake stocked in low-pH lakes. North American Journal of Fisheries Management, 12(3):460–464, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Staggs:1996:FAF

[SO96] Michael D. Staggs and Keith J. Otis. Factors affecting first-year growth of fishes in Lake Winnebago, Wisconsin. North American Journal of Fisheries Management, 16(3):608–618, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sorensen:1994:MBE

[Sor94] Peter W. Sorensen. Management briefs: Effects of electroshocking on the sexual behavior of goldfish and brook trout. North American Journal of Fisheries Management, 14(4):862–865, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stefferud:1996:LCF

[SP96] Jerome A. Stefferud and David L. Propst. A lightweight, constant-flow device for dispensing liquid piscicides into streams in remote areas. North American Journal of Fisheries Management, 16(1):228–230, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spencer:1993:FIS

[Spe93] Paul D. Spencer. Factors influencing satisfaction of anglers on Lake Miltona, Minnesota. North American Journal of Fisheries Management, 13(2):201–209, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepard:1996:FIR

[SRCIW96] Bradley B. Shepard, Jim Robison-Cox, Susan C. Ireland, and Robert G. White. Factors influencing retention of visible implant tags by westslope cutthroat trout inhabiting headwater streams of Montana. North American Journal of Fisheries Management, 16(4):913–920, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarnecchia:1997:CTW

[SRL97] Dennis L. Scarnecchia, L. Fred Ryckman, and Jason Lee. Capturing and tagging wild age-0 and age-1 paddlefish in a Great Plains reservoir. North American Journal of Fisheries Management, 17(3):800–802, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1991:OFL

[SS91] David R. Smith and John F. Smeltzer. Occurrence of fishing without a license on the Cache la Poudre River, Colorado. North American Journal of Fisheries Management, 11(1):109– 112, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spencer:1992:EPF

[SS92] Paul D. Spencer and George R. Spangler. Effect that providing fishing information has on angler expectations and satisfaction. North American Journal of Fisheries Management, 12(2):379– 385, May 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarnecchia:1997:IEC

[SS97a] Dennis L. Scarnecchia and Phillip A. Stewart. Implementation and evaluation of a catch-and-release fishery for paddlefish. North American Journal of Fisheries Management, 17(3):795–799, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:1997:BHR

[SS97b] D. J. Schill and R. L. Scarpella. Barbed hook restrictions in catch-and-release trout fisheries: a social issue. North American Journal of Fisheries Management, 17(4):873–881, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shasteen:1997:LEA

[SS97c] Scott P. Shasteen and Robert J. Sheehan. Laboratory evaluation of artificial swim bladder deflation in largemouth bass: Potential benefits for catch-and-release fisheries. North American Journal of Fisheries Management, 17(1):32–37, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stefanik:1999:DSE

[SS99] Elliott L. Stefanik and Mark B. Sandheinrich. Differences in spawning and emergence phenology between stocked and wild populations of brown trout in Southwestern Wisconsin streams. North American Journal of Fisheries Management, 19(4):1112– 1116, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1999:EIA

[SSA99] Barry D. Smith, Thomas F. Shardlow, and A. W. Argue. An error identification approach to forecasting the harvest rate of one species in a multispecies fishery, illustrated with two salmon fisheries. North American Journal of Fisheries Management, 19 (2):421–435, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simon:1993:FPT

[SSD93] David C. Simon, Charles G. Scalet, and Jeff C. Dillon. Field performance of triploid and diploid rainbow trout in South Dakota ponds. North American Journal of Fisheries Management, 13 (1):134–140, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarnecchia:1996:PRP

[SSL96] Dennis L. Scarnecchia, Phillip A. Stewart, and Youngtaik Lim. Profile of recreational paddlefish snaggers on the Lower Yellowstone River, Montana. North American Journal of Fisheries Management, 16(4):872–879, November 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swanberg:1999:CSS

[SSM99] Tim R. Swanberg, David A. Schmetterling, and David H. McEvoy. Comparison of surgical staples and silk sutures for closing incisions in rainbow trout. North American Journal of Fisheries Management, 19(1):215–218, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepard:1997:SRE

[SSUL97] Bradley B. Shepard, Brian Sanborn, Linda Ulmer, and Danny C. Lee. Status and risk of extinction for westslope cutthroat trout in the Upper Missouri River Basin, Montana. North American Journal of Fisheries Management, 17(4):1158–1172, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanley:1992:BCC

[Sta92] R. D. Stanley. Bootstrap calculation of catch-per-unit-effort variance from trawl logbooks: Do fisheries generate enough observations for stock assessments? North American Journal of Fisheries Management, 12(1):19–27, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steinback:1999:REI

[Ste99] Scott R. Steinback. Regional economic impact assessments of recreational fisheries: An application of the IMPLAN modeling system to marine party and charter boat fishing in Maine. North American Journal of Fisheries Management, 19(3):724–736, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stahl:1996:FAS

[STS^{+}96] Thomas P. Stahl, Gary P. Thiede, Roy A. Stein, Edward M. Lewis, Milton R. Austin, and David A. Culver. Factors affecting survival of age-0 saugeye Stizostedion vitreum × S. canadense stocked in Ohio reservoirs. North American Journal of Fisheries Management, 16(2):378–387, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stables:1990:ERE

[STT^{+}90] T. Brock Stables, Gary L. Thomas, Steven L. Thiesfeld, Gilbert B. Pauley, and Michael A. Wert. Effects of reservoir enlargement and other factors on the yield of wild rainbow and cutthroat trout in Spada Lake, Washington. North American Journal of Fisheries Management, 10(3):305–314, August 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stuecheli:1991:TBS

[Stu91] Ken Stuecheli. Trapping bias in sampling crayfish with baited funnel traps. North American Journal of Fisheries Management, 11(2):236–239, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Szendrey:1995:EFE

[SW95] Thomas A. Szendrey and David H. Wahl. Effect of feeding experience on growth, vulnerability to predation, and survival of esocids. North American Journal of Fisheries Management, 15 (3):610–620, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Szendrey:1996:SSS

[SW96] Thomas A. Szendrey and David H. Wahl. Size-specific survival and growth of stocked muskellunge: Effects of predation and prey availability. North American Journal of Fisheries Management, 16(2):395–402, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Somers:1997:SME

[SW97] Ian Somers and You-Gan Wang. A simulation model for evaluating seasonal closures in Australia’s multispecies northern prawn fishery. North American Journal of Fisheries Management, 17 (1):114–130, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swink:1999:EEB

[Swi99] William D. Swink. Effectiveness of an electrical barrier in blocking a sea lamprey spawning migration on the Jordan River, Michigan. North American Journal of Fisheries Management, 19(2):

397–405, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Santucci:1994:GMH

[SWS94] Victor J. Santucci, Jr., David H. Wahl, and Ted W. Storck. Growth, mortality, harvest, and cost-effectiveness of stocked channel catfish in a small impoundment. North American Journal of Fisheries Management, 14(4):781–789, November 1994. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sylvia:1994:MIF

[Syl94] Gilbert Sylvia. Market information and fisheries management: a multiple-objective analysis. North American Journal of Fisheries Management, 14(2):278–290, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thorn:1997:RTM

[TAL^{+}97] William C. Thorn, Charles S. Anderson, William E. Lorenzen, Deserae L. Hendrickson, and James W. Wagner. A review of trout management in Southeast Minnesota streams. North American Journal of Fisheries Management, 17(4):860–872, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:1992:LTO

[Tay92] David M. Taylor. Long-term observations on movements of tagged male snow crabs in Bonavista Bay, Newfoundland. North American Journal of Fisheries Management, 12(4):777–782, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:1996:PAH

[TB96] Kenneth E. Tiffan and Eric P. Bergersen. Performance of antimycin in high-gradient streams. North American Journal of Fisheries Management, 16(2):465–468, May 1996. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:1997:RAF

[TB97a] Daniel A. Thompson and H. Lee Blankenship. Regeneration of adipose fins given complete and incomplete clips. North American Journal of Fisheries Management, 17(2):467–469, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Turek:1997:CTM

[TB97b] Stephen M. Turek and Michael T. Brett. Comment: Trout mortality from baited barbed and barbless hooks. North American Journal of Fisheries Management, 17(3):807, August 1997. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trometer:1999:CAF

[TB99] E. S. Trometer and W-D. N. Busch. Changes in age-0 fish growth and abundance following the introduction of zebra mussels Dreissena polymorpha in the Western Basin of Lake Erie. North American Journal of Fisheries Management, 19(2):604–609, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Teisl:1993:AOR

[TBF93] Mario F. Teisl, Kevin J. Boyle, and Owen C. Fenderson. Angler opinions regarding management options to balance openwater and ice fishing effort in Maine. North American Journal of Fisheries Management, 13(2):353–359, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:1997:IBT

[TBN97] Kevin G. Thompson, Eric P. Bergersen, and R. Barry Nehring. Injuries to brown trout and rainbow trout induced by capture with pulsed direct current. North American Journal of Fisheries Management, 17(1):141–153, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:1997:LTE

[TBNB97] Kevin G. Thompson, Eric P. Bergersen, R. Barry Nehring, and David C. Bowden. Long-term effects of electrofishing on growth and body condition of brown trout and rainbow trout. North American Journal of Fisheries Management, 17(1):154– 159, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Teisl:1996:CAA

[TBR96] Mario F. Teisl, Kevin J. Boyle, and Brian Roe. Conjoint analysis of angler evaluations of Atlantic salmon restoration on the Penobscot River, Maine. North American Journal of Fisheries Management, 16(4):861–871, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tipping:1996:SEA

[TG96] Jack M. Tipping and Gary J. Gilhuly. Survival of electroanesthetized adult steelhead and eggs of fall Chinook salmon. North American Journal of Fisheries Management, 16(2):469–472, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Twedt:1992:SDL

[TGF92] Daniel J. Twedt, W. Clell Guest, and Bobby W. Farquhar. Selective dipnetting of largemouth bass during electrofishing. North American Journal of Fisheries Management, 12(3):609–611, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tillma:1998:RAH

[TGM98] Jeff S. Tillma, Christopher S. Guy, and Chris S. Mammoliti. Relations among habitat and population characteristics of spotted bass in Kansas streams. North American Journal of Fisheries Management, 18(4):886–893, November 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:1990:ISM

[TH90] Thomas Thompson and Wayne A. Hubert. Influence of survey method on estimates of statewide fishing activity. North American Journal of Fisheries Management, 10(1):111–113, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tipping:1993:UMW

[TH93] Jack M. Tipping and Jackie R. Heinricher. Use of magnetic wire tag locations to mark tiger muskellunge. North American Journal of Fisheries Management, 13(1):190–193, February 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Timmons:1995:RAS

[TH95] Tom J. Timmons and Mark H. Howell. Retention of anchor and spaghetti tags by paddlefish, catfishes, and buffalo fishes. North American Journal of Fisheries Management, 15(2):504–506, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thomas:1999:CLS

[TH99] Michael V. Thomas and Robert C. Haas. Capture of lake sturgeon with setlines in the St. Clair River, Michigan. North American Journal of Fisheries Management, 19(2):610–612, May 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tipping:1991:MBH

[Tip91] Jack M. Tipping. Management briefs: Heritability of age at maturity in steelhead. North American Journal of Fisheries Management, 11(1):105–108, February 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thurow:1997:DSS

[TLR97] Russell F. Thurow, Danny C. Lee, and Bruce E. Rieman. Distribution and status of seven native salmonids in the interior Columbia River Basin and portions of the Klamath River and Great Basins. North American Journal of Fisheries Management, 17(4):1094–1110, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tabor:1996:EDA

[TLW96] Roger Tabor, Chris Luecke, and Wayne Wurtsbaugh. Effects of Daphnia availability on growth and food consumption of rainbow trout in two Utah reservoirs. North American Journal of Fisheries Management, 16(3):591–599, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tarrant:1993:ERB

[TMBH93] Michael A. Tarrant, Michael J. Manfredo, Peter B. Bayley, and Richard Hess. Effects of recall bias and nonresponse bias on selfreport estimates of angling participation. North American Journal of Fisheries Management, 13(2):217–222, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thedinga:1994:DSS

[TMJ^{+}94] John F. Thedinga, Michael L. Murphy, Scott W. Johnson, J. Michael Lorenz, and K V. Koski. Determination of salmonid smolt yield with rotary-screw traps in the Situk River, Alaska, to predict effects of glacial flooding. North American Journal of Fisheries Management, 14(4):837–851, November 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Townsend:1994:ERU

[TP94] Ralph E. Townsend and Samuel G. Pooley. Effort reduction under limited entry: Attrition versus fractional licenses in the Hawaii longline fleet. North American Journal of Fisheries Management, 14(2):297–306, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Teirney:1992:ACA

[TR92] L. D. Teirney and J. Richardson. Attributes that characterize angling rivers of importance in New Zealand, based on angler use and perceptions. North American Journal of Fisheries Management, 12(4):693–702, November 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:1996:EDR

[TR96] Paul D. Thompson and Frank J. Rahel. Evaluation of depletionremoval electrofishing of brook trout in small Rocky Mountain streams. North American Journal of Fisheries Management, 16 (2):332–339, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:1998:EAB

[TR98] Paul D. Thompson and Frank J. Rahel. Evaluation of artificial barriers in small Rocky Mountain streams for preventing the upstream movement of brook trout. North American Journal of Fisheries Management, 18(1):206–210, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trippel:1993:RFM

[Tri93] Edward A. Trippel. Relations of fecundity, maturation, and body size of lake trout, and implications for management in Northwestern Ontario lakes. North American Journal of Fisheries Management, 13(1):64–72, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thurow:1996:CDS

[TS96] Russell E. Thurow and Daniel J. Schill. Comparison of day snorkeling, night snorkeling, and electrofishing to estimate bull trout abundance and size structure in a second-order Idaho stream. North American Journal of Fisheries Management, 16 (2):314–323, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyus:1999:EIT

[TSKS99] Harold M. Tyus, Wayne C. Starnes, Catherine A. Karp, and James F. Saunders III. Effects of invasive tissue collection on rainbow trout, razorback sucker, and bonytail chub. North American Journal of Fisheries Management, 19(3):848–855, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tabor:1993:PJS

[TSP93] Roger A. Tabor, Rip S. Shively, and Thomas P. Poe. Predation on juvenile salmonids by smallmouth bass and northern squawfish in the Columbia River near Richland, Washington. North American Journal of Fisheries Management, 13(4):831–838, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:1992:MAH

[TW92] Matthew J. Taylor and Karl R. White. A meta-analysis of hooking mortality of nonanadromous trout. North American Journal of Fisheries Management, 12(4):760–767, November 1992. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:1997:RTM

[TW97] Matthew J. Taylor and Karl R. White. Response: Trout mortality from baited barbed and barbless hooks. North American Journal of Fisheries Management, 17(3):808–809, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiersch:1997:FCH

[TWF^{+}97] Terrence R. Tiersch, William R. Wayman, Chester R. Figiel, Jr., Owen T. Gorman, J. Holt Williamson, and Gary J. Carmichael. Field collection, handling, and storage of sperm of the endangered razorback sucker. North American Journal of Fisheries Management, 17(1):167–173, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Unwin:1997:MBA

[UHL97] M. J. Unwin, J. T. Hill, and D. H. Lucas. Management briefs: An accurate and efficient method for reading sequentially coded wire tags aided by a PC-Based data capture program. North American Journal of Fisheries Management, 17(2):446–450, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Unmuth:1999:EMH

[UHP99] Jean M. L. Unmuth, Michael J. Hansen, and Thomas D. Pellett. Effects of mechanical harvesting of Eurasian watermilfoil on largemouth bass and bluegill populations in Fish Lake, Wisconsin. North American Journal of Fisheries Management, 19(4): 1089–1098, November 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Uphoff:1993:DSB

[Uph93] James H. Uphoff, Jr. Determining striped bass spawning stock status from the presence or absence of eggs in ichthyoplankton survey data. North American Journal of Fisheries Management, 13(4):645–656, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Uphoff:1997:UEP

[Uph97] James H. Uphoff, Jr. Use of egg presence–absence to derive probability-based management criteria for Upper Chesapeake Bay striped bass. North American Journal of Fisheries Management, 17(3):663–676, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Verrill:1995:EEB

[VB95] Donovan D. Verrill and Charles R. Berry, Jr. Effectiveness of an electrical barrier and lake drawdown for reducing common carp and bigmouth buffalo abundances. North American Journal of Fisheries Management, 15(1):137–141, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDenAvyle:1995:CCR

[VBM^{+}95] Michael J. Van Den Avyle, Jeff Boxrucker, Paul Michaletz, Bruce Vondracek, and Gene R. Ploskey. Comparison of catch rate, length distribution, and precision of six gears used to sample reservoir shad populations. North American Journal of Fisheries Management, 15(4):940–955, November 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vondracek:1995:AWT

[VD95] Bruce Vondracek and Donald J. Degan. Among- and withintransect variability in estimates of shad abundance made with hydroacoustics. North American Journal of Fisheries Management, 15(4):933–939, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

vandenAvyle:1990:TSS

[vdAE90] Michael J. van den Avyle and James W. Evans. Temperature selection by striped bass in a Gulf of Mexico coastal river system. North American Journal of Fisheries Management, 10(1):58–66, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanZee:1996:CCO

[VHW^{+}96] Brian E. Van Zee, Tracy D. Hill, David W. Willis, Lewis F. Brown, James B. Reynolds, Norman G. Sharber, and John Sharber. Comment: Clarification of the outputs from a coffelt VVP-15 electrofisher. North American Journal of Fisheries Management, 16(2):477–478, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanOffelen:1993:SGM

[VKS93] Henry K. Van Offelen, Charles C. Krueger, and Carl L. Schofield. Survival, growth, movement, and distribution of two brook trout strains stocked into small Adirondack streams. North American Journal of Fisheries Management, 13(1):86–95, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDenAvyle:1995:EGN

[VPB95] Michael J. Van Den Avyle, Gene R. Ploskey, and Phillip W. Bettoli. Evaluation of gill-net sampling for estimating abundance and length frequency of reservoir shad populations. North American Journal of Fisheries Management, 15(4):898–917, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Viola:1995:MRA

[VS95] Arthur E. Viola and Mark L. Schuck. A method to reduce the abundance of residual hatchery steelhead in rivers. North American Journal of Fisheries Management, 15(2):488–493, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDoornik:1999:ASP

[VWT99] Donald M. Van Doornik, Gary A. Winans, and David J. Teel. Allozyme studies of Pacific salmonids with nonlethal sampling of fin tissue. North American Journal of Fisheries Management, 19 (3):678–686, August 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wahl:1999:ECE

[Wah99] David H. Wahl. An ecological context for evaluating the factors influencing muskellunge stocking success. North American Journal of Fisheries Management, 19(1):238–248, February 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walsh:1992:SDS

[Wal92] Stephen J. Walsh. Size-dependent selection at the footgear of a groundfish survey trawl. North American Journal of Fisheries Management, 12(3):625–633, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waters:1992:APP

[Wat92] Thomas F. Waters. Annual production, production/biomass ratio, and the ecotrophic coefficient for management of trout in streams. North American Journal of Fisheries Management, 12 (1):34–39, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waite:1992:HCR

[WB92a] Ian R. Waite and Roger A. Barnhart. Habitat criteria for rearing steelhead: a comparison of site-specific and standard curves for use in the instream flow incremental methodology. North American Journal of Fisheries Management, 12(1):40–46, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weathers:1992:SBS

[WB92b] Kenneth C. Weathers and Mark B. Bain. Smallmouth bass in the shoals reach of the Tennessee River: Population characteristics and sport fishery. North American Journal of Fisheries Management, 12(3):528–537, August 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weisberg:1993:EFF

[WB93] Stephen B. Weisberg and William H. Burton. Enhancement of fish feeding and growth after an increase in minimum flow below the Conowingo Dam. North American Journal of Fisheries Management, 13(1):103–109, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Withler:1997:SIP

[WBMM97] Ruth E. Withler, Terry D. Beacham, Tobi J. Ming, and Kristina M. Miller. Species identification of Pacific salmon by means of a major histocompatibility complex gene. North American Journal of Fisheries Management, 17(4):929–938, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Westrheim:1992:SDD

[WBPQ92] S. Jergen Westrheim, William H. Barss, Ellen K. Pikitch, and Lawrence F. Quirollo. Stock delineation of Dover sole in the California–British Columbia region, based on tagging studies conducted during 1948–1979. North American Journal of Fisheries Management, 12(1):172–181, February 1992. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:1997:RAT

[WBYO97] David L. Ward, Raymond R. Boyce, Franklin R. Young, and Frederick E. Olney. A review and assessment of transportation studies for juvenile Chinook salmon in the Snake River. North American Journal of Fisheries Management, 17(3):652–662, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:1994:MOA

[WD94] Gene R. Wilde and Robert B. Ditton. A management-oriented approach to understanding diversity among largemouth bass anglers. North American Journal of Fisheries Management, 14 (1):34–40, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Warren:1995:DTS

[WD95] W. G. Warren and J. B. Dempson. Does temporal stratification improve the accuracy of mark–recapture estimates of smolt production? A case study based on the Conne River, Newfoundland. North American Journal of Fisheries Management, 15(1): 126–136, February 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waldman:1991:MBL

[WDM91] John R. Waldman, Dennis J. Dunning, and Mark T. Mattson. Management briefs: Long-term retention of anchor tags and internal anchor tags by striped bass. North American Journal of Fisheries Management, 11(2):232–234, May 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watson:1993:CST

[WDR93] Reg A. Watson, David J. Die, and Victor R. Restrepo. Closed seasons and tropical penaeid fisheries: a simulation including fleet dynamics and uncertainty. North American Journal of Fisheries Management, 13(2):326–336, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weaver:1993:MME

[WF93] Thomas M. Weaver and John J. Fraley. A method to measure emergence success of westslope cutthroat trout fry from varying substrate compositions in a natural stream channel. North American Journal of Fisheries Management, 13(4):817–822, November 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilkerson:1997:SBD

[WF97] Michael L. Wilkerson and William L. Fisher. Striped bass distribution, movements, and site fidelity in Robert S. Kerr Reservoir, Oklahoma. North American Journal of Fisheries Management, 17(3):677–686, August 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walters:1997:CCR

[WFB97] Jody P. Walters, Tom D. Fresques, and Scott D. Bryan. Comparison of creel returns from rainbow trout stocked at two sizes. North American Journal of Fisheries Management, 17(2):474– 476, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wenburg:1995:PVI

[WG95] John K. Wenburg and Gordon W. George. Placement of visible implant tags in the anal fin of wild coastal cutthroat trout. North American Journal of Fisheries Management, 15(4):874– 877, November 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiens:1996:RSW

[WGB96] Jennifer R. Wiens, Christopher S. Guy, and Michael L. Brown. A revised standard weight (WS) equation for spotted bass. North American Journal of Fisheries Management, 16(4):958– 959, November 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1991:DES

[WGM91] David W. Willis, Christopher S. Guy, and Brian R. Murphy. Development and evaluation of a standard weight (WS) equation for yellow perch. North American Journal of Fisheries Management, 11(3):374–380, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walters:1993:RIS

[WGR93] Carl Walters, R. D. Goruk, and Donald Radford. Rivers Inlet sockeye salmon: An experiment in adaptive management. North American Journal of Fisheries Management, 13(2):253–262, May 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winton:1994:LSC

[WH94] John Winton and Ray Hilborn. Lessons from supplementation of Chinook salmon in British Columbia. North American Journal of Fisheries Management, 14(1):1–13, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walker:1996:AAS

[WH96] Randal L. Walker and Peter B. Heffernan. Adult age-specific embryonic viability of the northern quahog. North American Journal of Fisheries Management, 16(3):633–639, August 1996. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watson:1997:FAD

[WH97] Greg Watson and T. W. Hillman. Factors affecting the distribution and abundance of bull trout: An investigation at hierarchical scales. North American Journal of Fisheries Management, 17(2): 237–252, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whittier:1998:EFS

[WH98] Thomas R. Whittier and Robert M. Hughes. Evaluation of fish species tolerances to environmental stressors in lakes in the Northeastern United States. North American Journal of Fisheries Management, 18(2):236–252, May 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whaley:1991:ITC

[Wha91] Roy A. Whaley. An improved technique for cleaning fish scales. North American Journal of Fisheries Management, 11(2):234– 236, May 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winkle:1990:RBB

[WHR90] Paul L. Winkle, Wayne A. Hubert, and Frank J. Rahel. Relations between brook trout standing stocks and habitat features in Beaver ponds in Southeastern Wyoming. North American Journal of Fisheries Management, 10(1):72–79, February 1990. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wichert:1995:EIS

[Wic95] Gordon A. Wichert. Effects of improved sewage effluent management and urbanization on fish associations of Toronto streams. North American Journal of Fisheries Management, 15(2):440– 456, May 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:1999:SDA

[Wil99] John G. Williams. Stock dynamics and adaptive management of habitat: An evaluation based on simulations. North American Journal of Fisheries Management, 19(2):329–341, May 1999. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilderbuer:1998:AFP

[WKG98] T. K. Wilderbuer, R. F. Kappenman, and D. R. Gunderson. Analysis of fishing power correction factor estimates from a trawl comparison experiment. North American Journal of Fisheries Management, 18(1):11–18, February 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whalen:1998:SGU

[WL98] Kevin G. Whalen and George W. Labar. Survival and growth of unfed and fed Atlantic salmon fry stocked in a Vermont tributary of the Connecticut River. North American Journal of Fisheries Management, 18(4):931–935, November 1998. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:1997:LGE

[WLC97] Richard N. Williams, Robb F. Leary, and Kenneth P. Currens. Localized genetic effects of a long-term hatchery stocking program on resident rainbow trout in the Metolius River, Oregon. North American Journal of Fisheries Management, 17(4):1079– 1093, November 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walters:1991:HDI

[WLJ91] Debra A. Walters, William E. Lynch, Jr., and David L. Johnson. How depth and interstice size of artificial structures influence fish attraction. North American Journal of Fisheries Management, 11(3):319–329, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:1998:DEH

[WLK98] Lizhu Wang, John Lyons, and Paul Kanehl. Development and evaluation of a habitat rating system for low-gradient Wisconsin streams. North American Journal of Fisheries Management, 18 (4):775–785, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weathers:1997:EOP

[WN97] Kenneth C. Weathers and Mike I. Newman. Effects of organizational procedures on mortality of largemouth bass during summer tournaments. North American Journal of Fisheries Management, 17(1):131–135, February 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:1999:SVC

[WN99] Scott M. Ward and Robert M. Neumann. Seasonal variations in concentrations of mercury in axial muscle tissue of largemouth bass. North American Journal of Fisheries Management, 19 (1):89–96, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:1994:IWD

[WNFK94] David L. Ward, Anthony A. Nigro, Ruth A. Farr, and Christopher J. Knutsen. Influence of waterway development on migrational characteristics of juvenile salmonids in the Lower Willamette River, Oregon. North American Journal of Fisheries Management, 14(2):362–371, May 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Webb:1991:ELB

[WO91] Mark A. Webb and Richard A. Ott, Jr. Effects of length and bag limits on population structure and harvest of white crappies in three Texas reservoirs. North American Journal of Fisheries Management, 11(4):614–622, November 1991. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weise:1998:ICB

[WP98a] Jerry G. Weise and Tiina A. Pajos. Intraspecific competition between larval sea lamprey year-classes as Salem Creek was recolonized, 1990–1994, after a lampricide application. North American Journal of Fisheries Management, 18(3):561–568, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welsh:1998:ISS

[WP98b] Stuart A. Welsh and Sue A. Perry. Influence of spatial scale on estimates of substrate use by benthic darters. North American Journal of Fisheries Management, 18(4):954–959, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wallace:1996:CCC

[WPE96] John R. Wallace, Ellen K. Pikitch, and Daniel L. Erickson. Can changing cod end mesh size and mesh shape affect the nearshore trawl fishery off the West Coast of the United States? North American Journal of Fisheries Management, 16(3):530–539, August 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walker:1993:PDM

[WPTW93] Susan H. Walker, Mark W. Prout, William W. Taylor, and Scott R. Winterstein. Population dynamics and management of lake whitefish stocks in Grand Traverse Bay, Lake Michigan. North American Journal of Fisheries Management, 13(1):73–85, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:1998:DAF

[WRD98] Gene R. Wilde, Robin K. Riechers, and Robert B. Ditton. Differences in attitudes, fishing motives, and demographic characteristics between tournament and nontournament black bass anglers in Texas. North American Journal of Fisheries Management, 18(2):422–431, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wells:1991:DYY

[WRDY91] Alan W. Wells, Donna M. Randall, Dennis J. Dunning, and John R. Young. Dispersal of young-of-the-year hatchery striped bass in the Hudson River. North American Journal of Fisheries Management, 11(3):381–392, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wright:1992:GSR

[Wri92] Sam Wright. Guidelines for selecting regulations to manage openaccess fisheries for natural populations of anadromous and resident trout in stream habitats. North American Journal of Fisheries Management, 12(3):517–527, August 1992. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walker:1991:INQ

[WS91] Randal L. Walker and Stuart A. Stevens. Increasing northern quahog harvest by using closed shellfishing areas in coastal Georgia. North American Journal of Fisheries Management, 11(3): 267–276, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:1999:CTH

[WSF99] John G. Williams, Terence P. Speed, and William F. Forrest. Comment: Transferability of habitat suitability criteria. North American Journal of Fisheries Management, 19(2):623–625, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:1996:APS

[WSL96] Lizhu Wang, Timothy D. Simonson, and John Lyons. Accuracy and precision of selected stream habitat estimates. North American Journal of Fisheries Management, 16(2):340–347, May 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:1998:CTB

[WSOM98] Gene R. Wilde, Dan W. Strickland, Kenneth G. Ostrand, and Maurice I. Muoneke. Characteristics of Texas black bass fishing tournaments. North American Journal of Fisheries Management, 18(4):972–977, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wasowicz:1994:NTR

[WV94] Anthony Wasowicz and Richard A. Valdez. A nonlethal technique to recover gut contents of roundtail chub. North American Journal of Fisheries Management, 14(3):656–658, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waldman:1996:SBH

[WV96] John R. Waldman and Victor J. Vecchio. Selected biocharacteristics of hatchery-reared striped bass captured in New York ocean waters. North American Journal of Fisheries Management, 16 (1):14–23, February 1996. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welch:1993:PRA

[WvdABD93] Timothy J. Welch, Michael J. van den Avyle, Robert K. Betsill, and Elizabeth M. Driebe. Precision and relative accuracy of striped bass age estimates from otoliths, scales, and anal fin rays and spines. North American Journal of Fisheries Management, 13(3):616–620, August 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilson:1993:DCB

[WW93] Harold T. Wilson and Stephen B. Weisberg. Design considerations for beach seine surveys of striped bass. North American Journal of Fisheries Management, 13(2):376–382, May 1993. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:1994:CBW

[WW94] Lizhu Wang and Ray J. White. Competition between wild brown trout and hatchery greenback cutthroat trout of largely wild parentage. North American Journal of Fisheries Management, 14(3):475–487, August 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiley:1993:ASH

[WWSF93a] Robert W. Wiley, Roy A. Whaley, Joe B. Satake, and Mark Fowden. Assessment of stocking hatchery trout: a Wyoming perspective. North American Journal of Fisheries Management, 13(1): 160–170, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiley:1993:EPT

[WWSF93b] Robert W. Wiley, Roy A. Whaley, Joe B. Satake, and Mark Fowden. An evaluation of the potential for training trout in hatcheries to increase poststocking survival in streams. North American Journal of Fisheries Management, 13(1):171–177, February 1993. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:1998:ESP

[WWTB98] John D. Williams, Kirk O. Winemiller, Donald C. Taphorn, and Luis Balbas. Ecology and status of piscivores in guri, an oligotrophic tropical reservoir. North American Journal of Fisheries Management, 18(2):274–285, May 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waldman:1999:CAA

[WYL^{+}99] John R. Waldman, John R. Young, B. Paul Lindsay, Robert E. Schmidt, and Helena Andreyko. A comparison of alternative approaches to discriminate larvae of striped bass and white perch. North American Journal of Fisheries Management, 19(2):470– 481, May 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yoshiyama:1998:HAD

[YFM98] Ronald M. Yoshiyama, Frank W. Fisher, and Peter B. Moyle. Historical abundance and decline of Chinook salmon in the Central Valley region of California. North American Journal of Fisheries Management, 18(3):487–521, August 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:1999:SRC

[YHN99] Kyle A. Young, Scott G. Hinch, and Tom G. Northcote. Status of resident coastal cutthroat trout and their habitat twenty-five years after riparian logging. North American Journal of Fisheries Management, 19(4):901–911, November 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:1991:SMS

[YHW91] Michael K. Young, Wayne A. Hubert, and Thomas A. Wesche. Selection of measures of substrate composition to estimate survival to emergence of salmonids and to detect changes in stream substrates. North American Journal of Fisheries Management, 11(3):339–346, August 1991. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zabel:1997:MTT

[ZA97] Richard W. Zabel and James J. Anderson. A model of the travel time of migrating juvenile salmon, with an application to Snake River spring Chinook salmon. North American Journal of Fisheries Management, 17(1):93–100, February 1997. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zale:1994:MBE

[ZB94] Alexander V. Zale and Mark B. Bain. Management briefs: Estimating tag-reporting rates with postcards as tag surrogates. North American Journal of Fisheries Management, 14(1):208– 211, February 1994. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zigler:1995:MBP

[ZD95] Steven J. Zigler and Michael R. Dewey. Management briefs: Phototaxis of larval and juvenile northern pike. North American Journal of Fisheries Management, 15(3):651–653, August 1995. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zigler:1999:DMH

[ZDK99] Steven J. Zigler, Michael R. Dewey, and Brent C. Knights. Diel movement and habitat use by paddlefish in navigation Pool 8 of the Upper Mississippi River. North American Journal of Fisheries Management, 19(1):180–187, February 1999. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhang:1990:YPR

[ZG90] Chang Ik Zhang and Donald R. Gunderson. Yield per recruit and management recommendations for Alaska plaice in the Eastern Bering Sea. North American Journal of Fisheries Management, 10(1):1–10, February 1990. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zweifel:1999:BIB

[ZHR99] Richard D. Zweifel, Robert S. Hayward, and Charles F. Rabeni. Bioenergetics insight into black bass distribution shifts in Ozark border region streams. North American Journal of Fisheries Management, 19(1):192–197, February 1999. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zheng:1998:IFS

[ZM98] Jie Zheng and Ole A. Mathisen. Inseason forecasting of Southeastern Alaska pink salmon abundance based on sex ratios and commercial catch and effort data. North American Journal of Fisheries Management, 18(4):872–885, November 1998. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zorn:1995:RBH

[ZS95] Troy G. Zorn and Paul W. Seelbach. The relation between habitat availability and the short-term carrying capacity of a stream reach for smallmouth bass. North American Journal of Fisheries Management, 15(4):773–783, November 1995. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhou:1997:MER

[ZS97] Shijie Zhou and Thomas C. Shirley. A model expressing the relationship between catch and soak time for trap fisheries. North American Journal of Fisheries Management, 17(2):482–487, May 1997. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).